A Complete Bibliography of Publications in Computational Statistics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 January 2024 Version 1.08

Title word cross-reference

(L_{2}) [1567]. 2 [122]. 2 × 2 [757]. 3 [636, 122]. 8 [1359, 1358]. _{α} [296, 462]. α [1399, 887]. ARMA(p, q) [515]. B [438]. X̄ [947]. C_{L} [81]. C_{pmk} [1524]. D [881, 1179, 841]. ℓ_{0} [1485, 832]. ℓ_{1} [853]. F [1055, 227, 563, 13]. G [1146, 1291, 1534, 341, 57]. I_{s}(x) [653]. I × J [142]. K [810, 1191, 756, 807, 395, 445, 578, 341, 813, 1426, 581, 867, 910, 235, 419, 420, 718, 717, 1200]. k_{2} > 2 [408]. K_{ϕ} [827]. L [1555, 409, 564]. L^{2} [1172]. L_{1} [666, 56]. L_{1/2} [1462, 1463]. L_{2} [394]. L_{p} [6, 14]. M [671, 109, 353, 1330, 41, 724]. n [456]. P [276, 342, 916, 1505, 232, 1399, 562, 1349, 456, 469, 221]. P (X&lt; Y) [1150]. P (Y &lt; X) [736]. πps [318]. Q [30]. r [724].

R^{2} [260]. ρ [1326]. S [986, 171]. t [1470, 117, 586, 87, 262, 1392, 1401, 752, 1181, 540, 1189, 1583]. τ [343, 142, 1326]. U [171]. ε [938, 1580].

-and- [341]. -bumps [756]. -D [122]. -dimensional [841]. -discrepancy [1567]. -discrimination [6, 14]. -distances [1172]. -distribution [752, 13]. -divergence [827]. -estimates [353]. -estimation [109]. -estimator [986, 564]. -estimators [1330]. -fold [1426]. -independent [581]. -inverse [57]. -means [395, 578, 910, 419, 420, 1200, 810, 1191]. -median [666]. -model [117, 87]. -nearest [813]. -normal [1189]. -optimal [881, 1291, 1179]. -parameter [1359, 1358]. -permanents [887]. -procedure [30].

-process [1401]. -records [807, 867]. -sample [445, 235]. -spline [276]. -stable [1399]. -surrogates [832]. -tempered [1399]. -test [1055]. -torus [1349]. -values [916, 1505, 232, 562, 469, 221, 342]. -Wright [671].

11 [112]. 1998- [174].

2.0 [1548]. 2.0.0 [338]. 2.1.0 [351]. 2001 [263]. 2002 [371]. 2003 [1083]. 2004 [379]. 2006 [591]. 2010 [688]. 2011 [808, 815]. 2013 [1235]. 2015 [1140]. 2016 [1405]. 2017 [1545, 1215]. 2018 [1614]. 21 [508]. 250th [558].

3 [176]. 3-7908-1134-3 [176].

4 [172]. 4/1999 [177]. 40al [584]. 40estimators [1050]. 40hidden [872].

532 [177].

71 [508].

97g [117]. 97h [112].

ability [680]. abnormal [592]. Abrupt [1116, 1138]. absence [843]. Accelerated [1526, 207, 63, 1403, 1557, 952, 1370, 1194]. Acceleration [497, 1105, 938, 1580, 584]. accept [765]. accept-reject [765]. acceptance [515, 1298]. acceptance-rejection [1298]. access [521]. accident [1406]. account [1013]. Accuracy [1615, 1055, 1278, 1620, 1320, 1616, 172, 858]. Accurate [736]. ACD [950]. achievement [377]. achievements [1439]. acoustic [1205]. activity [875, 948]. actuarial [779]. acyclic [1208, 740]. adaptation [355, 775]. adapted [1576]. Adaptive [1578, 141, 853, 805, 1348, 1263, 1554, 586, 243, 206, 721, 1101, 1424, 554, 1311, 1071, 900, 1391, 969, 925, 256, 339, 1560, 1602, 895, 1037]. Added [231]. Adding [306, 687, 324]. additional [12]. Additive [1147, 1442, 1375, 1193, 1314, 276, 1197, 382, 429, 1003, 1002, 309, 49]. Additive-multiplicative [1147]. additivity [47]. Adelchi [173]. adjoint [1092]. adjoint-state [1092]. adjust [1407]. Adjusted [260, 1117, 1380, 1584, 644]. Adjusting [1584, 699]. adjustment [469]. Adrian [173]. Advanced [1355, 699]. advances [1491, 159]. adventures [466]. advice [985]. advice-seeking [985]. affected [309]. affine [1365, 1100]. Africa [1316]. AFT [1357]. after [1407, 1011, 897]. against [881, 877, 1421]. Age [247, 622]. Aggregate [322, 507]. aggregated [786, 848, 1431]. aggregation [1211, 1591, 1387]. agreement [616]. AIC [1129]. Aided [247, 686]. AIDS [1160]. Ajax [454]. Albright [263]. Algebraic [701, 32, 132]. Algorithm [185, 220, 1456, 319, 768, 110, 756, 896, 1299, 1484, 134, 711, 765, 833, 1627, 1626, 1282, 384, 1377, 1161, 1481, 635, 1494, 130, 1232, 1138, 554, 154, 1040, 141, 6, 565, 1485, 170, 1391, 1498, 938, 1580, 198, 967, 1579, 1604, 1170, 1223, 1602, 50, 1519, 1191, 29, 131, 1180, 803, 1014, 862, 1306, 895, 702, 1570, 599, 1292, 1243, 1438, 171, 1267, 1370, 1281, 1340, 682, 584, 1192, 1262, 419, 1162, 753, 497, 762, 1469, 889, 977, 228, 978, 1004, 1450, 868, 221]. algorithmic [1051]. Algorithms [209, 109, 259, 360, 619, 842, 240, 48, 1470, 1254, 1017, 190, 343, 666, 600, 642, 34, 382, 332, 113, 553, 555, 1445, 1355, 846, 1455, 981, 651, 380, 420, 49, 14, 495]. alignment [1322]. Alive [229]. alleviate [874]. alternating [1580]. Alternative [505, 856, 1421]. alternatives [104, 52, 621, 877, 105]. America [592]. American [1405, 815, 1235, 591, 1614, 1545]. among [1337, 1010, 356, 1600, 985, 1326]. among-group [356]. amplitudes [1414]. Analyses [231, 1407, 1053, 886, 855, 304]. Analysing [125, 608]. Analysis [1195, 175, 1408, 1590, 246, 1393, 236, 204, 207, 173, 240, 253, 223, 293, 252, 790, 1012, 1436, 709, 17, 1395, 335, 364, 1175, 825, 729, 1108, 442, 1185, 1244, 121, 875, 948, 1476, 541, 851, 787, 1403, 331, 716, 1357, 1232, 1424, 643, 556, 446, 677, 1562, 1373, 542, 835, 400, 489, 1202, 1613, 195, 132, 1603, 398, 569, 127, 752, 365, 432, 1429, 1526, 664, 951, 1204, 1279, 374, 1580, 1303, 158, 35, 1103, 626, 818, 696, 747, 901, 1426, 1588, 725, 30, 1634, 449, 1242, 1033, 771, 1032, 64, 1570, 1410, 137, 162, 101, 163, 1520]. analysis [80, 417, 649, 633, 720, 847, 397, 922, 1323, 119, 1483, 524, 828, 694, 120, 783, 668, 939, 957, 1102, 1249, 640, 1517, 1319, 237, 358, 414, 977, 1029, 519, 769, 1612, 675, 379]. analytical [5]. analytics [638, 1275]. analyzer [1220]. Analyzing [218, 1269, 774]. ANCOVA [1021]. Andreas [478]. angular [8, 146]. ANN [768]. Annealing [619, 1484, 572, 1267, 775]. annealing-based [1484]. ANOM [563]. anomaly [1216, 1283]. ANOVA [563, 1074, 935, 27]. antigenic [1057]. Antony [340]. apple [1075]. apples [1075]. Applicable [919]. Application [26, 434, 1282, 399, 261, 312, 1555, 612, 1389, 856, 1533, 1130, 83, 1287, 1449, 733, 1499, 1316, 1562, 81, 892, 936, 132, 770, 1229, 1305, 429, 923, 334, 1156, 1042, 1570, 57, 1118, 885, 403, 380, 1160, 1394, 549, 1469, 769, 985, 940]. Applications [245, 212, 186, 1171, 1146, 1271, 1457, 1470, 1525, 825, 3, 605, 928, 766, 975, 1347, 1535, 93, 863, 827, 1103, 1606, 782, 1629, 1325, 895, 1113, 1431, 1093, 1233, 894, 1612, 371]. applied [1587, 1594, 432, 173]. Applying [1522]. Appointment [1277]. appraisal [104]. Approach [220, 292, 173, 63, 1087, 1407, 1538, 1466, 1433, 976, 1378, 1351, 672, 450, 1549, 729, 1108, 608, 873, 308, 859, 949, 1283, 347, 118, 643, 556, 378, 701, 946, 481, 1289, 1315, 854, 491, 1402, 1317, 798, 1620, 1391, 1142, 530, 158, 462, 1608, 1616, 200, 1437, 373, 100, 930, 162, 440, 685, 1174, 865, 1553, 1539, 1167, 1517, 918, 784, 1061, 1329, 1464, 1599, 112]. approaches [144, 972, 1316, 1502, 602, 573, 1245, 557, 1229, 650, 377, 1248, 1205]. Approximate [1592, 1186, 268, 949, 241, 1404, 805, 1519, 1106, 1453, 1213, 764]. approximated [1207]. approximately [774]. Approximation [267, 225, 262, 721, 11, 165, 102, 572, 598, 535, 169, 10, 1388, 653, 1422, 49, 348, 1]. Approximations [233, 704, 1175, 153, 1028, 280, 353, 342, 597, 1575]. approximator [762]. Arabic [713]. arbitrarily [71]. arbitrary [1007, 753]. Archimedean [1540, 1454, 824]. Area [292, 723, 818, 862, 100, 112, 107, 883]. areas [788]. ARFIMA [315, 407]. arithmetic [1100, 1289]. ARMA [73, 928, 92, 20, 128]. ARMA-scheme [20]. arrival [791]. art [401]. article [112]. Ascent [1290]. Aspects [627, 728, 1415, 18, 872, 553, 500, 56]. Assay [817]. assess [1241]. Assessing [1188, 909, 1006, 269, 21, 390]. assessment [143, 799, 1393, 77, 1461, 1410, 616, 1051]. asset [826]. assistance [1290]. associated [861]. Association [213, 214, 900, 1257, 1256, 486, 1454, 1497, 1258, 1255, 1260, 1259, 1530, 1517, 1405, 1614, 815, 1545, 1235, 591]. Assuming [266]. assumption [336, 1040, 801, 47]. Asymmetric [1400, 188, 1112, 789, 1521, 655, 1385, 963, 1517]. asymmetry [1173]. Asymptotic [534, 680, 795, 44, 88, 1330, 535]. asymptotically [1592]. Asymptotics [710]. attach [529]. attachment [1236, 1242, 1241]. attempt [1241]. Attractors [690]. attributes [1261]. auditory [1204]. augmentation [981]. autobinomial [1008]. autocorrelated [515, 849]. Autocorrelation [189, 328, 881, 1166, 561]. autocovariance [130]. Automated [316, 143]. Automatic [337, 488, 1459, 487, 1488]. Automatically [1037]. Autoregressive [270, 1230, 603, 739, 1476, 787, 1105, 660, 1482, 1327, 34, 1091, 1226, 1442, 1033, 1520, 1447, 1163, 1162]. auxiliary [1361, 483, 333, 1450]. available [12]. average [543, 603, 787, 130, 1400, 884, 34, 820, 140]. averaged [883]. averaging [59, 1038, 926, 1184, 1600, 1167]. avoiding [1116]. Award [458]. aware [640]. Azzalini [173].

B [465, 379, 185]. B-splines [185]. B. [538]. Backfitting [310]. backward [1080]. badly [1535]. Bagging [1252]. balanced [1484, 384, 1473]. balancing [887]. bands [470, 1506, 1390]. Bandwidth [372, 1066, 148, 781, 324, 422, 99, 956, 1253, 72, 1111, 177]. bank [1276]. banks [1276]. bar [294]. Bartlett [761]. barycentric [856]. Based [185, 988, 292, 912, 225, 1044, 611, 41, 205, 958, 1406, 1480, 1159, 1272, 124, 984, 1188, 728, 896, 807, 1216, 860, 1484, 1251, 436, 1550, 1230, 494, 1185, 121, 875, 948, 1276, 1376, 1432, 139, 37, 1219, 451, 972, 507, 1080, 1357, 1631, 701, 946, 1072, 601, 165, 1278, 532, 1086, 1141, 811, 809, 178, 1557, 1611, 251, 167, 573, 99, 560, 1312, 1027, 1391, 893, 269, 244, 353, 157, 1181, 585, 1296, 463, 1616, 1076, 581, 707, 1060, 1428, 1588, 1092, 285, 1568, 1437, 903, 210, 306, 1371, 1184, 882, 1442, 1008, 618, 990, 579, 1368, 1472]. based [1409, 1534, 1524, 1328, 1149, 1339, 865, 1332, 1553, 1261, 1539, 380, 1121, 1203, 1176, 1067, 549, 961, 1573, 237, 1280, 883, 1199, 978, 1035, 1326, 1450, 1387, 959, 1390, 1132, 1234, 394, 192, 193, 727, 874, 1229]. based-line [1573]. Bases [319]. basic [7]. Basis [273, 853, 485, 934, 461, 1398, 676]. basket [716]. Bayes [1407, 541, 1054, 1059, 135, 795, 461, 1452, 1111]. Bayesian [881, 63, 706, 1271, 791, 1501, 1221, 111, 1475, 110, 1264, 1265, 1395, 1351, 490, 1595, 182, 1392, 729, 1108, 955, 1185, 608, 826, 603, 949, 1352, 1476, 851, 474, 1219, 1105, 953, 1518, 1499, 1053, 1057, 1166, 1316, 1404, 1424, 656, 1026, 1594, 854, 181, 950, 1024, 295, 341, 1502, 1427, 1603, 1031, 1317, 42, 1179, 427, 1425, 1597, 560, 1027, 1397, 1429, 1516, 1592, 1393, 1038, 956, 572, 1225, 664, 1186, 374, 1109, 1598, 277, 1566, 180, 1103, 805, 1474, 1500, 204, 1058, 1602, 1220, 1426, 207, 1428, 553, 1224, 1519, 1568, 725, 952, 1504, 1184, 1156, 997]. Bayesian [1033, 485, 1263, 1106, 95, 1520, 773, 67, 53, 417, 1600, 582, 1503, 1295, 1341, 1318, 652, 1451, 1110, 1165, 96, 1396, 1340, 1593, 996, 1394, 1462, 1463, 504, 1453, 1167, 898, 957, 1517, 330, 907, 1029, 1268, 1269, 657, 1450, 379]. Bayesianbetareg [955]. BB [942]. bccp [1483]. be [233, 1182]. Beauchamp [878]. before [1407, 1011]. before-after [1407]. behavior [893, 1274, 224]. behavioral [1030]. Behaviour [359, 552]. belief [1416]. belongs [770]. benchmark [643, 764]. Benchmarking [797]. benefit [618]. Benefits [1204]. bent [1049]. Beran [865]. Beran-based [865]. Berge [584]. Berger [477]. Bergsma [962]. Berlin [925, 287]. Bernoulli [712]. Bernstein [1087, 1497]. best [1425, 888, 1129, 978, 1299]. best-subset [978]. beta [756, 1595, 955, 759, 866, 1099, 730, 1104, 1183, 1396]. better [902, 529]. between [402, 241, 44, 972, 367, 463, 65, 616, 533, 258]. beyond [542]. BGM [613]. Bi [1411]. Bi-level [1411]. Bias [288, 1407, 477, 829, 415, 86, 346, 1447, 763]. bias-optimized [86]. biased [876, 855]. big [1279, 1295, 1061]. bilateral [930]. Biliary [266]. Bilinear [220]. Biloxi [1237]. BINAR [1229]. Binary [589, 1206, 219, 1418, 464, 1188, 1097, 1401, 608, 365, 1155, 136, 551, 1207, 375, 1452, 1249, 718, 717, 774]. binary-choice [1155]. Bingham [841, 1359, 1358]. Binomial [225, 871, 1185, 733, 1515, 1536, 742, 1418, 135, 852, 1629, 1413]. biology [538].

Biomarker [693, 692, 968]. biomarkers [723]. Birds [817]. Birnbaum [1593, 957, 1111]. Birnbaum-Saunders [1111]. Birth [225]. birthday [558]. BITE [295]. Bivariate [1087, 1542, 505, 16, 608, 548, 1497, 792, 1419, 1418, 843, 1439, 994, 1508, 1530, 571, 840, 1149, 1084, 823]. BL [283]. Black [284, 1118]. black-box [1118]. block [916, 1481, 1423, 822, 1116, 1292]. block-bootstrap [822]. BMI [1244]. Bolfarine [117]. bond [1394]. Book [418, 466, 465, 489, 176, 467, 431, 263, 379, 538, 537, 173, 539, 371]. Boolean [690, 814]. boosted [818]. Boosting [972, 655, 809, 691, 1145, 1004, 1554]. Bootstrap [866, 1074, 315, 544, 1460, 1249, 802, 359, 1251, 1522, 563, 605, 507, 804, 602, 1115, 798, 361, 686, 116, 737, 439, 353, 166, 1379, 74, 346, 1250, 373, 822, 1116, 538, 1553, 659, 1099]. bootstrap-based [507]. Bootstrapping [202, 257, 1190, 711]. bootstraps [68]. bordered [57]. both [700]. bound [1180]. Boundary [239, 1433, 1417]. bounded [1230, 190, 1327]. bounds [68, 142, 1575]. Bowman [173]. box [1118, 21]. Box-Cox [21]. Boxplot [665]. BP [818]. branch [1340]. brand [336]. breakdown [89]. breaking [1056]. breaks [796]. bridge [1394]. bridge-randomized [1394]. Brownian [1501]. BSFM [111]. bucket [1211]. Buckley [876]. bug [1083]. build [713, 638]. Buja [478]. Bump [212]. bumps [756]. Burr [860, 1228, 1552]. bus [1409]. business [521, 1275].

C [531, 533]. C/C [531]. C2 [122]. Caching [528]. Calculate [318]. Calculating [1119]. Calculation [44, 343, 1178, 659]. calculator [1353]. calculus [1489]. Calibration [76, 117, 87]. Can [529, 691, 203, 1182, 299]. cancer [1587]. Canny [1368]. Canonical [1073, 352, 1564, 675]. capability [1524, 1398]. car [1406]. cardinality [1012, 1565]. care [770]. Carlin [379]. Carlo [538, 1078, 1005, 1313, 81, 1311, 602, 178, 361, 572, 374, 1187, 708, 535, 698, 743, 912, 1156, 931, 10, 652, 753, 1035, 484]. CARMA [941]. Carrier [186]. CART [1448]. cascades [442]. Case [370, 791, 300, 1632, 1287, 844, 909, 811, 593, 132, 127, 1245, 1420, 1323]. case-control [811, 1245]. Cashflow [285]. catastrophe [412]. Categorical [238, 1287, 1050, 35, 1556, 901, 1601, 1048, 66]. categorization [620]. Cauchy [194, 734, 564]. causal [884, 741, 42, 922]. cause [1581]. cause-specific [1581]. causes [952]. CCD [1622]. CD [371]. CD-ROM [371]. cell [894]. cells [696]. CEM [1445]. censored [916, 133, 1544, 958, 1264, 728, 1094, 807, 1039, 1017, 1185, 1376, 876, 98, 1590, 954, 1424, 792, 1609, 664, 1109, 1181, 1343, 1076, 581, 703, 1519, 867, 1380, 1362, 1534, 571, 609, 633, 720, 847, 1363, 647, 1420, 1332, 783, 1479, 1507, 1411, 793]. censoring [1544, 1094, 1532, 1381, 1101, 1473, 1557, 492, 1552, 1362, 840, 1465, 865, 996, 1051, 1531]. Census [215]. Central [592, 1007, 546]. certain [20, 2, 784]. cervical [1587]. Chain [374, 1005, 178, 743, 1156, 652, 344]. chains [1325, 1625, 1455]. Challenge [1614, 1405, 1545]. challenges [624]. Challenging [731]. Chance [418, 966]. Change [1397, 242, 780, 1444, 1011, 1138, 799, 575, 572, 1116, 1372, 712, 1139, 979, 630]. change-point [1138, 799, 575, 572, 979]. Change-Points [1397]. changepoint [1363]. changepoints [1318]. Changes [215, 592, 93, 149]. Chapman [379, 371]. characteristic [205, 312, 1435, 777, 426]. characterization [1480, 40, 734, 568, 48]. Characterizations [1515, 1189, 49]. characterized [1548]. chart [577, 820, 536, 947]. Charts [294, 1369, 1529, 1385]. checking [1496]. checks [179]. Chernoff [1575]. Chi [1013, 1354, 169, 1421]. chi-distributed [1354]. Chi-square [1013, 1421]. child [769]. China [875, 948]. Chinese [1276, 985]. Choice [186, 1135, 336, 464, 656, 738, 422, 1155, 969, 455, 899, 590, 1145, 51, 985]. Cholesky [1063]. Choosing [1069]. Christian [263]. Christians [985]. chunks [528]. CIR [1388]. circadian [542]. circle [708]. circles [107]. circular [665, 708, 777]. cities [1237]. claims [412]. Class [259, 916, 1079, 825, 902, 1026, 1043, 1266, 782, 1439, 822, 142, 546, 1064, 1233, 968]. Classes [222, 1532]. Classical [1109, 234, 181, 95]. Classification [611, 1210, 493, 1243, 226, 768, 755, 937, 1448, 797, 873, 451, 936, 398, 427, 888, 1598, 136, 180, 617, 845, 433, 1588, 903, 210, 435, 32, 812, 1345, 56, 693, 126]. classifier [507]. classifiers [620, 813, 506, 1144]. classifying [151, 907]. Clicks [1240]. Client [244]. Client/ [244]. climate [592, 303]. clinic [1277]. clinical [477]. cliques [1240]. close [329]. Closed [1287, 1432, 1226, 1388]. Closed-form [1287]. closeness [943, 867]. clouds [1621]. cluster [23, 1232, 425, 621, 1616, 908, 1570, 994, 1409, 304, 761, 519, 774]. cluster-based [1616, 1409]. cluster-weighted [994]. clustered [1454, 303, 235, 847, 1249, 1507]. clustergrams [304]. Clustering [396, 1343, 1630, 994, 1302, 1205, 984, 1188, 833, 393, 1169, 1157, 1283, 1219, 395, 507, 678, 1635, 1334, 560, 1342, 810, 617, 1030, 1191, 910, 1184, 1250, 1472, 1451, 1548, 1261, 420, 126, 1200, 1234, 1045, 394]. clustering-based [1261]. Clusters [1201, 1334, 637, 683, 54, 1190]. ClustGeo [1169]. co [1635]. co-clustering [1635]. COBS [138]. Code [524, 335, 364, 528, 531]. codes [1301]. coding [423]. coefficient [479, 1026, 1611, 1364, 969, 598, 818, 923, 1000, 890, 931, 1064, 922, 761, 1324, 699, 980, 959, 1531]. coefficients [232, 44, 798, 821, 19, 1439, 1568, 1581, 1441, 53, 472, 995, 1083, 784]. cointegration [802, 475, 698, 469]. collecting [621, 176]. College [1237]. collinear [1343]. collinearity [874, 1285]. column [142, 869]. COM [424]. combination [1549, 508, 385]. combinations [723, 404]. combine [1192]. Combined [512, 209, 294, 425, 1003, 1002, 1588]. Combining [964, 210, 937, 47]. Commemorating [558]. Comments [1257, 1256, 1258, 1259, 81]. Common [929, 1389, 792, 1027, 1598, 1152]. Community [1239, 1622, 1046, 1236, 1242, 1241, 1511, 1261, 1240]. compactly [676]. Companion [431]. Comparative [1024, 1102, 117, 464, 87, 352, 1114, 969, 1248, 582, 1601]. compare [312, 387]. Comparing [1632, 563, 1336, 857, 581, 861, 1315, 122, 551, 1022, 1596, 1390]. Comparison [321, 976, 595, 241, 187, 1337, 271, 214, 492, 557, 265, 1445, 231, 184, 1075, 829, 563, 1209, 972, 1028, 92, 315, 666, 181, 935, 602, 742, 361, 1379, 342, 698, 703, 377, 597, 1335, 685, 1413]. Comparisons [240, 52, 747, 1569, 408]. comperative [140]. competing [1185, 1557, 952, 597, 1362, 1465, 1149, 965, 1319]. competitive [336, 455]. complete [792, 1520, 1323]. complete-case-analysis [1323]. completely [309]. complex [554, 805, 634]. complexity [618, 137, 1081]. Component [202, 240, 1012, 1436, 1175, 556, 400, 1613, 1391, 1148, 1580, 158, 778, 356, 993, 449, 1478, 1477, 649, 939, 764, 358]. components [991, 476, 1389, 678, 831, 149, 4, 1018, 64, 1073, 1174, 651, 830, 504, 1443]. Componentwise [355]. Composite [880, 963, 968, 1364, 1247, 1355, 546, 1081, 1124]. Compositional [1383, 677, 939]. compound [450, 892]. compounding [732, 732]. Comprehensive [1387, 1013]. COMPSTAT [174, 458]. Comput [117, 112]. Computation [363, 1354, 504, 871, 1404, 438, 879, 1592, 313, 1186, 1154, 805, 1519, 724, 131, 1007, 1106, 1414, 1274, 1575, 1453, 783, 513, 962, 705, 1036, 1136]. Computational [728, 468, 1491, 187, 508, 6, 254, 18, 500, 551, 449, 56, 174, 1051, 1470, 1444, 982, 436, 278, 975, 132, 467, 409, 1415, 553, 1461, 1437, 919, 506, 162, 114, 1081]. Computationally [540, 1538, 587]. Computations [365, 137, 42]. Computer [247, 111, 767, 31, 1178, 622]. Computing [1055, 542, 244, 562, 926, 1490, 646, 1294, 1267, 36, 222, 258, 760, 1322, 614, 532, 1384, 561, 34, 332, 732, 555, 576, 495, 1309, 221]. concentrated [648]. concentration [850]. concepts [467]. concerning [206]. concordance [846, 1121, 453]. condense [813]. condition [1284, 1158, 1496]. Conditional [318, 1105, 366, 1610, 1253, 562, 535, 906, 822, 1171, 349]. conditionals [894]. conditioned [511]. Conditions [430]. Confidence [585, 503, 971, 916, 1272, 860, 826, 470, 1536, 1337, 1095, 1611, 575, 363, 1320, 1415, 38, 356, 1506, 904, 801, 426, 1339, 1553, 753, 348, 750, 1390]. confounders [380]. confounders- [380]. confounding [1056, 843]. conic [1413]. connect [1217]. connected [982, 8]. connectivity [533]. Conquer [319, 1293]. consensus [54]. consequences [1535]. conservativeness [942]. considerations [449]. considered [1615]. considering [63, 1278]. Consistency [1238, 51, 1420]. Consistent [1487, 311, 257]. consistently [1123]. constant [879, 204]. Constrained [1020, 682, 1211, 138, 810, 345, 513]. Constraint [380]. Constraint-based [380]. constraints [1169, 324, 852, 966, 894, 735, 940]. construct [1254, 753]. construction [1290, 198, 848, 659]. consumption [1546]. containing [1339]. Contaminated [189, 92]. Context [1016, 788]. Context-specific [1016]. contiguous [1456]. Contingency [297, 118, 496, 1408, 1293, 819, 562, 894, 1517, 344]. Continuous [268, 984, 1351, 1521, 1435, 500, 732, 1490, 1625, 1421, 842, 775]. continuous-discrete [500]. continuous-time [1625]. continuously [490, 822]. continuum [114]. Contour [767]. contoured [679, 776]. Contours [255]. Contrasts [326]. Contribution [422]. control [1369, 1559, 1529, 378, 811, 900, 1245, 1385, 820, 1292, 536, 616]. Controlling [1569]. controls [857, 408]. controversial [1244]. convergence [143, 1558, 565, 1580, 14, 549]. conversion [1177]. Convex [404, 459, 1604, 862, 1451, 762]. Convexity [966]. convolution [1273, 140]. convolution-type [140]. Conway [1487]. coordinate [1481, 1579]. coordinates [124, 637]. Copula [842, 668, 604, 1087, 711, 1454, 1497, 1502, 1557, 1568, 1149, 800, 1319]. copula-based [1557, 1568]. copulas [1419, 824]. core [889, 889]. Cornell [878]. corrected [1334]. Correcting [970]. Correction [1627, 508, 339, 1478, 1463, 1572, 1359, 1175, 1505, 872]. Correlated [266, 1401, 388, 276, 1521, 1179, 983, 1089, 1292, 987, 1452, 757]. correlation [826, 308, 1376, 44, 37, 448, 1137, 1611, 1620, 598, 1170, 1459, 1564, 1032, 1073, 36, 761, 784, 1029, 675]. correlations [964, 352, 1011, 814, 45]. correspondence [335, 364]. cost [1406, 1529, 1148, 1178]. cost-effectiveness [1178]. cost-optimal [1529]. count [770, 1342, 983, 956, 1182, 1441, 987, 1153]. Counting [819, 1024, 1328]. counts [1352, 1327]. Coupling [1106]. course [560]. covariables [462]. Covariance [1509, 44, 679, 257, 835, 34, 19, 845, 30, 131, 1014, 1423, 426, 962]. covariate [505, 477, 1595, 1376, 309, 1380, 1323]. covariate-adjusted [1380]. Covariates [238, 133, 1375, 920, 655, 1364, 1245, 1148, 1288, 1605, 1247, 996, 1160, 1452, 1021]. coverage [356]. covering [1514, 1520]. Cox [1444, 450, 21, 1377, 972, 274, 1606, 1145, 1192, 644, 745]. Cox-models [274, 1145]. CoxBoost [972]. crash [1410]. crash-safety [1410]. crashes [1409]. CRC [379, 371]. Creasy [1225]. Creating [465]. Credit [279, 1576, 607, 826, 929, 1325]. Crisp [420]. Criteria [296, 464, 1005, 990, 1121]. criterion [554]. critical [39, 726]. CRL [294]. Cross [296, 547, 1550, 1392, 501, 81, 1426, 758, 1634, 32, 1065, 1308, 1121, 51]. cross-classification [32]. cross-entropy [1550]. Cross-validated [547, 501]. cross-validation [1392, 81, 1426, 758, 1634, 1065, 1308]. cross-validatory [1121, 51]. crossing [514]. cubic [75]. Cumulants [258, 534]. Cumulative [300, 134, 332, 908]. cure [896, 1587, 1296, 1606, 1042, 1126, 1443, 793]. cured [1498]. current [1147]. curse [731]. curvature [1021]. curve [1538, 1307, 76, 1634, 582, 1176]. Curves [278, 1486, 300, 1432, 1114, 723, 1186, 1588, 618, 1201]. curvilinear [85]. cut [1244]. cutoff [965]. CXXR [627]. cyclical [872].

D [379, 636, 122]. D. [539]. DAG [1053]. DAG-models [1053]. daily [1264, 927, 1618]. Dandelion [886]. Dassios [962]. Data [266, 1195, 212, 1614, 209, 261, 325, 1545, 189, 322, 208, 1235, 251, 248, 219, 1297, 272, 327, 216, 591, 379, 173, 231, 293, 639, 252, 916, 133, 1238, 611, 958, 1406, 612, 1457, 1528, 300, 709, 567, 1365, 301, 124, 984, 479, 1264, 1265, 1039, 443, 1527, 1130, 1361, 1246, 1351, 1196, 1017, 434, 83, 1097, 21, 721, 442, 1401, 1282, 399, 844, 1185, 589, 608, 873, 121, 393, 859, 1376, 1432, 876, 1456, 541, 902, 1125, 815, 1219, 451, 395, 445, 1403, 1357, 527, 594, 1631, 289, 1424, 1347, 1074, 677, 1373, 1635, 854, 578]. data [1040, 1628, 400, 489, 935, 1202, 1278, 1454, 1497, 1591, 1502, 1621, 1010, 936, 1031, 642, 641, 573, 396, 127, 427, 99, 560, 1342, 363, 983, 1046, 156, 1526, 956, 969, 1418, 452, 664, 1186, 1279, 1498, 1109, 1142, 455, 1181, 199, 531, 1574, 158, 35, 923, 1147, 1343, 637, 915, 1058, 1002, 1063, 334, 901, 1588, 1519, 1191, 1344, 1437, 848, 1412, 772, 354, 302, 122, 303, 1349, 416, 1423, 447, 1042, 1380, 1630, 771, 552, 1032, 160, 64, 456, 618, 435, 930, 579, 994, 1441, 235, 931, 981, 333, 1295, 1447, 171, 22, 689, 616, 1302, 539, 1431]. data [571, 609, 633, 720, 847, 1465, 1363, 397, 1542, 1370, 647, 1345, 1496, 588, 1339, 606, 1163, 1323, 119, 1332, 125, 1160, 1394, 420, 375, 828, 987, 120, 299, 441, 1452, 783, 1167, 939, 1249, 307, 640, 1596, 126, 965, 66, 330, 1319, 1275, 1479, 1153, 1573, 414, 907, 977, 1029, 1269, 1507, 769, 985, 1200, 968, 1061, 1329, 1450, 774, 1411, 823, 793, 897, 1111, 1446, 394, 1405, 176, 371, 1003]. data-based [99, 618]. data-fitting [901]. datasets [191, 1510, 1460]. DAVIS [251, 306]. Day [239, 927]. Death [418, 906]. deaths [1130]. Decision [263, 291, 1188, 1299, 1550, 651]. deck [5]. decking [157]. Decomposition [1431, 932, 974, 1063]. deep [1218]. Deepwater [816, 817, 818]. default [607, 826]. defaults [929]. defectives [1398]. definite [676]. definition [1525]. degenerate [1182]. degree [1270]. deleting [687]. deletion [759]. demand [770]. demographic [300]. Denoising [499, 1528, 1469]. dense [1379]. densities [756, 1069, 312, 892, 167, 1038, 1306, 1346]. Density [984, 60, 322, 1273, 703, 685, 71, 1252, 206, 445, 695, 491, 422, 492, 99, 360, 1253, 342, 637, 708, 1383, 334, 1428, 583, 1490, 1514, 140, 1308, 1551, 1111]. Density-based [984]. depend [655]. Dependence [1084, 1087, 1015, 372, 1412, 670]. dependency [1316]. dependent [1128, 645, 1307, 1557, 1312, 1148, 1303, 1047, 579, 1581, 1328, 1452, 861, 1479, 1280, 1612]. depicting [1069]. Depth [1631, 241, 451, 936, 673, 1082, 1630, 1267].

Depth-based [1631]. Derivation [293]. Derivative [186, 768]. Derivatives [285, 1092]. descent [1481, 841, 1579]. description [152]. Descriptive [392, 911]. Design [246, 1148, 820, 916, 881, 1369, 1632, 3, 1522, 946, 742, 1500, 1567, 806, 623, 482]. designed [132, 998, 160]. Designing [1385, 967]. Designs [318, 1291, 1123, 1056, 1135, 554, 1290, 701, 1179, 1495, 131, 1292, 101, 546, 1332, 1553, 1390, 1134]. destructive [1042]. detect [744, 1621, 1089, 94, 683]. Detecting [93, 425, 818, 282, 606, 595, 24, 62, 1057, 1418, 630, 54]. Detection [1314, 149, 1397, 411, 1216, 79, 1138, 446, 1006, 1046, 872, 1333, 354, 1380, 649, 1318, 979, 1363, 1261, 1160, 1622]. detector [1368]. determinants [1276]. Determinating [103]. determination [1456, 1402, 1023, 1281]. determine [134]. Determining [469, 965]. Deterministic [1486, 103, 313, 180, 997, 469, 1390]. detrending [1088]. Detroit [1237]. Developing [263]. development [875, 948, 1276, 848, 624, 521]. developments [1574]. deviance [580]. Diagnosis [1285, 1277, 602, 1115]. Diagnostic [988, 1632, 1347, 1482, 1006, 551]. Diagnostics [94, 479, 759, 37, 370, 1120, 95, 414, 1037]. diagonalization [1282]. dialog [121]. diamond [1141]. Dicing [418]. difference [407, 1067, 1132]. difference-based [1132]. differenced [128]. differences [1379, 94, 65, 146]. Differencing [931]. Different [321, 1229, 1327, 1439, 1096, 404]. differential [1559, 1080, 1177, 909, 1317, 1001, 743, 403, 1262]. differentiation [1459]. diffusion [1337, 1090, 1280]. diffusions [652]. Dimension [1048, 991, 1509, 902, 1591, 1402, 1379, 845, 650, 433, 1301, 990, 1281, 1205]. dimensional [829, 1282, 1157, 1377, 1376, 79, 1635, 1350, 1628, 355, 550, 641, 969, 841, 311, 1604, 1608, 747, 1266, 910, 1118, 1145, 1262, 992, 1203, 1213, 1304, 1537, 907, 1029, 1269, 968, 1464]. dimensionality [1254, 620, 145, 731]. dimensions [594, 1100]. Direct [226, 1458, 1054, 51]. direct-path [1054]. directed [1208, 740]. direction [91]. directional [1527, 1062, 1074, 1630, 66]. Directions [195, 596]. Dirichlet [1365]. Discarding [240]. discharge [818]. discontinuous [413]. Discovered [213]. Discovering [786]. discovery [466, 946, 689, 693]. discrepancy [1567, 899, 978]. Discrete [1303, 268, 1457, 134, 1128, 1627, 1626, 1513, 1502, 1487, 363, 1415, 332, 462, 1459, 1426, 500, 732, 1514, 1233, 985]. discretely [754]. discretization [1011]. discretized [656]. Discriminant [17, 677, 236, 1278, 162, 163, 397]. discriminate [1010]. discrimination [896, 6, 50, 14]. Discriminative [833]. Discussion [144, 254, 179]. disease [1352, 595, 610, 551, 683, 615, 968]. diseases [460, 634]. dispersed [1487]. dispersion [704, 1229]. display [122]. dissimilarities [49]. dissimilarity [964, 402]. Distance [236, 320, 1525, 436, 1376, 834, 396, 654, 107, 1261, 1203, 604, 793, 394]. distance-based [436]. distances [1172, 11]. distress [1278]. Distributed [1467, 16, 1321, 1354, 552, 1446]. Distribution [1480, 227, 366, 1636, 188, 842, 225, 1593, 264, 763, 1555, 1044, 63, 41, 1227, 1270, 1457, 1272, 982, 1055, 871, 1039, 1433, 1533, 134, 1527, 1550, 262, 1119, 1532, 765, 243, 671, 1228, 1088, 1185, 733, 1515, 98, 1513, 1590, 1494, 954, 289, 115, 574, 648, 804, 1374, 892, 141, 669, 1273, 752, 1312, 97, 684, 1320, 1386, 203, 415, 1418, 658, 841, 1109, 1020, 1154, 1103, 1506, 1098, 1512, 1076, 70, 707, 864, 943, 1412, 970, 597, 732, 715, 882, 194, 1335, 1552, 13, 169, 702, 1113, 1362, 773, 406, 1183, 1503, 629, 1530, 1023, 647, 1332]. distribution [753, 90, 513, 957, 480, 1359, 1358, 348, 564, 750, 760, 221]. Distribution-free [1480, 243, 597, 221]. Distributions [266, 740, 1486, 389, 1146, 1271, 1470, 1079, 1112, 196, 308, 412, 679, 953, 944, 975, 1316, 1360, 1492, 1562, 341, 792, 1384, 850, 1336, 1043, 1025, 1027, 1386, 290, 1415, 1385, 353, 1181, 1434, 782, 777, 316, 1060, 1568, 725, 732, 776, 1490, 1514, 1421, 417, 590, 1534, 22, 1341, 653, 349, 1189, 428, 1233, 632, 376, 66, 659, 1084, 1583]. Distributome [975]. disturbances [933]. divergence [496, 851, 827, 1027, 961]. divergence-based [1027]. diverging [1375]. Divide [319, 1293]. Divide-and-Conquer [319, 1293]. divided [1591]. do [76, 123]. document [1596]. Documents [229, 528]. DOLDA [1266]. domain [1279, 55]. dominance [1315, 91]. Don’t [123]. Dose [326]. dot [559]. Double [1435, 739, 1471, 945]. double-regularized [1471]. double-truncation [945]. Doubly [1612, 958, 1357, 1373, 1591, 1412, 571, 609, 720, 847, 514]. doubly-censored [847]. doubly-robust [1373]. doubly-truncated [1357]. down [1236]. Download [218]. Drawing [216]. DRGP [988]. driven [1297]. Drivers [1242]. driving [1408]. dropout [1475, 608, 357]. dropouts [1329]. drowsy [1408]. dual [1030]. dual-view [1030]. due [1068]. duration [421]. Durbin [228]. Duxbury [263]. Dynamic [215, 1198, 875, 1513, 288, 253, 394, 1436, 1221, 1125, 1499, 528, 613, 1603, 664, 199, 256, 339, 785, 1504, 302, 1045, 948]. dynamical [103, 311, 1371].

e-E-insensitive [870]. e-learning [713]. EBLUP [544]. ecological [615]. ecology [100, 112]. Economic [1369, 1406]. economical [788]. EDF [98]. edge [79, 1368]. edge-preserving [79]. edges [687]. EDI [156]. Edition [379, 538]. editor [509]. Editorial [1195, 150, 520, 919, 391]. eds [467]. Educational [231]. EEG [689]. Effect [38, 817, 85, 444, 1244, 766, 1316, 663, 1122]. Effective [326, 465, 1081]. effectiveness [1178]. Effects [266, 239, 1064, 755, 1246, 1401, 1125, 62, 1056, 884, 1402, 1031, 183, 473, 1131, 1439, 580, 1126, 773, 931, 537, 47, 1201, 996, 1160, 1268, 1467]. efficiency [1276, 878, 203, 135, 553, 410, 555, 981, 1323]. Efficient [1470, 722, 871, 645, 1392, 187, 1100, 1373, 792, 1311, 254, 1154, 1104, 1098, 1063, 724, 1504, 934, 1274, 1575, 783, 962, 705, 659, 1538, 1377, 1481, 332, 1579, 540, 1170, 1602, 1301, 785, 346, 702, 1292, 762, 587, 564]. effort [147]. Efron [116]. eigenvector [6, 14]. either [91]. electrical [1143]. electricity [399, 927, 926]. Elizabeth [1316]. ellipse [146]. ellipses [106, 107]. elliptical [1542, 1523]. elliptically [679, 776]. ellipticity [1419]. emerging [908]. emission [82, 1205]. Empirical [541, 660, 325, 569, 473, 616, 745, 205, 1395, 711, 829, 1404, 312, 1417, 439, 1154, 135, 974, 1510, 970, 461, 1339, 380, 659, 644, 1124, 940, 5]. emulation [1301]. encoding [1565]. endogenous [980]. endowed [837]. energy [1527, 1546]. engagement [1239]. England [300]. enhanced [1232]. Ensemble [695, 1556, 611, 974]. Entropy [288, 1550]. environment [875, 948, 532, 122]. Environmental [272, 1491]. epidemiological [434]. epsilon [1228, 497]. equal [129]. Equality [988, 605, 679, 1114, 798, 821, 1085]. equation [284, 1317, 1176]. equations [1559, 1080, 1137, 1177, 983, 1001, 1003, 1002, 743, 30, 403, 1093]. equilibrium [1054]. Equivalents [326]. ergodicity [1340]. Errata [112]. Erratum [117, 948, 1099, 177, 1003, 718]. Error [202, 226, 989, 937, 68, 1112, 1376, 1456, 1586, 181, 1245, 1393, 166, 1310, 1569, 618, 1064, 1163, 1214, 33, 1160, 700, 1122, 918]. error-free [1456]. error-in-covariates [1245]. Errors [325, 292, 264, 223, 222, 276, 1482, 1179, 1059, 1442, 849, 1600, 800, 963, 996, 1162, 1324]. errors-in-variables [1324]. essential [1285]. EST [423]. estimate [1123, 312, 1317, 1253, 1306, 1214]. estimated [98, 990, 1443, 947]. Estimates [1406, 937, 422, 1592, 34, 360, 353, 342, 135, 990, 404, 51, 1452, 61, 1190]. Estimating [612, 232, 667, 1177, 189, 1334, 259, 147, 1330, 726, 1126, 681, 146, 800, 242, 894, 192, 193, 1444, 1538, 1128, 1244, 190, 62, 554, 1137, 884, 983, 585, 1003, 1002, 544, 1346, 114, 349, 1049, 1450, 484]. Estimation [1555, 41, 186, 1246, 1401, 1430, 1276, 1482, 239, 1454, 322, 1071, 274, 311, 1288, 1605, 852, 334, 265, 1349, 912, 267, 288, 1478, 1477, 472, 842, 1388, 1163, 761, 7, 1324, 1573, 348, 1531, 1087, 1271, 1528, 791, 9, 505, 905, 110, 896, 1039, 1314, 697, 1533, 1097, 645, 1230, 754, 1627, 1626, 468, 767, 739, 1432, 1481, 1125, 1559, 421, 451, 1518, 1307, 118, 109, 891, 446, 448, 284, 677, 315, 1373, 517, 950, 695, 1400, 60, 1374, 1613, 438, 341, 491, 628, 850, 945, 84, 492, 1043, 752, 1597, 99, 149, 824, 684, 557, 1155, 956]. estimation [1391, 415, 572, 737, 4, 1173, 1598, 256, 339, 1001, 1579, 1018, 1104, 1512, 1000, 1063, 1434, 1150, 362, 707, 1428, 743, 781, 906, 1437, 583, 1412, 40, 631, 899, 776, 416, 1445, 997, 1335, 771, 1116, 1008, 45, 57, 1441, 1113, 773, 1625, 67, 503, 849, 582, 1158, 413, 1447, 685, 1530, 1341, 571, 609, 1465, 880, 1110, 1345, 82, 588, 1332, 1331, 1189, 428, 33, 700, 1231, 369, 89, 1067, 1122, 1213, 604, 764, 330, 971, 1280, 1523, 1199, 735, 1151, 1488, 1624, 959, 999, 1132, 793, 1111, 25]. estimations [329]. Estimator [295, 15, 1287, 206, 1578, 257, 372, 1350, 1417, 1286, 313, 370, 166, 346, 1014, 680, 986, 1207, 161, 510, 654, 12, 1019, 1479, 564, 763, 940, 71]. Estimators [187, 254, 1633, 496, 1525, 1079, 710, 1252, 671, 139, 1086, 827, 893, 1059, 1330, 1568, 534, 795, 53, 140, 224, 794, 1508, 840, 1249, 604, 705, 549, 1551]. Eulerian [642]. Evaluating [1439]. Evaluation [801, 10, 1078, 1322, 1024, 669, 1273, 1279, 899, 1284, 764, 5]. event [427, 1526, 618]. events [689]. evidence [932]. evoked [1204]. evolution [1057]. Evolutionary [313]. evolving [1045]. EWMA [577]. Exact [68, 331, 166, 1023, 28, 134, 679, 1494, 556, 1290, 879, 363, 1386, 737, 332, 777, 764, 344, 1061]. Exactly [1636]. examination [1128]. examining [619]. example [737]. examples [582]. Excel [1083, 424, 858, 253, 263]. Exchangeable [168, 1097, 1423, 774]. exciting [1091]. exit [1417]. exome [1142]. Expansion [233, 853, 403]. expansions [680]. expectation [1627, 1626, 1311, 977]. expectation-maximization [1627, 1626, 977]. expectations [1294]. expected [791, 459, 1354, 463]. expectile [1353]. expectiles [1198, 951]. expenditure [1547, 1546]. experience [6]. experiment [122, 546]. experimental [1290]. Experiments [246, 1135, 643, 132, 967, 1500, 160]. expert [389]. explanatory [1287, 1310, 12]. explicitly [645]. exploration [300, 816, 1236, 1618]. explorations [90]. Explorative [302]. Exploratory [643, 886, 694, 120]. Exploring [215, 799, 183, 1408, 975, 1237, 1240]. Expo [815, 1235, 591, 1614]. exponent [311]. Exponential [97, 1227, 505, 1272, 1094, 792, 945, 1429, 1320, 951, 1109, 1020, 1212, 857, 29, 943, 1380, 417, 590, 1530, 1458, 66, 1199, 1111]. exponential-geometric [1227]. exponentiality [1075]. exponentially [1107, 820, 1042]. exponents [103]. Extended [1610, 323, 500, 634, 375]. Extending [522, 1129, 502, 537]. extensible [530]. Extension [399, 1228, 1191, 1382, 1617]. Extensions [234, 180, 1332]. external [1238]. Extracting [741, 457]. Extraction [211, 383, 1281]. Extreme [175, 1555, 1283, 1316, 574, 1594, 1086, 623].

F [1257, 1258, 538]. F. [489]. Factor [1402, 769, 1198, 1276, 1613, 752, 1303, 1566, 927, 901, 1220, 886, 114, 694, 1213, 1523]. factorial [1632, 701]. Factorization [328, 1044]. factorizations [838]. factorize [130]. factorized [1392]. factors [929, 1054, 1567, 1239, 47, 965, 1523]. fails [282]. failure [1544, 1403, 954, 1526, 1096, 952, 1534, 1149, 761, 1194, 1612]. fallacy [1369]. familial [365]. families [29, 417, 590, 634]. Family [287, 1146, 1271, 1384, 945, 200, 952, 1534, 659]. familywise [1569]. Faraway [537]. Fast [343, 561, 1082, 1092, 1414, 846, 1561, 1593, 384, 1161, 1028, 81, 1493, 1485, 862, 599, 653, 1262, 221]. faster [1306]. fastWKendall [1170]. fat [672]. fauna [389]. favorable [1374, 588]. fdANOVA [1202]. FDR [942]. Feature [1376, 617, 1377, 1608, 1008, 1281, 1537]. features [1031, 1058, 1030, 1565, 812, 482]. feedback [67]. Fernandes [371]. Ferraty [489]. fetal [1130]. fewer [1379]. FGM [1149]. fidelity [767]. Fields [175, 1263, 1599]. Fieller [1225]. FIGARCH [569]. filament [696]. file [25]. filling [1486]. Filter [328, 545]. filtered [491]. filtering [853, 500, 743]. filters [153]. Finance [175]. Financial [217, 1449, 928, 289, 1278, 1010]. Finding [1142, 151, 1615, 1292]. Finite [756, 153, 1360, 290, 224, 96, 727, 60, 824, 855, 1445, 726, 510, 417, 590, 1458]. finite-range [855]. Finite-sample [290, 727]. firm [1394]. first [791, 954, 1482, 1417, 407, 1207, 1534, 91]. first-exit [1417]. first-order [1482, 1207]. Fisher [1550, 833, 648, 850, 841, 113, 653, 428, 323, 513, 1359, 1358]. Fisher-Bingham [1359]. Fit [323, 1480, 21, 494, 444, 78, 1381, 1521, 648, 719, 1621, 1338, 1127, 194, 1552, 1421, 647, 90]. fits [94]. fitted [991]. Fitting [273, 289, 264, 375, 182, 550, 429, 708, 901, 1348, 772, 1438, 106, 107, 588, 700]. five [557]. Fixed [292, 838, 1, 837, 85, 1456, 1125, 1320, 933, 931, 430]. fixed-accuracy [1320]. fixed-effects [1125]. fixed-length [1456]. Fixed-rank [838, 837]. flexibility [806]. Flexible [1543, 1181, 683, 896, 1595, 1015, 1434]. flow [1165]. focal [786]. focus [972]. fold [1426]. follow [989]. forecast [1620, 1619, 1616, 926]. Forecasting [144, 927, 1167, 1221, 1251, 875, 948, 928, 1315, 1297]. forecasts [660, 1615, 1618, 1617]. forest [1298]. forests [1535]. form [1287, 50, 571, 1388, 1090]. Format [327]. forms [348]. formula [1369, 940]. formulas [863, 1007]. Fortran [318]. forward [301]. foundation [161]. Four [1254, 1237]. Fourier [542, 203, 1414]. fraction [1587, 1498, 1126, 1398]. fractional [1501, 754, 284, 701, 1165]. fractionally [787, 128]. frailty [1403, 357, 1110, 1319]. frailty-copula [1319]. frame [836]. framework [454, 1220, 1263, 502, 521]. free [768, 1480, 243, 1456, 1208, 597, 1455, 221]. Freight [186]. French [584]. frequencies [1414]. frequency [1013, 86, 771, 55]. Frequentist [1059, 1316, 1024, 1393]. Friendly [246]. Frommlet [1258]. Frontier [223, 111, 374, 997, 1136]. full [1632, 1015]. full-factorial [1632]. Fully [1065, 1053, 488, 487]. Function [189, 233, 256, 1555, 1087, 134, 1378, 1578, 470, 130, 312, 1611, 914, 432, 1354, 1155, 415, 1506, 147, 1076, 334, 1412, 899, 1171, 413, 571, 840, 1420, 632, 753, 762, 339]. function-based [1611]. function-on-function [1378, 1578, 914]. Functional [1195, 450, 873, 445, 448, 261, 921, 923, 649, 293, 1200, 144, 709, 920, 443, 1196, 1595, 1198, 442, 1282, 444, 451, 766, 660, 1631, 1347, 1489, 1197, 446, 489, 935, 1202, 1620, 1204, 1301, 1588, 1047, 1310, 449, 447, 771, 64, 435, 1302, 1341, 1345, 420, 441, 1048, 1573, 358, 1199, 999]. functionals [612, 1092, 82]. functioning [909]. Functions [273, 328, 205, 1119, 1088, 445, 1454, 936, 1070, 956, 585, 1154, 1030, 708, 1383, 777, 814, 1629, 40, 346, 970, 306, 416, 1490, 1630, 485, 934, 1118, 80, 461, 1065, 1308, 7, 676, 1390]. furnace [606]. fused [762]. Fusion [1192, 620, 689]. future [943]. Fuzzy [330, 395, 1570, 420].

G [1271, 1257, 1258, 176, 1433, 1533, 725]. g-and-h [1433, 1533]. G/ [725]. GA [1335]. gain [1500]. Galea [117]. Gamma [725, 756, 896, 605, 1590, 804, 1273, 1336, 374, 1098, 715, 1335, 1110, 763, 1136, 760, 294]. GARCH [1297, 672, 1449, 949, 1352, 511, 726, 283, 549]. gas [606]. Gauss [108, 170, 113, 29, 267]. Gauss-Hermite [267]. Gauss-Seidel [113]. Gaussian [1607, 1026, 695, 892, 260, 1502, 1010, 97, 1386, 329, 1301, 966, 416, 1250, 1356, 348, 1554]. GCV [960]. Gelman [379]. gene [613, 811, 889]. General [273, 281, 225, 911, 243, 1492, 1043, 35, 782, 32, 1037, 1406]. general-purpose [1037]. Generalised [267]. Generalization [188, 944, 29]. generalizations [1515]. Generalized [790, 1607, 1366, 1001, 988, 901, 288, 1133, 675, 916, 1544, 1457, 896, 1287, 1401, 115, 1137, 892, 561, 1557, 1485, 983, 1429, 658, 473, 77, 1127, 993, 1003, 1002, 1434, 785, 715, 1335, 702, 1362, 537, 1503, 105, 629, 1346, 1149, 1110, 647, 957, 705, 587, 959, 868, 192, 193]. Generating [388, 319, 764, 1469]. Generation [550, 982, 308, 337, 701, 946, 486, 740, 539, 623]. generator [46]. generics [502]. genes [691, 889]. Genetic [220, 600, 1391, 382, 619, 474, 1232, 811, 967, 895, 1570, 634, 1004]. genome [1257, 1256, 1258, 1255, 1260, 1259]. genome-wide [1257, 1256, 1258, 1255, 1260, 1259]. genomic [1403]. Genshiro [381]. Gentle [467, 298]. genuine [151]. Geoadditive [277]. geocoded [1447]. geographical [908]. geology [93]. Geometric [1621, 1340, 1227, 17, 589, 573, 894]. geometrical [48]. Geometrically [998]. geometries [837]. geometry [132, 162]. geophysical [83]. George [371]. Germ [169]. GEVA [573]. GGobi [522]. Gibbs [1246, 879, 350, 207]. Gini [1437, 1431]. given [543, 317, 131, 142, 107]. Glaciers [593]. GLMM [170]. global [517, 1604, 1451, 1021]. global-local [1451]. GLS [1088]. GMM [727, 822]. GMM-based [727]. GO [1350]. Gompertz [958]. good [1133]. Goodman [142]. Goodness [494, 1381, 648, 719, 194, 90, 1480, 78, 1621, 1338, 1552, 1421, 647]. Goodness-of-fit [494, 1381, 648, 719, 194, 90, 1480, 78, 1338, 1552, 1421]. GPM [1596]. GPU [1579]. GPU-parallel [1579]. GQL [1229]. GQL-based [1229]. gradient [684, 841, 1187, 1371, 1554]. Gram [1304]. Graph [216, 812, 1261]. Graph sampler [1053]. Graphic [466]. Graphical [1492, 817, 816, 593, 390, 1046, 1236, 425, 199, 650, 1016, 842, 119, 237]. graphically [594]. Graphics [215, 1481, 498, 120, 623, 523]. Graphs [641, 1222, 1208, 740, 559, 812, 465]. greater [1574]. Greedy [1342]. greyvalue [696]. greyvalue-oriented [696]. grinding [1141]. GRM [1450]. Group [991, 742, 139, 741, 1516, 356, 1474, 1302, 947]. group-specific [1302]. Group-wise [991]. grouped [1375, 854, 1186, 1437, 1243, 375]. grouping [22]. groups [35, 1344, 861]. growth [953, 1176]. GS [889]. GSDAR [1485]. GSDMM [1596]. guarantee [1485]. guaranteed [1055]. Guide [318]. Gumbel [41, 98, 1530].

H [117, 379, 1122, 1433, 1533]. H-relative [1122]. habitat [389]. half [997]. half-normal [997]. halfspace [1267]. Hall [379, 371]. Hall/CRC [379, 371]. Hamiltonian [1187, 912, 1035]. Hanafi-Wold [1366]. Hand [539]. Handbook [467]. handle [706]. handles [1299]. handling [1367]. Hands [614, 809]. Hands-on [614, 809]. Hard [1498]. Härdle [467]. harmonics [102]. Hastings [1223, 1340]. Hausdorff [396]. having [1148]. Hawaii [923]. Hawkes [1612]. Hazard [204, 1444, 470, 1581, 1479]. hazards [1375, 463, 1147]. health [1499, 770, 418]. healthcare [1457]. heavy [1112, 334, 1344, 1341, 588, 1111]. heavy-tailed [1112, 334, 1341]. Hellinger [793]. heplots [526]. heritability [831]. Hermite [267, 403]. Hessian [195, 596]. heterogeneity [421, 744, 922, 634, 1176]. Heterogeneous [1397, 825, 1333, 1192, 1467]. Heteroscedastic [1112, 1146, 1374, 361, 4, 1021]. heteroscedasticity [518, 857, 1152, 861]. Heteroscedastie [264]. heteroskedastic [698]. heteroskedasticity [739, 257]. heuristic [308, 1138, 1335]. heuristics [619, 549]. HGLM [357]. hhsmm [1623]. hidden [1623, 468, 501, 891, 199, 1231]. Hierarchical [1505, 1521, 1255, 307, 1407, 1527, 1169, 1057, 687, 183, 610, 636, 683, 1472, 304, 1548, 1257, 1256, 1258, 1260, 1259]. hierarchically [1454]. Higgins [431]. High [829, 1157, 747, 456, 1203, 1509, 1282, 902, 1635, 1350, 1628, 355, 641, 102, 969, 1604, 988, 1608, 637, 1266, 1565, 1118, 1145, 1262, 992, 89, 1213, 1304, 907, 1029, 1269, 968, 1464]. High-dimensional [829, 1157, 747, 1282, 1635, 1628, 969, 1604, 1608, 1266, 1145, 1262, 992, 1213, 907, 1029, 1269, 968, 1464]. high-order [102]. higher [736, 814, 583, 1007]. higher-order [736, 814, 583, 1007]. highest [1490, 1514]. highly [125, 564]. histograms [778]. Hjorth [1424]. HMMs [1576]. Holonomic [841]. homogeneity [205, 548, 1386, 877, 801, 746, 961]. homogeneous [1024]. homology [1540]. homoscedasticity [729, 1194]. Horizon [816, 817, 818]. Horvitz [554]. hospital [612]. hot [157, 5]. hot-deck [5]. House [287]. Household [1546]. households [1547]. Howard [466]. hubs [813]. hull [459, 862, 285]. Hunting [212]. hurdle [1499]. Hybrid [686, 1539, 1544, 1623, 1424, 1557, 1425, 1297]. Hydrology [175]. Hyper [347]. Hyper-rectangular [347]. hyperbolas [106]. hyperbolic [629]. hyperparameters [135]. hypoelliptic [1559]. hypotheses [359, 657]. hypothesis [526, 1072, 804, 1393, 198, 843, 1200, 1464]. hysteretic [1352].

i.i.d [726]. ICC [1249]. identifiable [835]. Identification [73, 128, 722, 405, 116, 872, 988, 908, 1263, 1052, 1467, 1583]. Identifying [326, 1070, 925, 813, 92, 573, 691, 634]. identity [761]. IDW [1539]. ignorable [1475, 1450]. II [133, 728, 1094, 807, 1185, 1381, 1590, 1424, 1473, 116, 1393, 867, 1552, 1362, 783]. III [1228, 314]. IIR [328]. ill [511]. ill-conditioned [511]. illness [906]. illness-death [906]. Illustrations [173]. image [81, 696, 1570, 406]. imagery [369]. images [46]. imaging [892]. imbalances [477]. immersive [122]. Impact [817, 909, 156, 818, 1052]. implement [502]. Implementation [182, 941, 1338, 286, 285, 108, 844, 1082, 653, 623, 869]. Implementing [436, 350, 1313]. Implied [286, 457]. Importance [268, 1188, 645, 1100, 1422]. important [691]. imposed [32]. Improve [233, 937, 203, 691]. Improved [860, 1404, 821, 661, 1184, 1113, 794, 225, 1553, 1270, 1494, 40]. Improvement [727, 872, 1323]. improvements [278]. Improving [1193, 324, 1278, 938, 990, 981, 369]. Imputation [259, 1361, 1522, 1161, 1288, 1605, 1310, 1248]. imputing [1549]. Inaccurate [1083]. INAR [1251, 1607, 1512, 864]. incidence [1594, 1023]. including [1413]. Inclusion [318, 554]. income [1547, 1431]. Incomplete [238, 191, 1631, 832]. incompletely [1447]. incorporating [448, 1208]. increasing [410]. independence [722, 1492, 562, 1016, 1084]. Independent [1105, 1273, 581, 703, 1345]. Index [921, 254, 988, 751, 1476, 1086, 1155, 1127, 1247, 1000, 1437, 631, 142, 1008, 1382, 1524, 1431, 1064, 865, 963, 1051, 1398, 1584]. indexes [1305, 930]. indicators [488, 487, 1531]. indices [719]. indirect [1263]. individual [580, 1319]. INDSCAL [584]. induced [475]. induced-order [475]. induction [1550, 651]. industrial [1456]. industry [930]. inertia [366]. infection [612]. infectious [460]. Inference [1557, 274, 1362, 1259, 1593, 644, 706, 1227, 1270, 917, 1470, 1576, 1501, 1172, 829, 1449, 1352, 902, 736, 1028, 1101, 656, 1026, 1257, 835, 950, 1473, 1256, 1321, 1258, 1311, 1502, 178, 1031, 170, 1109, 1560, 1091, 1131, 1058, 317, 1510, 1076, 1325, 1504, 1371, 1042, 1052, 1106, 579, 1255, 1260, 1534, 1458, 1149, 1165, 1340, 283, 380, 1462, 1463, 504, 987, 757, 883, 344]. Inferences [1152, 993, 632, 90, 1531]. inferential [1229]. Inferring [814]. infinite [1566, 1118]. infinite-dimensional [1118]. inflated [1044, 770, 1288, 1104, 1328]. Influence [989, 479, 1611, 219, 432, 893, 255, 370, 77, 698, 619, 553, 406, 1145]. Influential [37]. information [722, 896, 457, 1361, 448, 312, 452, 540, 570, 1500, 1567, 1577, 483, 333, 689, 1151, 539]. information-based [896]. informative [608, 741, 1329]. infrastructure [975, 641]. inhomogeneous [1026, 1070]. initial [990, 480]. initialization [1184]. injuries [1410]. injury [1406]. INLA [615]. innovation [1230, 864]. innovations [1226]. inputs [1118]. insensitive [870]. instability [1334]. installation [911]. instantaneous [771]. instrument [617]. Insurance [175, 1287]. integer [1352, 701, 1091]. integer-valued [1352, 1091]. Integrated [936, 1036, 787, 470, 166]. integrating [635]. Integration [209, 1054]. Integrative [1403]. intelligence [1275]. Intelligent [218]. Intensity [295, 791, 110, 557]. inter [366, 927, 1203]. inter-point [1203]. inter-zone [927]. Interacting [527, 855]. Interaction [1564, 1378, 119, 869]. Interactive [215, 121, 635, 127, 643, 1402, 1046, 498, 1178, 1242, 119, 125, 120, 623, 523]. intercept [178, 1134]. Interclass [398]. interdependence [1326]. interest [1280]. Interface [253, 249, 237]. intermediate [696]. internal [530]. internals [627]. Internet [247]. interpolants [856]. Interpolation [1264, 751]. interrelations [975]. Interval [401, 1150, 700, 567, 505, 392, 399, 844, 393, 395, 400, 1278, 396, 158, 778, 356, 610, 904, 848, 1570, 426, 720, 1465, 397, 1542, 1420, 1452, 1051, 971, 394]. interval-censored [1420]. interval-data-based [1278]. interval-valued [392, 778, 1542]. intervals [916, 958, 751, 728, 807, 860, 826, 866, 1099, 1536, 1289, 1095, 1384, 1611, 575, 1320, 1415, 38, 585, 820, 926, 801, 1472, 503, 1339, 1553, 753]. Interview [298, 340, 381, 478]. Intra [239, 366, 927, 930, 761]. intra-cluster [761]. intra-day [927]. intra-industry [930]. intraclass [784]. intractable [1599]. Introducing [974]. Introduction [208, 119, 542]. Invariance [407]. Inverse [1607, 189, 260, 108, 1039, 710, 859, 892, 1417, 97, 1386, 88, 845, 57, 760]. inversed [104]. inverses [332]. invertibility [1117]. investigate [1305]. Investigating [336, 1229]. Investigation [1250, 1083]. iPlots [623]. irregular [1528, 823]. IRT [183]. ISBN [176]. ISE [988]. isotonic [1123]. isotonizing [1014]. issue [1195, 452, 919, 119]. Issues [1034, 982, 436, 182, 468, 506]. item [909, 183, 534, 680, 795]. items [909]. Iterative [499, 1138, 1180, 1262, 783, 918, 959].

J [431, 379, 538, 537]. J. [467, 537]. Jackknife [503, 1250]. jackknifed [1207]. Jacobi [113, 36]. James [431, 298, 876]. James-type [876]. Jasp [306, 576]. Java [532, 251, 306, 533]. Java-based [251, 306]. Java/R [532]. Java/R-based [532]. JavaStat [532]. JavaStatSoft [482]. Joint [1182, 1160, 790, 1377, 1473, 1427, 1031, 1526, 1439, 1568, 1096, 1460, 996, 1462, 1463, 1319, 971, 750, 968, 1329]. joint-quantile [1427]. jointly [577]. judgment [1077]. Jump [1455, 1318, 1458, 1280]. jumps [1388]. just [123].

Kalman [500, 743]. Kalman-filtering [500]. Kaplan [370, 1479, 940]. kappa [801]. KDE [1297]. Keeping [229]. Kendall [343, 1611, 1170, 1326]. Kernel [102, 903, 1581, 173, 1087, 1254, 422, 1487, 492, 99, 956, 415, 1066, 1253, 703, 1371, 163, 440, 413, 514, 1345, 51, 1218, 1269, 1551, 1353, 1111]. Kernel-based [903]. kernels [1487, 1560]. KGode [1371]. Khmaladze [1338]. Kitagawa [381]. KLERC [1353]. KLIMT [635]. KmL [578]. knot [998, 1428]. knots [461]. knowing [632]. knowledge [1279, 1239]. known [1513, 1336, 1490, 1421, 146, 977, 1084]. Kolmogorov [1506]. kriging [1254]. Kronecker [26]. Kruskal [142]. Kullback [978]. Kumaraswamy [1534]. Kumaraswamy- [1534]. kurtosis [232]. KyPlot [252].

Labeled [1589]. lack [444, 1127]. lack-of-fit [444, 1127]. Lag [129, 714, 670]. Lag-length [129]. lagged [1067]. Lagrangian [1353]. land [925]. Landau [196]. Language [250, 713]. languages [626]. Laplace [508, 1360, 115, 203, 188, 385, 1507]. LARCH [569]. large [1265, 1481, 405, 1591, 1163, 638, 307]. large-scale [1481, 638, 307]. lasso [1471, 1579, 714, 674, 762, 1193, 481]. Late [748]. Latent [721, 825, 434, 1627, 1626, 799, 438, 749, 1030, 1156, 1569, 1478, 1477, 1149, 380, 1176, 968, 484]. lattice [60, 1133, 502]. lava [749]. lava-package [749]. law [791, 1597]. laws [1399]. layout [636]. Lazy [674]. LDA [1596]. learner [885]. Learning [1144, 713, 1216, 1217, 834, 635, 154, 1208, 1311, 525, 1547, 1142, 540, 570, 1619, 210, 1565, 1295, 1192, 1041]. Least [993, 40, 106, 49, 15, 1378, 1097, 1177, 1374, 878, 1366, 565, 719, 432, 437, 1580, 1368, 53, 849, 654, 1093, 71]. Least-squares [106, 1177, 53, 849, 1093]. leave [1392, 1634]. leave-one-out [1392]. leave-pair-out [1634]. left [412, 633, 840, 1465, 1420]. left-truncated [412, 633]. Leibler [978]. lemma [57]. length [876, 1456, 129, 1554]. length-biased [876]. lengths [1340]. Leone [1146]. Let [1616]. letter [509]. level [1069, 1056, 811, 1071, 360, 1178, 1411]. levels [543, 446]. Levene [877, 552]. Levene-type [877]. leverage [15, 988, 533]. Levinson [228]. Lévy [941, 684, 1560, 1164]. Leybourne [670]. libeemd [974]. Library [253]. Life [207, 63, 1557, 1609, 952, 1483]. life-testing [1483]. lifetime [896, 1017, 944, 1042, 330, 1051, 977]. lifetimes [1094]. like [1346]. Likelihood [1357, 259, 511, 219, 684, 1042, 1149, 268, 283, 632, 1280, 1136, 706, 896, 1395, 697, 645, 1287, 182, 501, 679, 736, 972, 1404, 1321, 628, 850, 945, 752, 1597, 824, 1386, 1107, 1253, 610, 1091, 463, 1131, 708, 1510, 1434, 777, 581, 1092, 30, 715, 1250, 39, 45, 67, 1508, 1388, 1339, 1214, 1332, 987, 513, 1213, 604, 659, 784, 763, 968, 644, 959, 1036, 1124, 745, 940, 1599]. likelihood- [896]. Likelihood-based [1357, 1149, 1280, 972, 1339]. likelihoods [1294]. Lima [1594]. limit [1006]. Limitations [1502]. limited [767]. limits [363, 1160, 348]. Lindley [1271, 1270, 1457, 1039, 954, 1473, 1332]. Lindley-G [1271]. line [1228, 700, 1573, 1049]. lineage [696]. Linear [325, 749, 177, 255, 292, 272, 1423, 267, 397, 238, 1464, 916, 1375, 920, 9, 1221, 26, 1119, 1287, 844, 1578, 928, 1476, 401, 153, 508, 75, 109, 687, 1026, 1489, 1482, 701, 526, 1289, 383, 1350, 138, 1115, 561, 723, 1031, 1485, 1364, 874, 1516, 425, 888, 658, 598, 473, 1020, 77, 148, 1018, 1127, 1604, 988, 197, 993, 1606, 1608, 1003, 1002, 1085, 743, 94, 933, 385, 1445, 1016, 822, 1569, 95, 57, 162, 440, 537, 472, 413, 105, 1145, 1341, 652, 720, 82, 731, 33, 1093, 1162, 48, 898, 1324, 1309, 587, 1199, 1583, 1124, 999, 1132, 1134]. linear [71]. linearity [709]. linearized [1286]. lines [861]. Linguistic [211]. Link [253, 1155, 1629, 632]. Linked [636]. Linking [245]. Liquidity [239]. LIR [844]. literature [93]. Liu [1159, 1471, 1019]. LMS [48]. load [399]. loadings [1012]. Local [443, 230, 949, 273, 177, 145, 1118, 413, 797, 1119, 527, 372, 1070, 42, 148, 1474, 1451, 731, 674, 1111]. locally [1062, 256, 339]. location [1244, 139, 1357, 446, 1384, 1598, 857, 781, 952, 564, 1554]. location-scale [1357, 1384, 781]. Locations [242, 1143]. log [1633, 1146, 1271, 1079, 26, 227, 687, 804, 1107, 370, 585, 463, 852, 952, 1016, 1113, 1413, 647, 950]. log-ACD [950]. log-binomial [852]. log-gamma [804]. log-likelihood [1107, 463]. log-linear [26, 687, 1016]. log-location-scale [952]. log-logistic [1146, 1271, 1113]. log-normal [1633, 647]. log-rank [370, 585]. log-symmetric [1079]. Logistic [265, 1120, 881, 1146, 1271, 1430, 1286, 1598, 1605, 818, 70, 50, 1207, 1113, 794, 1019, 699, 1446]. Logit [336, 985, 566, 1028, 1573, 735]. loglinear [305]. lognormal [1010]. logspectra [387]. long [730, 579, 404]. long-range [579]. long-term [730]. Longitudinal [325, 1475, 479, 1246, 1351, 83, 1401, 589, 608, 1219, 1373, 578, 1031, 1179, 915, 1058, 1003, 1002, 1063, 1160, 1394, 504, 1573, 414, 1329, 1134]. longitudinal-survival [1058]. look [1011]. Lorenz [1186]. Lorenzen [1369]. loss [1533, 1287, 885, 33, 762, 1199]. loss-based [1199]. lot [1398]. Low [84, 902, 1100, 20, 838, 910, 1398]. low-dimensional [910]. low-rank [838]. Low-storage [84]. LTS [48]. Lyapunov [103, 311].

M [117, 1257, 1258, 725]. M/ [725].

machine [1217, 241, 154, 525, 1142, 1296, 1565, 870, 1295, 514, 630]. made [410]. main [699]. maintenance [521]. Makeham [882]. making [530]. Malliavin [1489]. manifold [837, 165, 1572, 1571]. manipulation [360]. Manly [538]. Mann [555]. MANOVA [58, 1464]. manufacturing [127, 1279]. many [1339]. map [925]. mapping [610, 615]. Maps [215, 354]. Mardia [232]. marginal [308, 824, 462, 801, 1294, 140, 1452]. marginalization [1493]. marginals [879, 430, 894]. marginals/conditionals [894]. marked [490]. Market [281, 516, 716, 455, 1622]. markets [929]. Markov [1623, 1222, 468, 501, 949, 1005, 319, 891, 1562, 178, 374, 872, 743, 906, 1325, 1156, 1625, 1455, 1318, 652, 1458, 1496, 1231, 344, 1599]. Markov/semi [1623]. Markov/semi-Markov [1623].

Markovchart [1529]. Marquardt [29]. MARS [185, 461]. martingale [1338]. Mass [233, 956, 410]. massive [1510, 1446]. matched [1245]. matching [550, 1415, 1371, 691]. Mathematica [172]. matrices [837, 964, 1389, 679, 835, 836, 828, 1]. matrix [1044, 1509, 1282, 308, 1481, 257, 34, 1066, 1579, 50, 838, 1014, 1577, 57, 32, 426, 1458, 36, 1093, 27, 1624, 1]. maximal [1459]. maximization [1627, 1626, 1311, 966, 715, 977]. maximize [723]. Maximum [945, 259, 752, 1253, 1434, 288, 268, 1332, 1406, 697, 645, 1287, 628, 850, 1597, 708, 142, 45, 67, 1508, 1214, 513, 1213, 604, 763, 968]. maximum-likelihood [604]. maxLik [628]. Maxwell [1487]. may [233]. mboost [972, 809]. MCA [366, 910]. MCMC [143, 355, 341, 1104, 1602, 553, 1037]. MCS [976]. MD*ReX [245]. MDS [54]. Mean [242, 68, 450, 949, 190, 93, 280, 1027, 1610, 577, 969, 737, 166, 1379, 845, 1152, 416, 1116, 510, 1537, 918, 940, 25]. mean-conditional [1610]. mean-variance [1537]. means [1633, 16, 563, 605, 395, 1009, 578, 810, 455, 1191, 910, 346, 1022, 1478, 1477, 10, 419, 420, 1200]. Measure [236, 494, 851, 909, 1621, 893, 848, 1472, 846]. measured [1006, 1160]. Measurement [264, 1112, 1376, 1456, 1586, 1310, 1064, 1163, 1214]. measurements [611, 1475, 1143]. Measures [239, 260, 989, 1423, 961]. Measuring [692, 1326]. mechanism [529, 1056]. Median [1211, 320, 666, 1385, 80, 1508, 705]. medians [18]. meets [525]. Meier [370, 1479, 940]. Meixner [684]. melanoma [1042]. melt [593]. memorial [748]. memory [1369, 569, 404]. memory-type [1369]. mesenchymal [696]. Meta [1319, 531, 1239, 783]. Meta-analysis [1319, 783]. meta-data [531]. meta-knowledge [1239]. Metadata [155, 159]. Method [297, 992, 1444, 1188, 937, 1532, 942, 563, 605, 79, 876, 62, 1313, 1631, 655, 195, 1384, 390, 1046, 409, 425, 737, 471, 1253, 1343, 135, 1089, 886, 1092, 69, 1263, 1472, 461, 114, 1214, 349, 682, 584, 1452, 513, 957, 1081, 1469, 971, 1035, 959, 745, 484, 5, 221, 116]. methodological [1415]. methodologies [878]. methodology [1257, 1258, 1255, 1260, 1259]. Methods [1244, 274, 265, 303, 1496, 837, 611, 896, 1632, 797, 393, 595, 1209, 241, 24, 1404, 1114, 315, 1373, 517, 324, 296, 667, 178, 492, 31, 374, 1288, 1605, 1588, 551, 1310, 912, 1335, 1565, 1008, 538, 502, 616, 10, 1346, 1331, 1233, 764, 699, 968, 1036, 467]. metric [834, 1540]. Metropolis [141, 1223, 806, 1034, 1455, 1340]. microarray [541, 967, 616]. microdata [430]. Microsoft [263, 858, 1083]. milk [466]. Milledgeville [1237]. Miner [218]. minimal [917, 100, 112]. Minimax [82]. Minimization [226, 990]. minimize [918]. Minimizing [1005, 1536]. Minimum [326, 827, 793, 496, 1525, 241, 701, 604]. minimum-distance [604]. Mining [212, 209, 208, 371, 124, 900, 307]. minPtest [811]. misclassification [1234]. misclassifications [241]. Mises [1550, 648, 850, 653, 428, 1527, 702, 1503]. mislabeling [655]. Missing [333, 1299, 1361, 1364, 1288, 1605, 540, 570, 1247, 309, 1460, 1248, 171, 1323, 1324, 1450, 1624, 1531, 675]. missingness [1246, 1310]. misspecification [881, 733]. Mixed [325, 267, 249, 881, 755, 1401, 1607, 1028, 1056, 561, 1031, 1179, 1495, 473, 429, 993, 1445, 161, 994, 773, 537, 1420, 996, 1160, 275, 1268, 1309, 587, 1583]. mixed-effects [755, 1401, 1031, 773, 1160]. mixed-level [1056]. mixed-type [994]. Mixing [306, 1034, 1205, 513]. Mixture [133, 266, 321, 264, 222, 1443, 566, 697, 1017, 1449, 1157, 953, 854, 695, 1342, 1391, 888, 1296, 540, 570, 1030, 1344, 1168, 725, 903, 1250, 1520, 417, 590, 685, 1370, 783, 668]. Mixtures [1220, 1365, 756, 1360, 1586, 1043, 824, 1181, 1018, 1434, 1445, 1189]. MKPLS [440]. MLP [1539]. MMDS [54]. modal [1435, 795]. Mode [583, 438, 974, 886, 519]. Model [1582, 586, 1251, 209, 261, 921, 1591, 260, 1141, 809, 1397, 285, 899, 1600, 1121, 33, 293, 217, 222, 1044, 881, 1375, 751, 505, 905, 1221, 1444, 479, 1264, 896, 1466, 1265, 1094, 566, 860, 1527, 73, 1130, 85, 1246, 87, 83, 21, 1595, 1198, 729, 1108, 1401, 501, 1578, 589, 607, 1157, 875, 948, 928, 1377, 851, 1219, 395, 1499, 1057, 972, 1403, 507, 660, 1357, 481, 613, 1278, 1374, 1613, 730, 738, 178, 1557, 941, 183, 560, 874, 1429, 1516, 1038, 1229, 951, 1303, 473, 35, 357, 356, 1288, 180, 872, 1296, 988, 147, 650, 993, 1147, 1247, 1606, 1131]. model [135, 852, 128, 1266, 1439, 1224, 781, 906, 597, 309, 1184, 1250, 373, 997, 1569, 1042, 1033, 1380, 714, 1052, 1008, 57, 618, 32, 994, 137, 1441, 1126, 1362, 537, 1534, 794, 1508, 1328, 423, 1110, 654, 56, 282, 1420, 1462, 1463, 375, 1176, 1231, 1452, 7, 61, 1167, 1122, 1213, 521, 549, 1019, 1319, 1573, 1049, 1280, 414, 907, 1268, 1269, 480, 587, 883, 1190, 1194, 978, 1583, 1134, 793, 1234, 1045, 1136, 71, 117]. Model-based [1251, 809, 875, 948, 1219, 972, 507, 560, 1184, 1176, 1234]. Modelers [263]. Modeling [672, 1196, 356, 559, 922, 336, 1406, 1543, 1227, 1270, 1533, 1527, 490, 1287, 589, 474, 1011, 1166, 1594, 1366, 565, 719, 427, 1526, 1181, 429, 1344, 1168, 1372, 1367, 1542, 606, 630, 1160, 968, 1329]. Modelling [924, 412, 1316, 613, 1497, 272, 930, 231, 1300, 389, 920, 1475, 78, 438, 1558, 913, 357, 447, 1182, 441]. Models [266, 202, 321, 186, 220, 1535, 325, 254, 1327, 204, 265, 267, 288, 268, 323, 238, 223, 336, 133, 1146, 612, 1623, 515, 464, 9, 111, 755, 662, 1128, 976, 825, 1251, 26, 1017, 1287, 182, 1627, 1626, 468, 1392, 721, 1112, 1491, 512, 1449, 603, 739, 789, 949, 1352, 1476, 759, 1125, 1381, 421, 787, 1105, 953, 1053, 1028, 75, 276, 687, 1521, 1489, 891, 1138, 1197, 1482, 526, 315, 1337, 1289, 1315, 854, 1350, 950, 695, 1400, 1024, 799, 1311, 1502, 561, 738, 749, 1402, 1603, 1031, 1485, 569, 511, 752, 1179, 827, 1364, 770]. models [1312, 1155, 1526, 658, 4, 1173, 374, 382, 2, 562, 274, 19, 199, 256, 473, 1020, 77, 1566, 805, 818, 1018, 1104, 1127, 540, 570, 463, 923, 1608, 1030, 1003, 1000, 1002, 927, 1220, 1226, 1085, 843, 890, 553, 94, 933, 864, 1092, 1168, 785, 1504, 970, 903, 779, 1349, 1445, 912, 1016, 1156, 407, 1442, 1423, 670, 822, 1585, 726, 1106, 1581, 1520, 773, 1438, 931, 849, 65, 417, 590, 224, 842, 1447, 105, 1145, 1341, 633, 720, 1064, 1149, 1370, 403, 114, 1201, 922, 800, 283, 865, 1163, 963, 1192, 380, 33, 996, 1162, 305, 668, 275, 898]. models [1596, 1153, 1324, 1443, 1090, 1523, 1309, 869, 985, 1199, 980, 735, 1554, 644, 959, 1124, 868, 1411, 484, 999, 1132, 1531, 129, 537, 339]. modern [623]. modest [530]. Modified [1525, 113, 1183, 1339, 1019, 784, 1130, 851, 1385, 960, 1063, 882, 552, 1368, 1508, 571]. modified-Weibull [1183]. module [911]. Moment [671, 403, 1079, 1125, 1086, 1001, 822, 1452]. moment-type [1079]. Moments [629, 258, 135, 1180, 882, 13, 1007, 376]. monitor [577]. Monitoring [483, 717, 153, 1529, 460, 712, 718]. monotone [772, 1420]. monotonic [796]. Monte [374, 538, 1078, 1005, 1313, 81, 1311, 602, 178, 361, 572, 1187, 708, 535, 698, 743, 912, 1156, 931, 10, 652, 753, 1035, 484]. Monte-Carlo [81]. Mori [467]. mortgages [386]. Most [1363, 1069]. motion [1501]. mountains [593]. Moving [1436, 543, 603, 787, 130, 1400, 34, 820, 140]. moving-average [603]. MR1392535 [117]. MR1421635 [112]. MRI [1282]. MSE [544]. MSRL [885]. mTEXO [1212]. much [1574]. Müller [176]. Multi [662, 319, 621, 920, 1632, 1361, 1266, 973, 440, 885, 1496]. multi-case [1632]. multi-class [1266]. multi-functional [920]. multi-kernel [440]. multi-loss [885]. Multi-objective [621]. multi-reader [1632]. Multi-regime [662]. multi-stage [973]. multi-state [1496]. Multi-way [319]. Multiblock [434]. multiclass [635]. multicollinearity [1286]. multicomponent [1534]. Multidimensional [234, 297, 93, 18, 637, 1156, 1468]. multifractality [855]. multigrid [1440]. Multilevel [265, 182]. multilinear [405]. Multimodal [1500]. Multinomial [336, 427, 566, 607, 863, 752, 827, 1342, 907, 735, 484]. Multiple [759, 1289, 259, 1397, 1034, 354, 1310, 242, 780, 364, 44, 37, 1403, 181, 799, 1031, 874, 425, 572, 198, 1288, 1560, 1605, 626, 1058, 1223, 901, 806, 1569, 646, 1033, 1139, 1267, 1318, 979, 783, 889]. multiple-output [646]. multiple-try [1223]. multiplicative [1147, 1122]. multiplicity [378]. multiresolution [1306, 1032]. Multiscale [284]. multistage [1522, 978]. multistate [612]. Multivariate [1607, 273, 878, 1086, 1547, 188, 776, 235, 226, 72, 1436, 1470, 108, 1351, 672, 232, 765, 21, 1491, 399, 388, 1244, 308, 1283, 24, 421, 1015, 594, 115, 1521, 526, 666, 517, 1541, 1202, 390, 1603, 752, 1020, 820, 1154, 540, 570, 596, 1220, 354, 1349, 1423, 1007, 45, 773, 1302, 654, 349, 731, 1462, 1463, 420, 987, 1231, 1573, 1583, 745, 71]. Music [810]. Mutable [625]. mutation [1232, 895]. mutual [722]. MVP [498].

Nadarajah [508]. naive [427]. Nakagami [892]. Nakagami-generalized [892]. Naomi [465]. National [1618]. navigational [641]. Near [1291, 679, 21, 777]. Near-exact [679, 777]. nearest [157, 813, 1539]. Negative [225, 1044, 733, 1515, 1629]. neglected [1068]. neighbor [1539]. neighboring [553]. neighbour [157, 813]. Neighbourhood [788]. nested [464, 359, 1454, 1301, 546]. Network [211, 217, 1046, 1187, 696, 1426, 1262, 1443]. networks [1543, 722, 310, 690, 42, 1600, 1511, 1218, 1045]. Neural [1187, 211, 217, 1543, 310, 1443]. Newton [837, 1485, 868]. next [623]. next-generation [623]. Neyman [730]. NHPP [791]. No [444, 766, 117, 663, 112, 47]. node [1261]. nodelink [636]. noise [545, 43, 55]. noises [1193]. noisy [1119, 1432, 317, 823]. Nomclust [1548]. nominal [1548]. Non [1563, 204, 1030, 272, 796, 105, 514, 1044, 1271, 1543, 1436, 515, 1475, 359, 984, 1392, 928, 949, 109, 542, 613, 1024, 1054, 1540, 1068, 1229, 951, 374, 598, 1474, 1604, 619, 743, 906, 416, 822, 162, 1248, 440, 1414, 1285, 304, 652, 1345, 1450, 774]. Non- [105]. non-Archimedean [1540]. non-Bayesian [1271, 374]. Non-constant [204]. non-continuous [984]. non-convex [1604]. Non-crossing [514]. non-equilibrium [1054]. non-essential [1285]. non-factorized [1392]. non-Fourier [542, 1414]. non-Gaussian [416]. non-hierarchical [304]. non-homogeneous [1024]. non-ignorable [1475, 1450]. Non-Linear [272, 928, 109, 743, 822, 440, 652]. non-local [1474]. non-Markov [906]. Non-monotonic [796]. non-negative [1044]. non-nested [359]. non-normal [951, 598]. Non-parametric [1563, 1030, 1543, 1248, 1345, 774]. non-perfect [619]. non-sequential [515]. non-stationarity [949]. non-stationary [1436, 359, 1563, 613, 1068, 1229]. noncentral [1055, 169, 760]. noncentral- [1055]. noncentrality [760]. nonignorable [1246, 1323]. Noninformative [1025, 656, 1026]. noniterative [495]. Nonlinear [755, 268, 630, 662, 3, 1352, 1337, 8, 165, 76, 878, 170, 2, 1580, 1126, 403, 1531]. nonlinearity [606]. nonnegative [1111]. nonnormal [1011]. Nonparametric [807, 261, 1095, 239, 326, 431, 274, 272, 906, 1412, 571, 609, 1465, 408, 959, 336, 1087, 110, 755, 411, 1265, 83, 1166, 378, 518, 1393, 1038, 360, 471, 1288, 1461, 758, 537, 510, 1346, 840, 82, 1205, 513, 1067, 659, 1443, 489]. nonparametrics [28]. Nonsingular [1050]. nonsmooth [1308]. nonstationary [662]. nonuniform [337, 909]. nonzero [734]. norm [839]. Normal [266, 289, 1397, 1636, 598, 223, 1633, 1221, 1475, 765, 1392, 1105, 508, 1360, 1494, 656, 1482, 93, 669, 1043, 1027, 1592, 888, 38, 951, 821, 570, 993, 1434, 1220, 1152, 1226, 707, 1168, 597, 385, 997, 773, 10, 647, 408, 36, 1189, 348, 1464, 1136, 71]. normal-gamma [1136]. Normal-Pareto [289]. normalised [290]. normality [21, 1541, 20, 390, 1393, 661]. normalizing [879]. normally [16, 552]. Northwest [300]. note [567, 1222, 1549, 276, 687, 575, 53, 590, 653, 1049, 414]. notes [54]. notification [1130]. notions [451]. NOx [446]. NPMLE [1420]. nuclear [1118]. null [657]. number [1375, 501, 1244, 241, 1334, 1574, 1018, 650, 862, 1284, 796, 1158, 1145, 1174, 1469, 1523]. numbers [196, 740]. Numerical [206, 1088, 535, 349, 1080, 597, 80]. Numerically [1007].

Object [250, 152, 237]. object-based [237]. Objective [1108, 1027, 1429, 1516, 1225, 1103, 621, 533]. objects [1548, 625]. Oblique [1550]. observational [691]. observations [392, 656, 832, 20, 778, 317, 1292, 1339, 12, 694, 1190, 1624]. observed [754, 1559, 1154, 1577, 1625, 282]. obtained [178, 732]. Obtaining [777, 1382]. odd [1146, 1271]. ODE [1092]. Oil [817, 816, 818]. OLS [1088]. omnibus [744, 545, 661]. One [16, 1308, 519, 1392, 729, 79, 935, 577, 311, 356, 1182, 579, 1065, 571, 1464]. one-dimensional [79, 311]. One-mode [519]. one-sample [579]. One-sided [16, 1308, 1065]. one-way [935, 356, 1464]. Online [383, 287, 1613, 1311, 1007]. Open [525, 521]. open-access [521]. Open-source [525]. OpenCL [1105]. OpenML [1217]. operating [1435]. operations [524]. opinion [389]. opinions [1238]. Optimal [1272, 1361, 1559, 378, 946, 22, 1067, 1132, 1446, 881, 3, 1291, 1244, 1529, 1135, 1589, 1179, 1426, 101, 1362, 774]. Optimization [321, 279, 3, 1550, 1277, 1141, 900, 874, 1610, 838, 1348, 1335, 1118, 1413, 775, 638, 1572, 1571, 1144]. optimized [134, 1210, 637, 86]. optimizing [973]. Optimum [1056, 131]. option [280]. optional [725]. options [457, 607]. order [807, 736, 1211, 475, 574, 1482, 601, 667, 20, 1487, 983, 102, 1459, 128, 814, 583, 943, 867, 882, 670, 1207, 13, 1007, 702, 849, 91, 1388, 28, 783, 1067, 718, 717, 1090]. Ordered [1009, 877]. ordering [597]. ordinal [1466, 335, 721, 1219, 803, 493, 1394]. Ordinary [496, 1317]. organizing [354, 480]. orientation [769]. Oriented [278, 250, 1279, 696, 252]. Ornstein [1505, 754, 1561, 1165]. Oronsay [126]. orthant [669, 36, 1479]. Orthogonal [107, 1389, 701, 1402, 584]. orthogonalization [1304]. Oscillating [1501]. Other [175, 466]. OU-processes [1399]. our [539]. outbreaks [595]. outcomes [1521]. Outlier [189, 1380, 446, 116, 872, 354, 649]. Outliers [151, 1314, 24, 92, 149, 425, 1212]. Outliers- [151]. outpatient [1277]. outperforms [1565]. output [1156, 646]. outputs [1301]. over-dispersion [1229]. overall [937, 1276]. Overcoming [1558]. overlapping [519]. overview [437, 441].

P [489, 1428, 1251]. P-splines [1428]. P2P [1576]. Package [1212, 1623, 754, 1217, 955, 1169, 1322, 1307, 1529, 527, 526, 1202, 628, 809, 460, 749, 642, 941, 531, 974, 172, 973, 732, 779, 1371, 1548, 1483, 992, 624, 1300]. packages [972]. PADM [1610]. pair [1634, 1523]. pair-elliptical [1523]. paired [331, 1495, 769]. PairViz [642]. Pairwise [1131]. panel [1125, 1115, 1402, 664, 930, 1441, 931, 1363]. panic [123]. Papers [299]. ParallAX [124]. Parallel [185, 576, 1309, 124, 1222, 1481, 614, 123, 31, 1579, 1096, 1007]. parallelizing [524]. Parameter [1039, 754, 442, 1635, 204, 743, 428, 1457, 1272, 1433, 860, 645, 468, 671, 1559, 1536, 677, 1373, 1613, 884, 850, 900, 1320, 1391, 329, 1109, 585, 1098, 1063, 943, 1629, 1250, 534, 1335, 1284, 1052, 1106, 1520, 1113, 712, 590, 1530, 734, 653, 33, 604, 1359, 1358, 564, 763, 1035, 760, 192, 193]. Parameter-wise [1635]. parameterized [1603]. Parameters [222, 1375, 41, 505, 896, 1430, 98, 1454, 1317, 1336, 1025, 102, 1386, 409, 382, 857, 1349, 795, 1508, 146, 1214, 51, 753, 1231, 305, 947, 750]. Parametric [411, 1543, 917, 709, 1017, 1198, 563, 605, 739, 1499, 75, 1562, 1563, 798, 1393, 1608, 1030, 1126, 1248, 773, 1438, 931, 1345, 1553, 419, 1090, 774]. parent [769]. parent-child [769]. parents [1459]. Pareto [318, 1480, 697, 289, 1051, 750]. Partial [920, 1367, 1378, 1366, 565, 719, 723, 432, 437, 1180, 551, 1274, 28, 999]. Partially [631, 1375, 1476, 1559, 1482, 835, 1031, 1364, 1003, 1002, 1577, 1625, 1341, 1192, 1324, 980, 1124, 1132, 1531]. particle [900, 1348, 1504, 126]. particularly [82]. partition [678]. Partitioned [1452, 921, 1321]. Partitioning [209, 347, 1589]. Partitions [271, 1013, 1071]. past [791]. Path [438, 1054, 1366, 1558, 565, 719, 1367]. patient [1529, 1319]. pattern [398]. Patterns [328, 490, 542, 52, 1621, 573, 1546, 482]. PBoostGA [1004]. PC [803]. PCA [144, 165, 832, 1342, 1204, 1281]. PDE [285]. Pearson [1013]. Penalised [506, 599]. Penalized [1528, 566, 914, 219, 1364, 824, 580, 1441, 496, 1395, 913, 429, 1604, 1442, 1355, 685, 514, 1214, 1394, 1213, 823]. penalizing [796]. penalties [230]. penalty [1462, 1463]. Percentage [292]. Percentile [707, 68]. percentiles [262]. Perfect [317, 1077, 619]. perform [1535]. Performance [768, 1078, 1193, 1632, 182, 1315, 1502, 1619, 1184, 822, 691, 96, 1331, 1176, 1051]. performing [554, 974, 1569, 889]. Periodic [1226, 1563]. periodically [1089]. periods [282]. Permanents [887]. Permutation [1495, 1574, 654, 444, 1074, 361, 235, 1171, 10]. permutations [46, 1096, 1294]. persistent [1540]. personal [622]. perspective [1295]. perspectives [1527]. Peru [1594]. Peskun [981]. pharmacokinetic [1041]. Phase [855, 1312]. phylogenetic [1340]. Physica [176]. physical [788]. physics [1305]. Pickands [1087]. piece [598]. piece-wise [598]. pipeline [522]. Pitman [943, 867]. pivot [971]. placement [1428]. PlackettLuce [1300]. PLAD [439]. plaid [1157]. plan [1148, 1524]. Planar [23]. plans [515, 1398]. platform [1217]. Playfair [558]. plot [1435, 886, 1284]. plots [574, 990]. PLS [436, 438, 1558, 433, 435]. PLSR [440]. Plug [471, 72]. Plug-in [471, 72]. plumbing [523]. Plus [173, 335, 364, 121]. point [1444, 402, 490, 1138, 799, 52, 1621, 575, 1592, 572, 872, 317, 1168, 1133, 1139, 979, 1203, 89, 657]. Points [278, 1397, 242, 15, 780, 1244, 59, 988, 630]. Poisson [112, 318, 1457, 110, 982, 871, 1130, 1607, 851, 1024, 1487, 1179, 233, 1303, 1288, 1512, 1168, 1042, 100, 1023, 1281, 323, 375]. Poisson-generalized [1457]. Polasek [748]. policies [1407]. polychoric [1011, 45]. polygon [86]. Polynomial [1628, 329, 999, 1062, 1549, 1586, 372, 491, 228]. polynomials [772]. polyserial [45]. polytomous [909]. Pooled [917, 462, 693, 783]. pooling [312]. Population [376, 563, 474, 600, 387, 510, 472, 1176, 1134]. populations [1473, 888, 38, 1598, 1152]. Port [1316]. Portfolio [279, 826, 1610]. portmanteau [78]. positive [765, 1521, 676]. positron [82]. possibilistic [1367]. possible [410]. possibly [1038]. Post [1156]. Post-processing [1156]. Posterior [179, 178, 1107, 223, 851, 183, 90]. posteriors [495]. Potential [988, 1406, 423]. potentials [1204]. poverty [1594]. Power [942, 1209, 802, 496, 791, 1055, 1006, 1597, 951, 1020, 1178, 707, 724, 1171, 1111]. power-divergence [496]. power-exponential [1111]. power-normal [707]. powerful [1115]. pp [177]. Practical [960, 347, 1414]. Practice [187, 489]. praise [922]. pre [712]. pre-change [712]. precipitation [1615]. precision [1481, 1579, 1624]. Precomputing [1035]. preconditioner [1440]. predict [282]. Predicting [1248, 1143]. Prediction [1544, 958, 1094, 191, 610, 287, 217, 1193, 807, 1401, 866, 1099, 296, 1278, 1498, 926, 1460, 618, 885, 1539]. Predictive [1406, 179, 813, 510]. predictor [1052, 1048]. Predictors [1586, 1435, 1050, 943]. preference [769]. preferences [985]. Preliminary [1173, 1022, 1393]. presence [133, 392, 1532, 1381, 1286, 1148, 1212, 551, 1310, 1126, 1151]. Presentation [200]. preserving [79, 410]. presmoothing [492]. prespecified [1012]. Press [263, 1582]. Pretest [1633]. prevalence [968]. price [855, 926, 403]. Prices [287, 927]. pricing [607, 280]. primary [485]. primer [615]. Principal [278, 400, 195, 158, 778, 240, 939, 1012, 991, 476, 1436, 1389, 1175, 678, 831, 556, 1613, 1580, 596, 449, 64, 1073, 649, 1174, 358]. Principles [179]. Prior [67, 1518, 656, 1026, 448, 1567, 1568, 406, 590, 469, 1268]. priors [586, 1025, 1027, 1474, 1451, 1396, 1029]. probabilistic [1293, 42, 893, 180]. Probabilities [318, 554, 655, 669, 863, 332, 781, 906, 36, 894]. Probability [574, 233, 1234, 826, 1513, 975, 118, 946, 695, 956, 356, 334, 782, 966, 410, 1490, 753, 975]. probit [752, 1131, 843, 907, 1269, 484]. problem [1416, 243, 139, 445, 1009, 1177, 8, 935, 1225, 100, 112, 235, 1139]. problems [768, 411, 710, 81, 1417, 550, 1558, 398, 361, 874, 1305, 1604, 29, 966, 171, 775, 7, 764]. Procedural [250]. procedure [751, 1196, 1366, 811, 1320, 938, 30, 979, 1051, 657, 774]. Procedures [184, 378, 92, 181, 81, 129, 1418, 858]. Proceedings [174, 688, 808, 1140, 1215]. Process [294, 225, 323, 989, 1365, 402, 1505, 386, 450, 754, 1401, 589, 594, 130, 155, 1597, 577, 357, 1512, 1301, 1168, 100, 112, 483, 536, 1524, 1165, 1388, 375, 1398, 947, 1612]. processes [543, 515, 1230, 284, 92, 1562, 1399, 613, 879, 684, 1327, 34, 1091, 317, 1164, 1356, 1478, 1477, 140, 472, 1561, 1274, 1318, 1458, 1102, 717, 718]. processing [1481, 156, 1156]. Procrustes [790]. product [26, 316, 559, 1440]. production [155]. products [409, 1051, 1398]. Professor [748]. profile [784]. Profiles [218]. profiling [442, 399]. prognostic [965]. program [111, 974, 1178]. Programming [288, 701, 138, 1413, 894]. Programs [318]. Progress [99]. progressive [1544, 1094, 807, 1185, 1590, 1101, 1424, 1473, 1557, 781, 867, 1051]. progressively [954, 1552, 1362, 1534]. Projection [297, 451, 118, 1072, 1402, 673, 1305, 646, 682]. projection-based [451]. projections [1509]. propagation [1230]. propensity [1209]. properties [1271, 1457, 1525, 727, 660, 975, 1366, 569, 1043, 18, 437, 1305, 795, 1263, 161, 64, 1093]. property [1208]. Proportion [1151]. proportional [463]. proportions [742].

Proposal [775, 141, 530]. proposals [1223]. proposed [1369]. Prospective [908]. Prospects [624]. proteomics [885]. pruned [1139]. Pruning [213]. Pseudo [196, 1179, 1602, 45, 1596, 1077, 561, 1004]. Pseudo-Bayesian [1179, 1602]. pseudo-boosting [1004]. Pseudo-document [1596]. pseudo-models [561]. pseudo-samples [1077]. PSIR [296]. PSO [1335]. publishing [521]. purpose [911, 1037]. Pursuit [297, 118, 1305]. Putting [1236].

QR [932]. Quadratic [1093, 66, 1290, 50, 348]. Quadrature [267, 721, 203]. quadrivariate [669]. qualitative [324]. qualitatively [138]. Quality [640, 616]. Quality-aware [640]. Quantal [186]. quantification [1432, 901, 1047, 1285]. Quantifying [389]. Quantile [876, 362, 1162, 1351, 1062, 739, 1161, 1518, 1594, 1427, 84, 1364, 1609, 664, 904, 1247, 915, 1003, 1002, 1063, 1428, 1629, 926, 646, 1355, 503, 514, 681, 880, 995, 1201, 963, 996, 1160, 1394, 1462, 1463, 1081, 1049, 1124, 868, 1531]. quantile-based [1428]. quantile-specific [1518]. quantiles [1119, 951, 646, 169]. quantization [1144]. quantum [1229, 1602]. Quasi [1091, 1321, 1577, 959]. quasi-information [1577]. Quasi-likelihood [1091, 1321, 959]. queue [725]. queueing [557]. Quicksort [1634]. quotient [837].

R [529, 537, 1623, 754, 1217, 955, 1169, 787, 1322, 972, 1307, 1529, 614, 526, 454, 1202, 628, 809, 460, 338, 351, 525, 1385, 530, 531, 522, 886, 973, 732, 779, 1371, 1372, 1618, 1548, 1483, 992, 624, 524, 1300, 533, 625, 1153, 869, 424]. R-based [532]. R-mode [886]. R-package [955]. R-packages [972]. radial [1419, 676]. rain [1615]. rainfall [1264]. Rand [1584]. Random [266, 9, 982, 1298, 1293, 992, 1509, 1246, 1532, 450, 196, 308, 190, 902, 508, 1360, 1026, 337, 548, 1535, 481, 1072, 1289, 46, 141, 550, 1010, 178, 945, 167, 1354, 329, 356, 350, 1247, 1131, 1439, 69, 168, 385, 1047, 862, 309, 855, 1263, 1472, 472, 840, 734, 568, 1122, 1469, 1326, 774, 1134, 1531, 1599, 1045]. random-effects [1246]. Randomization [538]. randomized [916, 477, 1430, 1394]. randomly [1109]. randomness [52, 619]. range [855, 579]. Rank [1159, 916, 837, 802, 964, 1211, 1611, 370, 585, 838, 882, 1572, 1571, 1]. Rank-based [1159]. ranked [1555, 1095, 742, 1385, 1076, 1150, 1330, 1508, 1530, 1332, 1331, 1553, 769, 1151]. Ranking [711, 1077, 1210, 1347, 409, 1004, 1387]. rankings [1300]. ranks [711]. rarefaction [1538]. Rasch [909, 7]. Rasch-model [7]. rate [896, 1569, 1042, 895, 1443, 1280, 793]. Rates [226, 937, 1513, 181, 1393, 1023, 404]. rating [825, 1370]. ratings [1410, 1394]. ratio [1416, 679, 1025, 684, 1386, 1107, 777, 581, 39, 1023, 757, 957]. ratio-of-uniforms [957]. rational [856]. ratios [38]. Rayleigh [673, 1362]. RBF [211]. re [938]. re-starting [938]. reader [1632]. Real [1279, 1228, 1511]. Real-manufacturing-oriented [1279]. reality [302, 122]. realized [912]. Really [254]. receiver [1435]. recognition [1570]. recommendations [1574]. reconstructing [423]. reconstruction [1631, 406]. record [41, 860]. records [1272, 807, 867, 430]. recovery [81]. rectangle [52]. rectangular [347]. Recurrent [928, 1526]. recursion [228]. Recursions [879]. Recursive [863, 1345, 1231, 1470, 810, 57, 1139, 940]. reduce [1081]. Reduced [910, 1468, 1447, 396, 1572, 1571]. Reduced-bias [1447]. reduced-rank [1572, 1571]. reducible [1177]. Reducing [763, 415]. Reduction [234, 991, 1254, 1040, 620, 1610, 845, 433, 145, 1301, 990, 1205, 1048, 889]. reference [1585, 504, 1464]. refined [911]. reflections [622]. regime [662]. region [811]. region-level [811]. regions [1272, 411, 459, 786, 191, 1337, 1490, 1514, 646, 1460, 423, 750, 823]. Regression [266, 187, 260, 219, 157, 277, 177, 255, 1333, 292, 720, 238, 184, 989, 1087, 1146, 1270, 108, 104, 905, 1159, 443, 1128, 1130, 1378, 1351, 1062, 436, 434, 932, 1595, 1587, 230, 444, 399, 512, 844, 1578, 955, 1430, 859, 928, 1607, 876, 1161, 759, 241, 1123, 1518, 972, 766, 1166, 1080, 75, 1026, 1521, 1197, 518, 1289, 383, 1471, 8, 1493, 1374, 488, 1502, 1427, 914, 1068, 1364, 874, 998, 956, 425, 893, 1609, 1173, 664, 471, 1066, 1050, 439, 88, 437, 2, 1181, 148, 429, 585, 1288, 1560, 1605, 1018, 1104, 1604, 1129, 904, 845, 1247, 1608, 915]. regression [747, 1063, 1266, 1085, 1224, 933, 1568, 1629, 926, 1442, 986, 997, 1120, 1207, 646, 1380, 485, 934, 580, 1355, 599, 95, 1581, 137, 1126, 101, 1438, 53, 1367, 849, 537, 510, 870, 1600, 794, 413, 1382, 487, 1065, 1508, 615, 1413, 885, 1341, 847, 979, 1064, 514, 681, 880, 995, 1542, 96, 654, 1201, 82, 963, 1083, 630, 731, 992, 996, 1160, 1162, 1394, 1462, 1463, 48, 674, 61, 898, 1048, 1067, 1122, 1537, 861, 918, 1019, 1081, 1049, 699, 1507, 1572, 1571, 1199, 1061, 644, 1124, 745, 868, 1353, 999, 1134, 1446, 192, 193]. Regression-based [157, 1080]. regressions [359, 1119, 866, 1099, 1586, 313, 596]. regRSM [992]. regular [60, 800]. Regularised [236]. regularity [1035]. Regularization [136, 710, 839, 79, 1471, 890, 1284]. Regularized [831, 1565, 1543, 1471, 1485, 1266]. regulations [875, 948]. regulatory [613]. reinforcement [210]. Reisensburg [688, 808, 1140, 1215]. reject [1576, 765]. rejection [1298, 316, 1455]. rejection-free [1455]. Rejoinder [1260]. Related [1195, 284, 892, 70, 966]. Relations [258]. relationship [741]. relationships [516, 927]. relative [1416, 852, 1122]. relevance [630]. relevant [1057, 1069]. Reliability [207, 555, 1534, 1555, 1128, 1312, 1320, 269, 1148, 1478, 1477]. remote [527]. removals [1185]. renewal [1478, 1477]. repair [917]. reparameterization [3]. repeated [611, 1423, 80]. replicated [1112]. replication [47]. Replicationfree [101]. Reply [509]. repositories [527]. Representation [248, 23, 1389, 1383, 1434]. representations [1230, 1164, 559, 1479]. Reproducible [626]. Requirements [120]. Resampling [1076, 811, 31]. rescaling [1522]. Researcher [458]. residual [1609]. Residuals [270, 289, 726]. resistant [15]. resolution [485, 934]. resources [767]. respect [61]. Response [186, 670, 1430, 658, 596, 534, 680, 795, 1601, 1324]. responses [1041]. restarting [1580]. restricted [1395, 1220, 57, 1019]. restriction [1433, 32]. result [88]. Results [122, 68, 206, 81]. retirement [1499]. return [516]. returns [672, 928]. Reusable [651]. reversion [280]. Review [418, 466, 465, 489, 176, 467, 431, 263, 201, 538, 537, 173, 539, 74, 379, 830, 371]. revisions [153]. revisited [126]. Revisiting [772]. RGCCTranslationUnit [531]. rhythmic [542]. richness [1538]. Richter [431]. Ridge [735, 989, 1286, 1568, 1120, 1207, 794, 1158, 1382, 12, 918, 1061]. Riemannian [837, 838]. right [1017, 1532, 876, 1381, 1615, 581, 703, 1060, 633, 840, 1332, 783, 1411]. right-censored [1017, 876, 633, 1332, 1411]. Ripley [52]. Risk [212, 418, 976, 281, 279, 516, 826, 1276, 1408, 602, 1610, 852, 1297, 1325, 681, 757]. risk-return-volatility [516]. risks [1185, 1557, 597, 1362, 1465, 1149, 965, 1319]. RMSE [1536]. RMSE-minimizing [1536]. Robbins [465]. Robust [352, 1230, 1125, 451, 1374, 832, 429, 1566, 1018, 988, 1000, 1224, 1356, 987, 1523, 1199, 1124, 1134, 1555, 1528, 1527, 494, 78, 139, 395, 677, 1373, 601, 383, 6, 1385, 960, 94, 758, 986, 599, 137, 1065, 649, 1341, 96, 14, 861, 1329, 1583]. Robustness [881, 320, 1228, 878]. ROC [1307, 1114, 723, 1634, 1144]. ROC-optimization [1144]. Rojas [117]. role [1014, 771]. ROM [371]. root [434, 1088, 1563, 1010, 686, 1107, 670]. roots [1115, 1236, 426]. rotations [550]. roughness [230]. rounded [59]. Row [869, 142]. Row-column [869]. Rubin [379]. Rule [211, 143, 486, 900, 210, 539, 1345]. rule-based [210]. Rules [213, 214, 392, 724]. run [557]. running [1505]. runs [740, 724, 718, 717, 947].

S [508, 891, 181, 1050, 263, 379, 173, 335, 364, 121]. S-estimation [891]. S-Plus [173, 335, 364, 121]. SA [1335]. Saddlepoint [704, 916, 353, 342, 225, 597]. safety [1410, 1118]. SAGE [427]. salesman [1009]. SamP2CeT [1178]. Sample [288, 1153, 1416, 104, 243, 459, 727, 902, 139, 445, 1069, 1562, 361, 1495, 1393, 290, 969, 38, 18, 1574, 1379, 1178, 346, 862, 746, 579, 235, 224, 1171, 1023, 96, 1203, 961, 793, 25]. Sampler [207, 344]. samplers [350]. samples [1077, 359, 728, 1484, 388, 308, 98, 1095, 792, 178, 1206, 116, 1333, 1212, 1076, 581, 703, 631, 726, 1534, 693, 1390]. Sampling [318, 190, 1636, 268, 1555, 1486, 911, 515, 1222, 1246, 645, 1522, 384, 1105, 1513, 1293, 1009, 1494, 1100, 742, 1320, 1385, 820, 1150, 1330, 316, 362, 785, 1508, 1524, 1530, 1165, 404, 1422, 1332, 1331, 1553, 376, 1398, 1151]. Sanov [1453]. SAS [431, 371]. SASSC [635]. satisfy [20]. saturated [488, 487]. Saunders [1593, 957, 1111]. Savage [408]. SAVE [650]. scalability [291]. scalable [1265, 1007]. Scale [1032, 860, 1481, 139, 1357, 446, 1208, 1384, 1336, 1043, 1386, 888, 1434, 781, 952, 638, 307, 1551, 1554]. scale-free [1208]. scale-shape [1043, 1434]. scaled [1579]. Scaling [234, 1156, 1468]. scan [350, 683]. scans [740]. scatter [1344]. scenario [1610]. scheduling [1277]. scheme [1544, 1080, 1101, 1473, 20, 785, 1362, 1165]. schemes [806, 724]. Schimek [1257, 1258]. Schmidt [1304]. Scholes [284]. school [1439]. score [1209, 517, 1622]. scores [825, 408]. Scoring [323, 1576, 113, 513]. Scott [431]. screening [829, 1377, 1376, 1564, 1304, 1537, 897]. Scripting [250]. SDEs [1214]. Search [619, 768, 301, 1425, 1391, 480]. Seasonal [272, 1314, 1088, 1563, 923, 1167]. seasonally [1117]. Seasons [227]. Second [983, 379, 849, 1270, 1135, 1487, 725, 761, 1067, 1090]. second-degree [1270]. Second-order [983, 849, 1487, 1067, 1090]. see [299]. seeking [985]. Seemingly [277, 1224, 1600]. segmental [953]. segmentation [152, 799, 1418, 810, 696]. Seidel [170, 113]. selected [895]. Selecting [501, 199, 1484, 650, 461, 1174]. Selection [477, 220, 2, 177, 463, 288, 217, 1375, 1582, 464, 586, 1407, 1466, 833, 729, 1108, 1157, 603, 789, 1005, 1125, 1518, 716, 1197, 372, 1137, 1373, 517, 324, 129, 1374, 884, 488, 738, 1425, 770, 99, 1516, 1253, 621, 382, 692, 473, 148, 1474, 1608, 1434, 890, 781, 973, 373, 485, 934, 1585, 1052, 618, 137, 224, 842, 487, 1601, 1121, 33, 72, 898, 1153, 907, 1029, 1268, 1269, 587, 1572, 1571, 1190, 1551, 980, 978, 1004, 1554, 1411, 897, 1111, 1446]. Selectiongain [973]. Selective [486]. selectors [1066]. self [290, 1091, 354, 480]. self-exciting [1091]. self-normalised [290]. self-organizing [354, 480]. SEM [1445]. semantics [529]. Semi [1576, 739, 1562, 773, 709, 1351, 1198, 1499, 678, 1608, 1126, 1438, 931, 1319, 1199]. semi-competing [1319]. semi-continuous [1351]. semi-functional [1199]. Semi-Markov [1562, 1623]. Semi-parametric [739, 773, 709, 1198, 1499, 1608, 1126, 1438, 931]. semi-partition [678]. Semi-supervised [1576]. semicontinuous [968]. Semiparametric [1245, 1155, 913, 416, 633, 847, 980, 1017, 315, 956, 1296, 147, 843, 1439, 930, 105, 681, 1280, 1411, 793]. semiparametrics [919]. Senn [418]. sensibility [1008]. sensitive [947]. Sensitivity [835, 80, 358, 1175, 195, 1102]. sensory [1562]. sep [371]. separate [16]. Separated [222]. separately [1605]. separations [902]. Sequence [328, 1456, 712, 1612]. sequences [93, 867, 1460, 812, 423]. sequencing [1142]. Sequential [515, 1320, 712, 1282, 139, 742, 1524, 979]. Series [215, 220, 187, 270, 217, 144, 768, 751, 1436, 662, 1314, 1216, 78, 1166, 744, 1010, 1620, 1148, 1506, 923, 1117, 1164, 864, 855, 714, 1032, 1073, 1328, 282, 1167, 237]. Server [244]. Service [1275, 725, 1618]. Set [988, 1555, 1289, 1095, 1071, 742, 1385, 813, 1076, 1150, 1330, 1508, 1530, 1332, 1331, 1553, 430, 126, 1151]. SETAR [373]. sets [1457, 1069, 360, 1133, 1163, 889]. setting [127]. setting- [127]. several [144, 563, 605, 679, 92, 1336, 1027, 1386, 35, 857, 1152, 1085]. severity [1406, 412]. SGoF [942]. Shanker [1130]. Shape [204, 1189, 345, 1043, 1025, 1098, 1434, 763]. shapelet [1216]. shapes [1336]. Shapiro [1541]. shared [1110]. Sharp [142, 206]. Shiny [1618]. shooting [986]. short [557, 653, 1596, 1280]. short-term [1280]. shown [737]. Shrinkage [905, 1633, 586, 437, 346, 547, 1451, 1396, 1268]. shrinking [600]. sibling [331]. Side [947, 735]. sided [1227, 16, 669, 1065, 1308]. Sierpiński [1486]. sign [579, 962]. signal [383, 762]. signals [786, 1205]. signature [1096, 977]. signed [1107]. significance [555]. significant [543]. similar [1038]. similarities [367]. similarity [1008, 1472]. Simon [1103]. Simple [280, 332, 840, 1281, 937, 1094, 364, 262, 844, 1053, 1313, 130, 314, 830, 762, 348]. simplex [480]. Simplification [750]. Simplified [421, 1577, 211]. simulate [46]. Simulated [619, 268, 1433, 1484, 645, 1267, 775, 1214]. simulating [69]. Simulation [402, 177, 292, 1164, 911, 372, 792, 1141, 1107, 148, 1098, 317, 535, 1060, 1445, 702, 1561, 546, 568, 604, 1596]. Simulations [186, 1015, 698, 168, 380]. simulators [767]. Simultaneous [1475, 470, 904, 1390, 1337, 1506, 1281, 1572, 1571]. simultaneously [1605]. Single [921, 254, 259, 287, 751, 1476, 181, 1141, 1155, 1204, 1127, 1247, 1064, 865, 963]. single-index [751, 1476, 1155, 1127, 1064, 865, 963]. single-trial [1204]. singular [1]. Sinkhorn [887]. SIR [104, 296, 462, 650]. sites [1057]. size [902, 701, 600, 1591, 38, 1574, 1379, 1178, 1023, 126]. SiZer [368]. sizes [746, 774]. Skew [951, 1470, 1221, 1475, 1228, 953, 1360, 1482, 1541, 1043, 888, 570, 1434, 1220, 1226, 773, 22, 1189]. skew- [1470, 1189]. skew-normal [1221, 1475, 1482, 1043, 1220, 1226, 773]. skew-normality [1541]. skewed [136, 1060, 1191, 1344, 1102]. skewness [232, 494, 1360, 951, 734]. skip [1398]. skip-lot [1398]. Sliced [104, 108, 859, 88, 845]. slices [650]. slicing [104, 462]. slightly [786]. slightly-aggregated [786]. slope [80]. Small [456, 104, 359, 18, 1142, 1098, 631, 1249]. Smallest [1514]. Smirnov [1506]. smooth [936, 631, 346, 659]. Smoothed [1609, 166, 64]. smoother [197]. smoothers [200]. Smoothing [59, 1002, 173, 268, 184, 275, 676, 823, 443, 1578, 79, 75, 314, 884, 138, 1487, 925, 382, 960, 368, 703, 758, 1504, 652, 1440, 33, 345, 51, 1488, 192, 193, 1003]. Smoothness [219, 1070]. smoothROCtime [1307]. social [899]. socio [788]. socio-economical [788]. soft [1210]. Software [246, 201, 1128, 1202, 269, 1255, 638, 120, 764, 521, 1257, 1258, 1260, 1259]. sojourn [1562]. solution [23, 1244, 1139, 27]. solutions [1093, 48]. solvers [1413]. Solving [634, 30]. somatic [1142]. Some [464, 24, 1072, 52, 64, 54, 1233, 359, 134, 766, 561, 1415, 1574, 1239, 597, 39, 161, 22, 430, 1084]. somewhat [1379]. Soul [1241, 1237, 1239, 1240]. source [525, 1431]. South [1316]. Space [268, 293, 1486, 347, 799, 1311, 1391, 910, 447, 908, 1032, 1106, 1520, 846, 480, 1035]. space-filling [1486]. space-time [908, 846]. spaced [1528]. spaces [641]. Spain [386]. Spanish [516]. Sparse [1012, 839, 834, 678, 1613, 836, 828, 1218, 1269, 1571, 1624, 1254, 833, 786, 1057, 556, 1208, 1566, 1579, 890, 1600, 1451, 1281, 830, 1596, 1029, 1572]. sparsity [1040]. Spatial [1448, 176, 1620, 544, 490, 1230, 1169, 1476, 474, 1009, 594, 1594, 553, 1461, 788, 100, 112, 1447, 1617, 1163, 1539, 1102, 1573, 1488]. spatial-functional [1573]. spatially [1264, 1070]. Spatio [1372, 1265, 924, 1491, 1460]. Spatio-temporal [1372, 1265, 924, 1491, 1460]. Spearman [1326]. Special [452, 1195, 945, 919, 119]. species [1538]. Specific [281, 1518, 1016, 1581, 1302]. Specification [658, 601]. specified [308]. speckled [369]. spectra [1488]. Spectral [812, 1261, 328, 710, 1428, 1284]. Spectrometric [261, 766]. spectrum [472]. speed [1316]. Speeding [1580]. sphere [841, 1359, 1358]. spherical [933]. spherically [1603]. Spill [817, 816, 817]. Spline [184, 1578, 75, 276, 314, 429, 1606, 960, 1348, 506, 1441, 1440, 999, 192, 193]. spline-based [192, 193]. splines [856, 185, 230, 998, 913, 663, 1383, 368, 1428, 1442, 345, 676, 1488, 823]. split [221]. split-up [221]. splits [1210]. splitting [897]. sports [1418]. spot [926]. sppmix [1168].

Spreadsheet [245]. Spurious [1068, 151]. square [1010, 1013, 862, 169, 1421]. square-root [1010]. squared [297, 68, 44, 166, 33, 918, 1199]. squares [15, 1378, 1097, 1549, 1177, 878, 1366, 565, 719, 432, 437, 1580, 993, 40, 599, 1368, 53, 849, 106, 1093, 49, 71]. Stability [476, 291, 1174, 61, 692, 1190]. stabilization [511]. stabilizing [8]. stable [1399, 69, 168, 776, 1007, 1561, 568, 672]. stable-GARCH [672]. stacking [1218]. stage [1430, 1135, 1320, 1604, 973]. stages [1277, 1205]. Standard [245, 325, 1494, 574, 76, 878, 1059, 800]. Standardized [19, 726]. Standardizing [271]. Standards [222]. star [637]. starting [938]. stashR [527]. StatDataML [327]. State [268, 293, 501, 401, 1311, 1556, 781, 1092, 447, 1106, 1496, 480, 1237]. state-space [1311, 447, 1106]. States [1547, 1546, 1409]. Stationarity [410, 949]. stationary [543, 1436, 359, 468, 1563, 613, 1068, 1229, 1379]. Statist [117, 112]. statistic [1582, 679, 1386, 1013, 777, 661, 801, 683, 579]. Statistical [1405, 856, 1172, 175, 1527, 213, 1449, 1614, 787, 815, 1101, 1545, 1473, 1235, 248, 244, 250, 696, 1510, 1060, 327, 1325, 591, 253, 757, 249, 252, 728, 1128, 710, 1244, 595, 902, 1322, 155, 1321, 532, 1141, 1547, 1485, 31, 156, 269, 498, 626, 172, 858, 970, 210, 306, 1371, 895, 1263, 1295, 616, 1458, 1093, 1233, 521, 576]. Statistics [508, 208, 247, 229, 174, 622, 713, 807, 134, 392, 3, 475, 574, 601, 467, 969, 1459, 943, 867, 882, 919, 13, 408, 1121, 783, 164, 221, 431]. Statistics- [208]. status [1147]. Stein [1532, 1014]. step [1094, 594, 1557, 224, 1554]. step-length [1554]. step-stress [1094, 1557]. Stephen [418]. steps [1145]. Stern [379]. Stewart [1382]. Stochastic [1017, 1489, 259, 223, 111, 386, 789, 1559, 1080, 1177, 280, 1603, 572, 951, 374, 913, 1001, 743, 597, 912, 1106, 137, 1274, 91, 403, 977, 1136]. stock [672, 516, 403, 1622]. stopping [143]. storage [84]. storm [125]. straight [700]. strangers [533]. strata [152]. Strategies [321, 231, 440]. strategy [706, 1035]. stratified [1206, 463]. streakiness [1418]. stream [859]. strength [1312, 1429, 1320, 1534]. stress [1094, 1557, 1312, 1429, 1320, 1534]. stress-strength [1312, 1429, 1320, 1534]. Strong [1479, 1388]. structural [1193, 117, 87, 149, 1142, 796, 380, 1176, 757]. Structure [1208, 43, 281, 386, 474, 1015, 1137, 455, 1089, 553, 131, 1423, 32, 1511, 830, 1084]. structured [738, 19, 898]. structures [636, 30]. student [1439, 1392, 1181, 1164]. Student- [1392, 1181]. Student-Lévy [1164]. studentized [361, 575, 353, 270]. studies [1632, 766, 1257, 1258, 811, 1245, 691, 1255, 1260, 1259, 1160, 504, 1256]. Study [187, 892, 300, 464, 402, 1175, 352, 1499, 331, 1114, 372, 315, 367, 195, 593, 132, 127, 969, 439, 148, 1445, 1008, 1248, 931, 140, 582, 616, 1601, 546, 1102, 604, 177]. sub [546, 634]. sub-experiment [546]. sub-types [634]. subgraph [1045]. subgroup [1333, 1239, 1583]. subiterations [113]. Subject [264, 1012, 1465, 700]. subjects [1439]. subpopulation [1408]. subsample [1446]. subsamples [61]. subsampling [1050]. subsequences [1456]. subsequent [829]. Subset [220, 378, 603, 1129, 978]. Subsets [240, 37, 819, 1425]. Subspace [992, 667, 832]. subspaces [1302]. subtracting [1193]. successive [362]. sufficient [991, 990]. suitability [47]. sum [1274]. summarizing [594]. summary [612, 990]. sums [290]. super [885]. superposed [1478, 1477]. Supervised [570, 1588, 1576, 936, 1266, 1565]. Support [154, 263, 995, 928, 241, 1296, 506, 1372, 870, 681, 880]. supported [676]. surface [670]. surfaces [908]. surrogates [832, 1206]. surveillance [908, 460]. survey [1238]. surveys [333]. survival [133, 1587, 1244, 1381, 1424, 1454, 1497, 738, 1498, 1343, 1058, 1504, 779, 582, 885, 571, 840, 1465, 1496, 1420, 1483, 965, 793]. survivors [730]. SVD [1493]. SVG [454]. SVR [1297]. SVR-GARCH-KDE [1297]. swap [607, 929]. swappability [430]. swarm [900, 1348]. switching [949, 313, 440]. Symbolic [152, 163, 398, 158, 221]. symmetric [1079, 1442, 7, 735]. Symmetrizing [197]. symmetry [1078, 1419, 777, 1503]. synthetic [1599]. System [251, 248, 249, 713, 905, 667, 31, 557, 1173, 1148, 530, 855, 1143, 576, 977, 1406]. Systems [263, 911, 917, 1576, 121, 401, 103, 405, 311, 1096, 306, 1371, 638, 469]. Szego [228].

T [791]. Table

[297, 1408, 118, 562, 47, 757, 344]. Tables [319, 39, 496, 1293, 819, 1013, 142, 894, 1517]. Tackling [239]. tail [569, 631, 883]. tail-index [631]. tailed [1112, 334, 1341, 588, 1111]. tailored [1223]. tails [672, 951, 1344]. tale [1237]. Taligent [498]. talk [1616, 533]. target [543, 1466, 1565]. Targeted [884]. Targeting [1054]. taxonomy [1409]. Tchakaloff [1291]. Teaching [247]. technique [1430, 1610, 415]. Techniques [173, 315, 692, 1127, 931, 503, 164]. technologies [164]. temperature [924]. Tempered [1626, 1399, 1561, 1627]. temporal [1265, 924, 1491, 594, 561, 1460, 1372]. Ten [584]. tensor [1440]. term [730, 1280]. terms [989, 469]. Test [733, 387, 205, 1055, 26, 1196, 243, 563, 949, 679, 1347, 548, 601, 1541, 744, 1010, 1557, 1386, 370, 1338, 1173, 1148, 1020, 1127, 1379, 777, 597, 555, 661, 882, 877, 552, 1292, 683, 1421, 1171, 1524, 1203, 1021, 1090, 889, 1194, 1084]. test-control [1292]. Testing [709, 780, 1433, 21, 605, 475, 20, 663, 1085, 843, 746, 65, 91, 1509, 16, 1381, 75, 1114, 804, 1400, 602, 811, 1419, 686, 1027, 1386, 969, 1148, 198, 1574, 1212, 583, 952, 1552, 726, 1483, 1051, 657, 1200, 1326, 1464]. Tests [1077, 207, 55, 704, 916, 63, 802, 1075, 1078, 1480, 359, 16, 1532, 494, 444, 1088, 78, 727, 139, 98, 766, 1313, 1492, 1074, 526, 1072, 518, 1562, 1563, 935, 52, 361, 1419, 1495, 1393, 1225, 658, 1415, 821, 698, 864, 551, 194, 670, 1022, 39, 235, 1503, 10, 654, 28, 469, 961]. text [1596]. th [718, 717]. Their [286, 359, 1276, 168, 889]. them [1182, 299]. theorem [981, 1453]. theoretical [278]. theory [1283, 489, 836, 899, 776, 534, 680, 795, 549]. there [861, 1021]. thermodynamic [1054]. thick [1574]. third [538]. Thompson [554]. Three [328, 1435, 550, 1502, 781, 519]. three-dimensional [550]. three-modal [1435]. three-state [781]. three-way [519]. Threshold [789, 603, 739, 1435, 1615, 620, 888, 1091, 1382, 1583]. Thresholds [543, 1033]. tie [1151]. tight [560, 1575]. tilted [1107]. timbre [810]. Time [215, 1216, 220, 1312, 274, 270, 217, 144, 768, 751, 1436, 791, 505, 662, 1314, 78, 1403, 1166, 1307, 1562, 744, 1115, 1010, 1620, 560, 102, 1526, 1148, 818, 1506, 923, 864, 855, 908, 714, 1032, 618, 1581, 1441, 846, 1625, 53, 1073, 1328, 1149, 282, 1452, 1167, 237, 1280, 1194, 1612]. Time-dependent [1312, 1307, 1581, 1452, 1280, 1612]. time-heterogeneity [744]. time-series [1010]. time-to-event [618]. time-varying [1115, 102, 1441, 53]. times [1544, 1096]. title [354]. Tobit [953]. tolerance [411, 1384]. tomography [1602, 82]. Tool [212, 252, 124, 1053, 650]. tools [17, 1408, 67, 125]. topic [1266, 1596]. Topics [1195, 367]. topological [161]. Topp [1146]. torus [1349]. total [1276]. totals [142]. tour [642]. tourism [923]. tours [1305]. trace [839]. tracking [832]. trade [930]. traditional [1565]. training [768, 813]. traits [365]. trajectories [1176]. transduction [154]. transfers [410]. Transformation [274, 1344, 1158, 883, 8, 203, 1338, 633, 720, 1411]. Transformation-based [883]. transformations [1365, 21, 197]. transition [1399, 781, 906]. transmission [1143]. traveling [1009]. treatment [62, 1529, 747, 1467]. tree [611, 1299, 1505, 1550, 607, 1588, 1243, 651, 985]. tree-based [611, 1588]. treemap [636]. treemap-nodelink [636]. Trees [286, 291, 1188, 1448, 152, 1298, 399, 347, 1210, 1071, 818, 493, 683, 1176, 49]. Trellis [502]. trend [589, 853]. trends [1115, 1461]. trial [1204]. trials [477, 1178]. triangulations [823]. Trimmed [15, 459, 599]. Trinomial [286]. trout [466]. Truck [186]. true [1410, 632]. Truncated [568, 223, 791, 1533, 412, 1357, 1412, 609, 633, 720, 734, 1233]. truncation [945, 840, 1465, 1420]. Try [1034, 1223, 806, 1267]. Tukey [1082]. tuned [1037]. Tuning [1551]. turning [872]. tutorial [614, 809, 690]. Tweedie [1044, 982, 1015]. Twitter [1596]. Two [837, 1227, 518, 1562, 296, 1629, 1346, 961, 495, 699, 1416, 496, 1078, 1457, 1272, 728, 243, 671, 1108, 1430, 767, 139, 594, 1135, 1277, 1114, 1473, 669, 741, 742, 361, 1495, 1025, 1393, 1320, 1391, 969, 1109, 1598, 1379, 1604, 993, 387, 1178, 1301, 943, 867, 306, 551, 1182, 702, 140, 224, 1171, 47, 1023, 56, 1332, 1205, 1203, 1517, 861, 1021, 1551, 1390, 793]. two-component [1391, 993]. two-group [741]. two-level [1178]. Two-parameter [1629, 1457, 1272, 671]. Two-sample [961, 1416, 243, 139, 361, 1495, 1393, 1171, 1203, 793]. two-scale [1551]. Two-sided [1227, 669]. two-stage [1430, 1135, 1320, 1604]. two-step [594]. two-way [496, 47, 56, 1517]. Type [181, 133, 1369, 1079, 876, 1590, 1473, 46, 730, 1312, 1393, 1552, 877, 994, 140, 1051, 728, 1094, 807, 1039, 1381, 1424, 867, 1362, 1185, 783]. Type-I [1039]. type-II [1590, 1473, 1552, 1381, 1424, 1362, 1185, 783]. types [1403, 634].

U.S. [1619]. Uhlenbeck [1505, 754, 1561, 1165]. Ultra [1304]. Ultra-high [1304]. ultrahigh [1377, 1376, 1537]. ultrahigh-dimensional [1377, 1376, 1537]. Ultrametric [11, 1505]. ultrasound [892]. unbalanced [83, 1276, 356, 993, 32]. uncensored [1332]. uncertainties [516]. Uncertainty [1432, 1047, 706, 1250, 1367, 1370]. Unconstrained [1389]. uncover [1176]. underlying [632]. understand [902]. understanding [1547]. unequal [129, 746]. Unified [1537, 1470]. Uniform [1567, 1486, 909]. uniformity [1172]. uniforms [957]. unimodal [756]. unit [1270, 1088, 1563, 1115, 686, 862, 670]. unit-improved [1270]. United [1547, 1546, 1409]. unitizing [197]. units [1544, 1481]. univariate [73, 1202, 1490]. university [1387]. unknown [1206, 956, 712]. unobserved [421, 149, 843, 1143]. Unrelated [277, 1224, 1600]. unreplicated [541]. Unsupervised [1619, 1041, 755, 1188]. Unwin [340]. updated [822]. Updating [12, 42, 1556]. Usage [1350, 155]. Use [3, 1040, 236, 1335, 75, 8, 960, 777, 1588, 1414, 54]. used [1632, 1175, 233]. User [246, 249, 252, 1030, 237]. User-oriented [252]. Using [713, 902, 31, 577, 1305, 845, 207, 725, 263, 1585, 423, 323, 371, 328, 1555, 611, 1087, 1582, 1543, 911, 964, 515, 1077, 1509, 722, 1264, 896, 1527, 976, 1361, 1246, 1119, 1532, 721, 876, 1481, 1518, 1403, 1529, 527, 1026, 475, 1100, 517, 1471, 129, 1374, 832, 1497, 341, 809, 1071, 1487, 396, 1592, 425, 1418, 925, 439, 598, 938, 1580, 913, 1303, 357, 610, 872, 637, 708, 852, 1297, 113, 1076, 535, 698, 1164, 1428, 1092, 1168, 354, 1371, 926, 912, 1184, 1250, 1116, 1570, 990, 137, 1126, 483, 773, 503, 981, 1618, 510, 1600, 502, 559]. using [615, 1530, 885, 1341, 1388, 654, 408, 1261, 693, 125, 1162, 345, 497, 1596, 348, 676, 1583, 823]. usual [563]. utilizing [927, 689].

V [477]. valid [377, 1452]. validated [501, 547]. Validation [296, 864, 1392, 81, 741, 1426, 1461, 758, 1634, 1065, 1308, 1511]. validatory [1121, 51]. Value [976, 231, 1283, 1316, 574, 1086, 602, 1354, 1610, 925, 1279, 1297, 1426, 1570, 681, 769]. Value-at-Risk [976, 602, 1610, 1297]. valued [392, 1352, 395, 778, 1091, 1542]. Values [175, 916, 41, 1299, 860, 134, 1505, 232, 562, 1154, 342, 1460, 39, 726, 1248, 469, 965, 675, 221]. Values- [175]. Vance [1369]. VAR [129, 1315]. VAR-models [129]. variability [567, 924, 573, 483, 536]. variability-based [573]. Variable [1375, 716, 1197, 517, 738, 770, 890, 1601, 264, 234, 897, 1188, 833, 1627, 1626, 829, 721, 1157, 789, 1005, 1518, 766, 1373, 1374, 749, 1354, 998, 1516, 820, 1156, 1569, 1585, 898, 1304, 907, 1029, 1268, 1269, 1572, 1571, 980, 1004, 1411, 484]. Variables [240, 1398, 665, 1466, 16, 1128, 1287, 190, 548, 438, 1273, 741, 1006, 329, 2, 598, 199, 1574, 901, 803, 385, 1047, 1310, 1182, 493, 1243, 1478, 1477, 1158, 734, 568, 1548, 12, 700, 694, 1324, 980, 348, 1450, 71, 508]. Variance [95, 25, 9, 729, 1108, 1456, 679, 8, 1202, 577, 737, 4, 38, 356, 845, 1330, 416, 1116, 1438, 504, 1067, 1537, 764, 1390]. variance-covariance [679, 845]. variance-stabilizing [8]. Variances [1397, 1026, 535, 698, 877, 746]. variants [1142]. variate [982, 337]. variation [798, 821, 803]. Variational [652, 1422, 1028, 1560]. variogram [377]. various [1075, 595, 241, 1024]. varying [505, 479, 1264, 1115, 1364, 102, 818, 1000, 890, 1441, 53, 931, 1064, 995, 922, 1324, 980, 959, 1531]. varying-coefficient [890, 922, 1324, 959]. Vavilov [196]. VBA [263]. vector [928, 241, 154, 1400, 511, 969, 938, 1580, 1296, 506, 870, 681, 880, 995, 630, 524, 1144, 497]. vectors [1379, 1556, 69, 168, 939, 1326]. vehicle [1410]. verification [551]. Verlag [176]. Version [338, 351]. versions [359]. Versus [259, 411, 878, 1597]. vertexes [862]. VI [639]. via [1416, 586, 1505, 1549, 767, 451, 1293, 1404, 1493, 1334, 138, 1425, 1107, 853, 1288, 1579, 1605, 663, 1127, 462, 1564, 1063, 1434, 703, 1588, 814, 583, 970, 1250, 618, 994, 1458, 979, 514, 1451, 630, 234, 761, 1453, 957, 1304, 1037, 1572, 1571, 1200, 1488, 644, 1124]. Vieu [489]. view [1257, 296, 1258, 1173, 1030, 1255, 1260, 1259]. VIF [874, 979]. VIF-based [874]. vine [800]. violated [801]. violation [58]. virtual [302, 122]. virus [1057]. vision [1407, 1276]. Visions [164]. Visual [214, 638, 126, 466, 902, 307, 640]. Visualisation [456, 301, 452, 299]. visualisation-What [299]. Visualising [453]. Visualization [932, 251, 455, 216, 252, 1283, 635, 191, 642, 425, 886, 302, 306, 122, 303, 502]. Visualizing [592, 526, 637, 304, 305, 692, 636]. visually [1305]. Viz [639]. volatility [1227, 516, 789, 1563, 1278, 280, 1115, 1603, 1068, 951, 913, 912, 282, 1090]. Volpi [1083]. VOM [423]. voter [855]. vs [122].

W [176, 467, 173]. Wainer [466]. Wald [26, 727]. walk [141]. walks [1010]. warehouse [160]. warm [593]. Wavelet [167, 872, 1528, 873, 569, 547, 485, 934, 1396, 1469]. Wavelet-based [167]. way [496, 729, 1108, 319, 1347, 935, 356, 47, 56, 1517, 519, 1464]. ways [533]. Weather [1618, 1620, 1619, 1616, 1617]. web [975, 269, 248, 334]. web-based [269]. WebGIS [454]. Weibull [63, 1532, 1185, 730, 1597, 1025, 1391, 1629, 1042, 1183, 1319]. Weibull-power [1597]. Weight [320]. Weighted [186, 1097, 1247, 915, 933, 320, 1453, 706, 144, 1062, 1408, 59, 1364, 585, 820, 1170, 1567, 1250, 1042, 714, 994, 1368, 1356, 959, 1124, 25]. weights [1206, 925, 197, 933, 1007, 783, 1]. Weka [525]. Well [222, 563, 1084]. well-known [1084]. where [770]. which [151, 1535, 655, 723, 282]. white [55, 285]. whitening [1628]. Whitney [555]. wide [1257, 1256, 1258, 1255, 1260, 1259]. widely [1479]. width [38]. Wiener [169]. wiki [713]. wild [1115]. Wilk [1541]. William [558]. win [875, 948]. wind [1527, 1316]. wise [991, 1635, 598]. within [1026, 502, 1553, 521]. within-class [1026]. without [632, 897]. Wold [1366]. Wolfgang [748]. word [1322]. word.alignment [1322]. Working [1137, 531]. workshop [452]. workstation [122]. world [1387, 539]. wrap [1567]. wrap-around [1567]. wrapped [1349]. Wright [671].

X [1552, 294]. X-bar [294]. XGobi [122]. XII [860]. XML [327]. XploRe [245, 286, 218].

York [1622]. Young [458]. Yuima

[754, 941]. Yule [1103].

zag [694]. zero [1044, 1407, 1130, 851, 770, 1288, 1328, 1339, 1083, 1233, 757]. zeroinflated [1044, 770, 1288]. zero-modified [1130, 851]. zero-truncated [1233]. Zig [694]. Zig-zag [694]. zone [927].

References

deLeeuw:1984:FRM

[1] J. de Leeuw. Fixed rank matrix approximation with singular weights matrices. Computational Statistics Quarterly, 1 (1):3–12, 1984. CODEN CSQUEM. ISSN 0723-712X.

Kredler:1984:SVC

[2] Ch. Kredler. Selection of variables in certain nonlinear regression models. Computational Statistics Quarterly, 1 (1):13–27, 1984. CODEN CSQUEM. ISSN 0723-712X.

Bohning:1984:URN

[3] D. Böhning. Use of reparameterization in nonlinear optimization with applications to statistics and optimal design. Computational Statistics Quarterly, 1 (1):29–43, 1984. CODEN CSQUEM. ISSN 0723-712X.

Kleffe:1984:VCE

[4] J. Kleffe. On variance components estimation in heteroscedastic models. Computational Statistics Quarterly, 1 (3):179–204, 1984. CODEN CSQUEM. ISSN 0723-712X.

vanPraag:1984:HDM

[5] B. M. S. van Praag and B. Wesselman. The hot-deck method: an analytical and empirical evaluation. Computational Statistics Quarterly, 1(3):205– 231, 1984. CODEN CSQUEM. ISSN 0723-712X.

Haussler:1984:CEE

[6] W. M. Häussler. Computational experience with an eigenvector algorithm for robust L_{p}-discrimination. Computational Statistics Quarterly, 1(3):233– 244, 1984. CODEN CSQUEM. ISSN 0723-712X.

Verhelst:1984:EPR

[7] N. D. Verhelst, C. A. W. Glas, and A. van der Sluis. Estimation problems in the Rasch-model: the basic symmetric functions. Computational Statistics Quarterly, 1(3):245–262, 1984. CO-DEN CSQUEM. ISSN 0723-712X.

Gerisch:1984:NRP

[8] W. Gerisch and K. H. Müller. On a nonlinear regression problem connected with the use of the angular transformation as a variance-stabilizing transformation. Computational Statistics Quarterly, 1(4):271–294, 1984. CO-DEN CSQUEM. ISSN 0723-712X.

Aragon:1984:RVL

[9] Y. Aragon. Random variance linear models: estimation. Computational Statistics Quarterly, 1(4):295– 309, 1984. CODEN CSQUEM. ISSN 0723-712X.

Schrage:1984:EPT

[10] C. Schrage. Evaluation of permutation tests by means of normal approximation or Monte Carlo methods. Computational Statistics Quarterly, 1(4):325– 332, 1984. CODEN CSQUEM. ISSN 0723-712X.

Degens:1985:UAD

[11] P. O. Degens. Ultrametric approximation to distances. Computational Statistics Quarterly, 2(1):93–101, 1985. CODEN CSQUEM. ISSN 0723-712X.

Trenkler:1985:URE

[12] D. Trenkler, G. Trenkler, and B. Schipp. Updating the ridge estimator when additional explanatory variables or observations are available. Computational Statistics Quarterly, 2 (2):135–141, 1985. CODEN CSQUEM. ISSN 0723-712X.

Patil:1985:MOS

[13] S. A. Patil, K. Raghunandanan, and Ru Ying Lee. On the moments of order statistics of the F -distribution. Computational Statistics Quarterly, 2(3):285– 302, 1985. CODEN CSQUEM. ISSN 0723-712X.

Watson:1985:CEA

[14] G. A. Watson. On the convergence of eigenvector algorithms for robust l_{p}- discrimination. Computational Statistics Quarterly, 2(4):307–314, 1985. CO-DEN CSQUEM. ISSN 0723-712X.

Antoch:1985:TLS

[15] Jarom´ ır Antoch and J. Jurečková. Trimmed least squares estimator resistant to leverage points. Computational Statistics Quarterly, 2(4):329– 339, 1985. CODEN CSQUEM. ISSN 0723-712X.

Beedgen:1985:OST

[16] Rainer Beedgen. One-sided tests for a separate testing of means of bivariate normally distributed variables. Computational Statistics Quarterly, 2(4):341– 351, 1985. CODEN CSQUEM. ISSN 0723-712X.

Baufays:1986:DAG

[17] P. Baufays and J.-P. Rasson. Discriminant analysis with geometric tools. Computational Statistics Quarterly, 3 (1):1–19, 1986. CODEN CSQUEM. ISSN 0723-712X.

Kosfeld:1986:CAS

[18] R. Kosfeld. Computational aspects and small sample properties of multidimensional medians. Computational Statistics Quarterly, 3(1):21–36, 1986. CO-DEN CSQUEM. ISSN 0723-712X.

Laake:1986:SCC

[19] P. Laake. Standardized coefficients in covariance structured models. Computational Statistics Quarterly, 3(1):37– 48, 1986. CODEN CSQUEM. ISSN 0723-712X.

Heuts:1986:TNW

[20] R. M. J. Heuts and S. Rens. Testing normality when observations satisfy a certain low order ARMA-scheme. Computational Statistics Quarterly, 3 (1):49–60, 1986. CODEN CSQUEM. ISSN 0723-712X.

Bozdogan:1986:TMF

[21] H. Bozdogan and D. E. Ramirez. Testing for model fit: assessing and Box-Cox transformations of multivariate data to “near” normality. Computational Statistics Quarterly, 3(3):127– 150, 1986. CODEN CSQUEM. ISSN 0723-712X.

Schader:1986:OGD

[22] M. Schader and F. Schmid. Optimal grouping of data from some skew distributions. Computational Statistics Quarterly, 3(3):151–159, 1986. CO-DEN CSQUEM. ISSN 0723-712X.

Ader:1986:PRC

[23] H. J. Adèr. Planar representation of a cluster solution. Computational Statistics Quarterly, 3(3):161–175, 1986. CO-DEN CSQUEM. ISSN 0723-712X.

Cleroux:1986:SMD

[24] R. Cléroux, J.-M. Helbling, and N. Ranger. Some methods of detecting multivariate outliers. Computational Statistics Quarterly, 3(3):177– 195, 1986. CODEN CSQUEM. ISSN 0723-712X.

tenCate:1987:VEM

[25] A. ten Cate. Variance estimation of the mean of a weighted sample file. Computational Statistics Quarterly, 3(4):197– 205, 1987. CODEN CSQUEM. ISSN 0723-712X.

Bonett:1987:AKP

[26] D. G. Bonett and J. A. Woodward. Application of the Kronecker product and Wald test in log-linear models. Computational Statistics Quarterly, 3(4):235– 243, 1987. CODEN CSQUEM. ISSN 0723-712X.

Wang:1987:SMA

[27] C. Ming Wang. The solution matrix in ANOVA. Computational Statistics Quarterly, 3(4):257–263, 1987. CO-DEN CSQUEM. ISSN 0723-712X.

Streitberg:1988:ENP

[28] B. Streitberg and J. Röhmel. Exact nonparametrics for partial order tests. Computational Statistics Quarterly, 4 (1):23–41, 1988. CODEN CSQUEM. ISSN 0723-712X.

Messean:1988:GGM

[29] A. Messéan. A generalization of Gauss–Marquardt algorithm for exponential families problems. Computational Statistics Quarterly, 4(2):79–88, 1988. CODEN CSQUEM. ISSN 0723-712X.

Mukherjee:1988:PSL

[30] B. N. Mukherjee and S. S. Maiti. Qprocedure for solving likelihood equations in the analysis of covariance structures. Computational Statistics Quarterly, 4(2):105–128, 1988. CODEN CSQUEM. ISSN 0723-712X.

Kaufman:1988:UPC

[31] L. Kaufman, P. K. Hopke, and P. J. Rousseeuw. Using a parallel computer system for statistical resampling methods. Computational Statistics Quarterly, 4(2):129–141, 1988. CODEN CSQUEM. ISSN 0723-712X.

Prieto:1988:ASR

[32] J. M. Prieto and J. M. Caridad. Algebraic structure of the restriction matrix imposed on a general unbalanced cross-classification model. Computational Statistics Quarterly, 4(3):173– 190, 1988. CODEN CSQUEM. ISSN 0723-712X.

Terasvirta:1988:MSS

[33] T. Teräsvirta, G. Yi, and G. Judge. Model selection, smoothing and parameter estimation in linear models under squared error loss. Computational Statistics Quarterly, 4(3):191– 205, 1988. CODEN CSQUEM. ISSN 0723-712X.

Klein:1989:ACC

[34] A. Klein and G. Mélard. On algorithms for computing the covariance matrix of estimates in autoregressive moving average processes. Computational Statistics Quarterly, 5(1):1–9, 1989. CODEN CSQUEM. ISSN 0723-712X.

Lee:1989:GMA

[35] S.-Y. Lee and W.-Y. Poon. A general model for analysis of categorical data in several groups. Computational Statistics Quarterly, 5(1):11–25, 1989. CO-DEN CSQUEM. ISSN 0723-712X.

Sun:1989:CNO

[36] H.-J. Sun and C. Asano. Computing normal orthant probabilities with a Jacobi correlation matrix. Computational Statistics Quarterly, 5(1):27–41, 1989. CODEN CSQUEM. ISSN 0723-712X.

Cleroux:1989:ISD

[37] R. Cléroux, J.-M. Helbling, and N. Ranger. Influential subsets diagnostics based on multiple correlation. Computational Statistics Quarterly, 5 (2):99–117, 1989. CODEN CSQUEM. ISSN 0723-712X.

Klein:1990:ESS

[38] S. W. Klein. Effect of sample size on width of confidence intervals for variance and variance ratios in normal populations. Computational Statistics Quarterly, 5(3):171–179, 1990. CO-DEN CSQUEM. ISSN 0723-712X.

Parsian:1990:TCV

[39] A. Parsian and N. Nematolahi. Tables of critical values for some likelihood ratio tests. Computational Statistics Quarterly, 5(3):181–192, 1990. CO-DEN CSQUEM. ISSN 0723-712X.

Muller:1990:LSC

[40] P. Müller. Least squares characterization of improved estimation functions. Computational Statistics Quarterly, 5 (3):203–216, 1990. CODEN CSQUEM. ISSN 0723-712X.

Ahsanullah:1990:EPG

[41] M. Ahsanullah. Estimation of the parameters of the Gumbel distribution based on the m record values. Computational Statistics Quarterly, 5(3):231– 239, 1990. CODEN CSQUEM. ISSN 0723-712X.

Jensen:1990:BUC

[42] F. V. Jensen, S. L. Lauritzen, and K. G. Olesen. Bayesian updating in causal probabilistic networks by local computations. Computational Statistics Quarterly, 5(4):269–282, 1990. CO-DEN CSQUEM. ISSN 0723-712X.

Ohrvik:1990:SN

[43] J. Öhrvik. Structure or noise. Computational Statistics Quarterly, 6(1):1–20, 1990. CODEN CSQUEM. ISSN 0723-712X.

Chun:1990:CAC

[44] C. S. Y. Chun, O. J. Dunn, and R. I. Jennrich. Calculation of the asymptotic covariance between squared multiple correlation coefficients. Computational Statistics Quarterly, 6(1):21–30, 1990. CODEN CSQUEM. ISSN 0723-712X.

Poon:1990:PML

[45] W.-Y. Poon, S.-Y. Lee, and P. M. Bentler. Pseudo maximum likelihood estimation of multivariate polychoric and polyserial correlations. Computational Statistics Quarterly, 6(1):41–53, 1990. CODEN CSQUEM. ISSN 0723-712X.

Granville:1990:NTR

[46] V. Granville and J.-P. Rasson. A new type of random permutations generator to simulate random images. Computational Statistics Quarterly, 6(1):55–64, 1990. CODEN CSQUEM. ISSN 0723-712X.

Seyedsadr:1991:SAA

[47] M. Seyedsadr. On suitability of additivity assumption for combining effects of two factors in a two-way table with no replication. Computational Statistics Quarterly, 6(2):127–137, 1991. CO-DEN CSQUEM. ISSN 0723-712X.

Tichavsky:1991:AGC

[48] P. Tichavský. Algorithms for and geometrical characterization of solutions in the LMS and the LTS linear regression. Computational Statistics Quarterly, 6 (2):139–151, 1991. CODEN CSQUEM. ISSN 0723-712X.

Vach:1991:LSA

[49] W. Vach and P. O. Degens. Least squares approximation of additive trees to dissimilarities—characterizations and algorithms. Computational Statistics Quarterly, 6(3):203–218, 1991. CO-DEN CSQUEM. ISSN 0723-712X.

Mathes:1991:QLD

[50] H. Mathes and C. Lü. A quadratic logistic discrimination algorithm in matrix form. Computational Statistics Quarterly, 6(4):269–280, 1991. CO-DEN CSQUEM. ISSN 0723-712X.

Tutz:1991:CCV

[51] G. Tutz. Consistency of crossvalidatory choice of smoothing parameters for direct kernel estimates. Computational Statistics Quarterly, 6(4):295– 314, 1991. CODEN CSQUEM. ISSN 0723-712X.

Hermans:1988:SNT

[52] M. Hermans, J.-P. Rasson, and D. Weverbergh. Some new tests of randomness for point patterns in a rectangle, with some comparisons under Ripley’s alternatives. Computational Statistics Quarterly, 4(4):247–258, 1988/89. CO-DEN CSQUEM. ISSN 0723-712X.

Rodrigues:1988:NBL

[53] J. Rodrigues and H. Bolfarine. A note on Bayesian least-squares estimators of time-varying regression coefficients. Computational Statistics Quarterly, 4(4):259–265, 1988/89. CODEN CSQUEM. ISSN 0723-712X.

Telegdi:1988:SNM

[54] L. Telegdi. Some notes on MMDS and the use of MDS for detecting consensus clusters. Computational Statistics Quarterly, 4(4):267–280, 1988/89. CO-DEN CSQUEM. ISSN 0723-712X.

Schlittgen:1988:TWN

[55] R. Schlittgen. Tests for white noise in the frequency domain. Computational Statistics Quarterly, 4(4):281– 288, 1988/89. CODEN CSQUEM. ISSN 0723-712X.

Sposito:1988:CAT

[56] V. A. Sposito and M. B. Tirol. Computational aspects of the two-way classification model under L_{1}. Computational Statistics Quarterly, 4(4):289– 298, 1988/89. CODEN CSQUEM. ISSN 0723-712X.

Pordzik:1992:LIB

[57] P. R. Pordzik. A lemma on g-inverse of the bordered matrix and its application to recursive estimation in the restricted linear model. Computational Statistics, 7(1):31–37, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Amey:1992:MUV

[58] A. K. A. Amey and A. K. Gupta. MANOVA under violation. Computational Statistics, 7(1):39–57, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Hardle:1992:SWA

[59] W. K. Härdle and D. W. Scott. Smoothing by weighted averaging of rounded points. Computational Statistics, 7(2):97–128, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Granville:1992:DEF

[60] V. Granville and J.-P. Rasson. Density estimation on a finite regular lattice. Computational Statistics, 7(2): 129–136, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Visek:1992:SRM

[61] J. Á. V´ ıšek. Stability of regression model estimates with respect to subsamples. Computational Statistics, 7(2):183–203, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Dargahi-Noubary:1992:MED

[62] G. R. Dargahi-Noubary and A. Nanthakumar. A method for estimating and detecting treatment effects. Computational Statistics, 7(3): 301–308, 1992. CODEN CSTAEB.

ISSN 0943-4062 (print), 1613-9658 (electronic).

Achcar:1992:BAA

[63] J. A. Achcar and F. Louzada Neto. A Bayesian approach for accelerated life tests considering the Weibull distribution. Computational Statistics, 7 (4):355–369, 1992. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Pezzulli:1993:SPS

[64] S. Pezzulli and B. W. Silverman. Some properties of smoothed principal components analysis for functional data. Computational Statistics, 8(1):1–16, 1993. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Rubio:1993:TDB

[65] A. M. Rubio, L. Z. Aguilar, and J. Á. V´ ıšek. Testing for differences between models. Computational Statistics, 8(1):57–70, 1993. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Wood:1993:QED

[66] Andrew T. A. Wood. Quadratic exponential distributions for categorical and directional data. Computational Statistics, 8(2):141–159, 1993. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Robert:1993:PFB

[67] Christian P. Robert. Prior feedback: Bayesian tools for maximum likelihood estimation. Computational Statistics, 8(4):279–294, 1993. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Banks:1993:ERB

[68] David L. Banks. Exact results and bounds for the mean squared error of percentile bootstraps. Computational Statistics, 8(4):347–362, 1993. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Modarres:1994:MSS

[69] R. Modarres and J. P. Nolan. A method for simulating stable random vectors. Computational Statistics, 9(1):11–19, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Mallows:1994:NDR

[70] C. L. Mallows. A new distribution related to the logistic. Computational Statistics, 9(1):21–24, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

vanderGenugten:1994:DLS

[71] B. B. van der Genugten. Density of the least squares estimator in the multivariate linear model with arbitrarily normal variables. Computational Statistics, 9(1):25–39, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Wand:1994:MPB

[72] M. P. Wand and M. C. Jones. Multivariate plug-in bandwidth selection. Computational Statistics, 9(2):97–116, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Berlinet:1994:IUA

[73] Alain Berlinet and Christian Francq. Identification of a univariate ARMA model. Computational Statistics, 9(2):

117–133, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Manteiga:1994:BR

[74] Wenceslao González Manteiga, José Manuel Prada Sánchez, and Juan Romo. The bootstrap — a review. Computational Statistics, 9(3): 165–205, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Djojosugito:1994:UCS

[75] Rianto A. Djojosugito. On the use of cubic spline smoothing for testing parametric linear regression models. Computational Statistics, 9(3):213– 231, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Gruet:1994:CNS

[76] Marie-Anne Gruet and Emmanuel Jolivet. Calibration with a nonlinear standard curve: how to do it? Computational Statistics, 9(4):249–276, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Lee:1994:IAG

[77] Andy H. Lee and Yer Van Hui. On influence assessment in generalized linear models. Computational Statistics, 9(4):277–285, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Chan:1994:PGF

[78] W. S. Chan. On portmanteau goodness-of-fit tests in robust time series modelling. Computational Statistics, 9(4):301–310, 1994. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Cheng:1995:RMO

[79] Bing Cheng and Jim Kay. A regularization method for one-dimensional edge detection and edge-preserving smoothing. Computational Statistics, 10(1): 53–69, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Rousseeuw:1995:SFN

[80] Peter J. Rousseeuw, Christophe Croux, and Ola Hössjer. Sensitivity functions and numerical analysis of the repeated median slope. Computational Statistics, 10(1):71–90, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Girard:1995:FMC

[81] D. A. Girard. The fast Monte-Carlo cross-validation and C_{L} procedures: comments, new results and application to image recovery problems. Computational Statistics, 10(3):205–258, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). With comments and a reply by the author.

Stander:1995:MEL

[82] J. Stander and B. W. Silverman. Minimax estimation of linear functionals, particularly in nonparametric regression and positron emission tomography. Computational Statistics, 10(3):259–283, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Boularan:1995:NMU

[83] J. Boularan, L. Ferré, and P. Vieu. A nonparametric model for unbalanced longitudinal data with application to geophysical data. Computational Statistics, 10(3):285–298, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Hurley:1995:LSQ

[84] C. Hurley and R. Modarres. Lowstorage quantile estimation. Computational Statistics, 10(4):311–325, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Besse:1995:FEC

[85] P. C. Besse and F. Ferraty. A fixed effect curvilinear model. Computational Statistics, 10(4):339–351, 1995. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Minnotte:1996:BOF

[86] Michael C. Minnotte. The biasoptimized frequency polygon. Computational Statistics, 11(1):35–48, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Bolfarine:1996:SCC

[87] Heleno Bolfarine and Manuel Galea Rojas. On structural comparative calibration under a t-model. Computational Statistics, 11(1):63–85, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). See erratum [117].

Kotter:1996:ARS

[88] Thomas Kötter. An asymptotic result for sliced inverse regression. Computational Statistics, 11(2):113–136, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Visek:1996:HBP

[89] Jan Ámos V´ ıšek. On high breakdown point estimation. Computational Statistics, 11(2):137–146, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Vynckier:1996:GFI

[90] C. Vynckier, J. Beirlant, and L. Tierney. Goodness-of-fit inferences through posterior distribution explorations. Computational Statistics, 11(2):147– 163, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Schmid:1996:TFO

[91] Friedrich Schmid and Mark Trede. Testing for first order stochastic dominance in either direction. Computational Statistics, 11(2):165–173, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Flak:1996:CSP

[92] Th. Flak, W. Schmid, and R. Sigmund. A comparison of several procedures for identifying outliers in contaminated ARMA processes. Computational Statistics, 11(2):175–195, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Gupta:1996:DCM

[93] A. K. Gupta and Jie Chen. Detecting changes of mean in multidimensional normal sequences with applications to literature and geology. Computational Statistics, 11(3):211–221, 1996.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

McKean:1996:DDD

[94] Joseph W. McKean, Joshua D. Naranjo, and Simon J. Sheather. Diagnostics to detect differences in robust fits of linear models. Computational Statistics, 11(3):223–243, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Polasek:1996:VDC

[95] Wolfgang Polasek. Variance diagnostics for classical and Bayesian linear regression. Computational Statistics, 11(3):245–268, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Smith:1996:FSP

[96] Michael Smith, Simon Sheather, and Robert Kohn. Finite sample performance of robust Bayesian regression. Computational Statistics, 11(3):269– 301, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Kanefuji:1996:EIG

[97] Koji Kanefuji and Kosei Iwase. Exponential inverse Gaussian distribution. Computational Statistics, 11(3):315– 326, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Coronel-Brizio:1996:ETG

[98] Hector F. Coronel-Brizio and Federico O’Reilly. EDF tests for the Gumbel distribution with estimated parameters from censored samples. Computational Statistics, 11(3):327–335, 1996.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Jones:1996:PDB

[99] M. C. Jones, J. S. Marron, and S. J. Sheather. Progress in data-based bandwidth selection for kernel density estimation. Computational Statistics, 11(3):337–381, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Pfeifer:1996:MAP

[100] D. Pfeifer, H.-P. Bäumer, and U. Schleier. The “minimal area” problem in ecology: a spatial Poisson process approach. Computational Statistics, 11(4):415–428, 1996. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). See errata [112].

Rasch:1997:ROD

[101] Dieter A. M. K. Rasch, Eligius M. T. Hendrix, and Eric P. J. Boer. Replicationfree optimal designs in regression analysis. Computational Statistics, 12(1):19–52, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Katkovnik:1997:KHO

[102] Vladimir Katkovnik. Kernel high-order approximation of time-varying parameters of harmonics. Computational Statistics, 12(1):67–91, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Delecroix:1997:DLE

[103] M. Delecroix, D. Guégan, and G. Léorat. Determinating Lyapunov exponents in deterministic dynamical systems. Computational Statistics, 12(1):93–107, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Aragon:1997:SIR

[104] Yves Aragon and Jérôme Saracco. Sliced inversed regression (SIR): an appraisal of small sample alternatives to slicing. Computational Statistics, 12(1):109–130, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Schimek:1997:NSA

[105] Michael G. Schimek. Non- and semiparametric alternatives to generalized linear models. Computational Statistics, 12(2):173–191, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Spath:1997:LSF

[106] Helmuth Späth. Least-squares fitting of ellipses and hyperbolas. Computational Statistics, 12(3):329–341, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Spath:1997:ODF

[107] Helmuth Späth. Orthogonal distance fitting by circles and ellipses with given area. Computational Statistics, 12(3):343–354, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Aragon:1997:GIM

[108] Yves Aragon. A Gauss implementation of multivariate sliced inverse regression. Computational Statistics, 12 (3):355–372, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Edlund:1997:ANL

[109] Ove Edlund, Håkan Ekblom, and Kaj Madsen. Algorithms for non-linear M-estimation. Computational Statistics, 12(3):373–383, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Arjas:1997:ANB

[110] Elja Arjas and Juha Heikkinen. An algorithm for nonparametric Bayesian estimation of a Poisson intensity. Computational Statistics, 12(3):385–402, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Arickx:1997:BCP

[111] Frans Arickx, Jan Broeckhove, Muriel Dejonghe, and Julien van den Broeck. BSFM: a computer program for Bayesian stochastic frontier models. Computational Statistics, 12(3):403– 421, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Pfeifer:1997:EAM

[112] D. Pfeifer, H.-P. Bäumer, and U. Schleier. Errata to the article: The ‘minimal area’ problem in ecology: a spatial Poisson process approach [Comput. Statist. 11 (1996), no. 4, 415–428; MR1421635 (97h:62046)]. Computational Statistics, 12(3):429–430, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). See [100].

Ma:1997:MFS

[113] Jun Ma and H. Malcolm Hudson. Modified Fisher scoring algorithms using Jacobi or Gauss-Seidel subiterations.

Computational Statistics, 12(4):467– 479, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Sjostedt:1997:CME

[114] S. Sjöstedt and A. Barrlund. A computational method for estimating continuum factor models. Computational Statistics, 12(4):481–495, 1997. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Ernst:1998:MGL

[115] Michael D. Ernst. A multivariate generalized Laplace distribution. Computational Statistics, 13(2):227–232, 1998. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Jimenez-Gamero:1998:EMI

[116] M. D. Jiménez-Gamero, J. Muñoz-Garc´ ıa, A. Muñoz-Reyes, and R. Pino-Mej´ ıas. On Efron’s Method II with identification of outlier bootstrap samples. Computational Statistics, 13(3): 301–318, 1998. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Anonymous:1998:ESC

[117] Anonymous. Erratum: On structural comparative calibration under a t-model [Comput. Statist. 11 (1996), no. 1, 63–85; MR1392535 (97g:62117)] by H. Bolfarine and M. Galea Rojas. Computational Statistics, 13 (3):423–424, 1998. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). See [87].

Dong:1998:CTP

[118] Jianping Dong and Renfang Jiang. Contingency table probability estimation — a projection pursuit approach. Computational Statistics, 13(4):425– 445, 1998. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

Swayne:1999:ISI

[119] Deborah F. Swayne and Sigbert Klinke. Introduction to the special issue on interactive graphical data analysis: What is interaction? Computational Statistics, 14(1):1–6, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022700.

Unwin:1999:RIG

[120] Antony Unwin. Requirements for interactive graphics software for exploratory data analysis. Computational Statistics, 14(1):7–22, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022706.

Chao:1999:IDA

[121] Edward C. Chao. Interactive data analysis based on dialog systems in S-PLUS. Computational Statistics, 14(1):23–38, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022703.

Nelson:1999:XVC

[122] Laura Nelson, Dianne Cook, and Carolina Cruz-Neira. XGobi vs the C2:

Results of an experiment comparing data visualization in a 3-D immersive virtual reality environment with a 2-D workstation display. Computational Statistics, 14(1):39–51, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022704.

Inselberg:1999:DPL

[123] Alfred Inselberg. Don’t panic … just do it in parallel! Computational Statistics, 14(1):53–77, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022705.

Avidan:1999:PDM

[124] Tova Avidan and Shlomo Avidan. ParallAX — a data mining tool based on parallel coordinates. Computational Statistics, 14(1):79–89, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022707.

Theus:1999:ASD

[125] Martin Theus. Analysing storm data using highly interactive tools. Computational Statistics, 14(1):91–108, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022708.

Wilhelm:1999:VCC

[126] Adalbert F. X. Wilhelm, Edward J. Wegman, and Jürgen Symanzik. Visual clustering and classification: The Oronsay particle size data set revisited. Computational Statistics, 14 (1):109–146, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022701.

James:1999:IDA

[127] David A. James. Interactive data analysis in a manufacturing setting- a case study. Computational Statistics, 14 (1):147–159, March 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022702.

Lyhagen:1999:IOF

[128] Johan Lyhagen. Identification of the order of a fractionally differenced ARMA model. Computational Statistics, 14(2):161–169, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022709.

Gredenhoff:1999:LLS

[129] Mikael Gredenhoff and Sune Karlsson. Lag-length selection in VAR-models using equal and unequal lag-length procedures. Computational Statistics, 14(2):171–187, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022710.

Ehlgen:1999:SAF

[130] Jürgen Ehlgen. A simple algorithm to factorize the autocovariance function of a moving average process. Computational Statistics, 14(2):189–195, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/PL00022711.

Muller:1999:ACO

[131] Werner G. Müller and Andrej Pázman. An algorithm for the computation of optimum designs under a given covariance structure. Computational Statistics, 14(2):197–211, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050013.

Holliday:1999:ACA

[132] Tim Holliday, Giovanni Pistone, Eva Riccomagno, and Henry P. Wynn. The application of computational algebraic geometry to the analysis of designed experiments: a case study. Computational Statistics, 14(2):213–231, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800050014.

Achcar:1999:MMT

[133] Jorge Alberto Achcar and Gilberto de Araújo Pereira. Mixture models for type II censored survival data in the presence of covariates. Computational Statistics, 14(2):233–250, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800050015.

Beninel:1999:OAD

[134] Farid Beninel and François Husson. An optimized algorithm to determine the values of the exact cumulative distribution function of some discrete statistics. Computational Statistics, 14(2):251–261, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050016.

Lu:1999:EMM

[135] Wang-Shu Lu. The efficiency of the method of moments estimates for hyperparameters in the empirical Bayes binomial model. Computational Statistics, 14(2):263–276, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050017.

Lee:1999:RSB

[136] Sauchi Stephen Lee. Regularization in skewed binary classification. Computational Statistics, 14(2):277–292, July 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800050018.

Qian:1999:CAR

[137] Guoqi Qian. Computations and analysis in robust regression model selection using stochastic complexity. Computational Statistics, 14(3): 293–314, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03500911.

He:1999:CQC

[138] Xuming He and Pin Ng. COBS: qualitatively constrained smoothing via linear programming. Computational Statistics, 14(3):315–337, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800050019.

Christmann:1999:GST

[139] Andreas Christmann. On group sequential tests based on robust location and scale estimators in the two-sample problem. Computational Statistics, 14 (3):339–353, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050020.

Saavedra:1999:CST

[140] Angeles Saavedra and Ricardo Cao. A comperative study of two convolutiontype estimators of the marginal density of moving average processes. Computational Statistics, 14(3):355– 373, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050021.

Haario:1999:APD

[141] Heikki Haario, Eero Saksman, and Johanna Tamminen. Adaptive proposal distribution for random walk Metropolis algorithm. Computational Statistics, 14(3):375–395, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800050022.

Piccarreta:1999:SBM

[142] Raffaella Piccarreta. Sharp bounds for the maximum of the Goodman– Kruskal τ index in a class of I × J tables with given row and column totals. Computational Statistics, 14(3): 397–418, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022712.

Chauveau:1999:ASR

[143] Didier Chauveau and Jean Diebolt. An automated stopping rule for MCMC convergence assessment. Computational Statistics, 14(3):419–442, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s001800050024.

Aguilera:1999:FTS

[144] Ana M. Aguilera, Francisco A. Ocaña, and Mariano J. Valderrama. Forecasting time series by functional PCA. discussion of several weighted approaches. Computational Statistics, 14 (3):443–467, September 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050025.

Marchette:1999:LDR

[145] David J. Marchette and Wendy L. Poston. Local dimensionality reduction. Computational Statistics, 14(4): 469–489, December 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050026.

Spath:1999:EPE

[146] Helmuth Späth. Estimating the parameters of an ellipse when angular differences are known. Computational Statistics, 14(4):491–500, December 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800050027.

Linden:1999:EEF

[147] Mikael Linden. Estimating effort function with semiparametric model. Computational Statistics, 14(4):501– 513, December 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050028.

Lee:1999:BSL

[148] Thomas C. M. Lee and Victor Solo. Bandwidth selection for local linear regression: A simulation study. Computational Statistics, 14 (4):515–532, December 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050029. See erratum

[177].

Kaiser:1999:DES

[149] Regina Kaiser. Detection and estimation of structural changes and outliers in unobserved components. Computational Statistics, 14(4):533– 558, December 1999. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050030.

Lauro:2000:E

[150] Carlo Lauro and Vincenzo Esposito. Editorial. Computational Statistics, 15(1):1–2, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03500910;   http://link.
springer.com/content/pdf/10.1007/
BF03500910.pdf.

Bartkowiak:2000:OFC

[151] Anna Bartkowiak and Adam Szustalewicz. Outliers- finding and classifying which genuine and which spurious. Computational Statistics, 15(1):3–12, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800050031.

Bravo:2000:SOD

[152] M. Carmen Bravo and José M. García-Santesmases. Symbolic object description of strata by segmentation trees. Computational Statistics, 15 (1):13–24, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050032.

Depoutot:2000:FAL

[153] Raoul Depoutot and Christophe Planas. Finite approximations to linear filters and the monitoring of revisions. Computational Statistics, 15(1):25–30, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050033.

Gammermann:2000:SVM

[154] A. Gammermann. Support vector machine learning algorithm and transduction. Computational Statistics, 15(1):31–39, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050034.

Grossmann:2000:MUS

[155] Wilfried Grossmann. Metadata usage in the statistical production process. Computational Statistics, 15 (1):41–51, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050035.

Keller:2000:IES

[156] Wouter J. Keller and Jelke Bethlehem. The impact of EDI on statistical data processing. Computational Statistics, 15(1):53–63, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050036.

Laaksonen:2000:RBN

[157] Seppo Laaksonen. Regression-based nearest neighbour hot decking. Computational Statistics, 15(1):65–71, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800050037.

Lauro:2000:PCA

[158] Carlo N. Lauro and Francesco Palumbo. Principal component analysis of interval data: a symbolic data analysis approach. Computational Statistics, 15(1):73–87, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800050038.

Papageorgiou:2000:RAM

[159] H. Papageorgiou, Maria Vardaki, and Fragkiskos Pentaris. Recent advances on metadata. Computational Statistics, 15(1):89–97, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050039.

Payne:2000:DWD

[160] R. W. Payne, S. A. Harding, J. A. Dhaliwal, and S. S. Dhaliwal. A data warehouse for designed experiments. Computational Statistics, 15 (1):99–108, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050040.

Petrakos:2000:TFS

[161] George Petrakos. The topological foundation and some properties of the mixed estimator. Computational Statistics, 15(1):109–114, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800050041.

Ragozini:2000:CGA

[162] Giancarlo Ragozini. A computational geometry approach for linear and non linear discriminant analysis. Computational Statistics, 15 (1):115–125, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800050042.

Rasson:2000:SKD

[163] Jean-Paul Rasson and Sandrine Lissoir. Symbolic kernel discriminant analysis. Computational Statistics, 15 (1):127–132, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050043.

Wegman:2000:VNT

[164] Edward J. Wegman. Visions: New techniques and technologies in statistics. Computational Statistics, 15 (1):133–144, March 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800050044.

Girard:2000:NPB

[165] Stéphane Girard. A nonlinear PCA based on manifold approximation. Computational Statistics, 15 (2):145–167, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800000025.

Lee:2000:EMM

[166] Dominic Lee and Carey Priebe. Exact mean and mean squared error of the smoothed bootstrap mean integrated squared error estimator. Computational Statistics, 15(2):169–181, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800000026.

Insua:2000:WBR

[167] David Rios Insua and Brani Vidakovic. Wavelet-based random densities. Computational Statistics, 15 (2):183–203, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000027.

Mohammadpour:2000:ESR

[168] Adel Mohammadpour and A. Reza Soltani. Exchangeable stable random vectors and their simulations. Computational Statistics, 15(2):205–217, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800000028.

Penev:2000:WGA

[169] Spiridon Penev and Tenko Raykov. A Wiener germ approximation of the noncentral chi square distribution and of its quantiles. Computational Statistics, 15(2):219–228, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000029.

Jiang:2000:NGS

[170] Jiming Jiang. A nonlinear Gauss– Seidel algorithm for inference about GLMM. Computational Statistics, 15(2):229–241, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000030.

Satten:2000:AMD

[171] Glen A. Satten and Somnath Datta. The S–U algorithm for missing data problems. Computational Statistics, 15(2):243–277, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000031.

McCullough:2000:AMS

[172] B. D. McCullough. The accuracy of Mathematica 4 as a statistical package. Computational Statistics, 15 (2):279–299, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022713.

Rossini:2000:BRA

[173] A. J. Rossini. Book review: Applied Smoothing Techniques for Data Analysis: The Kernel Approach with S-Plus Illustrations by Adrian W. Bowman and Adelchi Azzalini. Computational Statistics, 15(2):301–302, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800000033.

Theus:2000:CPC

[174] Martin Theus. COMPSTAT 1998proceedings in computational statistics. Computational Statistics, 15 (2):303–305, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000034.

Baumer:2000:SAE

[175] Hans-Peter Bäumer. Statistical analysis of extreme values- from insurance, finance, hydrology and other fields. Computational Statistics, 15 (2):307–311, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000035.

Ickstadt:2000:BRC

[176] Katja Ickstadt. Book review: Collecting Spatial Data by W. G. Müller Physica–Verlag (1998), ISBN 3-7908-1134-3. Computational Statistics, 15 (2):313–314, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000036.

Lee:2000:EBS

[177] Thomas Lee and Victor Solo. Erratum to: Bandwidth Selection for Local Linear Regression: A Simulation Study 4/1999, pp 515– 532. Computational Statistics, 15(2):A3, July 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03500912;   http://link.
springer.com/content/pdf/10.1007/
BF03500912.pdf. See [148].

Hoijtink:2000:PIR

[178] Herbert Hoijtink. Posterior inference in the random intercept model based on samples obtained with Markov chain Monte Carlo methods. Computational Statistics, 15(3):315–336, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800000037.

Berkhof:2000:PPC

[179] Johannes Berkhof, Iven van Mechelen, and Herbert Hoijtink. Posterior predictive checks: Principles and discussion. Computational Statistics, 15(3): 337–354, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800000038.

Leenen:2000:BPE

[180] Iwin Leenen, Iven Van Mechelen, and Andrew Gelman. Bayesian probabilistic extensions of a deterministic classification model. Computational Statistics, 15(3):355–371, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800000039.

Gelman:2000:TER

[181] Andrew Gelman and Francis Tuerlinckx. Type s error rates for classical and Bayesian single and multiple comparison procedures. Computational Statistics, 15(3):373–390, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800000040.

Browne:2000:IPI

[182] William J. Browne and David Draper. Implementation and performance issues in the Bayesian and likelihood fitting of multilevel models.

Computational Statistics, 15(3):391– 420, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000041.

Janssen:2000:EPH

[183] Rianne Janssen and Paul De Boeck. Exploring the posterior of a hierarchical IRT model for item effects. Computational Statistics, 15(3): 421–442, September 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022714.

Wand:2000:CRS

[184] M. P. Wand. A comparison of regression spline smoothing procedures. Computational Statistics, 15 (4):443–462, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000047.

Bakin:2000:PMA

[185] Sergey Bakin, Markus Hegland, and Michael R. Osborne. Parallel MARS algorithm based on B-splines. Computational Statistics, 15(4):463–484, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022715.

Bergkvist:2000:WDE

[186] Erik Bergkvist and Per Johansson. Weighted derivative estimation of quantal response models: Simulations and applications to choice of truck freight carrier. Computational Statistics, 15(4):485–510, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022716.

Delecroix:2000:RSE

[187] Michel Delecroix and Camelia Protopopescu. Are regression series estimators efficient in practice? a computational comparison study. Computational Statistics, 15(4):511– 529, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000045.

Kozubowski:2000:MAG

[188] Tomasz J. Kozubowski and Krzysztof Podgórski. A multivariate and asymmetric generalization of Laplace distribution. Computational Statistics, 15 (4):531–540, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/PL00022717.

Glendinning:2000:EIA

[189] R. H. Glendinning. Estimating the inverse autocorrelation function from outlier contaminated data. Computational Statistics, 15(4):541– 565, December 2000. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800000048.

Cheng:2001:SAE

[190] Jian Cheng. Sampling algorithms for estimating the mean of bounded random variables. Computational Statistics, 16(1):1–23, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100049.

Dybowski:2001:PRV

[191] Richard Dybowski and Peter R. Weller. Prediction regions for the visualization of incomplete datasets. Computational Statistics, 16(1):25–41, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/PL00022718.

vanderLinde:2001:ESPa

[192] Angelika van der Linde. Estimating the smoothing parameter in generalized spline-based regression. Computational Statistics, 16(1):43–71, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100051.

vanderLinde:2001:ESPb

[193] Angelika van der Linde. Estimating the smoothing parameter in generalized spline-based regression. Computational Statistics, 16(1):73–95, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100052.

Onen:2001:GFT

[194] Bora H. Onen, Dennis C. Dietz, Vincent C. Yen, and Albert H. Moore. Goodness-of-fit tests for the Cauchy distribution. Computational Statistics, 16(1):97–107, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100053.

Heng-Hui:2001:SSA

[195] Lue Heng-Hui. A study of sensitivity analysis on the method of principal Hessian directions. Computational Statistics, 16(1):109–130, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100054.

Chamayou:2001:PRN

[196] J.-F. Chamayou. Pseudo random numbers for the Landau and Vavilov distributions. Computational Statistics, 16(1):131–152, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100055.

Linton:2001:SUT

[197] Oliver Linton. Symmetrizing and unitizing transformations for linear smoother weights. Computational Statistics, 16(1):153–164, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100056.

Kwong:2001:ACM

[198] Koon-Shing Kwong. An algorithm for construction of multiple hypothesis testing. Computational Statistics, 16(1):165–171, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100057.

Lacruz:2001:SDG

[199] Beatriz Lacruz, Pilar Lasala, and Alberto Lekuona. Selecting dynamic graphical models with hidden variables from data. Computational Statistics, 16(1):173–194, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100058.

Marron:2001:PSF

[200] J. S. Marron and S. S. Chung. Presentation of smoothers: the family approach. Computational Statistics, 16 (1):195–207, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100059.

Polat:2001:SR

[201] Cihat Polat. Software review. Computational Statistics, 16(1):209–219, March 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s001800100060.

Andersson:2001:BEC

[202] Michael K. Andersson and Sune Karlsson. Bootstrapping error component models. Computational Statistics, 16 (2):221–231, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100061.

Kim:2001:FTC

[203] Jinhyo Kim and Sangwon Seo. Fourier transformation can improve quadrature efficiency of Laplace distribution. Computational Statistics, 16 (2):233–242, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100062.

Louzada-Neto:2001:BAH

[204] Francisco Louzada-Neto. Bayesian analysis for hazard models with nonconstant shape parameter. Computational Statistics, 16(2):243–254, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100063.

Alba:2001:HTB

[205] M. V. Alba, D. Barrera, and M. D. Jiménez. A homogeneity test based on empirical characteristic functions. Computational Statistics, 16 (2):255–270, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100064.

Butucea:2001:NRC

[206] Cristina Butucea. Numerical results concerning a sharp adaptive density estimator. Computational Statistics, 16(2):271–298, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100065.

Mattos:2001:BAR

[207] Néli Maria Costa Mattos and Hélio Santos dos Migon. A Bayesian analysis of reliability in accelerated life tests using Gibbs sampler. Computational Statistics, 16(2):299–312, July 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100066.

Hofmann:2001:DMS

[208] Heike Hofmann, Antony Unwin, and Adalbert Wilhem. Data mining and statistics- introduction. Computational Statistics, 16(3):317–321, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100069.

Conversano:2001:PAC

[209] Claudio Conversano, Francesco Mola, and Roberta Siciliano. Partitioning algorithms and combined model integration for data mining. Computational Statistics, 16(3):323–339, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100070.

Muruzabal:2001:CSR

[210] Jorge Muruzábal. Combining statistical and reinforcement learning in rulebased classification. Computational Statistics, 16(3):341–359, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100071.

PhD:2001:LRE

[211] Xiuju Fu Ph.D and Lipo Wang. Linguistic rule extraction from a simplified RBF neural network. Computational Statistics, 16(3):361–372, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800100072.

Becker:2001:BHR

[212] Ursula Becker and Ludwig Fahrmeir. Bump hunting for risk: a new data mining tool and its applications. Computational Statistics, 16(3): 373–386, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800100073.

Bruzzese:2001:SPD

[213] Dario Bruzzese and Cristina Davino. Statistical pruning of discovered association rules. Computational Statistics, 16(3):387–398, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800100074.

Hofmann:2001:VCA

[214] Heike Hofmann and Adalbert Wilhelm. Visual comparison of association rules. Computational Statistics, 16 (3):399–415, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800100075.

Andrienko:2001:ECC

[215] Natalia Andrienko, Gennady Andrienko, and Peter Gatalsky. Exploring changes in census time series with interactive dynamic maps and graphics. Computational Statistics, 16(3): 417–433, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100076.

Michailidis:2001:DVT

[216] George Michailidis and Jan de Leeuw. Data visualization through graph drawing. Computational Statistics, 16 (3):435–450, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800100077.

Virili:2001:NNM

[217] Francesco Virili and Bernd Freisleben. Neural network model selection for financial time series prediction. Computational Statistics, 16(3):451–463, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s001800100078.

Sofyan:2001:AXD

[218] Hizir Sofyan and Axel Werwatz. Analyzing XploRe download profiles with intelligent Miner. Computational Statistics, 16(3):465–479, September 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800100079.

Jernigan:2001:ISP

[219] Robert Jernigan and Julie O’Connell. Influence on smoothness in penalized likelihood regression for binary data. Computational Statistics, 16(4): 481–504, December 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s180-001-8326-z.

Chen:2001:SSB

[220] Cathy W. S. Chen, Tsai-Hung Cherng, and Berlin Wu. On the selection of subset bilinear time series models: a genetic algorithm approach. Computational Statistics, 16 (4):505–517, December 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s180-001-8327-9.

vandeWiel:2001:SAF

[221] Mark van de Wiel. The split-up algorithm: a fast symbolic method for computing p-values of distribution-free statistics. Computational Statistics, 16 (4):519–538, December 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s180-001-8328-6.

Wedel:2001:CSE

[222] Michel Wedel. Computing the standards errors of mixture model parameters with EM when classes are well separated. Computational Statistics, 16 (4):539–558, December 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s180-001-8329-3.

Tsionas:2001:PAS

[223] Efthymios G. Tsionas. Posterior analysis of stochastic frontier models with truncated normal errors. Computational Statistics, 16(4):559– 575, December 2001. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s180-001-8330-0.

Sainz:2002:FSB

[224] Ana Fernandez Sainz, Juan Rodriguez-Poo, and Inmaculada Villanua Martin. Finite sample behavior of two step estimators in selection models. Computational Statistics, 17(1):1–16, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800200087.

Smyth:2002:ISA

[225] Gordon K. Smyth and Heather M. Podlich. An improved saddlepoint approximation based on the negative binomial distribution for the general birth process. Computational Statistics, 17(1):17–28, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200088.

Rohl:2002:DME

[226] Michael C. Röhl, Claus Weihs, and Winfried Theis. Direct minimization of error rates in multivariate classification. Computational Statistics, 17(1):29–46, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200089.

Brown:2002:LDA

[227] Barry W. Brown, Floyd M. Spears, and Lawrence B. Levy. The log F: A distribution for all seasons. Computational Statistics, 17(1):47–58, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200098.

Yi:2002:LDA

[228] Seongbaek Yi and ByoungSeon Choi. Levinson–Durbin algorithm as a Szegő polynomial recursion. Computational Statistics, 17(1):59–64, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200090.

Sawitzki:2002:KSA

[229] Günther Sawitzki. Keeping statistics alive in documents. Computational Statistics, 17(1):65–88, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200091.

Cardot:2002:LRP

[230] Hervé Cardot. Local roughness penalties for regression splines. Computational Statistics, 17(1):89–102, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200092.

Spencer:2002:CMS

[231] Neil H. Spencer and Antony Fielding. A comparison of modelling strategies for value–added analyses of educational data. Computational Statistics, 17(1):103–116, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200093.

Bonett:2002:EVM

[232] Douglas G. Bonett, J. Arthur Woodward, and Robert L. Randall. Estimating p-values for Mardia’s coefficients of multivariate skewness and kurtosis. Computational Statistics, 17 (1):117–122, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200094.

Kathman:2002:EPM

[233] Steven J. Kathman. An expansion of the probability mass function that may be used to improve Poisson approximations. Computational Statistics, 17 (1):123–140, March 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200095.

Trosset:2002:ECM

[234] Michael W. Trosset. Extensions of classical multidimensional scaling via variable reduction. Computational Statistics, 17(2):147–163, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200099.

Rahnenfuhrer:2002:MPT

[235] Jörg Rahnenführer. Multivariate permutation tests for the k-sample problem with clustered data. Computational Statistics, 17(2):165–184, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800200100.

Koolaard:2002:UDM

[236] J. P. Koolaard, S. Ganesalingam, and C. R. O. Lawoko. The use of a distance measure in regularised discriminant analysis. Computational Statistics, 17(2):185–202, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200101.

Yamamoto:2002:OBG

[237] Yoshikazu Yamamoto and Junji Nakano. An object-based graphical user interface for time series analysis. Computational Statistics, 17 (2):203–213, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200102.

Toutenburg:2002:LRM

[238] Helge Toutenburg and Thomas Nittner. Linear regression models with incomplete categorical covariates. Computational Statistics, 17 (2):215–232, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200103.

Grammig:2002:TBE

[239] Joachim Grammig, Reinhard Hujer, and Stefan Kokot. Tackling boundary effects in nonparametric estimation of intra–day liquidity measures. Computational Statistics, 17 (2):233–249, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200104.

Silva:2002:DVP

[240] António Pedro Duarte Silva. Discarding variables in a principal component analysis: Algorithms for all–subsets comparisons. Computational Statistics, 17(2):251–271, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200105.

Christmann:2002:CBV

[241] Andreas Christmann, Paul Fischer, and Thorsten Joachims. Comparison between various regression depth methods and the support vector machine to approximate the minimum number of misclassifications. Computational Statistics, 17(2):273–287, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800200106.

Sullivan:2002:ELM

[242] Joe H. Sullivan. Estimating the locations of multiple change points in the mean. Computational Statistics, 17(2):289–296, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200107.

Buning:2002:ADF

[243] Herbert Büning. An adaptive distribution-free test for the general two-sample problem. Computational Statistics, 17(2):297–313, July 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200108.

Kleinow:2002:CSB

[244] Torsten Kleinow and Heiko Lehmann. Client/ server based statistical computing. Computational Statistics, 17 (3):315–328, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200109.

Aydinli:2002:MLX

[245] Gökhan Aydinli, Wolfgang Härdle, Torsten Kleinow, and Hizir Sofyan. MD*ReX: Linking XploRe to standard spreadsheet applications. Computational Statistics, 17(3):329– 341, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200110.

Kang:2002:UFS

[246] Gunseog Kang and Jung Jin Lee. A user friendly software for the design and analysis of experiments. Computational Statistics, 17(3):343– 353, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200111.

Lee:2002:CAT

[247] Jung Jin Lee, Gunseog Kang, and Kyung Soo Han. Computer aided teaching for statistics in Internet age. Computational Statistics, 17(3): 355–365, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200112.

Inoue:2002:SDR

[248] Tatsuki Inoue, Yumi Asahi, Hiroshi Yadohisa, and Yoshiro Yamamoto. A statistical data representation system on the Web. Computational Statistics, 17(3):367–378, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800200113.

Yamamoto:2002:MUI

[249] Yoshikazu Yamamoto, Junji Nakano, Takeshi Fujiwara, and Ikunori Kobayashi. A mixed user interface for a statistical system. Computational Statistics, 17 (3):379–393, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200114.

Kobayashi:2002:POO

[250] Ikunori Kobayashi, Takeshi Fujiwara, Junji Nakano, and Yoshikazu Yamamoto. A procedural and object– oriented statistical scripting language. Computational Statistics, 17(3):395– 410, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200115.

Huh:2002:DJB

[251] Moon Yul Huh and Kwangryeol Song. DAVIS: A Java-based data visualization system. Computational Statistics, 17(3):411–423, September 2002.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200116.

Yoshioka:2002:KUO

[252] Koichi Yoshioka. KyPlot — a useroriented tool for statistical data analysis and visualization. Computational Statistics, 17(3):425–437, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200117.

Takeuchi:2002:DLL

[253] Akinobu Takeuchi, Yadohisa Hiroshi, Kazunori Yamaguchi, Michiko Watanabe, and Chooichiro Asano. Dynamic link library for statistical analysis and its Excel interface. Computational Statistics, 17(3):439–452, September 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800200118.

Hristache:2002:EES

[254] Marian Hristache. Are efficient estimators in single–index models really efficient? a computational discussion. Computational Statistics, 17 (4):453–464, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200119.

Lengvarszky:2002:ICL

[255] Zsolt Lengvárszky and R. Webster West. Influence contours in linear regression. Computational Statistics, 17 (4):465–477, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200120.

Lang:2002:FEL

[256] Stefan Lang, Eva-Maria Pronk, and Ludwig Fahrmeir. Function estimation with locally adaptive dynamic models. Computational Statistics, 17(4): 479–499, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200121. See correction

[339].

Flachaire:2002:BHC

[257] Emmanuel Flachaire. Bootstrapping heteroskedasticity consistent covariance matrix estimator. Computational Statistics, 17(4):501–506, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800200122.

Zheng:2002:CRB

[258] Qi Zheng. Computing relations between moments and cumulants. Computational Statistics, 17(4):507– 515, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800200123.

Ip:2002:SVM

[259] Edward H. Ip. On single versus multiple imputation for a class of stochastic algorithms estimating maximum likelihood. Computational Statistics, 17 (4):517–524, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200124.

Heinzl:2002:ARM

[260] Harald Heinzl and Martina Mittlböck. Adjusted R^{2} measures for the inverse Gaussian regression model. Computational Statistics, 17(4):525– 544, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200125.

Ferraty:2002:FNM

[261] Frédéric Ferraty and Philippe Vieu. The functional nonparametric model and application to spectrometric data. Computational Statistics, 17(4):545– 564, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200126.

Bonett:2002:SAP

[262] Douglas G. Bonett. A simple approximation to the percentiles of the t distribution. Computational Statistics, 17 (4):565–568, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800200127.

Nedelman:2002:BRV

[263] Jerry R. Nedelman. Book review: VBA for Modelers: Developing Decision Support Systems Using Microsoft Excel, by S. Christian Albright Duxbury Press, 2001. Computational Statistics, 17(4):569–571, December 2002. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03500930.

Thamerus:2003:FMD

[264] Markus Thamerus. Fitting a mixture distribution to a variable subject to heteroscedastie measurement errors. Computational Statistics, 18 (1):1–17, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800300129.

Moerbeek:2003:CEM

[265] Mirjam Moerbeek, Gerard J. P. Van Breukelen, and Martijn P. F. Berger. A comparison of estimation methods for multilevel logistic models. Computational Statistics, 18(1):19–37, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800300130.

Achcar:2003:RMC

[266] Jorge Alberto Achcar, Vanderly Janeiro, and Josmar Mazucheli. Regression models for correlated biliary data with random effects assuming a mixture of normal distributions. Computational Statistics, 18(1):39–55, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800300131.

Pan:2003:GHQ

[267] Jianxin Pan and Robin Thompson. Gauss-Hermite quadrature approximation for estimation in generalised linear mixed models. Computational Statistics, 18(1):57–78, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300132.

Singer:2003:SML

[268] Hermann Singer. Simulated maximum likelihood in nonlinear continuous– discrete state space models: Importance sampling by approximate smoothing. Computational Statistics, 18(1):79–106, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300133.

Kitchen:2003:ARW

[269] A. M. Kitchen, R. Drachenberg, and J. Symanzik. Assessing the reliability of web-based statistical software. Computational Statistics, 18 (1):107–122, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300134.

Terpstra:2003:SAT

[270] Jeffrey T. Terpstra, Joseph W. McKean, and Kirk Anderson. Studentized autoregressive time series residuals. Computational Statistics, 18 (1):123–141, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300135.

Garczarek:2003:SCP

[271] Ursula Garczarek and Glaus Weihs. Standardizing the comparison of partitions. Computational Statistics, 18 (1):143–162, March 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800300136.

McMullan:2003:NLN

[272] A. McMullan, A. W. Bowman, and E. M. Scott. Non-linear and nonparametric modelling of seasonal environmental data. Computational Statistics, 18(2):167–183, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800300139.

Einbeck:2003:MLF

[273] Jochen Einbeck. Multivariate local fitting with general basis functions. Computational Statistics, 18 (2):185–203, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s001800300140.

Kvaloy:2003:EIN

[274] Jan Terje Kvaløy and Bo Henry Lindqvist. Estimation and inference in nonparametric Cox-models: Time transformation methods. Computational Statistics, 18(2):205–221, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s001800300141.

Wand:2003:SMM

[275] M. P. Wand. Smoothing and mixed models. Computational Statistics, 18 (2):223–249, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300142.

Durban:2003:NSA

[276] Maria Durbán and Iain D. Currie. A note on P -spline additive models with correlated errors. Computational Statistics, 18(2):251–262, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s001800300143.

Lang:2003:BGS

[277] Stefan Lang, Samson B. Adebayo, Ludwig Fahrmeir, and Winfried J. Steiner. Bayesian geoadditive seemingly unrelated regression. Computational Statistics, 18(2):263–292, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300144.

Delicado:2003:PCO

[278] Pedro Delicado and Mario Huerta. Principal curves of oriented points: theoretical and computational improvements. Computational Statistics, 18(2):293–315, July 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s001800300145.

Zagst:2003:POU

[279] Rudi Zagst, Jan Kehrbaum, and Bernd Schmid. Portfolio optimization under credit risk. Computational Statistics, 18(3):317–338, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03354601.

Hafner:2003:SAO

[280] Christian M. Hafner. Simple approximations for option pricing under mean reversion and stochastic volatility. Computational Statistics, 18(3): 339–353, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354602.

Platen:2003:SGS

[281] Eckhard Platen and Gerhard Stahl. A structure for general and specific market risk. Computational Statistics, 18(3):355–373, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03354603.

Stadie:2003:DPW

[282] Andreas Stadie. Detecting periods in which a time series model fails to predict the observed volatility. Computational Statistics, 18(3): 375–386, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354604.

Storti:2003:LIB

[283] Giuseppe Storti and Cosimo Vitale. Likelihood inference in BL–GARCH models. Computational Statistics, 18 (3):387–400, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354605.

Fernandez-Pascual:2003:MEP

[284] R. Fernández-Pascual, M. D. Ruiz-Medina, and J. M. Angulo. Multiscale estimation of processes related to the fractional black–Scholes equation. Computational Statistics, 18(3): 401–415, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354606.

Meyer:2003:PBI

[285] Sascha Meyer and Willi Schwarz. A PDE based implementation of the Hull & White model for cashflow derivatives. Computational Statistics, 18(3): 417–434, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354607.

Komorad:2003:ITT

[286] Karel Komorád. Implied trinomial trees and their implementation with XploRe. Computational Statistics, 18 (3):435–448, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354608.

Schulz:2003:OPB

[287] Rainer Schulz, Hizir Sofyan, Axel Werwatz, and Rodrigo Witzel. Online prediction of Berlin single–family house prices. Computational Statistics, 18 (3):449–462, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354609.

Papalia:2003:GME

[288] Rosa Bernardini Papalia. Generalized maximum entropy estimation of dynamic programming models with sample selection bias. Computational Statistics, 18(3):463–475, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03354610.

Ellis:2003:FPN

[289] Suria Ellis, Faans Steyn, and Hennie Venter. Fitting a Pareto–normal-Pareto distribution to the residuals of financial data. Computational Statistics, 18(3):477–491, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03354611.

Kim:2003:FSD

[290] Jeong-Ryeol Kim. Finite-sample distributions of self-normalised sums. Computational Statistics, 18(3):493– 504, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354612.

Aluja-Banet:2003:SSD

[291] Tomàs Aluja-Banet and Eduard Nafria. Stability and scalability in decision trees. Computational Statistics, 18(3): 505–520, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03354613.

M:2003:SEL

[292] Rahmatullah Imon A. H. M. Simulation of errors in linear regression: An approach based on fixed percentage area. Computational Statistics, 18 (3):521–531, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03354614.

Valderrama:2003:DSS

[293] Mariano J. Valderrama, Mónica Ortega-Moreno, Pedro González, and Ana M. Aguilera. Derivation of a state– space model by functional data analysis. Computational Statistics, 18(3): 533–546, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03354615.

Sim:2003:CXB

[294] C. H. Sim. Combined X-bar and CRL charts for the Gamma process. Computational Statistics, 18(3): 547–563, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF03354616.

Harkanen:2003:BBI

[295] Tommi Härkänen. BITE: A Bayesian intensity estimator. Computational Statistics, 18(3):565–583, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03354617.

Gannoun:2003:TCV

[296] Ali Gannoun and Jérôme Saracco. Two cross validation criteria for SIR_{α} and PSIR_{α} methods in view of prediction. Computational Statistics, 18(3): 585–603, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354618.

Ahn:2003:PPM

[297] Ju Sun Ahn, Heike Hofmann, and Dianne Cook. A projection pursuit method on the multidimensional squared contingency table. Computational Statistics, 18(3):605– 626, September 2003. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03354619.

Anonymous:2004:IJG

[298] Anonymous. Interview with James E. Gentle. Computational Statistics, 19(1):1–4, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915272.

Unwin:2004:PDV

[299] Antony Unwin and Martin Theus. Papers on data visualisation-what can we see in them? Computational Statistics, 19(1):5–8, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02915273.

Andrienko:2004:CCE

[300] Natalia Andrienko and Gennady Andrienko. Cumulative curves for exploration of demographic data: a case study of Northwest England. Computational Statistics, 19(1):9–28, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02915274.

Atkinson:2004:FSD

[301] Anthony C. Atkinson and Marco Riani. The forward search and data visualisation. Computational Statistics, 19(1):29–54, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02915275.

Nagel:2004:EDV

[302] Henrik R. Nagel and Erik Granum. Explorative and dynamic visualization of data in virtual reality. Computational Statistics, 19(1):55–73, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02915276.

Nocke:2004:MVC

[303] Thomas Nocke, Heidrun Schumann, and Uwe Böhm. Methods for the visualization of clustered climate data. Computational Statistics, 19 (1):75–94, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915277.

Schonlau:2004:VNH

[304] Matthias Schonlau. Visualizing nonhierarchical and hierarchical cluster analyses with clustergrams. Computational Statistics, 19(1):95–111, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02915278.

Valero-Mora:2004:VPL

[305] Pedro Valero-Mora, María F. Rodrigo, and Forrest W. Young. Visualizing parameters from loglinear models. Computational Statistics, 19(1): 113–135, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02915279.

Nakano:2004:AVF

[306] Junji Nakano, Moon Yul Huh, Yoshikazu Yamamoto, Takeshi Fujiwara, and Ikunori Kobayashi. Adding visualization functions of DAVIS to Jasp: Mixing two Java-based statistical systems. Computational Statistics, 19 (1):137–146, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02915280.

Ward:2004:HVD

[307] Matthew Ward, Wei Peng, and Xiaoning Wang. Hierarchical visual data mining for large-scale data. Computational Statistics, 19(1):147–158, February 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02915281.

Charmpis:2004:HAG

[308] Dimos C. Charmpis and Panayiota L. Panteli. A heuristic approach for the generation of multivariate random samples with specified marginal distributions and correlation matrix. Computational Statistics, 19(2):??, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892061.

Nittner:2004:AMA

[309] Thomas Nittner. The additive model affected by missing completely at random in the covariate. Computational Statistics, 19(2):??, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892060.

Ghosh:2004:BNN

[310] Anil Kumar Ghosh and Smarajit Bose. Backfitting neural networks. Computational Statistics, 19(2):??, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02892056.

Lardjane:2004:CLE

[311] Salim Lardjane. Consistent Lyapunov exponent estimation for onedimensional dynamical systems. Computational Statistics, 19(2):159–168, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892054.

Ferre:2004:AEC

[312] L. Ferré and J. Whittaker. Application of the empirical characteristic function to compare and estimate densities by pooling information. Computational Statistics, 19(2):169–192, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892055.

Karavas:2004:ECD

[313] Vassilios N. Karavas and L. Joe Moffitt. Evolutionary computation of a deterministic switching regressions estimator. Computational Statistics, 19(2):211–225, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892057.

Eubank:2004:SSS

[314] R. L. Eubank. A simple smoothing spline, III. Computational Statistics, 19(2):227–241, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892058.

Franco:2004:BTS

[315] Glaura C. Franco and Valderio A. Reisen. Bootstrap techniques in semiparametric estimation methods for ARFIMA models: A comparison study. Computational Statistics, 19(2):243–259, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892059.

Marrelec:2004:ARS

[316] G. Marrelec and H. Benali. Automated rejection sampling from product of distributions. Computational Statistics, 19(2):301–315, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02892062.

Lund:2004:PSI

[317] Jens Lund and Elke Thönnes. Perfect simulation and inference for point processes given noisy observations. Computational Statistics, 19(2):317–336, May 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02892063.

Aires:2004:GFP

[318] Nibia Aires. A guide to the Fortran programs to calculate inclusion probabilities for conditional Poisson sampling and Pareto πps sampling designs. Computational Statistics, 19(3): 337–345, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372100.

Dobra:2004:DCA

[319] Adrian Dobra and Seth Sullivant. A divide-and-conquer algorithm for generating Markov bases of multi-way tables. Computational Statistics, 19(3): 347–366, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372101.

Park:2004:RWW

[320] Dongryeon Park. Robustness weight by weighted median distance. Computational Statistics, 19(3):367–383, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/BF03372102.

Berchtold:2004:OMM

[321] André Berchtold. Optimization of mixture models: Comparison of different strategies. Computational Statistics, 19 (3):385–406, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372103.

Hazelton:2004:DEA

[322] Martin Hazelton. Density estimation from aggregate data. Computational Statistics, 19(3):407–423, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03372104.

Toscas:2004:UFS

[323] Peter J. Toscas and Malcolm J. Faddy. Using Fisher scoring to fit extended Poisson process models. Computational Statistics, 19(3):425–443, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03372105.

Futschik:2004:IBS

[324] Andreas Futschik and Brenton R. Clarke. Improving bandwidth selection methods by adding qualitative constraints. Computational Statistics, 19(3):445–453, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03372106.

Frees:2004:ESE

[325] Edward W. Frees and Chunfang Jin. Empirical standard errors for longitudinal data mixed linear models. Computational Statistics, 19(3): 455–475, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372107.

Jan:2004:NEC

[326] Show-Li Jan. Nonparametric equivalents of contrasts for identifying the minimum effective dose. Computational Statistics, 19(3):477– 491, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372108.

Meyer:2004:SXF

[327] David Meyer, Friedrich Leisch, Torsten Hothorn, and Kurt Hornik. Stat-DataML: An XML format for statistical data. Computational Statistics, 19 (3):493–509, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372109.

Yi:2004:SFA

[328] Seongbaek Yi and ByoungSeon Choi. Spectral factorization of the autocorrelation sequence of an IIR filter using patterns of three functions. Computational Statistics, 19(3): 511–519, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF03372110.

Kochkarev:2004:PPE

[329] Yuriy A. Kochkarev. Polynomial parameter estimations of close to Gaussian random variables. Computational Statistics, 19(3):521–523, September 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF03372111.

Wu:2004:FBE

[330] Hsien-Chung Wu. Fuzzy Bayesian estimation on lifetime data. Computational Statistics, 19(4):??, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753915.

Dinwoodie:2004:EAP

[331] I. H. Dinwoodie and Brenda MacGibbon. Exact analysis of a paired sibling study. Computational Statistics, 19 (4):525–534, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02753910.

Krzensk:2004:SEA

[332] Udo Krzensk. Simple and efficient algorithms for computing exact cumulative discrete probabilities and its inverses. Computational Statistics, 19(4):535–550, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753911.

Rueda:2004:MDA

[333] M. Rueda and S. González. Missing data and auxiliary information in surveys. Computational Statistics, 19(4): 551–567, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02753912.

Maiboroda:2004:EHT

[334] Rostislav E. Maiboroda and Natalia M. Markovich. Estimation of heavy-tailed probability density function with application to Web data. Computational Statistics, 19(4):569–592, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753913.

Beh:2004:PCO

[335] Eric J. Beh. S-PLUS code for ordinal correspondence analysis. Computational Statistics, 19(4):593–612, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753914.

Abe:2004:ICA

[336] Makoto Abe, Yasemin Boztug, and Lutz Hildebrandt. Investigating the competitive assumption of multinomial logit models of brand choice by nonparametric modeling. Computational Statistics, 19(4):635–657, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753916.

Espejo:2004:ANR

[337] Mariano Ruiz Espejo and Miguel Delgado Pineda. Automatic nonuniform random variate generation. Computational Statistics, 19(4):659–660, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753917.

Hornik:2004:RV

[338] Kurt Hornik and Friedrich Leisch. R Version 2.0.0. Computational Statistics, 19(4):661–664, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753918.

Lang:2004:CFE

[339] Stefan Lang, Eva-Maria Fronk, and Ludwig Fahrmeir. Correction to: Function estimation with locally adaptive dynamic models. Computational Statistics, 19(4):665, December 2004. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02753919;
 http://link.springer.com/content/

                                        pdf/10.1007/BF02753919.pdf. See

[256].

Anonymous:2005:IAU

[340] Anonymous. Interview with Antony Unwin. Computational Statistics, 20(1):1–5, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02736119.

Haynes:2005:BED

[341] Michele Haynes and Kerrie Mengersen. Bayesian estimation of g-and- k distributions using MCMC. Computational Statistics, 20(1):7–30, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02736120.

Lim:2005:SAV

[342] Hoi-Jeong Lim. Saddlepoint approximations to P -values for comparison of density estimates. Computational Statistics, 20(1):31–50, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02736121.

Christensen:2005:FAC

[343] David Christensen. Fast algorithms for the calculation of Kendall’s τ. Computational Statistics, 20(1):51–62, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02736122.

Yuan:2005:MCS

[344] Ao Yuan and Yimin Yang. A Markov chain sampler for contingency table exact inference. Computational Statistics, 20(1):63–80, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02736123.

Turlach:2005:SCS

[345] Berwin A. Turlach. Shape constrained smoothing using smoothing splines. Computational Statistics, 20 (1):81–104, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02736124.

Munoz-Reyes:2005:ESB

[346] A. Muñoz-Reyes, J. L. Moreno-Rebollo, M. D. Jiménez-Gamero, and J. Munoz-Garcia. An efficient shrinkage bootstrap bias estimator for smooth functions of sample means. Computational Statistics, 20(1):105–118, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02736125.

DeMacq:2005:HRS

[347] Isabelle De Macq and Leopold Simar. Hyper-rectangular space partitioning trees: A practical approach. Computational Statistics, 20(1):119–135, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02736126.

Zele:2005:ECL

[348] Mina Zele and Djani Juricić. Estimation of the confidence limits for the quadratic forms in normal variables using a simple Gaussian distribution approximation. Computational Statistics, 20(1):137–150, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02736127.

Sutzle:2005:NME

[349] Eric A. Sützle and Tomas Hrycej. Numerical method for estimating multivariate conditional distributions. Computational Statistics, 20(1):151–176, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02736128.

Levine:2005:IRS

[350] Richard A. Levine, Zhaoxia Yu, William G. Hanley, and John J. Nitao. Implementing random scan Gibbs samplers. Computational Statistics, 20 (1):177–196, March 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02736129.

Hornik:2005:RV

[351] Kurt Hornik and Friedrich Leisch. R Version 2.1.0. Computational Statistics, 20(2):197–202, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/BF02789699.

Branco:2005:RCC

[352] J. A. Branco, C. Croux, P. Filzmoser, and M. R. Oliveira. Robust canonical correlations: A comparative study. Computational Statistics, 20(2):203–229, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02789700.

Kuonen:2005:SAS

[353] Diego Kuonen. Saddlepoint approximations to Studentized bootstrap distributions based on M-estimates. Computational Statistics, 20(2):231–244, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02789701.

Nag:2005:MOD

[354] Ashok K. Nag, Amit Mitra, and Sharmishtha Mitra. Multiple outlier detection in multivariate data using self-organizing maps title. Computational Statistics, 20(2):245–264, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02789702.

Haario:2005:CAH

[355] Heikki Haario, Eero Saksman, and Johanna Tamminen. Componentwise adaptation for high dimensional MCMC. Computational Statistics, 20(2):265–273, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02789703.

Lee:2005:MCP

[356] Juneyoung Lee, André I. Khuri, and KeeWhan Kim. Modeling the coverage probability of a confidence interval on the among-group variance component in the unbalanced random one-way model. Computational Statistics, 20(2):275–294, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02789704.

Lee:2005:HMD

[357] Youngjo Lee, Maengseok Noh, and Keunkwan Ryu. HGLM modelling of dropout process using a frailty model. Computational Statistics, 20 (2):295–309, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02789705.

Yamanishi:2005:SAF

[358] Yoshihiro Yamanishi and Yutaka Tanaka. Sensitivity analysis in functional principal component analysis. Computational Statistics, 20 (2):311–326, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02789706.

Ayuda:2005:BSS

[359] María-Isabel Ayuda and Antonio Aznar. Behaviour in small samples of some tests of non-nested hypotheses in non-stationary regressions and their bootstrap versions. Computational Statistics, 20(2):327–347, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02789707.

Klemela:2005:AML

[360] Jussi Klemelä. Algorithms for manipulation of level sets of nonparametric density estimates. Computational Statistics, 20(2):349–368, June 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02789708.

Janssen:2005:MCC

[361] Arnold Janssen and Thorsten Pauls. A Monte Carlo comparison of studentized bootstrap and permutation tests for heteroscedastic two-sample problems. Computational Statistics, 20(3): 369–383, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741303.

Martinez-Miranda:2005:QEU

[362] M. D. Martínez-Miranda, M. Rueda-Garcia, A. Arcos-Cebrián, Y. Román-Montoya, and S. Gonzalez-Aguilera. Quantile estimation under successive sampling. Computational Statistics, 20 (3):385–399, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741304.

Kabaila:2005:CEC

[363] Paul Kabaila. Computation of exact confidence limits from discrete data. Computational Statistics, 20(3): 401–414, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741305.

Beh:2005:PCS

[364] Eric J. Beh. S-PLUS code for simple and multiple correspondence analysis. Computational Statistics, 20(3): 415–438, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02741306.

Joe:2005:CFA

[365] Harry Joe and A. H. M. Mahbub ul Latif. Computations for the familial analysis of binary traits. Computational Statistics, 20(3):439– 448, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/BF02741307.

Daunis-i-Estadella:2005:DII

[366] Josep Daunis i Estadella, Tomàs Aluja-Banet, and Santiago Thió-Henestrosa. Distribution of the inter and intra inertia in conditional MCA. Computational Statistics, 20(3):449–463, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02741308.

Gonzalez:2005:SSB

[367] Luis González, Francisco Velasco, and Rafael M. Gasca. A study of the similarities between topics. Computational Statistics, 20(3):465–479, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/BF02741309.

Marron:2005:SSS

[368] J. S. Marron and Jin Ting Zhang. SiZer for smoothing splines. Computational Statistics, 20(3):481–502, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02741310.

Vasconcellos:2005:IES

[369] Klaus L. P. Vasconcellos, Alejandro C. Frery, and Luciano B. Silva. Improving estimation in speckled imagery. Computational Statistics, 20(3): 503–519, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741311.

Kim:2005:CID

[370] Choongrak Kim and Whasoo Bae. Case influence diagnostics in the Kaplan– Meier estimator and the log-rank test. Computational Statistics, 20(3): 521–534, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741312.

Warner:2005:BRD

[371] John Warner. Book review: Data Mining Using SAS Applications by George Fernandes Chapman & Hall/ CRC, with sep. CD-ROM, 2002. Computational Statistics, 20(3):535– 537, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741313.

Francisco-Fernandez:2005:BSL

[372] Mario Francisco-Fernández and Juan M. Vilar-Fernández. Bandwidth selection for the local polynomial estimator under dependence: A simulation study. Computational Statistics, 20 (4):539–558, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741314.

Ohrvik:2005:SMS

[373] John Öhrvik and Gabriella Schoier. SE-TAR model selection — a bootstrap approach. Computational Statistics, 20 (4):559–573, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741315.

Kozumi:2005:BNB

[374] Hideo Kozumi and Xingyuan Zhang. Bayesian and non-Bayesian analysis of gamma stochastic frontier models by Markov Chain Monte Carlo methods. Computational Statistics, 20(4): 575–593, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741316.

Toscas:2005:FEP

[375] Peter J. Toscas and Malcolm J. Faddy. Fitting the extended Poisson process model to grouped binary data. Computational Statistics, 20(4):595–609, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02741317.

Vrbik:2005:PMS

[376] Jan Vrbik. Population moments of sampling distributions. Computational Statistics, 20(4):611–621, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02741318.

Menezes:2005:CAV

[377] Raquel Menezes, Pilar Garcia-Soidán, and Manuel Febrero-Bande. A comparison of approaches for valid variogram achievement. Computational Statistics, 20(4):623–642, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02741319.

Finos:2005:NNA

[378] Livio Finos and Luigi Salmaso. A new nonparametric approach for multiplicity control: Optimal subset procedures. Computational Statistics, 20 (4):643–654, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/BF02741320.

Nedelman:2005:BRB

[379] Jerry R. Nedelman. Book review: Bayesian Data Analysis, Second Edition by A. Gelman, J. B. Carlin, H. S. Stern, and D. B. Rubin, Chapman & Hall/CRC, 2004. Computational Statistics, 20(4):655–657, December 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/BF02741321.

Temme:2006:CBI

[380] Dirk Temme. Constraint-based inference algorithms for structural models with latent confounders- empirical application and simulations. Computational Statistics, 21(1):??, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0257-8.

Anonymous:2006:IGK

[381] Anonymous. Interview with Genshiro Kitagawa. Computational Statistics, 21(1):1–7, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0247-x.

Krause:2006:GAS

[382] Rüdiger Krause and Gerhard Tutz. Genetic algorithms for the selection of smoothing parameters in additive models. Computational Statistics, 21(1):9–31, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0248-9.

Gather:2006:OSE

[383] Ursula Gather, Karen Schettlinger, and Roland Fried. Online signal extraction by robust linear regression. Computational Statistics, 21(1):33–51, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0249-8.

Chauvet:2006:FAB

[384] Guillaume Chauvet and Yves Tillé. A fast algorithm for balanced sampling. Computational Statistics, 21(1):53–62, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0250-2.

Nadarajah:2006:LCN

[385] Saralees Nadarajah. On the linear combination of normal and Laplace random variables. Computational Statistics, 21(1):63–71, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0251-1. See correc-

tion [508].

Bouzas:2006:SSP

[386] P. R. Bouzas, A. M. Aguilera, M. J. Valderrama, and N. Ruiz-Fuentes. On the structure of the stochastic process of mortgages in Spain. Computational Statistics, 21(1):73–89, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0252-0.

Luengo:2006:TCT

[387] I. Luengo, C. N. Hernández, and P. Saavedra. Test to compare two population logspectra. Computational Statistics, 21(1):91–101, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0253-z.

Chakraborty:2006:GMC

[388] Arnab Chakraborty. Generating multivariate correlated samples. Computational Statistics, 21(1):103–119, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0254-y.

Al-Awadhi:2006:QEO

[389] Shafiqah A. Al-Awadhi and Paul H. Garthwaite. Quantifying expert opinion for modelling fauna habitat distributions. Computational Statistics, 21(1):121–140, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0255-x.

Holgersson:2006:GMA

[390] H. E. T. Holgersson. A graphical method for assessing multivariate normality. Computational Statistics, 21 (1):141–149, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0256-9.

Palumbo:2006:E

[391] Francesco Palumbo. Editorial. Computational Statistics, 21(2):183–185, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0258-7;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        006-0258-7.pdf.

Billard:2006:DSI

[392] L. Billard and E. Diday. Descriptive statistics for interval-valued observations in the presence of rules. Computational Statistics, 21(2):187–210, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0259-6.

Chavent:2006:NCM

[393] Marie Chavent, Francisco de A. T. de Carvalho, Yves Lechevallier, and Rosanna Verde. New clustering methods for interval data. Computational Statistics, 21(2):211–229, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0260-0.

deCarvalho:2006:DCI

[394] Francisco de A. T. de Carvalho, Paula Brito, and Hans-Hermann Bock. Dynamic clustering for interval data based on L_{2} distance. Computational Statistics, 21(2):231–250, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0261-z.

DUrso:2006:RFM

[395] Pierpaolo D’Urso and Paolo Giordani. A robust fuzzy k-means clustering model for interval valued data. Computational Statistics, 21(2):251–269, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-006-
                                        0262-y.

Irpino:2006:CRI

[396] Antonio Irpino and Valentino Tontodonato. Clustering reduced interval data using Hausdorff distance. Computational Statistics, 21(2):271–288, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0263-x.

Silva:2006:LDA

[397] António Pedro Duarte Silva and Paula Brito. Linear discriminant analysis for interval data. Computational Statistics, 21(2):289–308, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0264-9.

Ichino:2006:IAS

[398] Manabu Ichino and Shinya Ishikawa. Interclass analysis in symbolic pattern classification problems. Computational Statistics, 21(2):309–323, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0265-8.

Cariou:2006:EMR

[399] Véronique Cariou. Extension of multivariate regression trees to interval data. application to electricity load profiling. Computational Statistics, 21(2):325–341, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0266-7.

Gioia:2006:PCA

[400] Federica Gioia and Carlo N. Lauro. Principal component analysis on interval data. Computational Statistics, 21(2):343–363, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0267-6.

Corsaro:2006:ILS

[401] Stefania Corsaro and Marina Marino. Interval linear systems: the state of the art. Computational Statistics, 21 (2):365–384, ???? 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0268-5.

Bar-Hen:2006:SSD

[402] Avner Bar-Hen and Nicolas Picard. Simulation study of dissimilarity between point process. Computational Statistics, 21(3–4):??, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0008-x.

Singer:2006:MEH

[403] Hermann Singer. Moment equations and Hermite expansion for nonlinear stochastic differential equations with application to stock price models. Computational Statistics, 21(3– 4):385–397, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0001-4.

Souza:2006:CCL

[404] Leonardo R. Souza, Jeremy Smith, and Reinaldo C. Souza. Convex combinations of long memory estimates from different sampling rates. Computational Statistics, 21(3–4):399–413, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0002-3.

Franke:2006:ILM

[405] Jürgen Franke and Jochen Löhr. On the identification of large multilinear systems. Computational Statistics, 21(3–4):415–429, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-006-0003-
2.

Rootzen:2006:IPD

[406] Holger Rootzén and Jonny Olsson. On the influence of the prior distribution in image reconstruction. Computational Statistics, 21(3–4):431–444, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0004-1.

Olbermann:2006:IFD

[407] Barbara P. Olbermann, Sílvia R. C. Lopes, and Valdério A. Reisen. Invariance of the first difference in ARFIMA models. Computational Statistics, 21(3–4):445–461, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0005-0.

Solorzano:2006:NCC

[408] Eleanne Solorzano. Nonparametric comparisons with k_{2} > 2 controls using normal scores and Savage statistics. Computational Statistics, 21(3– 4):463–472, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0006-z.

Kim:2006:CMR

[409] Hea-Jung Kim. A computational method for ranking L products of parameters. Computational Statistics, 21(3–4):473–485, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-006-0007-
y.

Mira:2006:SPE

[410] Antonietta Mira. Stationarity preserving and efficiency increasing probability mass transfers made possible. Computational Statistics, 21(3– 4):509–522, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-006-0009-9.

Baillo:2006:PVN

[411] Amparo Baíllo and Antonio Cuevas. Parametric versus nonparametric tolerance regions in detection problems. Computational Statistics, 21(3– 4):523–536, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0010-3.

Chernobai:2006:MCC

[412] Anna Chernobai, Krzysztof Burnecki, Svetlozar Rachev, Stefan Trück, and Rafał Weron. Modelling catastrophe claims with left-truncated severity distributions. Computational Statistics, 21(3–4):537–555, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0011-2.

Sanchez-Borrego:2006:LLK

[413] I. R. Sánchez-Borrego, M. D. Martínez-Miranda, and A. González-Carmona. Local linear kernel estimation of the discontinuous regression function. Computational Statistics, 21(3– 4):557–569, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0014-z.

Yang:2006:NMD

[414] Kung Han Yang and Yue-Cune Chang. A note on model diagnostics in longitudinal data analysis. Computational Statistics, 21(3–4):571–587, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0015-y.

Kim:2006:BRT

[415] Choongrak Kim, Sungsoo Kim, Mira Park, and Hakbae Lee. A bias reducing technique in kernel distribution function estimation. Computational Statistics, 21(3–4):589–601, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0016-x.

Nott:2006:SEM

[416] David Nott. Semiparametric estimation of mean and variance functions for non-Gaussian data. Computational Statistics, 21(3–4):603–620, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0017-9.

Rufo:2006:BAF

[417] M. J. Rufo, J. Martín, and C. J. Pérez. Bayesian analysis of finite mixture models of distributions from exponential families. Computational Statistics, 21(3–4):621–637, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0018-8.

Berger:2006:BRS

[418] V. W. Berger. Book review: Stephen Senn: Dicing with Death: Chance, Risk, and Health. Computational Statistics, 21(3–4):639–641, December 2006. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-006-
0012-1.

Tarpey:2007:PMA

[419] Thaddeus Tarpey. A parametric k-means algorithm. Computational Statistics, 22(1):??, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0022-7.

Tokushige:2007:CFM

[420] Shuichi Tokushige, Hiroshi Yadohisa, and Koichi Inada. Crisp and fuzzy kmeans clustering algorithms for multivariate functional data. Computational Statistics, 22(1):1–16, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-006-
                                        0013-0.

Colby:2007:SEM

[421] Gordana Colby and Paul Rilstone. Simplified estimation of multivariate duration models with unobserved heterogeneity. Computational Statistics, 22(1):17–29, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-006-0019-7.

Horova:2007:CBC

[422] Ivana Horová and Jirí Zelinka. Contribution to the bandwidth choice for kernel density estimates. Computational Statistics, 22(1):31–47, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0020-9.

Shmilovici:2007:UMR

[423] Armin Shmilovici and Irad Ben-Gal. Using a VOM model for reconstructing potential coding regions in EST sequences. Computational Statistics, 22(1):49–69, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0021-8.

Baier:2007:ECR

[424] Thomas Baier and Erich Neuwirth. Excel :: COM :: R. Computational Statistics, 22(1):91–108, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0023-6.

Kim:2007:DMO

[425] Sung-Soo Kim and W. J. Krzanowski. Detecting multiple outliers in linear regression using a cluster method combined with graphical visualization. Computational Statistics, 22 (1):109–119, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0026-3.

Sakaori:2007:NCI

[426] Fumitake Sakaori, Takayuki Yamada, Akihisa Kawamura, and Takakazu Sugiyama. A new confidence interval for all characteristic roots of a covariance matrix. Computational Statistics, 22(1):121–131, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0028-1.

Jin:2007:MEN

[427] Xin Jin, Wengang Zhou, and Rongfang Bie. Multinomial event naive Bayesian modeling for SAGE data classification. Computational Statistics, 22(1):133–143, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0029-0.

Tanabe:2007:PEM

[428] Akihiro Tanabe, Kenji Fukumizu, Shigeyuki Oba, Takashi Takenouchi, and Shin Ishii. Parameter estimation for von Mises–Fisher distributions. Computational Statistics, 22 (1):145–157, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0030-7.

Lee:2007:RPR

[429] Thomas C. M. Lee and Hee-Seok Oh. Robust penalized regression spline fitting with application to additive mixed modeling. Computational Statistics, 22(1):159–171, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0031-6.

Takemura:2007:CSR

[430] Akimichi Takemura and Hisayuki Hara. Conditions for swappability of records in a microdata set when some marginals are fixed. Computational Statistics, 22(1):173–185, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-007-
0024-5.

Karlsson:2007:BRS

[431] Andreas Karlsson. Book review: A SAS Companion for Nonparametric Statistics, by Scott J. Richter and James J. Higgins. Computational Statistics, 22(1):187–188, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0032-5.

Johnson:2007:IFA

[432] Kjell Johnson and William Rayens. Influence function analysis applied to partial least squares. Computational Statistics, 22(2):??, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0037-0.

Liu:2007:PDR

[433] Yushu Liu and William Rayens. PLS and dimension reduction for classification. Computational Statistics, 22(2):189–208, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0039-y.

Bougeard:2007:MLR

[434] Stéphanie Bougeard, Mohamed Hanafi, and El Mostafa Qannari. Multiblock latent root regression. application to epidemiological data. Computational Statistics, 22(2):209–222, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-007-
                                        0036-1.

Preda:2007:PCF

[435] Cristian Preda, Gilbert Saporta, and Caroline Lévéder. PLS classification of functional data. Computational Statistics, 22(2):223–235, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0041-4.

Boj:2007:IPD

[436] Eva Boj, Aurea Grané, Josep Fortiana, and M. Mercè Claramunt. Implementing PLS for distance-based regression: computational issues. Computational Statistics, 22(2):237–248, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0035-2.

Kramer:2007:OSP

[437] Nicole Krämer. An overview on the shrinkage properties of partial least squares regression. Computational Statistics, 22(2):249–273, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0038-z.

Hanafi:2007:PPM

[438] Mohamed Hanafi. PLS path modelling: computation of latent variables with the estimation mode B. Computational Statistics, 22(2):275–292, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-007-
0042-3.

Kondylis:2007:ESP

[439] Athanassios Kondylis and Joe Whittaker. An empirical study of PLAD regression using the bootstrap. Computational Statistics, 22(2):307–321, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0034-3.

Renteria:2007:MAS

[440] Raúl Rentería, Ruy Milidiú, and Rafael Souza. MKPLS approach: switching strategies for the non-linear multikernel PLSR. Computational Statistics, 22(2):323–330, July 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0040-5.

Valderrama:2007:OMF

[441] Mariano J. Valderrama. An overview to modelling functional data. Computational Statistics, 22(3):331–334, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s00180-
007-0043-2.

Cao:2007:PCP

[442] Jiguo Cao and James O. Ramsay. Parameter cascades and profiling in functional data analysis. Computational Statistics, 22(3):335–351, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-007-
                                        0044-1.

Benhenni:2007:LSR

[443] K. Benhenni, F. Ferraty, M. Rachdi, and P. Vieu. Local smoothing regression with functional data. Computational Statistics, 22(3):353– 369, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0045-0.

Cardot:2007:NEL

[444] Hervé Cardot, Lubos Prchal, and Pascal Sarda. No effect and lack-of-fit permutation tests for functional regression. Computational Statistics, 22(3): 371–390, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0046-z.

Delicado:2007:FSP

[445] Pedro Delicado. Functional k-sample problem when data are density functions. Computational Statistics, 22(3): 391–410, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0047-y.

Febrero:2007:FAN

[446] Manuel Febrero, Pedro Galeano, and Wenceslao González-Manteiga. A functional analysis of NOx levels: location and scale estimation and outlier detection. Computational Statistics, 22(3): 411–427, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0048-x.

Ortega-Moreno:2007:SSM

[447] Mónica Ortega-Moreno and Manuel Escabias. On a state-space modelling for functional data. Computational Statistics, 22(3):429–438, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0049-9.

Fernandez-Alcala:2007:FEI

[448] R. M. Fernández-Alcalá, J. Navarro-Moreno, and J. C. Ruiz-Molina. Functional estimation incorporating prior correlation information. Computational Statistics, 22(3):439–447, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0050-3.

Ocana:2007:CCF

[449] Francisco A. Ocaña, Ana M. Aguilera, and Manuel Escabias. Computational considerations in functional principal component analysis. Computational Statistics, 22(3):449–465, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0051-2.

Bouzas:2007:FAR

[450] P. R. Bouzas, N. Ruiz-Fuentes, and F. M. Ocaña. Functional approach to the random mean of a compound Cox process. Computational Statistics, 22 (3):467–479, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0052-1.

Cuevas:2007:REC

[451] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. Robust estimation and classification for functional data via projection-based depth notions. Computational Statistics, 22(3): 481–496, September 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0053-0.

Klinke:2007:SIW

[452] Sigbert Klinke. Special issue: workshop data and information visualisation 2006. Computational Statistics, 22(4):497, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0056-x.

Vidmar:2007:VC

[453] Gaj Vidmar and Nino Rode. Visualising concordance. Computational Statistics, 22(4):499–509, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0057-9.

Fujino:2007:SNF

[454] Tomokazu Fujino. SVG+Ajax+R: a new framework for WebGIS. Computational Statistics, 22(4):511–520, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0058-8.

Kunz:2007:VCM

[455] Werner Kunz. Visualization of competitive market structure by means of choice data. Computational Statistics, 22(4):521–531, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-007-0059-
7.

Pittelkow:2007:VHP

[456] Y. E. Pittelkow and S. R. Wilson. Visualisation of “high p, small n” data. Computational Statistics, 22(4): 533–541, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0060-1.

Benko:2007:EII

[457] M. Benko, M. Fengler, W. Härdle, and M. Kopa. On extracting information implied in options. Computational Statistics, 22(4):543–553, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0061-0.

Leisch:2007:CYR

[458] Friedrich Leisch. Compstat 2006 Young Researcher Award. Computational Statistics, 22(4):555, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-007-
                                        0096-2.

Cascos:2007:ECH

[459] Ignacio Cascos. The expected convex hull trimmed regions of a sample. Computational Statistics, 22(4): 557–569, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0095-3.

Hohle:2007:RPM

[460] Michael Höhle. surveillance: an R package for the monitoring of infectious diseases. Computational Statistics, 22 (4):571–582, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0074-8.

Sakamoto:2007:MSB

[461] Wataru Sakamoto. MARS: selecting basis functions and knots with an empirical Bayes method. Computational Statistics, 22(4):583–597, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0075-7.

Liquet:2007:PMS

[462] Benoît Liquet and Jérôme Saracco. Pooled marginal slicing approach via SIR_{α} with discrete covariables. Computational Statistics, 22(4):599–617, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0078-4.

Liquet:2007:SBP

[463] Benoit Liquet, Jérôme Saracco, and Daniel Commenges. Selection between proportional and stratified hazards models based on expected loglikelihood. Computational Statistics, 22 (4):619–634, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0079-3.

Aparicio:2007:SSC

[464] Teresa Aparicio and Inmaculada Villanúa. Some selection criteria for nested binary choice models: a comparative study. Computational Statistics, 22(4):635–660, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0080-x.

Bradstreet:2007:BRN

[465] Thomas E. Bradstreet. Book review: Naomi B. Robbins: Creating more effective graphs. Computational Statistics, 22(4):661–663, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0064-x.

Bradstreet:2007:BRH

[466] Thomas E. Bradstreet. Book review: Howard Wainer: Graphic discovery: a trout in the milk and other visual adventures. Computational Statistics, 22 (4):665–667, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0065-9.

Karlsson:2007:BRJ

[467] Andreas Karlsson. Book review: J. E. Gentle, W. Härdle, and Y. Mori (eds): Handbook of computational statistics: concepts and methods. Computational Statistics, 22(4):669–670, December 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0086-4.

Bulla:2008:CIP

[468] Jan Bulla and Andreas Berzel. Computational issues in parameter estimation for stationary hidden Markov models. Computational Statistics, 23 (1):1–18, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0063-y.

Trenkler:2008:DVS

[469] Carsten Trenkler. Determining pvalues for systems cointegration tests with a prior adjustment for deterministic terms. Computational Statistics, 23(1):19–39, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0066-8.

Dudek:2008:SCB

[470] Anna Dudek, Maciej Goćwin, and Jacek Leśkow. Simultaneous confidence bands for the integrated hazard function. Computational Statistics, 23(1):41–62, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0067-7.

Kolacek:2008:PMN

[471] Jan Kolácek. Plug-in method for nonparametric regression. Computational Statistics, 23(1):63–78, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0068-6.

Saavedra:2008:EPS

[472] P. Saavedra, C. N. Hernández, I. Luengo, J. Artiles, and A. Santana. Estimation of population spectrum for linear processes with random coefficients. Computational Statistics, 23 (1):79–98, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0069-5.

Lavergne:2008:EMS

[473] Christian Lavergne, Marie-José Martinez, and Catherine Trottier. Empirical model selection in generalized linear mixed effects models. Computational Statistics, 23(1):99–109, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0071-y.

Corander:2008:BSM

[474] Jukka Corander, Jukka Sirén, and Elja Arjas. Bayesian spatial modeling of genetic population structure. Computational Statistics, 23(1):111–129, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0072-x.

Escribano:2008:TCU

[475] Alvaro Escribano, M. Teresa Santos, and Ana E. Sipols. Testing for cointegration using induced-order statistics. Computational Statistics, 23 (1):131–151, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0081-9.

Al-Ibrahim:2008:SPC

[476] A. H. Al-Ibrahim and Noriah M. Al-Kandari. Stability of principal components. Computational Statistics, 23 (1):153–171, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0082-8.

Bradstreet:2008:VBS

[477] Thomas E. Bradstreet. V. Berger: Selection bias and covariate imbalances in randomized clinical trials. Computational Statistics, 23(1):173–176, January 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0087-3.

Symanzik:2008:IAB

[478] Jürgen Symanzik. Interview with Andreas Buja. Computational Statistics, 23(2):177–184, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0113-0.

Bae:2008:IDV

[479] Whasoo Bae, Soonyoung Hwang, and Choongrak Kim. Influence diagnostics in the varying coefficient model with longitudinal data. Computational Statistics, 23(2):185–196, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0025-4.

Yano:2008:SOS

[480] Koiti Yano. A self-organizing state space model and simplex initial distribution search. Computational Statistics, 23(2):197–216, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-007-0027-
2.

Foster:2008:RMA

[481] Scott D. Foster, Arunas P. Verbyla, and Wayne S. Pitchford. A random model approach for the LASSO. Computational Statistics, 23(2):217–233, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0033-4.

Wei:2008:JDP

[482] Wen Hsiang Wei and Guan Jhih Chen. JavaStatSoft: design patterns and features. Computational Statistics, 23 (2):235–251, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0092-6.

Riaz:2008:MPV

[483] Muhammad Riaz. Monitoring process variability using auxiliary information. Computational Statistics, 23 (2):253–276, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0084-6.

Zhou:2008:MCE

[484] Xingcai Zhou and Xinsheng Liu. The Monte Carlo EM method for estimating multinomial probit latent variable models. Computational Statistics, 23 (2):277–289, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0091-7.

Park:2008:BSP

[485] Chun Gun Park, Hee-Seok Oh, and Hakbae Lee. Bayesian selection of primary resolution and wavelet basis functions for wavelet regression. Computational Statistics, 23(2):291–302, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0055-y.

Hahsler:2008:SAR

[486] Michael Hahsler, Christian Buchta, and Kurt Hornik. Selective association rule generation. Computational Statistics, 23(2):303–315, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0062-z.

Santos:2008:ASI

[487] Carlos Santos, David F. Hendry, and Soren Johansen. Automatic selection of indicators in a fully saturated regression. Computational Statistics, 23(2):317–335, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0054-z.

Hendry:2008:ASI

[488] David F. Hendry, Søren Johansen, and Carlos Santos. Automatic selection of indicators in a fully saturated regression. Computational Statistics, 23(2):337–339, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0112-1;   http:
//link.springer.com/content/pdf/
10.1007/s00180-008-0112-1.pdf.

Gonzalez-Manteiga:2008:BRF

[489] Wenceslao González-Manteiga. Book review: F. Ferraty and P. Vieu: Nonparametric functional data analysis: theory and practice. Computational Statistics, 23(2):341–342, April 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0111-2.

Bognar:2008:BMC

[490] Matthew A. Bognar. Bayesian modeling of continuously marked spatial point patterns. Computational Statistics, 23(3):??, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0073-9.

Heinzmann:2008:FPA

[491] Dominik Heinzmann. A filtered polynomial approach to density estimation. Computational Statistics, 23 (3):343–360, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0070-z.

Jacome:2008:CPM

[492] M. A. Jácome, I. Gijbels, and R. Cao. Comparison of presmoothing methods in kernel density estimation under censoring. Computational Statistics, 23(3):381–406, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0076-6.

Piccarreta:2008:CTO

[493] Raffaella Piccarreta. Classification trees for ordinal variables. Computational Statistics, 23(3):407–427, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0077-5.

Brys:2008:GFT

[494] Guy Brys, Mia Hubert, and Anja Struyf. Goodness-of-fit tests based on a robust measure of skewness. Computational Statistics, 23(3):429–442, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0083-7.

Yang:2008:TNA

[495] Jun Yang, Guohua Zou, and Yu Zhao. Two noniterative algorithms for computing posteriors. Computational Statistics, 23(3):443–453, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-007-0085-
5.

Alin:2008:OPM

[496] Aylin Alin and Serdar Kurt. Ordinary and penalized minimum powerdivergence estimators in two-way contingency tables. Computational Statistics, 23(3):455–468, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0088-2.

Wang:2008:AEA

[497] Mingfeng Wang, Masahiro Kuroda, Michio Sakakihara, and Zhi Geng. Acceleration of the EM algorithm using the vector epsilon algorithm. Computational Statistics, 23(3):469–486, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0089-1.

Law:2008:TMI

[498] Derek Law. Taligent MVP in interactive statistical graphics. Computational Statistics, 23(3):487–495, July 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0093-5.

Giles:2008:ID

[499] Kendall E. Giles, Michael W. Trosset, David J. Marchette, and Carey E. Priebe. Iterative denoising. Computational Statistics, 23(4):497–517, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0090-8.

Mazzoni:2008:CAC

[500] Thomas Mazzoni. Computational aspects of continuous-discrete extended Kalman-filtering. Computational Statistics, 23(4):519–539, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0094-4.

Celeux:2008:SHM

[501] Gilles Celeux and Jean-Baptiste Durand. Selecting hidden Markov model state number with cross-validated likelihood. Computational Statistics, 23 (4):541–564, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0097-1.

Sarkar:2008:ELU

[502] Deepayan Sarkar. Extending lattice: using generics and methods to implement new visualization methods within the Trellis framework. Computational Statistics, 23(4):565–572, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0098-0.

Roman-Montoya:2008:CIQ

[503] Y. Román-Montoya, M. Rueda, and A. Arcos. Confidence intervals for quantile estimation using jackknife techniques. Computational Statistics, 23(4):573–585, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0099-z.

Tsai:2008:CRB

[504] Miao-Yu Tsai and Chuhsing K. Hsiao. Computation of reference Bayesian inference for variance components in longitudinal studies. Computational Statistics, 23(4):587–604, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0100-x.

Arasan:2008:AIE

[505] Jayanthi Arasan and Mary Lunn. Alternative interval estimation for parameters of bivariate exponential model with time varying covariate. Computational Statistics, 23(4):605–622, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0101-9.

Ormerod:2008:PSS

[506] John T. Ormerod, M. P. Wand, and Inge Koch. Penalised spline support vector classifiers: computational issues. Computational Statistics, 23 (4):623–641, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-007-0102-8.

Dias:2008:BBA

[507] José G. Dias and Jeroen K. Vermunt. A bootstrap-based aggregate classifier for model-based clustering. Computational Statistics, 23(4):643–659, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-007-
                                        0103-7.

Diaz-Frances:2008:CLC

[508] Eloísa Díaz-Francés and José A. Montoya. Correction to On the linear combination of normal and Laplace random variables, by Nadarajah, S., Computational Statistics, 2006, 21, 63–71. Computational Statistics, 23 (4):661–666, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0137-5. See [385].

Nadarajah:2008:RLE

[509] Saralees Nadarajah. Reply to the letter to the editor. Computational Statistics, 23(4):667–668, October 2008. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0136-6.

Rueda:2009:PEF

[510] M. Rueda and I. R. Sánchez-Borrego. A predictive estimator of finite population mean using nonparametric regression. Computational Statistics, 24 (1):1–14, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0140-x.

Jerez:2009:LSI

[511] Miguel Jerez, José Casals, and Sonia Sotoca. Likelihood stabilization for ill-conditioned vector GARCH models. Computational Statistics, 24 (1):15–35, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-007-0104-6.

Castillo:2009:CRM

[512] Enrique Castillo, Carmen Castillo, Ali S. Hadi, and José M. Sarabia. Combined regression models. Computational Statistics, 24(1):37–66, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0105-5.

Wang:2009:CFS

[513] Yong Wang. The constrained Fisher scoring method for maximum likelihood computation of a nonparametric mixing distribution. Computational Statistics, 24(1):67–81, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-007-
0106-4.

Shim:2009:NCQ

[514] Jooyong Shim, Changha Hwang, and Kyung Ha Seok. Non-crossing quantile regression via doubly penalized kernel machine. Computational Statistics, 24(1):83–94, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0123-y.

Aminzadeh:2009:SNS

[515] M. S. Aminzadeh. Sequential and non-sequential acceptance sampling plans for autocorrelated processes using ARMA(p, q) models. Computational Statistics, 24(1):95–111, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0108-x.

Cao:2009:UAR

[516] Ricardo Cao, Alicia de las Heras, and Angeles Saavedra. The uncertainties about the relationships riskreturn-volatility in the Spanish stock market. Computational Statistics, 24 (1):113–126, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0141-9.

Fueda:2009:VSM

[517] Kaoru Fueda, Masaya Iizuka, and Yuichi Mori. Variable selection in multivariate methods using global score estimation. Computational Statistics, 24 (1):127–144, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0109-9.

Francisco-Fernandez:2009:TTH

[518] Mario Francisco-Fernández and Juan M. Vilar-Fernández. Two tests for heteroscedasticity in nonparametric regression. Computational Statistics, 24 (1):145–163, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0110-3.

Yokoyama:2009:OMT

[519] Satoru Yokoyama, Atsuho Nakayama, and Akinori Okada. One-mode three-way overlapping cluster analysis. Computational Statistics, 24(1): 165–179, February 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0114-z.

Murrell:2009:E

[520] Paul Murrell. Editorial. Computational Statistics, 24(2):181–182, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0157-9;   http://link.springer.
com/content/pdf/10.1007/s00180-
009-0157-9.pdf.

Wessa:2009:FSS

[521] Patrick Wessa. A framework for statistical software development, maintenance, and publishing within an open-access business model. Computational Statistics, 24(2):183–193, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0107-y.

Lawrence:2009:EGP

[522] Michael Lawrence, Hadley Wickham, Dianne Cook, Heike Hofmann, and Deborah F. Swayne. Extending the GGobi pipeline from R. Computational Statistics, 24(2):195–205, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0115-y.

Wickham:2009:PIG

[523] Hadley Wickham, Michael Lawrence, Dianne Cook, Andreas Buja, Heike Hofmann, and Deborah F. Swayne. The plumbing of interactive graphics. Computational Statistics, 24 (2):207–215, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0116-x.

Tierney:2009:CAP

[524] Luke Tierney. Code analysis and parallelizing vector operations in R. Computational Statistics, 24(2):217–223, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0117-9.

Hornik:2009:OSM

[525] Kurt Hornik, Christian Buchta, and Achim Zeileis. Open-source machine learning: R meets Weka. Computational Statistics, 24(2):225–232, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0119-7.

Fox:2009:VHT

[526] John Fox, Michael Friendly, and Georges Monette. Visualizing hypothesis tests in multivariate linear models: the heplots package for R. Computational Statistics, 24(2):233–246, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0120-1.

Eckel:2009:ILR

[527] Sandrah P. Eckel and Roger D. Peng. Interacting with local and remote data repositories using the stashR package. Computational Statistics, 24 (2):247–254, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0124-x.

Falcon:2009:CCC

[528] Seth Falcon. Caching code chunks in dynamic documents. Computational Statistics, 24(2):255–261, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0125-9.

Dalgaard:2009:CWB

[529] Peter Dalgaard. Can we have better semantics for r’s attach mechanism? Computational Statistics, 24(2):263–269, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0126-8.

Lang:2009:MPA

[530] Duncan Temple Lang. A modest proposal: an approach to making the internal R system extensible. Computational Statistics, 24(2):271–281, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0127-7.

Lang:2009:WMD

[531] Duncan Temple Lang. Working with meta-data from C/C++ code in R: the

                                        RGCCTranslationUnit package. Com-

putational Statistics, 24(2):283–293, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0128-6.

Harner:2009:JJR

[532] E. James Harner, Dajie Luo, and Jun Tan. JavaStat: a Java/R-based statistical computing environment. Computational Statistics, 24(2):295–302, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0131-y.

Urbanek:2009:HTS

[533] Simon Urbanek. How to talk to strangers: ways to leverage connectivity between R, Java and Objective C. Computational Statistics, 24 (2):303–311, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0132-x.

Ogasawara:2009:ACP

[534] Haruhiko Ogasawara. Asymptotic cumulants of the parameter estimators in item response theory. Computational Statistics, 24(2):313–331, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0118-8.

Mak:2009:NAC

[535] Tak K. Mak and Fassil Nebebe. Numerical approximation of conditional asymptotic variances using Monte Carlo simulation. Computational Statistics, 24(2):333–344, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0121-0.

Riaz:2009:PVC

[536] Muhammad Riaz and Ronald J. M. M. Does. A process variability control chart. Computational Statistics, 24 (2):345–368, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0122-z;   http:
//link.springer.com/content/pdf/
10.1007/s00180-008-0122-z.pdf.

Rosenblad:2009:BRJ

[537] Andreas Rosenblad. Book review: J. J. Faraway: Extending the linear model with R: generalized linear, mixed effects and nonparametric regression models. Computational Statistics, 24(2):369–370, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-009-0152-1.

Rosenblad:2009:BRB

[538] Andreas Rosenblad. Book review: B. F. J. Manly: Randomization, bootstrap and Monte Carlo methods in biology, third edition. Computational Statistics, 24(2):371–372, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-009-0150-3.

Shang:2009:BRD

[539] Han Lin Shang. Book review: D. Hand: Information generation (2007): how data rule our world. Computational Statistics, 24(2):373–374, May 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0151-2.

Lin:2009:CEL

[540] Tsung-I Lin, Hsiu J. Ho, and Pao S. Shen. Computationally efficient learning of multivariate t mixture models with missing information. Computational Statistics, 24(3):375–392, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-008-
0129-5.

Cho:2009:EBA

[541] HyungJun Cho, Jaewoo Kang, and Jae K. Lee. Empirical Bayes analysis of unreplicated microarray data. Computational Statistics, 24(3):393–408, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0133-9.

Fushing:2009:CCR

[542] Hsieh Fushing, Shu-Chun Chen, and How-Jing Lee. Computing circadian rhythmic patterns and beyond: introduction to a new non-Fourier analysis. Computational Statistics, 24 (3):409–430, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-008-0134-8.

Al-Awadhi:2009:TMA

[543] F. Al-Awadhi and A. R. Soltani. Thresholds of moving average of stationary processes for given on target significant levels. Computational Statistics, 24(3):431–440, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0135-7.

Molina:2009:BEM

[544] Isabel Molina, Nicola Salvati, and Monica Pratesi. Bootstrap for estimating the MSE of the spatial EBLUP. Computational Statistics, 24(3):441–458, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0138-4.

Morana:2009:ONF

[545] Claudio Morana. An omnibus noise filter. Computational Statistics, 24 (3):459–479, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0139-3.

Shen:2009:SSC

[546] Pao sheng Shen. A simulation study for a class of central composite designs with nested sub-experiment. Computational Statistics, 24(3):481–495, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-008-
                                        0142-8.

Oh:2009:CVW

[547] Hee-Seok Oh, Donghoh Kim, and Youngjo Lee. Cross-validated wavelet shrinkage. Computational Statistics, 24 (3):497–512, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-008-0143-7.

Fernandez:2009:HTB

[548] V. Alba Fernández, D. Barrera Rosillo, M. J. Ibáñez Pérez, and M. D. Jiménez Gamero. A homogeneity test for bivariate random variables. Computational Statistics, 24(3):513–531, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-008-
0144-6.

Winker:2009:CEB

[549] Peter Winker and Dietmar Maringer. The convergence of estimators based on heuristics: theory and application to a GARCH model. Computational Statistics, 24(3):533–550, August 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0145-5.

Habeck:2009:GTD

[550] Michael Habeck. Generation of threedimensional random rotations in fitting and matching problems. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0156-x;   http://link.springer.
com/content/pdf/10.1007/s00180-
009-0156-x.pdf.

Nofuentes:2009:CMC

[551] J. A. Roldán Nofuentes, J. D. Luna del Castillo, and P. Femia Marzo. Computational methods for comparing two binary diagnostic tests in the presence of partial verification of the disease. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0155-y.

Parra-Frutos:2009:BML

[552] Isabel Parra-Frutos. The behaviour of the modified Levene’s test when data are not normally distributed. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0154-z.

Mayrink:2009:CAB

[553] Vinicius Diniz Mayrink and Dani Gamerman. On computational aspects of Bayesian spatial models: influence of the neighboring structure in the efficiency of MCMC algorithms. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0153-0.

Fattorini:2009:AAE

[554] Lorenzo Fattorini. An adaptive algorithm for estimating inclusion probabilities and performing the Horvitz– Thompson criterion in complex designs. Computational Statistics, 24 (4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0149-9.

Nagarajan:2009:REA

[555] Niranjan Nagarajan and Uri Keich. Reliability and efficiency of algorithms for computing the significance of the Mann–Whitney test. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0148-x.

Farcomeni:2009:EAS

[556] Alessio Farcomeni. An exact approach to sparse principal component analysis. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0147-3.

Ke:2009:CFE

[557] Jau-Chuan Ke and Yunn-Kuang Chu. Comparison on five estimation approaches of intensity for a queueing system with short run. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-008-
0146-4.

Symanzik:2009:CWP

[558] Jürgen Symanzik, William Fischetti, and Ian Spence. Commemorating William Playfair’s 250th birthday. Computational Statistics, 24(4):??, December 2009. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0170-z;   http://link.springer.
com/content/pdf/10.1007/s00180-
009-0170-z.pdf.

Scheinerman:2010:MGU

[559] Edward R. Scheinerman and Kimberly Tucker. Modeling graphs using dot product representations. Computational Statistics, 25(1):1–16, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-009-
                                        0158-8.

Joo:2010:BMB

[560] Yongsung Joo, George Casella, and James Hobert. Bayesian model-based tight clustering for time course data. Computational Statistics, 25(1):17–38, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0159-7.

Hoef:2010:FCS

[561] Jay M. Ver Hoef, Josh M. London, and Peter L. Boveng. Fast computing of some generalized linear mixed pseudo-models with temporal autocorrelation. Computational Statistics, 25(1):39–55, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-009-0160-1.

Kuroda:2010:CVC

[562] Masahiro Kuroda, Hiroki Hashiguchi, and Shigakazu Nakagawa. Computing p-values in conditional independence models for a contingency table. Computational Statistics, 25(1):57–70, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0161-0.

Chang:2010:CSP

[563] Ching-Hui Chang, Nabendu Pal, Wooi Khai Lim, and Jyh-Jiuan Lin. Comparing several population means: a parametric bootstrap method, and its comparison with usual ANOVA F test as well as ANOM. Computational Statistics, 25(1):71–95, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0162-z.

Zhang:2010:HEE

[564] Jin Zhang. A highly efficient Lestimator for the location parameter of the Cauchy distribution. Computational Statistics, 25(1):97–105, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0163-y.

Henseler:2010:CPL

[565] Jörg Henseler. On the convergence of the partial least squares path modeling algorithm. Computational Statistics, 25(1):107–120, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0164-x;   http:
//link.springer.com/content/pdf/
10.1007/s00180-009-0164-x.pdf.

Bashir:2010:PMM

[566] Shaheena Bashir and Edward M. Carter. Penalized multinomial mixture logit model. Computational Statistics, 25(1):121–141, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0165-9;   http:
//link.springer.com/content/pdf/
10.1007/s00180-009-0165-9.pdf.

Antoch:2010:NVI

[567] Jaromír Antoch, Miroslav Brzezina, and Rafaelle Miele. A note on variability of interval data. Computational Statistics, 25(1):143–153, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0166-8.

Soltani:2010:TSR

[568] A. R. Soltani and A. Shirvani. Truncated stable random variables: characterization and simulation. Computational Statistics, 25(1):155–161, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0167-7.

Jach:2010:EWA

[569] Agnieszka Jach and Piotr Kokoszka. Empirical wavelet analysis of tail and memory properties of LARCH and FIGARCH models. Computational Statistics, 25(1):163–182, March 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0168-6.

Lin:2010:SLM

[570] Tzy-Chy Lin and Tsung-I Lin. Supervised learning of multivariate skew normal mixture models with missing information. Computational Statistics, 25(2):183–201, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0169-5.

Shen:2010:NEB

[571] Pao sheng Shen. Nonparametric estimation of the bivariate survival function for one modified form of doubly censored data. Computational Statistics, 25(2):203–213, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0171-y.

Kim:2010:BMC

[572] Jaehee Kim and Sooyoung Cheon. Bayesian multiple change-point estimation with annealing stochastic approximation Monte Carlo. Computational Statistics, 25(2):215–239, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0172-x.

Irigoien:2010:GGV

[573] Itziar Irigoien, Concepcion Arenas, Elena Fernández, and Francisco Mestres. GEVA: geometric variabilitybased approaches for identifying patterns in data. Computational Statistics, 25(2):241–255, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0173-9.

Fard:2010:PPO

[574] Mir Nabi Pirouzi Fard. Probability plots and order statistics of the standard extreme value distribution. Computational Statistics, 25(2):257–267, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0174-8.

Huskova:2010:NSC

[575] Marie Husková and Claudia Kirch. A note on Studentized confidence intervals for the change-point. Computational Statistics, 25(2):269–289, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0175-7.

Yamamoto:2010:PCS

[576] Yoshikazu Yamamoto, Junji Nakano, and Takeshi Fujiwara. Parallel computing in the statistical system Jasp. Computational Statistics, 25(2):291–298, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-009-
                                        0176-6.

Khoo:2010:UOE

[577] Michael Boon Chong Khoo, Zhang Wu, Chung-Ho Chen, and Kah Wai Yeong. Using one EWMA chart to jointly monitor the process mean and variance. Computational Statistics, 25(2):299–316, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-009-0177-5.

Genolini:2010:KML

[578] Christophe Genolini and Bruno Falissard. KmL: k-means for longitudinal data. Computational Statistics, 25(2):317–328, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0178-4.

Psaradakis:2010:IBO

[579] Zacharias Psaradakis. On inference based on the one-sample sign statistic for long-range dependent data. Computational Statistics, 25(2):329–340, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0179-3.

Perperoglou:2010:PRI

[580] Aris Perperoglou and Paul H. C. Eilers. Penalized regression with individual deviance effects. Computational Statistics, 25(2):341–361, June 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0180-x;   http://link.springer.
com/content/pdf/10.1007/s00180-
009-0180-x.pdf.

Martinez-Camblor:2010:CIR

[581] Pablo Martínez-Camblor. Comparing k-independent and right censored samples based on the likelihood ratio. Computational Statistics, 25(3): 363–374, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-009-0181-9.

Salinas:2010:BES

[582] Victor H. Salinas, José S. Romeo, and Alexis Peña. On Bayesian estimation of a survival curve: comparative study and examples. Computational Statistics, 25(3):375–389, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-009-
0182-8.

Minnotte:2010:MTH

[583] Michael C. Minnotte. Mode testing via higher-order density estimation. Computational Statistics, 25(3): 391–407, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0183-7.

Takane:2010:AMT

[584] Yoshio Takane, Kwanghee Jung, and Heungsun Hwang. An acceleration method for ten Berge et 40al.’s algorithm for orthogonal INDSCAL. (French) []. Computational Statistics, 25(3):409–428, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-010-0184-
6.

Lee:2010:CIR

[585] Seung-Hwan Lee. Confidence intervals for the regression parameter based on weighted log-rank estimating functions. Computational Statistics, 25(3): 429–440, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0185-5.

Armagan:2010:MSA

[586] Artin Armagan and Russell L. Zaretzki. Model selection via adaptive shrinkage with t priors. Computational Statistics, 25(3):441–461, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0186-4.

Yoshida:2010:CEM

[587] Takuma Yoshida, Masaru Kanba, and Kanta Naito. A computationally efficient model selection in the generalized linear mixed model. Computational Statistics, 25(3):463–484, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0187-3.

Stehlik:2010:FEF

[588] Milan Stehlík, Rastislav Potocký, Helmut Waldl, and Zdenek Fabián. On the favorable estimation for fitting heavy tailed data. Computational Statistics, 25(3):485–503, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0189-1.

Chan:2010:BGP

[589] Jennifer S. K. Chan and Doris Y. P. Leung. Binary geometric process model for the modeling of longitudinal binary data with trend. Computational Statistics, 25(3):505–536, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0190-8;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        010-0190-8.pdf.

Rufo:2010:NPP

[590] M. J. Rufo, J. Martín, and C. J. Pérez. A note on the prior parameter choice in finite mixture models of distributions from exponential families. Computational Statistics, 25(3): 537–550, September 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0188-2.

Murrell:2010:DEA

[591] Paul Murrell. The 2006 Data Expo of the American Statistical Association. Computational Statistics, 25(4): 551–554, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0207-3;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-010-0207-3.pdf.

Cho:2010:VAC

[592] Sang-Hoon Cho and Hyonho Chun. Visualizing abnormal climate changes in central America from 1995 to 2000. Computational Statistics, 25(4): 555–567, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0204-6.

Hobbs:2010:GMM

[593] J. Hobbs, H. Wickham, H. Hofmann, and D. Cook. Glaciers melt as mountains warm: a graphical case study. Computational Statistics, 25 (4):569–586, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0202-8.

Eden:2010:TSP

[594] Svetlana K. Eden, Angel Q. An, Jeffrey Horner, Cathy A. Jenkins, and Theresa A. Scott. A two-step process for graphically summarizing spatial temporal multivariate data in two dimensions. Computational Statistics, 25 (4):587–601, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0210-8.

Choi:2010:CVS

[595] Byeong Yeob Choi, Ho Kim, Un Yeong Go, Jong-Hyeon Jeong, and Jae Won Lee. Comparison of various statistical methods for detecting disease outbreaks. Computational Statistics, 25 (4):603–617, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0191-7.

Lue:2010:PHD

[596] Heng-Hui Lue. On principal Hessian directions for multivariate response regressions. Computational Statistics, 25 (4):619–632, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0192-6.

Murakami:2010:NCN

[597] Hidetoshi Murakami. A numerical comparison of the normal and some saddlepoint approximations to a distributionfree test for stochastic ordering in the competing risks model. Computational Statistics, 25(4):633–643, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0193-5.

Kugiumtzis:2010:NCC

[598] Dimitris Kugiumtzis and Efthymia Bora-Senta. Normal correlation coefficient of non-normal variables using piece-wise linear approximation. Computational Statistics, 25(4):645– 662, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0195-3.

Pitsoulis:2010:FAR

[599] L. Pitsoulis and G. Zioutas. A fast algorithm for robust regression with penalised trimmed squares. Computational Statistics, 25(4):663–689, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0196-2.

Hallam:2010:GAS

[600] Joshua W. Hallam, Olcay Akman, and Füsun Akman. Genetic algorithms with shrinking population size. Computational Statistics, 25(4):691–705, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0197-1.

Furno:2010:RTS

[601] Marilena Furno. A robust test of specification based on order statistics. Computational Statistics, 25(4): 707–723, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0199-z.

Herwartz:2010:CBM

[602] Helmut Herwartz and Israel Waichman. A comparison of bootstrap and Monte–Carlo testing approaches to value-at-risk diagnosis. Computational Statistics, 25(4):725–732, December 2010. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0194-4.

Chen:2011:BSS

[603] Cathy W. S. Chen, Feng Chi Liu, and Richard Gerlach. Bayesian subset selection for threshold autoregressive moving-average models. Computational Statistics, 26(1):1–30, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0198-0.

Weiss:2011:CPE

[604] Gregor Weiß. Copula parameter estimation by maximum-likelihood and minimum-distance estimators: a simulation study. Computational Statistics, 26(1):31–54, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0203-7.

Chang:2011:TES

[605] Ching-Hui Chang, Jyh-Jiuan Lin, and Nabendu Pal. Testing the equality of several gamma means: a parametric bootstrap method with applications. Computational Statistics, 26(1):55–76, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0209-1.

Stokes:2011:DMN

[606] Houston H. Stokes and Melvin Hinich. Detecting and modeling nonlinearity in the gas furnace data. Computational Statistics, 26(1):77–93, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0211-7.

Chang:2011:MTM

[607] Yi-Ping Chang, Ming-Chin Hung, and Yi-Chen Ko. A multinomial tree model for pricing credit default swap options. Computational Statistics, 26 (1):95–120, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0212-6.

Chan:2011:BAA

[608] Jennifer S. K. Chan and Wai Y. Wan. Bayesian approach to analysing longitudinal bivariate binary data with informative dropout. Computational Statistics, 26(1):121–144, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0213-5.

Shen:2011:NED

[609] Pao sheng Shen. Nonparametric estimation with doubly censored and truncated data. Computational Statistics, 26(1):145–157, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0214-4.

Lee:2011:PID

[610] Youngjo Lee, Myoungjin Jang, and Woojoo Lee. Prediction interval for disease mapping using hierarchical likelihood. Computational Statistics, 26 (1):159–179, March 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0215-3.

Adler:2011:CRM

[611] Werner Adler, Sergej Potapov, and Berthold Lausen. Classification of repeated measurements data using tree-based ensemble methods. Computational Statistics, 26(2):??, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0249-1.

Allignol:2011:ESF

[612] Arthur Allignol, Martin Schumacher, and Jan Beyersmann. Estimating summary functionals in multistate models with an application to hospital infection data. Computational Statistics, 26(2):181–197, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0200-x.

Grzegorczyk:2011:MNS

[613] Marco Grzegorczyk, Dirk Husmeier, and Jörg Rahnenführer. Modelling non-stationary dynamic gene regulatory processes with the BGM model. Computational Statistics, 26 (2):199–218, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0201-9.

Eugster:2011:HTP

[614] Manuel J. A. Eugster, Jochen Knaus, Christine Porzelius, Markus Schmidberger, and Esmeralda Vicedo. Handson tutorial for parallel computing with R. Computational Statistics, 26(2):219–239, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0206-4.

Schrodle:2011:PDM

[615] Birgit Schrödle and Leonhard Held. A primer on disease mapping and ecological regression using INLA. Computational Statistics, 26(2):241–258, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0208-2.

Schmidberger:2011:ESA

[616] Markus Schmidberger, Esmeralda Vicedo, and Ulrich Mansmann. Empirical study for the agreement between statistical methods in quality assessment and control of microarray data. Computational Statistics, 26(2):259–277, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0216-2.

Ligges:2011:FCI

[617] Uwe Ligges and Sebastian Krey. Feature clustering for instrument classification. Computational Statistics, 26(2):279–291, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0234-8.

Porzelius:2011:BDB

[618] Christine Porzelius, Martin Schumacher, and Harald Binder. The benefit of data-based model complexity selection via prediction error curves in time-to-event data. Computational Statistics, 26(2):293–302, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0236-6.

Maucher:2011:SHI

[619] M. Maucher, U. Schöning, and H. A. Kestler. Search heuristics and the influence of non-perfect randomness: examining genetic algorithms and simulated annealing. Computational Statistics, 26(2):303–319, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0237-5.

Kestler:2011:FTC

[620] Hans A. Kestler, Ludwig Lausser, Wolfgang Lindner, and Günther Palm. On the fusion of threshold classifiers for categorization and dimensionality reduction. Computational Statistics, 26(2):321–340, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0243-7.

Kraus:2011:MOS

[621] Johann M. Kraus, Christoph Müssel, Günther Palm, and Hans A. Kestler. Multi-objective selection for collecting cluster alternatives. Computational Statistics, 26(2):341–353, June 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0244-6.

Tjur:2011:SCA

[622] Tue Tjur. Statistics in the computer age: personal reflections. Computational Statistics, 26(3):371– 379, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0257-1.

Urbanek:2011:IEN

[623] Simon Urbanek. iPlots eXtreme: nextgeneration interactive graphics design and implementation of modern interactive graphics. Computational Statistics, 26(3):381–393, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0240-x.

Theussl:2011:PCR

[624] Stefan Theußl, Uwe Ligges, and Kurt Hornik. Prospects and challenges in R package development. Computational Statistics, 26(3):395–404, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0205-5.

Wickham:2011:MOR

[625] Hadley Wickham. Mutable objects in R. Computational Statistics, 26(3): 405–418, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0235-7.

Lenth:2011:RSA

[626] Russell Lenth and Søren Højsgaard. Reproducible statistical analysis with multiple languages. Computational Statistics, 26(3):419–426, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0245-5.

Runnalls:2011:ACI

[627] Andrew R. Runnalls. Aspects of CXXR internals. Computational Statistics, 26 (3):427–442, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0218-0.

Henningsen:2011:PMP

[628] Arne Henningsen and Ott Toomet. maxLik: A package for maximum likelihood estimation in R. Computational Statistics, 26(3):443–458, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0217-1.

Scott:2011:MGH

[629] David J. Scott, Diethelm Würtz, Christine Dong, and Thanh Tam Tran. Moments of the generalized hyperbolic distribution. Computational Statistics, 26 (3):459–476, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0219-z.

Tateishi:2011:NRM

[630] Shohei Tateishi and Sadanori Konishi. Nonlinear regression modeling and detecting change points via the relevance vector machine. Computational Statistics, 26(3):477–490, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0220-6.

Muller:2011:PST

[631] Samuel Müller and Houng Chhay. Partially smooth tail-index estimation for small samples. Computational Statistics, 26(3):491–505, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0221-5.

Tsou:2011:LIL

[632] Tsung-Shan Tsou. Likelihood inferences for the link function without knowing the true underlying distributions. Computational Statistics, 26(3): 507–519, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0222-4.

Shen:2011:SAT

[633] Pao sheng Shen. Semiparametric analysis of transformation models with left-truncated and right-censored data. Computational Statistics, 26(3): 521–537, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-010-0223-3.

Tayeb:2011:SGH

[634] Arafat Tayeb, Aurélie Labbe, Alexandre Bureau, and Chantal Mérette. Solving genetic heterogeneity in extended families by identifying subtypes of complex diseases. Computational Statistics, 26(3):539–560, September 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s00180-
010-0224-2.

Conversano:2011:IVM

[635] Claudio Conversano. Interactive visualization in multiclass learning: integrating the SASSC algorithm with KLIMT. Computational Statistics, 26(4):??, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0255-3.

Linsen:2011:LTT

[636] Lars Linsen and Sabine Behrendt. Linked treemap: a 3D treemapnodelink layout for visualizing hierarchical structures. Computational Statistics, 26(4):??, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-011-0272-
2.

Long:2011:VHD

[637] Tran Van Long and Lars Linsen. Visualizing high density clusters in multidimensional data using optimized star coordinates. Computational Statistics, 26(4):??, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0271-3.

Telea:2011:VSA

[638] Alexandru Telea and Lucian Voinea. Visual software analytics for the build optimization of large-scale software systems. Computational Statistics, 26(4):??, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0248-2;   http:
//link.springer.com/content/pdf/
10.1007/s00180-011-0248-2.pdf.

Wilhelm:2011:DVV

[639] Adalbert F. X. Wilhelm and Lars Linsen. Data viz VI. Computational Statistics, 26(4):??, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0278-9;   http:
//link.springer.com/content/pdf/
10.1007/s00180-011-0278-9.pdf.

Ward:2011:QAV

[640] Matthew Ward, Zaixian Xie, Di Yang, and Elke Rundensteiner. Qualityaware visual data analysis. Computational Statistics, 26(4):567–584, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-010-
0226-0.

Hurley:2011:GNI

[641] C. B. Hurley and R. W. Oldford. Graphs as navigational infrastructure for high dimensional data spaces. Computational Statistics, 26 (4):585–612, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0228-6.

Hurley:2011:ETA

[642] C. B. Hurley and R. W. Oldford. Eulerian tour algorithms for data visualization and the PairViz package. Computational Statistics, 26(4):613–633, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0229-5.

Eugster:2011:EAB

[643] Manuel J. A. Eugster and Friedrich Leisch. Exploratory analysis of benchmark experiments an interactive approach. Computational Statistics, 26 (4):699–710, December 2011. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-010-0227-z.

Zhao:2012:ICR

[644] Yichuan Zhao and Ali Jinnah. Inference for Cox’s regression models via adjusted empirical likelihood. Computational Statistics, 27(1):1–12, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-010-
                                        0225-1.

Brinch:2012:ESM

[645] Christian N. Brinch. Efficient simulated maximum likelihood estimation through explicitly parameter dependent importance sampling. Computational Statistics, 27(1):13–28, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0230-z.

Paindaveine:2012:CMO

[646] Davy Paindaveine and Miroslav Siman. Computing multiple-output regression quantile regions from projection quantiles. Computational Statistics, 27 (1):29–49, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0231-y.

Singh:2012:GLN

[647] Bhupendra Singh, K. K. Sharma, Shubhi Rathi, and Gajraj Singh. A generalized log-normal distribution and its goodness of fit to censored data. Computational Statistics, 27(1):51–67, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0233-9.

Figueiredo:2012:GFC

[648] Adelaide Maria Sousa Figueiredo. Goodness-of-fit for a concentrated von Mises–Fisher distribution. Computational Statistics, 27(1):69–82, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0238-4.

Sawant:2012:FOD

[649] Pallavi Sawant, Nedret Billor, and Hyejin Shin. Functional outlier detection with robust functional principal component analysis. Computational Statistics, 27(1):83–102, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0239-3.

Liquet:2012:GTS

[650] Benoît Liquet and Jérôme Saracco. A graphical tool for selecting the number of slices and the dimension of the model in SIR and SAVE approaches. Computational Statistics, 27(1):103–125, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0241-9.

Suknovic:2012:RCD

[651] Milija Suknovic, Boris Delibasic, Milos Jovanovic, Milan Vukicevic, Dragana Becejski-Vujaklija, and Zoran Obradovic. Reusable components in decision tree induction algorithms. Computational Statistics, 27 (1):127–148, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0242-8.

Shen:2012:VMC

[652] Yuan Shen, Dan Cornford, Manfred Opper, and Cedric Archambeau. Variational Markov chain Monte Carlo for Bayesian smoothing of non-linear diffusions. Computational Statistics, 27 (1):149–176, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0246-4.

Sra:2012:SNP

[653] Suvrit Sra. A short note on parameter approximation for von Mises–Fisher distributions: and a fast implementation of I_{s}(x). Computational Statistics, 27(1):177–190, March 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0232-x.

Sohn:2012:PTU

[654] Sooncheol Sohn, Byoung Cheol Jung, and Myoungshic Jhun. Permutation tests using least distance estimator in the multivariate regression model. Computational Statistics, 27 (2):191–201, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0247-3.

Hayashi:2012:BMA

[655] Kenichi Hayashi. A boosting method with asymmetric mislabeling probabilities which depend on covariates. Computational Statistics, 27(2):203–218, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0250-8.

Elster:2012:CNP

[656] Clemens Elster and Ignacio Lira. On the choice of a noninformative prior for Bayesian inference of discretized normal observations. Computational Statistics, 27(2):219–235, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0251-7.

Yin:2012:NBP

[657] Yuliang Yin. A new Bayesian procedure for testing point null hypotheses. Computational Statistics, 27 (2):237–249, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0252-6.

Klar:2012:STR

[658] Bernhard Klar and Simos G. Meintanis. Specification tests for the response distribution in generalized linear models. Computational Statistics, 27(2):251–267, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0253-5.

Worton:2012:ECS

[659] Bruce J. Worton. Efficient construction of a smooth nonparametric family of empirical distributions and calculation of bootstrap likelihood. Computational Statistics, 27(2):269–283, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0254-4.

Didericksen:2012:EPF

[660] Devin Didericksen, Piotr Kokoszka, and Xi Zhang. Empirical properties of forecasts with the functional autoregressive model. Computational Statistics, 27(2):285–298, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0256-2.

Nakagawa:2012:IOT

[661] Shigekazu Nakagawa, Hiroki Hashiguchi, and Naoto Niki. Improved omnibus test statistic for normality. Computational Statistics, 27(2):299–317, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0258-0.

Battaglia:2012:MRM

[662] Francesco Battaglia and Mattheos K. Protopapas. Multi-regime models for nonlinear nonstationary time series. Computational Statistics, 27 (2):319–341, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0259-z.

Li:2012:TNE

[663] Chin-Shang Li. Testing for no effect via splines. Computational Statistics, 27(2):343–357, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0260-6.

Kobayashi:2012:BAQ

[664] Genya Kobayashi and Hideo Kozumi. Bayesian analysis of quantile regression for censored dynamic panel data. Computational Statistics, 27(2):359–380, June 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0263-3.

Abuzaid:2012:BCV

[665] Ali H. Abuzaid, Ibrahim B. Mohamed, and Abdul G. Hussin. Boxplot for circular variables. Computational Statistics, 27(3):381–392, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-011-
                                        0261-5.

Fritz:2012:CAM

[666] Heinrich Fritz, Peter Filzmoser, and Christophe Croux. A comparison of algorithms for the multivariate L_{1}- median. Computational Statistics, 27 (3):393–410, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0262-4.

Garcia-Hiernaux:2012:ESO

[667] Alfredo García-Hiernaux, José Casals, and Miguel Jerez. Estimating the system order by subspace methods. Computational Statistics, 27(3):411– 425, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0264-2.

Vrac:2012:CAM

[668] M. Vrac, L. Billard, E. Diday, and A. Chédin. Copula analysis of mixture models. Computational Statistics, 27 (3):427–457, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0266-0.

Hayter:2012:ETS

[669] A. J. Hayter and Y. Lin. The evaluation of two-sided orthant probabilities for a quadrivariate normal distribution. Computational Statistics, 27 (3):459–471, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0267-z.

Otero:2012:RSM

[670] Jesús Otero and Jeremy Smith. Response surface models for the Leybourne unit root tests and lag order dependence. Computational Statistics, 27 (3):473–486, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0268-y.

Cahoy:2012:MET

[671] Dexter O. Cahoy. Moment estimators for the two-parameter M-Wright distribution. Computational Statistics, 27 (3):487–497, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0269-x.

Bonato:2012:MFT

[672] Matteo Bonato. Modeling fat tails in stock returns: a multivariate stable-GARCH approach. Computational Statistics, 27(3):499–521, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0270-4.

Hu:2012:RPD

[673] Yonggang Hu, Qiang Li, Yong Wang, and Yi Wu. Rayleigh projection depth. Computational Statistics, 27 (3):523–530, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0273-1.

Vidaurre:2012:LLL

[674] Diego Vidaurre, Concha Bielza, and Pedro Larrañaga. Lazy lasso for local regression. Computational Statistics, 27(3):531–550, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-011-0274-
                                        0.

vandeVelden:2012:GCC

[675] Michel van de Velden and Yoshio Takane. Generalized canonical correlation analysis with missing values. Computational Statistics, 27(3): 551–571, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0276-y;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-011-0276-y.pdf.

Zhang:2012:SSU

[676] Guoyi Zhang. Smoothing splines using compactly supported, positive definite, radial basis functions. Computational Statistics, 27(3):573–584, September 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0277-x.

Filzmoser:2012:DAC

[677] Peter Filzmoser, Karel Hron, and Matthias Templ. Discriminant analysis for compositional data and robust parameter estimation. Computational Statistics, 27(4):585–604, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0279-8.

Enki:2012:SPC

[678] Doyo G. Enki and Nickolay T. Trendafilov. Sparse principal components by semi-partition clustering. Computational Statistics, 27(4): 605–626, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0280-2.

Coelho:2012:NED

[679] Carlos A. Coelho and Filipe J. Marques. Near-exact distributions for the likelihood ratio test statistic to test equality of several variance-covariance matrices in elliptically contoured distributions. Computational Statistics, 27 (4):627–659, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0281-1.

Ogasawara:2012:AEA

[680] Haruhiko Ogasawara. Asymptotic expansions for the ability estimator in item response theory. Computational Statistics, 27(4):661–683, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0282-0.

Shim:2012:EVR

[681] Jooyong Shim, Yongtae Kim, Jangtaek Lee, and Changha Hwang. Estimating value at risk with semiparametric support vector quantile regression. Computational Statistics, 27(4): 685–700, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0283-z.

Takai:2012:CEA

[682] Keiji Takai. Constrained EM algorithm with projection method. Computational Statistics, 27(4):701–714, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0285-x.

Prates:2012:FSS

[683] Marcos O. Prates, Renato M. Assunção, and Marcelo A. Costa. Flexible scan statistic test to detect disease clusters in hierarchical trees. Computational Statistics, 27(4):715–737, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0286-9.

Kawai:2012:LRG

[684] Reiichiro Kawai. Likelihood ratio gradient estimation for Meixner distribution and Lévy processes. Computational Statistics, 27(4):739–755, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0288-7.

Schellhase:2012:DEC

[685] Christian Schellhase and Göran Kauermann. Density estimation and comparison with a penalized mixture approach. Computational Statistics, 27 (4):757–777, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0289-6.

Jentsch:2012:HBA

[686] C. Jentsch, J.-P. Kreiss, P. Mantalos, and E. Paparoditis. Hybrid bootstrap aided unit root testing. Computational Statistics, 27(4):779–797, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0290-0.

Edwards:2012:NAD

[687] David Edwards. A note on adding and deleting edges in hierarchical log-linear models. Computational Statistics, 27 (4):799–803, December 2012. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0287-8;   http:
//link.springer.com/content/pdf/
10.1007/s00180-011-0287-8.pdf.

Binder:2013:PR

[688] Harald Binder, Hans A. Kestler, and Matthias Schmid. Proceedings of Reisensburg 2010. Computational Statistics, 28(1):1–3, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0394-9;   http://link.springer.
com/content/pdf/10.1007/s00180-
013-0394-9.pdf.

Schels:2013:DEE

[689] Martin Schels, Stefan Scherer, Michael Glodek, Hans A. Kestler, Günther Palm, and Friedhelm Schwenker. On the discovery of events in EEG data utilizing information fusion. Computational Statistics, 28(1):5–18, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0292-y.

Hopfensitz:2013:ABN

[690] Martin Hopfensitz, Christoph Müssel, Markus Maucher, and Hans A. Kestler. Attractors in Boolean networks: a tutorial. Computational Statistics, 28 (1):19–36, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0324-2.

Reiser:2013:CMI

[691] Veronika Reiser, Christine Porzelius, Susanne Stampf, Martin Schumacher, and Harald Binder. Can matching improve the performance of boosting for identifying important genes in observational studies? Computational Statistics, 28(1):37–49, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0306-4.

Lausser:2013:MVS

[692] Ludwig Lausser, Christoph Müssel, Markus Maucher, and Hans A. Kestler. Measuring and visualizing the stability of biomarker selection techniques. Computational Statistics, 28 (1):51–65, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0284-y.

Telaar:2013:BDC

[693] Anna Telaar, Dirk Repsilber, and Gerd Nürnberg. Biomarker discovery: classification using pooled samples. Computational Statistics, 28 (1):67–106, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0302-0.

Unkel:2013:ZZE

[694] Steffen Unkel and Nickolay T. Trendafilov. Zig-zag exploratory factor analysis with more variables than observations. Computational Statistics, 28(1):107–125, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-011-0275-
z.

Glodek:2013:EGM

[695] Michael Glodek, Martin Schels, and Friedhelm Schwenker. Ensemble Gaussian mixture models for probability density estimation. Computational Statistics, 28(1):127–138, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0374-5.

Luck:2013:SAI

[696] Sebastian Lück, Alois Fichtl, Michaela Sailer, Helga Joos, Rolf E. Brenner, Paul Walther, and Volker Schmidt. Statistical analysis of the intermediate filament network in cells of mesenchymal lineage by greyvalue-oriented image segmentation. Computational Statistics, 28(1):139–160, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0265-1.

Bee:2013:MLE

[697] Marco Bee, Roberto Benedetti, and Giuseppe Espa. On maximum likelihood estimation of a Pareto mixture. Computational Statistics, 28 (1):161–178, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0291-z.

Maki:2013:IHV

[698] Daiki Maki. The influence of heteroskedastic variances on cointegration tests: A comparison using Monte Carlo simulations. Computational Statistics, 28(1):179–198, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-011-
                                        0293-x.

Yang:2013:TAM

[699] Ya-Wen Yang and Chong Yau Fu. Two advanced methods for adjusting the main coefficient in logistic regression. Computational Statistics, 28 (1):199–218, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0294-9.

Tsai:2013:IEF

[700] Jia-Ren Tsai. Interval estimation for fitting straight line when both variables are subject to error. Computational Statistics, 28(1):219–240, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0295-8.

Fontana:2013:AGM

[701] Roberto Fontana. Algebraic generation of minimum size orthogonal fractional factorial designs: an approach based on integer linear programming. Computational Statistics, 28(1):241–253, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-011-
0296-7.

Pfyffer:2013:ESA

[702] Samuel Pfyffer and Riccardo Gatto. An efficient simulation algorithm for the generalized von Mises distribution of order two. Computational Statistics, 28 (1):255–268, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-011-0297-6.

Martinez-Camblor:2013:DCI

[703] Pablo Martínez-Camblor and Jacobo de Uña-Álvarez. Density comparison for independent and right censored samples via kernel smoothing. Computational Statistics, 28(1): 269–288, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0298-5.

Abd-Elfattah:2013:SAT

[704] Ehab F. Abd-Elfattah. Saddlepoint approximations for tests of dispersion. Computational Statistics, 28(1): 289–306, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0299-4.

Wilde:2013:ECG

[705] Sarah L. Wilde and Scott D. Grimshaw. Efficient computation of generalized median estimators. Computational Statistics, 28(1):307– 317, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0300-2.

Agostinelli:2013:WSH

[706] Claudio Agostinelli and Luca Greco. A weighted strategy to handle likelihood uncertainty in Bayesian inference. Computational Statistics, 28 (1):319–339, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-011-0301-1.

Maruo:2013:PEB

[707] Kazushi Maruo and Masashi Goto. Percentile estimation based on the power-normal distribution. Computational Statistics, 28(1):341–356, February 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0303-7.

Lund:2013:MCM

[708] Ulric J. Lund. Monte Carlo maximum likelihood circle fitting using circular density functions. Computational Statistics, 28(2):393–411, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0307-3.

Aneiros-Perez:2013:TLS

[709] Germán Aneiros-Pérez and Philippe Vieu. Testing linearity in semiparametric functional data analysis. Computational Statistics, 28 (2):413–434, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0308-2.

Bissantz:2013:ASR

[710] Nicolai Bissantz and Hajo Holzmann. Asymptotics for spectral regularization estimators in statistical inverse problems. Computational Statistics, 28(2):435–453, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0309-1.

Blumentritt:2013:RRR

[711] Thomas Blumentritt and Oliver Grothe. Ranking ranks: a ranking algorithm for bootstrapping from the empirical copula. Computational Statistics, 28(2):455–462, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0310-8.

Ross:2013:SMB

[712] Gordon J. Ross, Dimitris K. Tasoulis, and Niall M. Adams. Sequential monitoring of a Bernoulli sequence when the pre-change parameter is unknown. Computational Statistics, 28 (2):463–479, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0311-7.

Ahmad:2013:UWB

[713] Taleb Ahmad, Wolfgang Härdle, Sigbert Klinke, and Shafiqah Alawadhi. Using wiki to build an e-learning system in statistics in the Arabic language. Computational Statistics, 28 (2):481–491, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0312-6.

Park:2013:LWL

[714] Heewon Park and Fumitake Sakaori. Lag weighted lasso for time series model. Computational Statistics, 28 (2):493–504, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0313-5.

Noufaily:2013:MLG

[715] Angela Noufaily and M. C. Jones. On maximization of the likelihood for the generalized gamma distribution. Computational Statistics, 28(2):505–517, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0314-4.

Dippold:2013:VSM

[716] Katrin Dippold and Harald Hruschka. Variable selection for market basket analysis. Computational Statistics, 28(2):519–539, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0315-3.

Weiss:2013:MTO

[717] Christian H. Weiß. Monitoring k th order runs in binary processes. Computational Statistics, 28(2):541–562, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-

0316-2. See erratum [718].

Weiss:2013:EMO

[718] Christian H. Weiß. Erratum to: Monitoring k th order runs in binary processes. Computational Statistics, 28(2):563, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0334-0;   http:
//link.springer.com/content/pdf/
10.1007/s00180-012-0334-0.pdf. See

Henseler:2013:GFI

[719] Jörg Henseler and Marko Sarstedt. Goodness-of-fit indices for partial least squares path modeling. Computational Statistics, 28(2):565–580, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0317-1;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        012-0317-1.pdf.

Shen:2013:RAI

[720] Pao sheng Shen. Regression analysis of interval censored and doubly truncated data with linear transformation models. Computational Statistics, 28(2):581–596, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0318-0.

Cagnone:2013:LVM

[721] Silvia Cagnone and Paola Monari. Latent variable models for ordinal data by using the adaptive quadrature approximation. Computational Statistics, 28(2):597–619, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0319-z.

Bacciu:2013:EII

[722] Davide Bacciu, Terence A. Etchells, Paulo J. G. Lisboa, and Joe Whittaker. Efficient identification of independence networks using mutual information. Computational Statistics, 28(2):621–646, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0320-6.

Hsu:2013:LCB

[723] Man-Jen Hsu and Huey-Miin Hsueh. The linear combinations of biomarkers which maximize the partial area under the ROC curves. Computational Statistics, 28(2):647–666, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0321-5.

Mehmood:2013:EPC

[724] Rashid Mehmood, Muhammad Riaz, and Ronald J. M. M. Does. Efficient power computation for r out of m runs rules schemes. Computational Statistics, 28(2):667–681, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0322-4;   http://link.springer.
com/content/pdf/10.1007/s00180-
012-0322-4.pdf.

Mohammadi:2013:UMG

[725] A. Mohammadi, M. R. Salehi-Rad, and E. C. Wit. Using mixture of gamma distributions for Bayesian analysis in an M/ G/ 1 queue with optional second service. Computational Statistics, 28(2):683–700, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0323-3;   http:
//link.springer.com/content/pdf/
10.1007/s00180-012-0323-3.pdf.

Perez-Rodriguez:2013:ECV

[726] Jorge V. Pérez-Rodríguez and Julián Andrada-Félix. Estimating critical values for testing the i.i.d. in standardized residuals from GARCH models in finite samples. Computational Statistics, 28(2):701–734, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0325-1.

Chen:2013:IFS

[727] Qihui Chen and Yu Ren. Improvement in finite-sample properties of GMM-based Wald tests. Computational Statistics, 28(2):735–749, April 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0326-0.

Balakrishnan:2013:CAS

[728] N. Balakrishnan, E. Beutner, and E. Cramer. Computational aspects of statistical intervals based on two Type–II censored samples. Computational Statistics, 28(3):893–917, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0335-z.

Cano:2013:BMS

[729] J. A. Cano, C. Carazo, and D. Salmerón. Bayesian model selection approach to the one way analysis of variance under homoscedasticity. Computational Statistics, 28(3):919–931, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0339-8.

Hashimoto:2013:NNT

[730] Elizabeth M. Hashimoto, Gauss M. Cordeiro, and Edwin M. M. Ortega. The new Neyman type a beta Weibull model with long-term survivors. Computational Statistics, 28(3):933–954, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0338-9.

Taylor:2013:CCD

[731] James Taylor and Jochen Einbeck. Challenging the curse of dimensionality in multivariate local linear regression. Computational Statistics, 28(3):955–976, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0342-0.

Nadarajah:2013:CRP

[732] Saralees Nadarajah, Bozidar V. Popović, and Miroslav M. Ristić. Compounding: an R package for computing continuous distributions obtained by compounding a continuous and a discrete distribution. Computational Statistics, 28(3):977–992, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0336-y.

Chua:2013:TMA

[733] K. C. Chua and S. H. Ong. Test of misspecification with application to negative binomial distribution. Computational Statistics, 28(3):993–1009, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0345-x.

Shirvani:2013:CTC

[734] A. Shirvani and A. R. Soltani. A characterization for truncated Cauchy random variables with nonzero skewness parameter. Computational Statistics, 28(3):1011–1016, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0340-2.

Zahid:2013:REM

[735] Faisal Maqbool Zahid and Gerhard Tutz. Ridge estimation for multinomial logit models with symmetric side constraints. Computational Statistics, 28(3):1017–1034, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0341-1.

Cortese:2013:AHO

[736] Giuliana Cortese and Laura Ventura. Accurate higher-order likelihood inference on P (Y &lt; X). Computational Statistics, 28(3):1035–1059, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0343-z.

Kisielinska:2013:EBM

[737] Joanna Kisielinska. The exact bootstrap method shown on the example of the mean and variance estimation. Computational Statistics, 28 (3):1061–1077, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0350-0;   http:
//link.springer.com/content/pdf/
10.1007/s00180-012-0350-0.pdf.

Hofner:2013:VSM

[738] Benjamin Hofner, Torsten Hothorn, and Thomas Kneib. Variable selection and model choice in structured survival models. Computational Statistics, 28 (3):1079–1101, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0337-x.

Chen:2013:SPQ

[739] Cathy W. S. Chen and Richard Gerlach. Semi-parametric quantile estimation for double threshold autoregressive models with heteroskedasticity. Computational Statistics, 28(3):1103–1131, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0346-9.

Inoue:2013:DNR

[740] Kiyoshi Inoue and Sigeo Aki. Distributions of numbers of runs and scans on directed acyclic graphs with generation. Computational Statistics, 28 (3):1133–1150, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0349-6.

Hung:2013:EIV

[741] Ying-Chao Hung and Neng-Fang Tseng. Extracting informative variables in the validation of two-group causal relationship. Computational Statistics, 28(3):1151–1167, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-012-0351-
                                        z.

Hussein:2013:GSC

[742] A. Hussein, H. A. Muttlak, and M. Saleh. Group sequential comparison of two binomial proportions under ranked set sampling design. Computational Statistics, 28(3):1169–1194, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0347-8.

Mbalawata:2013:PES

[743] Isambi S. Mbalawata, Simo Särkkä, and Heikki Haario. Parameter estimation in stochastic differential equations with Markov chain Monte Carlo and non-linear Kalman filtering. Computational Statistics, 28(3):1195–1223, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0352-y.

Guegan:2013:OTD

[744] Dominique Guégan and Philippe de Peretti. An omnibus test to detect time-heterogeneity in time series. Computational Statistics, 28(3):1225–1239, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0356-7.

Zheng:2013:ELM

[745] Ming Zheng and Wen Yu. Empirical likelihood method for multivariate Cox regression. Computational Statistics, 28(3):1241–1267, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0348-7.

Parra-Frutos:2013:THV

[746] I. Parra-Frutos. Testing homogeneity of variances with unequal sample sizes. Computational Statistics, 28 (3):1269–1297, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0353-x.

Lue:2013:HDR

[747] Heng-Hui Lue and Bing-Ran You. High-dimensional regression analysis with treatment comparisons. Computational Statistics, 28(3):1299–1317, June 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0357-6.

Mori:2013:MLP

[748] Yuichi Mori, Helmut Herwartz, and Jürgen Symanzik. A memorial for the Late Professor Wolfgang Polasek. Computational Statistics, 28 (4):1383–1384, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0431-8;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-013-0431-8.pdf.

Holst:2013:LLV

[749] Klaus Kähler Holst and Esben Budtz-Jørgensen. Linear latent variable models: the lava-package. Computational Statistics, 28(4):1385–1452, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0344-y.

Zhang:2013:SJC

[750] Jin Zhang. Simplification of joint confidence regions for the parameters of the Pareto distribution. Computational Statistics, 28(4):1453–1462, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0354-9.

Alonso:2013:SIM

[751] Andrés M. Alonso, Ana E. Sipols, and Silvia Quintas. A single-index model procedure for interpolation intervals in time series. Computational Statistics, 28(4):1463–1484, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0355-8.

Jiang:2013:MLE

[752] Jie Jiang, Xinsheng Liu, and Keming Yu. Maximum likelihood estimation of multinomial probit factor analysis models for multivariate t-distribution. Computational Statistics, 28(4):1485–1500, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0363-8.

Tyralis:2013:ACM

[753] Hristos Tyralis, Demetris Koutsoyiannis, and Stefanos Kozanis. An algorithm to construct Monte Carlo confidence intervals for an arbitrary function of probability distribution parameters. Computational Statistics, 28 (4):1501–1527, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0364-7.

Brouste:2013:PED

[754] Alexandre Brouste and Stefano M. Iacus. Parameter estimation for the discretely observed fractional Ornstein– Uhlenbeck process and the Yuima R package. Computational Statistics, 28 (4):1529–1547, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0365-6.

Azzimonti:2013:NNM

[755] Laura Azzimonti, Francesca Ieva, and Anna Maria Paganoni. Nonlinear nonparametric mixed-effects models for unsupervised classification. Computational Statistics, 28(4):1549–1570, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-012-
                                        0366-5.

Bagnato:2013:FMU

[756] Luca Bagnato and Antonio Punzo. Finite mixtures of unimodal beta and gamma densities and the k-bumps algorithm. Computational Statistics, 28 (4):1571–1597, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0367-4.

Wang:2013:SIR

[757] Shun-Fang Wang and Xue-Ren Wang. Statistical inference of risk ratio in a correlated 2 × 2 table with structural zero. Computational Statistics, 28 (4):1599–1615, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0368-3.

Morell:2013:RCV

[758] Oliver Morell, Dennis Otto, and Roland Fried. On robust crossvalidation for nonparametric smoothing. Computational Statistics, 28(4): 1617–1637, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0369-2.

Chien:2013:MDD

[759] Li-Chu Chien. Multiple deletion diagnostics in beta regression models. Computational Statistics, 28(4):1639–1661, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-012-
0370-9.

deOliveira:2013:CNG

[760] Izabela Regina Cardoso de Oliveira and Daniel Furtado Ferreira. Computing the noncentral gamma distribution, its inverse and the noncentrality parameter. Computational Statistics, 28 (4):1663–1680, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0371-8.

Tsou:2013:EIC

[761] Tsung-Shan Tsou and Wan-Chen Chen. Estimation of intra-cluster correlation coefficient via the failure of Bartlett’s second identity. Computational Statistics, 28(4):1681–1698, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0372-7.

Wang:2013:SEA

[762] Lichun Wang, Yuan You, and Heng Lian. A simple and efficient algorithm for fused lasso signal approximator with convex loss function. Computational Statistics, 28(4):1699–1714, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0373-6.

Zhang:2013:RBM

[763] Jin Zhang. Reducing bias of the maximum likelihood estimator of shape parameter for the gamma distribution. Computational Statistics, 28(4):

1715–1724, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0375-4.

Wensch:2013:EVC

[764] Jörg Wensch, Monika Wensch-Dorendorf, and Hermann H. Swalve. The evaluation of variance component estimation software: generating benchmark problems by exact and approximate methods. Computational Statistics, 28 (4):1725–1748, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0376-3.

Botts:2013:ARA

[765] Carsten Botts. An accept-reject algorithm for the positive multivariate normal distribution. Computational Statistics, 28(4):1749–1773, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0377-2.

Delsol:2013:NET

[766] Laurent Delsol. No effect tests in regression on functional variable and some applications to spectrometric studies. Computational Statistics, 28 (4):1775–1811, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0378-1.

Chen:2013:CET

[767] Ray-Bing Chen, Ying-Chao Hung, Weichung Wang, and Sung-Wei Yen. Contour estimation via two fidelity computer simulators under limited resources. Computational Statistics, 28 (4):1813–1834, August 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0380-7.

Ahmed:2013:PDF

[768] Shamsuddin Ahmed. Performance of derivative free search ANN training algorithm with time series and classification problems. Computational Statistics, 28(5):1881–1914, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0386-1.

Yu:2013:FAP

[769] Philip L. H. Yu, Paul H. Lee, and W. M. Wan. Factor analysis for paired ranked data with application on parent-child value orientation preference data. Computational Statistics, 28 (5):1915–1945, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0387-0.

Jochmann:2013:WBW

[770] Markus Jochmann. What belongs where? variable selection for zeroinflated count models with an application to the demand for health care. Computational Statistics, 28(5): 1947–1964, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0388-z.

Park:2013:RFD

[771] Minjeong Park, Sinsup Cho, and Hee-Seok Oh. The role of functional data analysis for instantaneous frequency estimation. Computational Statistics, 28 (5):1965–1987, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0389-y.

Murray:2013:RFM

[772] Kevin Murray, Samuel Müller, and Berwin A. Turlach. Revisiting fitting monotone polynomials to data. Computational Statistics, 28(5):1989– 2005, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0390-5.

Rikhtehgaran:2013:SPB

[773] Reyhaneh Rikhtehgaran and Iraj Kazemi. Semi-parametric Bayesian estimation of mixed-effects models using the multivariate skew-normal distribution. Computational Statistics, 28(5): 2007–2027, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0392-3.

Zhao:2013:AON

[774] Hui-Xiu Zhao and Jin-Guan Lin. An approximately optimal non-parametric procedure for analyzing exchangeable binary data with random cluster sizes. Computational Statistics, 28(5): 2029–2047, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0393-2.

Solonen:2013:PAS

[775] Antti Solonen. Proposal adaptation in simulated annealing for continuous optimization problems. Computational Statistics, 28(5):2049–2065, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0395-8.

Nolan:2013:MEC

[776] John P. Nolan. Multivariate elliptically contoured stable distributions: theory and estimation. Computational Statistics, 28(5):2067–2089, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0396-7.

Marques:2013:OEN

[777] Filipe J. Marques and Carlos A. Coelho. Obtaining the exact and nearexact distributions of the likelihood ratio statistic to test circular symmetry through the use of characteristic functions. Computational Statistics, 28(5): 2091–2115, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0398-5.

Le-Rademacher:2013:PCH

[778] J. Le-Rademacher and L. Billard. Principal component histograms from interval-valued observations. Computational Statistics, 28(5):2117–2138, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0399-4.

Nadarajah:2013:NRP

[779] S. Nadarajah and S. A. A. Bakar. A new R package for actuarial survival models. Computational Statistics, 28 (5):2139–2160, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0400-2.

Antoch:2013:TMC

[780] Jaromír Antoch and Daniela Jarusková. Testing for multiple change points. Computational Statistics, 28(5):2161– 2183, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0401-1.

Meira-Machado:2013:BSE

[781] Luís Meira-Machado, Javier Roca-Pardiñas, Ingrid Van Keilegom, and Carmen Cadarso-Suárez. Bandwidth selection for the estimation of transition probabilities in the locationscale progressive three-state model. Computational Statistics, 28(5):2185– 2210, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0402-0.

Mak:2013:GCP

[782] Tak K. Mak and Fassil Nebebe. On a general class of probability distributions and its applications. Computational Statistics, 28(5):2211–2230, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0403-z.

Volterman:2013:EIC

[783] William Volterman and N. Balakrishnan. Efficient iterative computation of mixture weights for pooled order statistics for meta-analysis of multiple type-II right censored data. Computational Statistics, 28(5):2231–2239, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0404-y.

Xiao:2013:MPL

[784] Yuanhui Xiao and Huayu Liu. Modified profile likelihood approach for certain intraclass correlation coefficients. Computational Statistics, 28 (5):2241–2265, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0405-x.

Migon:2013:ESS

[785] Helio S. Migon, Alexandra M. Schmidt, Romy E. R. Ravines, and João B. M. Pereira. An efficient sampling scheme for dynamic generalized models. Computational Statistics, 28(5): 2267–2293, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0406-9.

Chen:2013:DFR

[786] Shu-Chun Chen, Hsieh Fushing, and Chii-Ruey Hwang. Discovering focal regions of slightly-aggregated sparse signals. Computational Statistics, 28(5): 2295–2308, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0407-8.

Contreras-Reyes:2013:SAA

[787] Javier E. Contreras-Reyes and Wilfredo Palma. Statistical analysis of autoregressive fractionally integrated moving average models in R. Computational Statistics, 28(5):2309–2331, October 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0408-7.

Mlodak:2013:NSA

[788] Andrzej Młodak. Neighbourhood of spatial areas in the physical and socioeconomical context. Computational Statistics, 28(6):2379–2414, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0411-z.

Chen:2013:TVS

[789] Cathy W. S. Chen, Feng-Chi Liu, and Mike K. P. So. Threshold variable selection of asymmetric stochastic volatility models. Computational Statistics, 28(6):2415–2447, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-013-
0412-y.

Adachi:2013:GJP

[790] Kohei Adachi. Generalized joint Procrustes analysis. Computational Statistics, 28(6):2449–2464, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0413-x.

Aminzadeh:2013:BEE

[791] M. S. Aminzadeh. Bayesian estimation of the expected time of first arrival past a truncated time t: the case of NHPP with power law intensity. Computational Statistics, 28(6): 2465–2477, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0414-9.

He:2013:ESC

[792] Qinying He, H. N. Nagaraja, and Chunjie Wu. Efficient simulation of complete and censored samples from common bivariate exponential distributions. Computational Statistics, 28(6): 2479–2494, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0415-8.

Zhu:2013:MHD

[793] Yayuan Zhu, Jingjing Wu, and Xuewen Lu. Minimum Hellinger distance estimation for a two-sample semiparametric cure rate model with censored survival data. Computational Statistics, 28(6):2495–2518, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0416-7.

Saleh:2013:IRR

[794] A. K. Md. E. Saleh and B. M. Golam Kibria. Improved ridge regression estimators for the logistic regression model. Computational Statistics, 28(6): 2519–2558, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0417-6.

Ogasawara:2013:APB

[795] Haruhiko Ogasawara. Asymptotic properties of the Bayes modal estimators of item parameters in item response theory. Computational Statistics, 28(6):2559–2583, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-013-0418-
5.

Reschenhofer:2013:NMP

[796] Erhard Reschenhofer, David Preinerstorfer, and Lukas Steinberger. Nonmonotonic penalizing for the number of structural breaks. Computational Statistics, 28(6):2585–2598, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0419-4.

Bischl:2013:BLC

[797] Bernd Bischl, Julia Schiffner, and Claus Weihs. Benchmarking local classification methods. Computational Statistics, 28(6):2599–2619, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0420-y.

Jafari:2013:PBA

[798] Ali Akbar Jafari and Mohammad Reza Kazemi. A parametric bootstrap approach for the equality of coefficients of variation. Computational Statistics, 28(6):2621–2639, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0421-x.

Guedon:2013:ELS

[799] Yann Guédon. Exploring the latent segmentation space for the assessment of multiple change-point models. Computational Statistics, 28(6): 2641–2678, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0422-9.

Stober:2013:ESE

[800] Jakob Stöber and Ulf Schepsmeier. Estimating standard errors in regular vine copula models. Computational Statistics, 28(6):2679–2707, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-013-
                                        0423-8.

Parpia:2013:ECI

[801] Sameer Parpia, John J. Koval, and Allan Donner. Evaluation of confidence intervals for the kappa statistic when the assumption of marginal homogeneity is violated. Computational Statistics, 28(6):2709–2718, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0424-7.

Ahlgren:2013:PBT

[802] Niklas Ahlgren and Jan Antell. The power of bootstrap tests of cointegration rank. Computational Statistics, 28(6):2719–2748, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-013-0425-
                                        6.

Musella:2013:PAV

[803] Flaminia Musella. A PC algorithm variation for ordinal variables. Computational Statistics, 28(6):2749– 2759, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0426-5.

Gonzalez:2013:TLG

[804] Eduardo Gutiérrez González, José A. Villaseñor Alva, Olga Vladimirovna Panteleeva, and Humberto Vaquera Huerta. On testing the log-gamma distribution hypothesis by bootstrap.

Computational Statistics, 28(6):2761– 2776, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0427-4.

Lenormand:2013:AAB

[805] Maxime Lenormand, Franck Jabot, and Guillaume Deffuant. Adaptive approximate Bayesian computation for complex models. Computational Statistics, 28(6):2777–2796, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0428-3.

Martino:2013:FDM

[806] Luca Martino and Jesse Read. On the flexibility of the design of multiple try Metropolis schemes. Computational Statistics, 28(6):2797–2823, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0429-2.

Basiri:2013:NPI

[807] Elham Basiri and Jafar Ahmadi. Nonparametric prediction intervals for progressive Type–II censored order statistics based on k-records. Computational Statistics, 28(6):2825–2848, December 2013. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0430-9.

Binder:2014:PR

[808] Harald Binder, Hans A. Kestler, and Matthias Schmid. Proceedings of Reisensburg 2011. Computational Statistics, 29(1–2):1–2, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0475-9;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        013-0475-9.pdf.

Hofner:2014:MBB

[809] Benjamin Hofner, Andreas Mayr, Nikolay Robinzonov, and Matthias Schmid. Model-based boosting in R: a handson tutorial using the R package mboost. Computational Statistics, 29 (1–2):3–35, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-012-0382-5.

Krey:2014:MTS

[810] Sebastian Krey, Uwe Ligges, and Friedrich Leisch. Music and timbre segmentation by recursive constrained K-means clustering. Computational Statistics, 29(1–2):37–50, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-012-
                                        0358-5.

Hieke:2014:MRB

[811] Stefanie Hieke, Harald Binder, Alexandra Nieters, and Martin Schumacher. minPtest: a resampling based gene region-level testing procedure for genetic case-control studies. Computational Statistics, 29(1–2):51–63, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0391-4.

Schmidt:2014:SGF

[812] Miriam Schmidt, Günther Palm, and Friedhelm Schwenker. Spectral graph features for the classification of graphs and graph sequences. Computational Statistics, 29(1–2):65–80, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0381-6.

Lausser:2014:IPH

[813] Ludwig Lausser, Christoph Müssel, Alexander Melkozerov, and Hans A. Kestler. Identifying predictive hubs to condense the training set of knearest neighbour classifiers. Computational Statistics, 29(1–2):81–95, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-012-
0379-0.

Maucher:2014:IBF

[814] Markus Maucher, David V. Kracht, Steffen Schober, Martin Bossert, and Hans A. Kestler. Inferring Boolean functions via higher-order correlations. Computational Statistics, 29(1– 2):97–115, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-012-0385-2.

Cook:2014:DEA

[815] Dianne Cook. The 2011 data Expo of the American Statistical Association. Computational Statistics, 29(1– 2):117–119, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0474-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-013-0474-x.pdf.

Follett:2014:GED

[816] Lendie Follett, Ulrike Genschel, and Heike Hofmann. A graphical exploration of the deepwater horizon oil spill. Computational Statistics, 29(1– 2):121–132, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0432-7.

Tran:2014:EOS

[817] Tony Tran, Aida Yazdanparast, and Eric A. Suess. Effect of oil spill on birds: A graphical assay of the deepwater horizon oil Spill’s impact on birds. Computational Statistics, 29(1– 2):133–140, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0472-z.

Li:2014:DIA

[818] Tianxi Li, Chao Gao, Meng Xu, and Bala Rajaratnam. Detecting the impact area of BP deepwater horizon oil discharge: an analysis by time varying coefficient logistic models and boosted trees. Computational Statistics, 29(1– 2):141–157, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0449-y.

Fishman:2014:CSC

[819] George S. Fishman. Counting subsets of contingency tables. Computational Statistics, 29(1–2):159–187, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0442-5.

Lee:2014:DME

[820] Ming Ha Lee and Michael B. C. Khoo. Design of a multivariate exponentially weighted moving average control chart with variable sampling intervals. Computational Statistics, 29(1– 2):189–214, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0443-4.

Krishnamoorthy:2014:ITE

[821] K. Krishnamoorthy and Meesook Lee. Improved tests for the equality of normal coefficients of variation. Computational Statistics, 29(1– 2):215–232, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0445-2.

Ouysse:2014:PBB

[822] Rachida Ouysse. On the performance of block-bootstrap continuously updated GMM for a class of non-linear conditional moment models. Computational Statistics, 29(1– 2):233–261, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0447-0.

Zhou:2014:SND

[823] Lan Zhou and Huijun Pan. Smoothing noisy data for irregular regions using penalized bivariate splines on triangulations. Computational Statistics, 29(1–2):263–281, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-013-0448-
z.

Kauermann:2014:PML

[824] Göran Kauermann and Renate Meyer. Penalized marginal likelihood estimation of finite mixtures of Archimedean copulas. Computational Statistics, 29(1–2):283–306, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-013-0454-
1.

Bertrand:2014:HLC

[825] Aurélie M. E. Bertrand and Christian M. Hafner. On heterogeneous latent class models with applications to the analysis of rating scores. Computational Statistics, 29(1– 2):307–330, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0450-5.

Chang:2014:BCI

[826] Yi-Ping Chang and Chih-Tun Yu. Bayesian confidence intervals for probability of default and asset correlation of portfolio credit risk. Computational Statistics, 29(1–2):331–361, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0453-2.

Jimenez-Gamero:2014:MDE

[827] M. D. Jiménez-Gamero, R. Pino-Mejías, and A. Rufián-Lizana. Minimum K_{ϕ}-divergence estimators for multinomial models and applications. Computational Statistics, 29(1–2): 363–401, February 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0452-3.

Trendafilov:2014:SMD

[828] Nickolay Trendafilov, Martin Kleinsteuber, and Hui Zou. Sparse matrices in data analysis. Computational Statistics, 29(3–4):403–405, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0468-8;   http://link.springer.
com/content/pdf/10.1007/s00180-
013-0468-8.pdf.

Buhlmann:2014:HDV

[829] Peter Bühlmann and Jacopo Mandozzi. High-dimensional variable screening and bias in subsequent inference, with an empirical comparison. Computational Statistics, 29(3–4):407–430, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0436-3.

Trendafilov:2014:SSS

[830] Nickolay T. Trendafilov. From simple structure to sparse components: a review. Computational Statistics, 29 (3–4):431–454, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0434-5.

Fang:2014:RPC

[831] Yixin Fang, Yang Feng, and Ming Yuan. Regularized principal components of heritability. Computational Statistics, 29(3–4):455–465, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0444-3.

Hage:2014:RPS

[832] Clemens Hage and Martin Kleinsteuber. Robust PCA and subspace tracking from incomplete observations using ℓ_{0}-surrogates. Computational Statistics, 29(3–4):467–487, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-013-
                                        0435-4.

Bouveyron:2014:DVS

[833] Charles Bouveyron and Camille Brunet-Saumard. Discriminative variable selection for clustering with the sparse Fisher–EM algorithm. Computational Statistics, 29(3–4):489–513, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0433-6.

Choy:2014:SDM

[834] Tze Choy and Nicolai Meinshausen. Sparse distance metric learning. Computational Statistics, 29(3–4):515–528, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0437-2.

GSell:2014:SAI

[835] Max Grazier G’Sell, Shai S. Shen-Orr, and Robert Tibshirani. Sensitivity analysis for inference with partially identifiable covariance matrices. Computational Statistics, 29(3–4):529–546, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0451-4.

Krahmer:2014:SMF

[836] Felix Krahmer, Gitta Kutyniok, and Jakob Lemvig. Sparse matrices in frame theory. Computational Statistics, 29(3–4):547–568, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0446-1.

Absil:2014:TNM

[837] P.-A. Absil, Luca Amodei, and Gilles Meyer. Two Newton methods on the manifold of fixed-rank matrices endowed with Riemannian quotient geometries. Computational Statistics, 29 (3–4):569–590, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0441-6.

Mishra:2014:FRM

[838] Bamdev Mishra, Gilles Meyer, Silvère Bonnabel, and Rodolphe Sepulchre. Fixed-rank matrix factorizations and Riemannian low-rank optimization. Computational Statistics, 29 (3–4):591–621, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0464-z.

Chen:2014:STN

[839] Jianhui Chen and Jieping Ye. Sparse trace norm regularization. Computational Statistics, 29(3–4):623–639, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0440-7.

Shen:2014:SNE

[840] Pao sheng Shen. Simple nonparametric estimators of the bivariate survival function under random left truncation and right censoring. Computational Statistics, 29(3–4):641–659, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0455-0.

Koyama:2014:HGD

[841] Tamio Koyama, Hiromasa Nakayama, Kenta Nishiyama, and Nobuki Takayama. Holonomic gradient descent for the Fisher–Bingham distribution on the d-dimensional sphere. Computational Statistics, 29(3–4):661–683, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0456-z.

Salinas-Gutierrez:2014:CSG

[842] Rogelio Salinas-Gutiérrez, Arturo Hernández-Aguirre, and Enrique R. Villa-Diharce. Copula selection for graphical models in continuous estimation of distribution algorithms. Computational Statistics, 29(3–4):685–713, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0457-y.

Marra:2014:THA

[843] Giampiero Marra, Rosalba Radice, and Silvia Missiroli. Testing the hypothesis of absence of unobserved confounding in semiparametric bivariate probit models. Computational Statistics, 29 (3–4):715–741, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0458-x;   http:
//link.springer.com/content/pdf/
10.1007/s00180-013-0458-x.pdf.

Cattaneo:2014:ILC

[844] Marco E. G. V. Cattaneo and Andrea Wiencierz. On the implementation of LIR: the case of simple linear regression with interval data. Computational Statistics, 29(3–4):743–767, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0459-9.

Lindsey:2014:USM

[845] Charles D. Lindsey, Simon J. Sheather, and Joseph W. McKean. Using sliced mean variance-covariance inverse regression for classification and dimension reduction. Computational Statistics, 29(3–4):769–798, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-013-0460-
3.

Rey:2014:FAS

[846] Sergio J. Rey. Fast algorithms for a space-time concordance measure. Computational Statistics, 29(3–4):799–811, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0461-2.

Shen:2014:SRA

[847] Pao sheng Shen. Semiparametric regression analysis for clustered doubly-censored data. Computational Statistics, 29(3–4):813–828, June 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0462-1.

Mlodak:2014:CAM

[848] Andrzej Młodak. On the construction of an aggregated measure of the development of interval data. Computational Statistics, 29(5):895–929, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0469-7.

Rosadi:2014:SOL

[849] Dedi Rosadi and Shelton Peiris. Second-order least-squares estimation for regression models with autocorrelated errors. Computational Statistics, 29(5):931–943, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-013-0470-1.

Hornik:2014:MLE

[850] Kurt Hornik and Bettina Grün. On maximum likelihood estimation of the concentration parameter of von Mises–Fisher distributions. Computational Statistics, 29(5):945–957, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-013-
0471-0;   http://link.springer.
com/content/pdf/10.1007/s00180-
013-0471-0.pdf.

Conceicao:2014:ZMP

[851] Katiane S. Conceição, Marinho G. Andrade, and Francisco Louzada. On the zero-modified Poisson model: Bayesian analysis and posterior divergence measure. Computational Statistics, 29 (5):959–980, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0473-y.

Luo:2014:ERR

[852] Ji Luo, Jiajia Zhang, and Han Sun. Estimation of relative risk using a log-binomial model with constraints. Computational Statistics, 29 (5):981–1003, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0476-8.

Kim:2014:ABE

[853] Daeju Kim, Shuichi Kawano, and Yoshiyuki Ninomiya. Adaptive basis expansion via ℓ_{1} trend filtering. Computational Statistics, 29(5): 1005–1023, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0477-7.

Gau:2014:BAM

[854] Shiow-Lan Gau, Jean de Dieu Tapsoba, and Shen-Ming Lee. Bayesian approach for mixture models with grouped data. Computational Statistics, 29(5): 1025–1043, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-013-0478-6.

Niu:2014:PMA

[855] Hongli Niu and Jun Wang. Phase and multifractality analyses of random price time series by finite-range interacting biased voter system. Computational Statistics, 29(5):1045–1063, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0479-0.

Baker:2014:SAB

[856] Rose D. Baker and Dan Jackson. Statistical application of barycentric rational interpolants: an alternative to splines. Computational Statistics, 29 (5):1065–1081, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0480-7;   http:
//link.springer.com/content/pdf/
10.1007/s00180-014-0480-7.pdf.

Malekzadeh:2014:CEL

[857] A. Malekzadeh, M. Kharrati-Kopaei, and S. M. Sadooghi-Alvandi. Comparing exponential location parameters with several controls under heteroscedasticity. Computational Statistics, 29(5):1083–1094, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-014-0481-
6.

Melard:2014:ASP

[858] Guy Mélard. On the accuracy of statistical procedures in Microsoft Excel 2010. Computational Statistics, 29(5): 1095–1128, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0482-5.

Chavent:2014:SIR

[859] Marie Chavent, Stéphane Girard, Vanessa Kuentz-Simonet, Benoit Liquet, Thi Mong Ngoc Nguyen, and Jérôme Saracco. A sliced inverse regression approach for data stream. Computational Statistics, 29 (5):1129–1152, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0483-4.

Belaghi:2014:ICI

[860] R. Arabi Belaghi, M. Arashi, and S. M. M. Tabatabaey. Improved confidence intervals for the scale parameter of Burr XII model based on record values. Computational Statistics, 29 (5):1153–1173, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0484-3.

Wilcox:2014:CRR

[861] Rand R. Wilcox and Florence Clark. Comparing robust regression lines associated with two dependent groups when there is heteroscedasticity. Computational Statistics, 29(5):1175–1186, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0485-2.

Ng:2014:FAS

[862] Chi Tim Ng, Johan Lim, Kyeong Eun Lee, Donghyeon Yu, and Sujung Choi. A fast algorithm to sample the number of vertexes and the area of the random convex hull on the unit square. Computational Statistics, 29 (5):1187–1205, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0486-1.

Hayter:2014:RFM

[863] A. J. Hayter. Recursive formulas for multinomial probabilities with applications. Computational Statistics, 29 (5):1207–1219, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0487-0.

Meintanis:2014:VTI

[864] Simos G. Meintanis and Dimitris Karlis. Validation tests for the innovation distribution in INAR time series models. Computational Statistics, 29 (5):1221–1241, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0488-z.

Strzalkowska-Kominiak:2014:BBA

[865] Ewa Strzalkowska-Kominiak and Ricardo Cao. Beran-based approach for single-index models under censoring. Computational Statistics, 29(5): 1243–1261, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0489-y.

Espinheira:2014:BPI

[866] Patrícia L. Espinheira, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Bootstrap prediction intervals in beta regressions. Computational Statistics, 29 (5):1263–1277, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0490-5. See erra-

tum [1099].

Mirfarah:2014:PCR

[867] Elham Mirfarah and Jafar Ahmadi. Pitman closeness of k-records from two sequences to progressive Type–II censored order statistics. Computational Statistics, 29(5):1279–1300, October 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0491-4.

Zheng:2014:GNA

[868] Songfeng Zheng. A generalized Newton algorithm for quantile regression models. Computational Statistics, 29(6): 1403–1426, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0498-x.

Yee:2014:RCI

[869] Thomas W. Yee and Alfian F. Hadi. Row-column interaction models, with an R implementation. Computational Statistics, 29(6):1427–1445, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0499-9.

Safari:2014:ISV

[870] Amir Safari. An e-E-insensitive support vector regression machine.

Computational Statistics, 29(6):1447– 1468, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0500-7.

Barrett:2014:ECP

[871] Bruce E. Barrett and J. Brian Gray. Efficient computation for the Poisson binomial distribution. Computational Statistics, 29(6):1469–1479, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0501-6.

Li:2014:WIT

[872] Yushu Li and Simon Reese. Wavelet improvement in turning point detection using a 40hidden Markov model: from the aspects of cyclical identification and outlier correction. Computational Statistics, 29(6):1481–1496, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0502-5.

Chang:2014:FDC

[873] Chung Chang, Yakuan Chen, and R. Todd Ogden. Functional data classification: a wavelet approach. Computational Statistics, 29(6):1497–1513, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0503-4.

Jou:2014:VBO

[874] Yow-Jen Jou, Chien-Chia Liäm Huang, and Hsun-Jung Cho. A VIF-based optimization model to alleviate collinearity problems in multiple linear regression. Computational Statistics, 29(6): 1515–1541, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0504-3.

Chen:2014:DAA

[875] Shiyi Chen and Wolfgang K. Härdle. Dynamic activity analysis model-based win–win development forecasting under environment regulations in China. Computational Statistics, 29(6):1543– 1570, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0505-2. See erra-

tum [948].

Cheng:2014:QRR

[876] Jung-Yu Cheng and Shinn-Jia Tzeng. Quantile regression of right-censored length-biased data using the Buckley– James-type method. Computational Statistics, 29(6):1571–1592, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0507-0.

Pallmann:2014:LTT

[877] Philip Pallmann, Ludwig A. Hothorn, and Gemechis D. Djira. A levenetype test of homogeneity of variances against ordered alternatives. Computational Statistics, 29(6):1593–1608, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0508-z.

Guseo:2014:MNL

[878] Renato Guseo and Cinzia Mortarino. Multivariate nonlinear least squares: robustness and efficiency of standard versus Beauchamp and Cornell methodologies. Computational Statistics, 29(6):1609–1636, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0509-y.

Hardouin:2014:RME

[879] Cécile Hardouin and Xavier Guyon. Recursions on the marginals and exact computation of the normalizing constant for Gibbs processes. Computational Statistics, 29(6):1637–1650, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0510-5.

Shim:2014:CSV

[880] Jooyong Shim, Changha Hwang, and Kyungha Seok. Composite support vector quantile regression estimation. Computational Statistics, 29(6): 1651–1665, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0511-4.

Abebe:2014:RBO

[881] H. T. Abebe, F. E. S. Tan, G. J. P. Van Breukelen, and M. P. F. Berger. Robustness of Bayesian D-optimal design for the logistic mixed model against misspecification of autocorrelation. Computational Statistics, 29(6): 1667–1690, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0512-3.

Ogura:2014:RTB

[882] T. Ogura and H. Murakami. A rank test based on the moments of order statistics of the modified Makeham distribution. Computational Statistics, 29(6):1691–1711, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0513-2.

Yu:2014:TBM

[883] Wei Yu, Wangli Xu, and Lixing Zhu. Transformation-based model averaged tail area inference. Computational Statistics, 29(6):1713–1726, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0514-1.

Haggstrom:2014:TSP

[884] Jenny Häggström and Xavier de Luna. Targeted smoothing parameter selection for estimating average causal effects. Computational Statistics, 29(6): 1727–1748, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0515-0;   http:
//link.springer.com/content/pdf/
10.1007/s00180-014-0515-0.pdf.

Shah:2014:MLS

[885] Jasmit Shah, Somnath Datta, and Susmita Datta. A multi-loss super regression learner (MSRL) with application to survival prediction using proteomics. Computational Statistics, 29(6):1749– 1767, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0516-z.

Manukyan:2014:DPM

[886] Artür Manukyan, Erhan Çene, Ahmet Sedef, and Ibrahim Demir. Dandelion plot: a method for the visualization of R-mode exploratory factor analyses. Computational Statistics, 29(6): 1769–1791, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0518-x.

Sullivan:2014:PPS

[887] Francis Sullivan and Isabel Beichl. Permanents, α-permanents and Sinkhorn balancing. Computational Statistics, 29(6):1793–1798, December 2014. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-014-0506-
1.

Kim:2015:BLT

[888] Hea-Jung Kim. A best linear threshold classification with scale mixture of skew normal populations. Computational Statistics, 30(1):1–28, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0517-y.

Yang:2015:GCA

[889] Tae Young Yang. A GS–CORE algorithm for performing a reduction test on multiple gene sets and their core genes. Computational Statistics, 30(1):29–41, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0519-9.

Matsui:2015:VSV

[890] Hidetoshi Matsui and Toshihiro Misumi. Variable selection for varyingcoefficient models with the sparse regularization. Computational Statistics, 30(1):43–55, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0520-3.

Farcomeni:2015:EHM

[891] Alessio Farcomeni and Luca Greco. Sestimation of hidden Markov models. Computational Statistics, 30(1):57–80, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0521-2.

Gupta:2015:SCG

[892] Abhinav Gupta and Karmeshu. Study of compound generalized Nakagamigeneralized inverse Gaussian distribution and related densities: application to ultrasound imaging. Computational Statistics, 30(1):81–96, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0522-1.

Kim:2015:IMB

[893] Myung Geun Kim. Influence measure based on probabilistic behavior of regression estimators. Computational Statistics, 30(1):97–105, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0524-z.

Wang:2015:ECP

[894] Xinlei Wang, Johan Lim, Seung-Jean Kim, and Kyu S. Hahn. Estimating cell probabilities in contingency tables with constraints on marginals/conditionals by geometric programming with applications. Computational Statistics, 30(1):107–129, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0525-y.

Pereira:2015:GAA

[895] André G. C. Pereira and Bernardo B. de Andrade. On the genetic algorithm with adaptive mutation rate and selected statistical applications. Computational Statistics, 30(1):131–150, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-014-
                                        0526-x.

Balakrishnan:2015:EAE

[896] N. Balakrishnan and Suvra Pal. An EM algorithm for the estimation of parameters of a flexible cure rate model with generalized gamma lifetime and model discrimination using likelihoodand information-based methods. Computational Statistics, 30(1):151–189, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0527-9.

Zhu:2015:VSA

[897] Xiaoyi Zhu and Yuhong Yang. Variable selection after screening: with or without data splitting? Computational Statistics, 30(1):191–203, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0528-8.

Wang:2015:BSV

[898] Min Wang, Xiaoqian Sun, and Tao Lu. Bayesian structured variable selection in linear regression models. Computational Statistics, 30(1):205–229, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0529-7.

Neath:2015:MED

[899] Andrew A. Neath, Joseph E. Cavanaugh, and Adam G. Weyhaupt. Model evaluation, discrepancy function estimation, and social choice theory. Computational Statistics, 30 (1):231–249, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0532-z.

Indira:2015:ARM

[900] K. Indira and S. Kanmani. Association rule mining through adaptive parameter control in particle swarm optimization. Computational Statistics, 30(1):251–277, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0533-y.

Makino:2015:GDF

[901] Naomichi Makino. Generalized datafitting factor analysis with multiple quantification of categorical variables. Computational Statistics, 30 (1):279–292, March 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0536-8.

Chowdhury:2015:UVS

[902] Niladri Roy Chowdhury, Dianne Cook, Heike Hofmann, Mahbubul Majumder, Eun-Kyung Lee, and Amy L. Toth. Using visual statistical inference to better understand random class separations in high dimension, low sample size data. Computational Statistics, 30(2):293–316, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0534-x.

Murua:2015:KBM

[903] Alejandro Murua and Nicolas Wicker. Kernel-based mixture models for classification. Computational Statistics, 30(2):317–344, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0535-9.

Lim:2015:SCI

[904] Yaeji Lim and Hee-Seok Oh. Simultaneous confidence interval for quantile regression. Computational Statistics, 30(2):345–358, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0537-7.

Arashi:2015:SES

[905] Mohammad Arashi and Mahdi Roozbeh. Shrinkage estimation in system regression model. Computational Statistics, 30(2):359–376, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-014-0539-5.

Meira-Machado:2015:NEC

[906] Luís Meira-Machado, Jacobo de Uña-Álvarez, and Somnath Datta. Nonparametric estimation of conditional transition probabilities in a non-Markov illness-death model. Computational Statistics, 30(2):377–397, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0538-6.

Yang:2015:BVS

[907] Aijun Yang, Yunxian Li, Niansheng Tang, and Jinguan Lin. Bayesian variable selection in multinomial probit model for classifying high-dimensional data. Computational Statistics, 30 (2):399–418, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0540-z.

Paiva:2015:PST

[908] Thais Paiva, Renato Assunção, and Taynãna Simões. Prospective spacetime surveillance with cumulative surfaces for geographical identification of the emerging cluster. Computational Statistics, 30(2):419–440, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0541-y.

Golia:2015:AIU

[909] Silvia Golia. Assessing the impact of uniform and nonuniform differential item functioning items on Rasch measure: the polytomous case. Computational Statistics, 30(2):441–461, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0542-x.

Mitsuhiro:2015:RMC

[910] Masaki Mitsuhiro and Hiroshi Yadohisa. Reduced k-means clustering with MCA in a low-dimensional space. Computational Statistics, 30(2):463–475, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0544-8.

Aloui:2015:GPM

[911] Abdelouhab Aloui, Arezki Zioui, Megdouda Ourbih-Tari, and Amine Alioui. A general purpose module using refined descriptive sampling for installation in simulation systems. Computational Statistics, 30(2):477–490, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0545-7.

Nugroho:2015:ERS

[912] Didit B. Nugroho and Takayuki Morimoto. Estimation of realized stochastic volatility models using Hamiltonian Monte Carlo–based methods. Computational Statistics, 30(2):491–516, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-014-
                                        0546-6.

Langrock:2015:SSV

[913] Roland Langrock, Théo Michelot, Alexander Sohn, and Thomas Kneib. Semiparametric stochastic volatility modelling using penalized splines. Computational Statistics, 30(2):517– 537, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

                                        springer.com/article/10.1007/s00180-
                                        014-0547-5.

Ivanescu:2015:PFF

[914] Andrada E. Ivanescu, Ana-Maria Staicu, Fabian Scheipl, and Sonja Greven. Penalized function-on-function regression. Computational Statistics, 30(2):539–568, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0548-4.

Lu:2015:WQR

[915] Xiaoming Lu and Zhaozhi Fan. Weighted quantile regression for longitudinal data. Computational Statistics, 30(2):569–592, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0550-x.

Abd-Elfattah:2015:SVC

[916] Ehab F. Abd-Elfattah. Saddlepoint pvalues and confidence intervals for the class of linear rank tests for censored data under generalized randomized block design. Computational Statistics, 30(2):593–604, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0551-9.

Amini:2015:PPI

[917] Morteza Amini and Narayanaswamy Balakrishnan. Pooled parametric inference for minimal repair systems. Computational Statistics, 30(2):605–623, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0552-8.

Wong:2015:IAM

[918] Ka Yiu Wong and Sung Nok Chiu. An iterative approach to minimize the mean squared error in ridge regression. Computational Statistics, 30(2):625–639, June 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0557-y.

Okhrin:2015:ESI

[919] Ostap Okhrin and Stefan Trück. Editorial to the special issue on applicable semiparametrics of computational statistics. Computational Statistics, 30 (3):641–646, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0616-4;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-015-0616-4.pdf.

Aneiros:2015:PLM

[920] Germán Aneiros and Philippe Vieu. Partial linear modelling with multifunctional covariates. Computational Statistics, 30(3):647–671, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0568-8.

Goia:2015:PSF

[921] Aldo Goia and Philippe Vieu. A partitioned single functional index model. Computational Statistics, 30 (3):673–692, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0530-1.

Sperlich:2015:MHP

[922] Stefan Sperlich and Raoul Theler. Modeling heterogeneity: a praise for varying-coefficient models in causal analysis. Computational Statistics, 30 (3):693–718, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0581-y.

Liu:2015:FCS

[923] Xialu Liu, Zongwu Cai, and Rong Chen. Functional coefficient seasonal time series models with an application of Hawaii tourism data. Computational Statistics, 30(3):719–744, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0574-x.

Cao:2015:MST

[924] Xiaofeng Cao, Ostap Okhrin, Martin Odening, and Matthias Ritter. Modelling spatio-temporal variability of temperature. Computational Statistics, 30(3):745–766, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0561-2.

Kolbe:2015:IBL

[925] Jens Kolbe, Rainer Schulz, Martin Wersing, and Axel Werwatz. Identifying Berlin’s land value map using adaptive weights smoothing. Computational Statistics, 30(3):767–790, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0559-9.

Nowotarski:2015:CES

[926] Jakub Nowotarski and Rafał Weron. Computing electricity spot price prediction intervals using quantile regression and forecast averaging. Computational Statistics, 30(3):791– 803, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0523-0;   http:
//link.springer.com/content/pdf/
10.1007/s00180-014-0523-0.pdf.

Maciejowska:2015:FDE

[927] Katarzyna Maciejowska and Rafał Weron. Forecasting of daily electricity prices with factor models: utilizing intra-day and inter-zone relationships. Computational Statistics, 30(3): 805–819, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-014-0531-0;   http:
//link.springer.com/content/pdf/
10.1007/s00180-014-0531-0.pdf.

Chen:2015:RSV

[928] Shiyi Chen, Kiho Jeong, and Wolfgang K. Härdle. Recurrent support vector regression for a non-linear ARMA model with applications to forecasting financial returns. Computational Statistics, 30(3):821–843, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0543-9.

Chen:2015:CFC

[929] Cathy Yi-Hsuan Chen and Wolfgang Karl Härdle. Common factors in credit defaults swap markets. Computational Statistics, 30(3): 845–863, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0578-6.

Proenca:2015:MBI

[930] Isabel Proença and Horácio C. Faustino. Modelling bilateral intraindustry trade indexes with panel data: a semiparametric approach. Computational Statistics, 30(3):865– 884, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0556-z.

Rodriguez-Poo:2015:DTS

[931] Juan M. Rodriguez-Poo and Alexandra Soberón. Differencing techniques in semi-parametric panel data varying coefficient models with fixed effects: a Monte Carlo study. Computational Statistics, 30(3):885–906, September 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-014-
0549-3.

Braun:2015:VER

[932] W. John Braun. Visualization of evidence in regression with the QR decomposition. Computational Statistics, 30 (4):907–927, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0558-x.

Meermeyer:2015:WLR

[933] Martin Meermeyer. Weighted linear regression models with fixed weights and spherical disturbances. Computational Statistics, 30(4):929–955, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0572-z.

Park:2015:ERB

[934] Chun Gun Park and Inyoung Kim. Efficient resolution and basis functions selection in wavelet regression. Computational Statistics, 30(4):957–986, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0575-9.

Gorecki:2015:CTO

[935] Tomasz Górecki and Lukasz Smaga. A comparison of tests for the oneway ANOVA problem for functional data. Computational Statistics, 30(4): 987–1010, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0555-0;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-015-0555-0.pdf.

Hlubinka:2015:IDD

[936] Daniel Hlubinka, Irène Gijbels, Marek Omelka, and Stanislav Nagy. Integrated data depth for smooth functions and its application in supervised classification. Computational Statistics, 30(4):1011–1031, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0566-x.

Balakrishnan:2015:SMC

[937] Narayanaswamy Balakrishnan and Majid Mojirsheibani. A simple method for combining estimates to improve the overall error rates in classification. Computational Statistics, 30(4): 1033–1049, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0571-0.

Kuroda:2015:IVA

[938] Masahiro Kuroda, Zhi Geng, and Michio Sakakihara. Improving the vector ε acceleration for the EM algorithm using a re-starting procedure. Computational Statistics, 30(4):1051–1077, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0565-y.

Wang:2015:PCA

[939] Huiwen Wang, Liying Shangguan, Rong Guan, and Lynne Billard. Principal component analysis for compositional data vectors. Computational Statistics, 30(4):1079–1096, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-015-
                                        0570-1.

Zhou:2015:RFK

[940] Mai Zhou and Yifan Yang. A recursive formula for the Kaplan–Meier estimator with mean constraints and its application to empirical likelihood. Computational Statistics, 30(4): 1097–1109, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0567-9.

Iacus:2015:ILC

[941] Stefano M. Iacus and Lorenzo Mercuri. Implementation of Lévy CARMA model in Yuima package. Computational Statistics, 30(4):1111–1141, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0569-7.

Castro-Conde:2015:PFC

[942] Irene Castro-Conde and Jacobo de Uña-Álvarez. Power, FDR and conservativeness of BB–SGoF method. Computational Statistics, 30(4):1143– 1161, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0553-2.

MirMostafaee:2015:PCP

[943] S. M. T. K. MirMostafaee, Jafar Ahmadi, and Narjes Sadeghian. Pitman closeness of predictors of future order statistics for two parameter exponential distribution. Computational Statistics, 30(4):1163–1183, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0554-1.

Delgarm:2015:NGL

[944] Leila Delgarm and Mohammad Reza Zadkarami. A new generalization of lifetime distributions. Computational Statistics, 30(4):1185–1198, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0563-0.

Hu:2015:MLE

[945] Ya-Hsuan Hu and Takeshi Emura. Maximum likelihood estimation for a special exponential family under random double-truncation. Computational Statistics, 30(4):1199–1229, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0564-z.

Fontana:2015:ODG

[946] Roberto Fontana. Optimal design generation: an approach based on discovery probability. Computational Statistics, 30(4):1231–1244, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0562-1.

You:2015:SSG

[947] H. W. You, Michael B. C. Khoo, P. Castagliola, and Yanjing Ou. Side sensitive group runs X̄ chart with estimated process parameters. Computational Statistics, 30(4):1245–1278, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0573-y.

Chen:2015:EDA

[948] Shiyi Chen and Wolfgang K. Härdle. Erratum to: Dynamic activity analysis model-based win–win development forecasting under environment regulations in China. Computational Statistics, 30(4):1279, December 2015. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0592-8;   http://link.springer.
com/content/pdf/10.1007/s00180-
015-0592-8.pdf. See [875].

Chen:2016:LNS

[949] Cathy W. S. Chen, Sangyeol Lee, and Shu-Yu Chen. Local non-stationarity test in mean for Markov switching GARCH models: an approximate Bayesian approach. Computational Statistics, 31(1):1–24, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0624-4.

Gerlach:2016:BEI

[950] Richard Gerlach, Shelton Peiris, and Edward M. H. Lin. Bayesian estimation and inference for log-ACD models. Computational Statistics, 31(1):25–48, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0576-8.

Kobayashi:2016:SEP

[951] Genya Kobayashi. Skew exponential power stochastic volatility model for analysis of skewness, non-normal tails, quantiles and expectiles. Computational Statistics, 31(1):49–88, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0596-4.

Mukhopadhyay:2016:BAL

[952] Chiranjit Mukhopadhyay and Soumya Roy. Bayesian accelerated life testing under competing log-location-scale family of causes of failure. Computational Statistics, 31(1):89–119, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0602-x.

Dagne:2016:BSG

[953] Getachew A. Dagne. Bayesian segmental growth mixture tobit models with skew distributions. Computational Statistics, 31(1):121–137, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0620-8.

Dube:2016:PFF

[954] Madhulika Dube, Renu Garg, and Hare Krishna. On progressively first failure censored Lindley distribution. Computational Statistics, 31(1):139–163, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0622-6.

Cepeda-Cuervo:2016:BBR

[955] Edilberto Cepeda-Cuervo, Daniel Jaimes, Margarita Marín, and Javier Rojas. Bayesian beta regression with

Bayesianbetareg R-package. Com-

putational Statistics, 31(1):165–187, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0591-9.

Kiesse:2016:BEB

[956] Tristan Senga Kiessé, Nabil Zougab, and Célestin C. Kokonendji. Bayesian estimation of bandwidth in semiparametric kernel estimation of unknown probability mass and regression functions of count data. Computational Statistics, 31(1):189–206, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0627-1.

Wang:2016:BAB

[957] Min Wang, Xiaoqian Sun, and Chanseok Park. Bayesian analysis of Birnbaum–Saunders distribution via the generalized ratio-of-uniforms method. Computational Statistics, 31 (1):207–225, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0629-z.

Ali:2016:PIB

[958] S. F. Niazi Ali. Prediction intervals based on Gompertz doubly censored data. Computational Statistics, 31 (1):227–246, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0613-7.

Zhao:2016:NEG

[959] Yan-Yong Zhao, Jin-Guan Lin, and Xing-Fang Huang. Nonparametric estimation in generalized varyingcoefficient models based on iterative weighted quasi-likelihood method. Computational Statistics, 31(1):247– 268, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s00180-
015-0579-5.

Lukas:2016:PUR

[960] Mark A. Lukas, Frank R. de Hoog, and Robert S. Anderssen. Practical use of robust GCV and modified GCV for spline smoothing. Computational Statistics, 31(1):269–289, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0577-7.

Wornowizki:2016:TSH

[961] Max Wornowizki and Roland Fried. Two-sample homogeneity tests based on divergence measures. Computational Statistics, 31(1):291–313, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0633-3.

Weihs:2016:ECB

[962] Luca Weihs, Mathias Drton, and Dennis Leung. Efficient computation of the Bergsma–Dassios sign covariance. Computational Statistics, 31 (1):315–328, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0639-x.

Sun:2016:CQR

[963] Jing Sun. Composite quantile regression for single-index models with asymmetric errors. Computational Statistics, 31(1):329–351, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-016-0645-
                                        7.

Amerise:2016:CDM

[964] Ilaria L. Amerise and Agostino Tarsitano. Combining dissimilarity matrices by using rank correlations. Computational Statistics, 31(1):353–367, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0590-x.

Woo:2016:DCV

[965] Sook young Woo, Seonwoo Kim, and Jinheum Kim. Determining cutoff values of prognostic factors in survival data with competing risks. Computational Statistics, 31(1):369–386, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0582-x.

Minoux:2016:CGC

[966] Michel Minoux and Riadh Zorgati. Convexity of Gaussian chance constraints and of related probability maximization problems. Computational Statistics, 31(1):387–408, March 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0580-z.

Latif:2016:GAD

[967] A. H. M. Mahbub Latif and Edgar Brunner. A genetic algorithm for designing microarray experiments. Computational Statistics, 31(2):409–424, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0618-2.

Zhang:2016:CLM

[968] Bo Zhang, Wei Liu, Hui Zhang, Qihui Chen, and Zhiwei Zhang. Composite likelihood and maximum likelihood methods for joint latent class modeling of disease prevalence and highdimensional semicontinuous biomarker data. Computational Statistics, 31 (2):425–449, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0597-3.

Kim:2016:HDT

[969] Soeun Kim, Jae Youn Ahn, and Woojoo Lee. On high-dimensional two sample mean testing statistics: a comparative study with a data adaptive choice of coefficient vector. Computational Statistics, 31(2):451–464, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0605-7.

Munteanu:2016:CSM

[970] Alexander Munteanu and Max Wornowizki. Correcting statistical models via empirical distribution functions. Computational Statistics, 31(2):465–495, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0607-5;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        015-0607-5.pdf.

Xu:2016:CIE

[971] Wangli Xu, Wei Yu, and Zaixing Li. Confidence interval estimation by a joint pivot method. Computational Statistics, 31(2):497–511, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0585-7.

DeBin:2016:BCR

[972] Riccardo De Bin. Boosting in Cox regression: a comparison between the likelihood-based and the modelbased approaches with focus on the Rpackages CoxBoost and mboost. Computational Statistics, 31(2):513–531, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0642-2.

Mi:2016:SRP

[973] Xuefei Mi, H. Friedrich Utz, and Albrecht E. Melchinger. Selectiongain: an R package for optimizing multistage selection. Computational Statistics, 31(2):533–543, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-015-0583-
9.

Luukko:2016:IPL

[974] P. J. J. Luukko, J. Helske, and E. Räsänen. Introducing libeemd: a program package for performing the ensemble empirical mode decomposition. Computational Statistics, 31 (2):545–557, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0603-9.

Dinov:2016:PDW

[975] Ivo D. Dinov, Kyle Siegrist, Dennis K. Pearl, Alexandr Kalinin, and Nicolas Christou. Probability Distributome: a web computational infrastructure for exploring the properties, interrelations, and applications of probability distributions. Computational Statistics, 31(2):559–577, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0594-6.

Bernardi:2016:CVR

[976] Mauro Bernardi and Leopoldo Catania. Comparison of value-at-risk models using the MCS approach. Computational Statistics, 31(2):579–608, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0646-6.

Yang:2016:SEM

[977] Yandan Yang, Hon Keung Tony Ng, and Narayanaswamy Balakrishnan. A stochastic expectation-maximization algorithm for the analysis of system lifetime data with known signature. Computational Statistics, 31 (2):609–641, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0586-6.

Zhang:2016:MAB

[978] Tao Zhang and Joseph E. Cavanaugh. A multistage algorithm for bestsubset model selection based on the Kullback–Leibler discrepancy. Computational Statistics, 31(2):643–669, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0584-8.

Shi:2016:SMC

[979] Xiaoping Shi, Xiang-Sheng Wang, Dongwei Wei, and Yuehua Wu. A sequential multiple change-point detection procedure via VIF regression. Computational Statistics, 31 (2):671–691, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0587-5.

Yuan:2016:SVS

[980] Jinyi Yuan, Peixin Zhao, and Weiguo Zhang. Semiparametric variable selection for partially varying coefficient models with endogenous variables. Computational Statistics, 31 (2):693–707, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0601-y.

Roy:2016:IED

[981] Vivekananda Roy. Improving efficiency of data augmentation algorithms using Peskun’s theorem. Computational Statistics, 31(2):709–728, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0632-4.

Baccini:2016:RVG

[982] Alberto Baccini, Lucio Barabesi, and Luisa Stracqualursi. Random variate generation and connected computational issues for the Poisson–Tweedie distribution. Computational Statistics, 31(2):729–748, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0623-5.

Kalema:2016:SOG

[983] George Kalema, Geert Molenberghs, and Wondwosen Kassahun. Secondorder generalized estimating equations for correlated count data. Computational Statistics, 31(2):749–770, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0599-1.

Azzalini:2016:DBC

[984] Adelchi Azzalini and Giovanna Menardi. Density-based clustering with noncontinuous data. Computational Statistics, 31(2):771–798, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-016-0644-
8.

Yu:2016:LTM

[985] Philip L. H. Yu, Paul H. Lee, S. F. Cheung, Esther Y. Y. Lau, Doris S. Y. Mok, and Harry C. Hui. Logit tree models for discrete choice data with application to advice-seeking preferences among Chinese Christians. Computational Statistics, 31(2):799–827, June 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0588-4.

Ollerer:2016:SER

[986] Viktoria Öllerer, Andreas Alfons, and Christophe Croux. The shooting S-estimator for robust regression. Computational Statistics, 31(3): 829–844, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0593-7.

Tsou:2016:RLI

[987] Tsung-Shan Tsou. Robust likelihood inference for multivariate correlated count data. Computational Statistics, 31(3):845–857, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0589-3.

Lim:2016:DRG

[988] Hock Ann Lim and Habshah Midi. Diagnostic robust generalized potential based on index set equality (DRGP (ISE)) for the identification of high leverage points in linear model. Computational Statistics, 31 (3):859–877, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0662-6.

Acar:2016:IMR

[989] Tugba Söküt Açar and M. Revan Özkale. Influence measures in ridge regression when the error terms follow an Ar(1) process. Computational Statistics, 31(3):879–898, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0615-5.

Prendergast:2016:IES

[990] Luke A. Prendergast and Alan F. Healey. Improving estimated sufficient summary plots in dimension reduction using minimization criteria based on initial estimates. Computational Statistics, 31(3):899–922, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0614-6.

Adragni:2016:GWS

[991] Kofi P. Adragni, Elias Al-Najjar, Sean Martin, Sai K. Popuri, and Andrew M. Raim. Group-wise sufficient dimension reduction with principal fitted components. Computational Statistics, 31(3): 923–941, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0611-9.

Teisseyre:2016:RSM

[992] Paweł Teisseyre, Robert A. Kłopotek, and Jan Mielniczuk. Random subspace method for high-dimensional regression with the R package regRSM. Computational Statistics, 31(3):943– 972, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0658-2;   http:
//link.springer.com/content/pdf/
10.1007/s00180-016-0658-2.pdf.

Liu:2016:LSG

[993] Xuhua Liu, Xingzhong Xu, and Jan Hannig. Least squares generalized inferences in unbalanced twocomponent normal mixed linear model. Computational Statistics, 31(3):973– 988, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0604-8.

Punzo:2016:CBM

[994] Antonio Punzo and Salvatore Ingrassia. Clustering bivariate mixed-type data via the cluster-weighted model. Computational Statistics, 31(3):989– 1013, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0600-z.

Shim:2016:SVQ

[995] Jooyong Shim, Changha Hwang, and Kyungha Seok. Support vector quantile regression with varying coefficients. Computational Statistics, 31(3): 1015–1030, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0647-5.

Tian:2016:BJQ

[996] Yuzhu Tian, Er’qian Li, and Maozai Tian. Bayesian joint quantile regression for mixed effects models with censoring and errors in covariates. Computational Statistics, 31(3):1031–1057, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0659-1.

Ortega:2016:BEH

[997] Francisco J. Ortega and Jose M. Gavilan. Bayesian estimation of the half-normal regression model with deterministic frontier. Computational Statistics, 31(3):1059–1078, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-016-
                                        0648-4.

Kaishev:2016:GDV

[998] Vladimir K. Kaishev, Dimitrina S. Dimitrova, Steven Haberman, and Richard J. Verrall. Geometrically designed, variable knot regression splines. Computational Statistics, 31(3):1079– 1105, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0621-7.

Zhou:2016:PSE

[999] Jianjun Zhou, Zhao Chen, and Qingyan Peng. Polynomial spline estimation for partial functional linear regression models. Computational Statistics, 31(3):1107–1129, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-015-0636-
                                        0.

Lv:2016:REV

[1000] Jing Lv, Hu Yang, and Chaohui Guo. Robust estimation for varying index coefficient models. Computational Statistics, 31(3):1131–1167, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0595-5.

Laurini:2016:GME

[1001] Márcio Poletti Laurini and Luiz Koodi Hotta. Generalized moment estimation of stochastic differential equations. Computational Statistics, 31(3): 1169–1202, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0598-2.

Lv:2016:SCG

[1002] Jing Lv, Hu Yang, and Chaohui Guo. Smoothing combined generalized estimating equations in quantile partially linear additive models with longitudinal data. Computational Statistics, 31(3):1203–1234, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-

0612-8. See erratum [1003].

Lv:2016:ESC

[1003] Jing Lv, Hu Yang, and Chaohui Guo.

Erratum to: Smoothing combined generalized estimating equations in quantile partially linear additive models with longitudinal data. Computational Statistics, 31(3):1235, September 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0626-2;   http://link.springer.
com/content/pdf/10.1007/s00180-
015-0626-2.pdf. See [1002].

Zhang:2016:PPB

[1004] Chun-Xia Zhang, Jiang-She Zhang, and Sang-Woon Kim. PBoostGA: pseudo-boosting genetic algorithm for variable ranking and selection. Computational Statistics, 31(4):1237– 1262, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0652-8.

Chin:2016:MVS

[1005] Yen-Shiu Chin and Ting-Li Chen. Minimizing variable selection criteria by Markov chain Monte Carlo. Computational Statistics, 31(4):1263–1286, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0649-3.

Jia:2016:ADP

[1006] Bochao Jia, Yuan chin Ivan Chang, and Zhanfeng Wang. Assessing the diagnostic power of variables measured with a detection limit. Computational Statistics, 31(4):1287–1303, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0628-0.

Pebay:2016:NSS

[1007] Philippe Pébay, Timothy B. Terriberry, Hemanth Kolla, and Janine Bennett. Numerically stable, scalable formulas for parallel and online computation of higher-order multivariate central moments with arbitrary weights. Computational Statistics, 31(4):1305– 1325, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0637-z.

Pistonesi:2016:SSA

[1008] Silvina Pistonesi, Jorge Martinez, and Silvia M. Ojeda. A sensibility study of the autobinomial model estimation methods based on a feature similarity index. Computational Statistics, 31(4): 1327–1357, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0634-2.

Dickson:2016:OSS

[1009] Maria Michela Dickson and Yves Tillé. Ordered spatial sampling by means of the traveling salesman problem. Computational Statistics, 31(4): 1359–1372, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-015-0635-1.

Heymann:2016:TFT

[1010] Yuri Heymann. A test of financial time-series data to discriminate among lognormal, Gaussian and square-root random walks. Computational Statistics, 31(4):1373–1383, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0630-6.

Demirtas:2016:NLP

[1011] Hakan Demirtas, Robab Ahmadian, Sema Atis, Fatma Ezgi Can, and Ilker Ercan. A nonnormal look at polychoric correlations: modeling the change in correlations before and after discretization. Computational Statistics, 31(4): 1385–1401, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0653-7.

Adachi:2016:SPC

[1012] Kohei Adachi and Nickolay T. Trendafilov. Sparse principal component analysis subject to prespecified cardinality of loadings. Computational Statistics, 31(4):1403–1427, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0608-4.

Loisel:2016:PPC

[1013] Sébastien Loisel and Yoshio Takane. Partitions of Pearson’s chi-square statistic for frequency tables: a comprehensive account. Computational Statistics, 31(4):1429–1452, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0619-1.

Naul:2016:RIA

[1014] Brett Naul, Bala Rajaratnam, and Dario Vincenzi. The role of the isotonizing algorithm in Stein’s covariance matrix estimator. Computational Statistics, 31(4):1453–1476, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0672-4.

Cuenin:2016:SFM

[1015] Johann Cuenin, Bent Jørgensen, and Célestin C. Kokonendji. Simulations of full multivariate Tweedie with flexible dependence structure. Computational Statistics, 31(4):1477–1492, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0617-3.

Nyman:2016:CSI

[1016] Henrik Nyman, Johan Pensar, Timo Koski, and Jukka Corander. Contextspecific independence in graphical log-linear models. Computational Statistics, 31(4):1493–1512, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0606-6.

Bordes:2016:SEA

[1017] Laurent Bordes and Didier Chauveau. Stochastic EM algorithms for parametric and semiparametric mixture models for right-censored lifetime data. Computational Statistics, 31(4): 1513–1538, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0661-7.

Li:2016:REN

[1018] Meng Li, Sijia Xiang, and Weixin Yao. Robust estimation of the number of components for mixtures of linear regression models. Computational Statistics, 31(4):1539–1555, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0610-x.

Wu:2016:MRL

[1019] Jibo Wu. Modified restricted Liu estimator in logistic regression model.

Computational Statistics, 31(4):1557– 1567, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0609-3.

Leao:2016:CTL

[1020] Jeremias Leão, Francisco Cysneiros, Helton Saulo, and N. Balakrishnan. Constrained test in linear models with multivariate power exponential distribution. Computational Statistics, 31(4):1569–1592, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0650-x.

Wilcox:2016:AHG

[1021] Rand R. Wilcox. ANCOVA: a heteroscedastic global test when there is curvature and two covariates. Computational Statistics, 31(4):1593– 1606, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-015-0640-4.

Parra-Frutos:2016:PTW

[1022] I. Parra-Frutos. Preliminary tests when comparing means. Computational Statistics, 31(4):1607–1631, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0656-4.

Shan:2016:ESS

[1023] Guogen Shan. Exact sample size determination for the ratio of two incidence rates under the Poisson distribution. Computational Statistics, 31(4): 1633–1644, December 2016. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0654-6.

Grzegorczyk:2017:CEV

[1024] Marco Grzegorczyk and Mahdi Shafiee Kamalabad. Comparative evaluation of various frequentist and Bayesian nonhomogeneous Poisson counting models. Computational Statistics, 32(1):1–33, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0686-y;   http://link.springer.
com/content/pdf/10.1007/s00180-
016-0686-y.pdf.

Kang:2017:NPR

[1025] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Noninformative priors for the ratio of the shape parameters of two Weibull distributions. Computational Statistics, 32(1):35–50, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-015-
0631-5.

Elster:2017:BIU

[1026] Clemens Elster and Gerd Wübbeler.

Bayesian inference using a noninformative prior for linear Gaussian random coefficient regression with inhomogeneous within-class variances. Computational Statistics, 32(1):51–69, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-015-
                                        0641-3.

Kang:2017:OBT

[1027] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective Bayesian testing on the common mean of several normal distributions under divergencebased priors. Computational Statistics, 32(1):71–91, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0699-6.

Depraetere:2017:CVA

[1028] Nicolas Depraetere and Martina Vandebroek. A comparison of variational approximations for fast inference in mixed logit models. Computational Statistics, 32(1):93–125, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-015-
                                        0638-y.

Yang:2017:BVS

[1029] Aijun Yang, Xuejun Jiang, Lianjie Shu, and Jinguan Lin. Bayesian variable selection with sparse and correlation priors for high-dimensional data analysis. Computational Statistics, 32 (1):127–143, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0665-3.

Lumbreras:2017:NPC

[1030] Alberto Lumbreras, Julien Velcin, Marie Guégan, and Bertrand Jouve. Non-parametric clustering over user features and latent behavioral functions with dual-view mixture models. Computational Statistics, 32 (1):145–177, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0668-0.

Huang:2017:BIP

[1031] Yangxin Huang and Tao Lu. Bayesian inference on partially linear mixedeffects joint models for longitudinal data with multiple features. Computational Statistics, 32(1):179–196, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0671-5.

Pasanen:2017:SSM

[1032] Leena Pasanen and Lasse Holmström.

Scale space multiresolution correlation analysis for time series data. Computational Statistics, 32(1):197–218, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0670-6.

Pan:2017:BAM

[1033] Jiazhu Pan, Qiang Xia, and Jinshan Liu. Bayesian analysis of multiple thresholds autoregressive model. Computational Statistics, 32(1):219–237, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0673-3.

Martino:2017:IMT

[1034] L. Martino and F. Louzada. Issues in the multiple try Metropolis mixing. Computational Statistics, 32 (1):239–252, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0643-9.

Zhang:2017:PSH

[1035] Cheng Zhang, Babak Shahbaba, and Hongkai Zhao. Precomputing strategy for Hamiltonian Monte Carlo method based on regularity in parameter space. Computational Statistics, 32(1):253–279, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0683-1.

Zhao:2017:ILC

[1036] Zhenyu Zhao and Thomas A. Severini. Integrated likelihood computation methods. Computational Statistics, 32 (1):281–313, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0677-z.

Yang:2017:ATG

[1037] Jinyoung Yang and Jeffrey S. Rosenthal. Automatically tuned generalpurpose MCMC via new adaptive diagnostics. Computational Statistics, 32(1):315–348, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0682-2.

Ker:2017:BMA

[1038] Alan P. Ker and Yong Liu. Bayesian model averaging of possibly similar nonparametric densities. Computational Statistics, 32(1):349–365, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0700-4.

Basu:2017:PEI

[1039] Suparna Basu, Sanjay Kumar Singh, and Umesh Singh. Parameter estimation of inverse Lindley distribution for Type-I censored data. Computational Statistics, 32(1):367–385, March 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0704-0.

Ghosh:2017:UEA

[1040] Atanu Kumar Ghosh and Arnab Chakraborty. Use of EM algorithm for data reduction under sparsity assumption. Computational Statistics, 32(2):387–407, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0657-3.

Tomas:2017:ULP

[1041] Elson Tomás, Susana Vinga, and Alexandra M. Carvalho. Unsupervised learning of pharmacokinetic responses. Computational Statistics, 32(2):409–428, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0707-x.

Pal:2017:LID

[1042] Suvra Pal and N. Balakrishnan. Likelihood inference for the destructive exponentially weighted Poisson cure rate model with Weibull lifetime and an application to melanoma data. Computational Statistics, 32(2):429–449, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0660-8.

Jamalizadeh:2017:GCS

[1043] Ahad Jamalizadeh and Tsung-I Lin. A general class of scale-shape mixtures of skew-normal distributions: properties and estimation. Computational Statistics, 32(2):451–474, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0691-1.

Abe:2017:NNM

[1044] Hiroyasu Abe and Hiroshi Yadohisa. A non-negative matrix factorization model based on the zeroinflated Tweedie distribution. Computational Statistics, 32(2):475–499, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0689-8.

Zreik:2017:DRS

[1045] Rawya Zreik, Pierre Latouche, and Charles Bouveyron. The dynamic random subgraph model for the clustering of evolving networks. Computational Statistics, 32(2):501–533, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0655-5.

Kaplan:2017:IGM

[1046] Andee Kaplan, Heike Hofmann, and Daniel Nordman. An interactive graphical method for community detection in network data. Computational Statistics, 32(2):535–557, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0663-5.

Nanty:2017:UQF

[1047] Simon Nanty, Céline Helbert, Amandine Marrel, Nadia Pérot, and Clémentine Prieur. Uncertainty quantification for functional dependent random variables. Computational Statistics, 32(2):559–583, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0676-0.

Wang:2017:DRF

[1048] Guochang Wang. Dimension reduction in functional regression with categorical predictor. Computational Statistics, 32(2):585–609, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0675-1.

Yan:2017:NEB

[1049] Yanyang Yan, Feipeng Zhang, and Xiaoying Zhou. A note on estimating the bent line quantile regression model. Computational Statistics, 32 (2):611–630, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0711-9.

Koller:2017:NSR

[1050] Manuel Koller and Werner A. Stahel. Nonsingular subsampling for regression s 40estimators with categorical predictors. Computational Statistics, 32(2):631–646, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0679-x.

Wu:2017:CTA

[1051] Shu-Fei Wu and Jin-Yang Lu. Computational testing algorithmic procedure of assessment for lifetime performance index of Pareto products under progressive type I interval censoring. Computational Statistics, 32 (2):647–666, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0717-3.

Pfeiffer:2017:IMS

[1052] Ruth M. Pfeiffer, Andrew Redd, and Raymond J. Carroll. On the impact of model selection on predictor identification and parameter inference. Computational Statistics, 32 (2):667–690, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0690-2;   http:
//link.springer.com/content/pdf/
10.1007/s00180-016-0690-2.pdf.

Datta:2017:GST

[1053] Sagnik Datta, Ghislaine Gayraud, Eric Leclerc, and Frederic Y. Bois. Graph sampler: a simple tool for fully Bayesian analyses of DAGmodels. Computational Statistics, 32 (2):691–716, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0719-1.

Grzegorczyk:2017:TBF

[1054] Marco Grzegorczyk, Andrej Aderhold, and Dirk Husmeier. Targeting Bayes factors with direct-path nonequilibrium thermodynamic integration. Computational Statistics, 32 (2):717–761, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0721-7;   http:
//link.springer.com/content/pdf/
10.1007/s00180-017-0721-7.pdf.

Baharev:2017:CND

[1055] Ali Baharev, Hermann Schichl, and Endre Rév. Computing the noncentral- F distribution and the power of the F - test with guaranteed accuracy. Computational Statistics, 32(2):763–779, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0701-3;   http://link.springer.

                                        com/content/pdf/10.1007/s00180-
                                        016-0701-3.pdf.

Elsawah:2017:OMB

[1056] A. M. Elsawah and Hong Qin. Optimum mechanism for breaking the confounding effects of mixed-level designs. Computational Statistics, 32 (2):781–802, June 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0651-9.

Davies:2017:SHB

[1057] Vinny Davies, Richard Reeve, William T. Harvey, François F. Maree, and Dirk Husmeier. A sparse hierarchical Bayesian model for detecting relevant antigenic sites in virus evolution. Computational Statistics, 32(3): 803–843, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0730-6;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-017-0730-6.pdf.

Lu:2017:BIL

[1058] Tao Lu. Bayesian inference on longitudinal-survival data with multiple features. Computational Statistics, 32(3):845–866, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-016-0681-
                                        3.

Lee:2017:FSE

[1059] DongHyuk Lee, Raymond J. Carroll, and Samiran Sinha. Frequentist standard errors of Bayes estimators. Computational Statistics, 32(3): 867–888, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0710-x.

Maruo:2017:SSB

[1060] Kazushi Maruo, Takaharu Yamabe, and Yusuke Yamaguchi. Statistical simulation based on right skewed distributions. Computational Statistics, 32 (3):889–907, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0664-4.

Zhang:2017:EAR

[1061] Tonglin Zhang and Baijian Yang. An exact approach to ridge regression for big data. Computational Statistics, 32 (3):909–928, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0731-5.

Bocek:2017:WLP

[1062] Pavel Bocek and Miroslav Siman.

On weighted and locally polynomial directional quantile regression. Computational Statistics, 32(3):929– 946, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0708-9.

Lv:2017:EPE

[1063] Jing Lv and Chaohui Guo. Efficient parameter estimation via modified Cholesky decomposition for quantile regression with longitudinal data. Computational Statistics, 32(3):947– 975, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0714-6.

Shi:2017:EME

[1064] Jianhong Shi, Qian Yang, Xiongya Li, and Weixing Song. Effects of measurement error on a class of single-index varying coefficient regression models. Computational Statistics, 32(3): 977–1001, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0726-2.

Savchuk:2017:FRO

[1065] Olga Y. Savchuk and Jeffrey D. Hart. Fully robust one-sided cross-validation for regression functions. Computational Statistics, 32(3):1003–1025, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0713-7.

Kolacek:2017:BMS

[1066] Jan Kolácek and Ivana Horová. Bandwidth matrix selectors for kernel regression. Computational Statistics, 32(3): 1027–1046, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0709-3.

Wang:2017:OVE

[1067] WenWu Wang, Lu Lin, and Li Yu.

Optimal variance estimation based on lagged second-order difference in nonparametric regression. Computational Statistics, 32(3):1047–1063, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0666-2;   http://link.springer.
com/content/pdf/10.1007/s00180-
016-0666-2.pdf.

Jin:2017:SRD

[1068] Hao Jin, Si Zhang, and Jinsuo Zhang. Spurious regression due to neglected of non-stationary volatility. Computational Statistics, 32(3): 1065–1081, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0687-x.

Delicado:2017:CMR

[1069] Pedro Delicado and Philippe Vieu.

Choosing the most relevant level sets for depicting a sample of densities. Computational Statistics, 32(3):1083– 1113, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0746-y.

Jang:2017:ILS

[1070] Dongik Jang, Hee-Seok Oh, and Philippe Naveau. Identifying local smoothness for spatially inhomogeneous functions. Computational Statistics, 32(3):1115–1138, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0694-y.

Holmstrom:2017:ELS

[1071] Lasse Holmström, Kyösti Karttunen, and Jussi Klemelä. Estimation of level set trees using adaptive partitions. Computational Statistics, 32(3): 1139–1163, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0702-2.

Fraiman:2017:SHT

[1072] Ricardo Fraiman, Leonardo Moreno, and Sebastian Vallejo. Some hypothesis tests based on random projection. Computational Statistics, 32(3): 1165–1189, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0732-4.

Samadi:2017:CCP

[1073] S. Yaser Samadi, L. Billard, M. R. Meshkani, and A. Khodadadi. Canonical correlation for principal components of time series. Computational Statistics, 32(3):1191–1212, September 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0667-1.

Figueiredo:2017:BPT

[1074] Adelaide Figueiredo. Bootstrap and permutation tests in ANOVA for directional data. Computational Statistics, 32(4):1213–1240, December 2017.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-017-0739-
x.

Allison:2017:AAC

[1075] J. S. Allison, L. Santana, N. Smit, and I. J. H. Visagie. An ‘apples to apples’ comparison of various tests for exponentiality. Computational Statistics, 32(4):1241–1283, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0733-3.

Mahdizadeh:2017:RBI

[1076] M. Mahdizadeh and E. Strzalkowska-Kominiak. Resampling based inference for a distribution function using censored ranked set samples. Computational Statistics, 32(4):1285–1308, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0716-4.

Amiri:2017:TPJ

[1077] Saeid Amiri, Reza Modarres, and Silvelyn Zwanzig. Tests of perfect judgment ranking using pseudo-samples. Computational Statistics, 32(4):1309– 1322, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0698-7.

Allison:2017:MCE

[1078] James S. Allison and Charl Pretorius.

A Monte Carlo evaluation of the performance of two new tests for symmetry. Computational Statistics, 32(4): 1323–1338, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0680-4.

Balakrishnan:2017:MTE

[1079] N. Balakrishnan, Helton Saulo, Marcelo Bourguignon, and Xiaojun Zhu. On moment-type estimators for a class of log-symmetric distributions. Computational Statistics, 32(4):1339–1355, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0722-6.

Ding:2017:RBN

[1080] Deng Ding, Xiaofei Li, and Yiqi Liu. A regression-based numerical scheme for backward stochastic differential equations. Computational Statistics, 32(4): 1357–1373, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0763-x.

Wu:2017:EMR

[1081] Yanke Wu and Maozai Tian. An effective method to reduce the computational complexity of composite quantile regression. Computational Statistics, 32(4):1375–1393, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0749-8.

Liu:2017:FIT

[1082] Xiaohui Liu. Fast implementation of the Tukey depth. Computational Statistics, 32(4):1395–1410, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0697-8.

Sun:2017:IRC

[1083] H.-J. Sun, Kaoru Fukuda, and B. D.

McCullough. Inaccurate regression coefficients in Microsoft Excel 2003: an investigation of Volpi’s “zero bug”. Computational Statistics, 32(4):1411– 1421, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0764-9.

Zargar:2017:DST

[1084] M. Zargar, H. Jabbari, and M. Amini.

Dependence structure and test of independence for some well-known bivariate distributions. Computational Statistics, 32(4):1423–1451, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-016-
0696-9.

Marques:2017:TES

[1085] Filipe J. Marques, Carlos A. Coelho, and Paulo C. Rodrigues. Testing the equality of several linear regression models. Computational Statistics, 32(4):1453–1480, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-016-
                                        0703-1.

Heikkila:2017:MMB

[1086] Matias Heikkilä, Yves Dominicy, and Pauliina Ilmonen. Multivariate moment based extreme value index estimators. Computational Statistics, 32(4):1481–1513, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-016-0706-
                                        y.

Ahmadabadi:2017:BNE

[1087] Alireza Ahmadabadi and Burcu Hudaverdi Ucer. Bivariate nonparametric estimation of the Pickands dependence function using Bernstein copula with kernel regression approach. Computational Statistics, 32(4):1515–1532, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0750-2.

Castro:2017:NDF

[1088] Tomás del Barrio Castro, Andrii Bodnar, and Andreu Sansó. Numerical distribution functions for seasonal unit root tests with OLS and GLS detrending. Computational Statistics, 32(4): 1533–1568, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0688-9.

Mahmoudi:2017:NMD

[1089] Mohammad Reza Mahmoudi and Mohsen Maleki. A new method to detect periodically correlated structure. Computational Statistics, 32(4): 1569–1581, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0705-z.

Yan:2017:TPF

[1090] Tianshun Yan and Changlin Mei. A test for a parametric form of the volatility in second-order diffusion models. Computational Statistics, 32(4): 1583–1596, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0685-z.

Li:2017:QLI

[1091] Han Li, Kai Yang, and Dehui Wang.

Quasi-likelihood inference for selfexciting threshold integer-valued autoregressive processes. Computational Statistics, 32(4):1597–1620, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0748-9.

Melicher:2017:FDL

[1092] Valdemar Melicher, Tom Haber, and Wim Vanroose. Fast derivatives of likelihood functionals for ODE based models using adjoint-state method. Computational Statistics, 32(4):1621– 1643, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0765-8.

Tian:2017:QPL

[1093] Yongge Tian and Bo Jiang. Quadratic properties of least-squares solutions of linear matrix equations with statistical applications. Computational Statistics, 32(4):1645–1663, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-016-
                                        0693-z.

Basak:2017:PCE

[1094] Indrani Basak and N. Balakrishnan. Prediction of censored exponential lifetimes in a simple step-stress model under progressive Type II censoring. Computational Statistics, 32(4): 1665–1687, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-016-0684-0.

Ghosh:2017:NCI

[1095] Santu Ghosh, Arpita Chatterjee, and N. Balakrishnan. Nonparametric confidence intervals for ranked set samples. Computational Statistics, 32(4): 1689–1725, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0744-0.

Mohammadi:2017:JSP

[1096] Leila Mohammadi. The joint signature of parallel systems for different permutations of failure times. Computational Statistics, 32(4):1727–1746, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-017-0718-
2.

Bowman:2017:WLS

[1097] Dale Bowman and E. Olusegun George. Weighted least squares estimation for exchangeable binary data. Computational Statistics, 32(4):1747– 1765, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-016-0695-x.

Liu:2017:ESG

[1098] Chuanhai Liu, Ryan Martin, and Nick Syring. Efficient simulation from a gamma distribution with small shape parameter. Computational Statistics, 32(4):1767–1775, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-016-0692-
0.

Espinheira:2017:EBP

[1099] Patrícia L. Espinheira, Silvia L. P. Ferrari, and Francisco Cribari-Neto. Erratum to: Bootstrap prediction intervals in beta regressions. Computational Statistics, 32(4):1777, December 2017. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-017-0754-

y. See [866].

Everitt:2018:EIS

[1100] Richard G. Everitt. Efficient importance sampling in low dimensions using affine arithmetic. Computational Statistics, 33(1):1–29, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0729-z;   http://link.springer.
com/content/pdf/10.1007/s00180-
017-0729-z.pdf.

El-Din:2018:SIU

[1101] M. M. Mohie El-Din, A. R. Shafay, and M. Nagy. Statistical inference under adaptive progressive censoring scheme. Computational Statistics, 33(1):31–74, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0745-z.

Wang:2018:CSS

[1102] Jiangyan Wang, Miao Yang, and Anandamayee Majumdar. Comparative study and sensitivity analysis of skewed spatial processes. Computational Statistics, 33(1):75–98, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0741-3.

Leisen:2018:OBA

[1103] Fabrizio Leisen, Luca Rossini, and Cristiano Villa. Objective Bayesian analysis of the Yule–Simon distribution with applications. Computational Statistics, 33(1):99–126, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0735-1;   http://link.springer.
com/content/pdf/10.1007/s00180-
017-0735-1.pdf.

Li:2018:EME

[1104] Phillip Li. Efficient MCMC estimation of inflated beta regression models. Computational Statistics, 33 (1):127–158, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0747-x.

Cowles:2018:ISB

[1105] Mary Kathryn Cowles, Stephen Bonett, and Michael Seedorff. Independent sampling for Bayesian normal conditional autoregressive models with OpenCL acceleration. Computational Statistics, 33(1):159–177, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0752-0.

Picchini:2018:CSE

[1106] Umberto Picchini and Adeline Samson.

Coupling stochastic EM and approximate Bayesian computation for parameter inference in state-space models. Computational Statistics, 33 (1):179–212, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0770-y;   http:
//link.springer.com/content/pdf/
10.1007/s00180-017-0770-y.pdf.

Kharroubi:2018:PSE

[1107] Samer A. Kharroubi. Posterior simulation via the exponentially tilted signed root log-likelihood ratio. Computational Statistics, 33(1):213–234, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0772-9.

Cano:2018:OBM

[1108] J. A. Cano, C. Carazo, and D. Salmerón. Objective Bayesian model selection approach to the two way analysis of variance. Computational Statistics, 33 (1):235–248, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0727-1.

Krishna:2018:CBI

[1109] H. Krishna and N. Goel. Classical and Bayesian inference in two parameter exponential distribution with randomly censored data. Computational Statistics, 33(1):249–275, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0725-3.

Sidhu:2018:BEG

[1110] Sukhmani Sidhu, Kanchan Jain, and Suresh Kumar Sharma. Bayesian estimation of generalized gamma shared frailty model. Computational Statistics, 33(1):277–297, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0728-0.

Ziane:2018:BSP

[1111] Yasmina Ziane, Nabil Zougab, and Smail Adjabi. Birnbaum-Saunders power-exponential kernel density estimation and Bayes local bandwidth selection for nonnegative heavy tailed data. Computational Statistics, 33 (1):299–318, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0712-8.

Cao:2018:HRM

[1112] Chunzheng Cao, Mengqian Chen, Yahui Wang, and Jian Qing Shi. Heteroscedastic replicated measurement error models under asymmetric heavy-tailed distributions. Computational Statistics, 33(1):319–338, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0720-8.

Reath:2018:IPE

[1113] Joseph Reath, Jianping Dong, and Min Wang. Improved parameter estimation of the log-logistic distribution with applications. Computational Statistics, 33(1):339–356, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0738-y.

Fanjul-Hevia:2018:CSM

[1114] Arís Fanjul-Hevia and Wenceslao González-Manteiga. A comparative study of methods for testing the equality of two or more ROC curves. Computational Statistics, 33(1):357–377, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-017-
                                        0783-6.

Herwartz:2018:PWB

[1115] Helmut Herwartz and Yabibal M. Walle. A powerful wild bootstrap diagnosis of panel unit roots under linear trends and time-varying volatility. Computational Statistics, 33 (1):379–411, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0784-5.

Pestova:2018:ACM

[1116] Barbora Pestová and Michal Pesta. Abrupt change in mean using block bootstrap and avoiding variance estimation. Computational Statistics, 33 (1):413–441, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0785-4.

Lovcha:2018:ISA

[1117] Yuliya Lovcha, Alejandro Perez-Laborda, and Luis Gil-Alana. On the invertibility of seasonally adjusted series. Computational Statistics, 33 (1):443–465, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0715-5.

Roche:2018:LOB

[1118] Angelina Roche. Local optimization of black-box functions with high or infinite-dimensional inputs: application to nuclear safety. Computational Statistics, 33(1):467–485, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0751-1.

Bornkamp:2018:CQN

[1119] Björn Bornkamp. Calculating quantiles of noisy distribution functions using local linear regressions. Computational Statistics, 33(1):487–501, March 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0736-0.

Ozkale:2018:LRD

[1120] M. Revan Özkale, Stanley Lemeshow, and Rodney Sturdivant. Logistic regression diagnostics in ridge regression. Computational Statistics, 33 (2):563–593, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0755-x.

TenEyck:2018:MSC

[1121] Patrick Ten Eyck and Joseph E. Cavanaugh. Model selection criteria based on cross-validatory concordance statistics. Computational Statistics, 33(2):595–621, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0766-7.

Wang:2018:HRE

[1122] Zhanfeng Wang, Zhuojian Chen, and Zimu Chen. H-relative error estimation for multiplicative regression model with random effect. Computational Statistics, 33(2):623–638, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0798-7.

Colubi:2018:NDC

[1123] Ana Colubi, J. Santos Dominguez-Menchero, and Gil Gonzalez-Rodriguez. New designs to consistently estimate the isotonic regression. Computational Statistics, 33(2):639–658, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0792-0.

Zhao:2018:REL

[1124] Peixin Zhao and Xiaoshuang Zhou. Robust empirical likelihood for partially linear models via weighted composite quantile regression. Computational Statistics, 33(2):659–674, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0793-z.

Cizek:2018:REM

[1125] P. Cízek and M. Aquaro. Robust estimation and moment selection in dynamic fixed-effects panel data models. Computational Statistics, 33 (2):675–708, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0782-7;   http:
//link.springer.com/content/pdf/
10.1007/s00180-017-0782-7.pdf.

Ramires:2018:ENE

[1126] Thiago G. Ramires, Niel Hens, Gauss M. Cordeiro, and Edwin M. M. Ortega. Estimating nonlinear effects in the presence of cure fraction using a semi-parametric regression model. Computational Statistics, 33(2):709–730, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0781-8.

Li:2018:LFT

[1127] Chin-Shang Li and Minggen Lu. A lack-of-fit test for generalized linear models via single-index techniques. Computational Statistics, 33 (2):731–756, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0802-2.

Bergtold:2018:ERS

[1128] Jason S. Bergtold, Krishna P. Pokharel, Allen M. Featherstone, and Lijia Mo. On the examination of the reliability of statistical software for estimating regression models with discrete dependent variables. Computational Statistics, 33(2):757–786, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0776-5.

Liao:2018:EAB

[1129] J. G. Liao, Joseph E. Cavanaugh, and Timothy L. McMurry. Extending AIC to best subset regression. Computational Statistics, 33(2):787–806, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0797-8.

Bertoli:2018:ZMP

[1130] Wesley Bertoli, Katiane S. Conceição, Marinho G. Andrade, and Francisco Louzada. On the zero-modified Poisson–Shanker regression model and its application to fetal deaths notification data. Computational Statistics, 33(2):807–836, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0788-1.

Lo:2018:PLI

[1131] T.-F. Lo, P.-H. Ke, and W.-J. Tsay. Pairwise likelihood inference for the random effects probit model. Computational Statistics, 33(2):837–861, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0757-8.

Zhou:2018:ODB

[1132] Yuejin Zhou, Yebin Cheng, Wenlin Dai, and Tiejun Tong. Optimal differencebased estimation for partially linear models. Computational Statistics, 33 (2):863–885, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0786-3.

Qi:2018:GGL

[1133] Zong-Feng Qi, Xue-Ru Zhang, and Yong-Dao Zhou. Generalized good lattice point sets. Computational Statistics, 33(2):887–901, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-017-0737-
z.

Zhou:2018:RPD

[1134] Xiao-Dong Zhou, Yun-Juan Wang, and Rong-Xian Yue. Robust population designs for longitudinal linear regression model with a random intercept. Computational Statistics, 33(2):903–931, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0767-6.

Elsawah:2018:COS

[1135] A. M. Elsawah. Choice of optimal second stage designs in two-stage experiments. Computational Statistics, 33(2):933–965, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0778-3.

deAndrade:2018:LCN

[1136] Bernardo B. de Andrade and Geraldo S. Souza. Likelihood computation in the normal-gamma stochastic frontier model. Computational Statistics, 33(2):967–982, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0768-5.

Fu:2018:WCS

[1137] Liya Fu, Yangyang Hao, and You-Gan Wang. Working correlation structure selection in generalized estimating equations. Computational Statistics, 33(2):983–996, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0800-4.

Fasola:2018:HIA

[1138] Salvatore Fasola, Vito M. R. Muggeo, and Helmut Küchenhoff. A heuristic, iterative algorithm for change-point detection in abrupt change models. Computational Statistics, 33(2):997–1015, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0740-4.

Ruggieri:2018:PRS

[1139] Eric Ruggieri. A pruned recursive solution to the multiple change point problem. Computational Statistics, 33 (2):1017–1045, June 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0756-9.

Kestler:2018:PR

[1140] Hans A. Kestler, Bernd Bischl, and Matthias Schmid. Proceedings of Reisensburg 2014–2015. Computational Statistics, 33(3):1125–1126, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0823-x;   http://link.springer.
com/content/pdf/10.1007/s00180-
018-0823-x.pdf.

Herbrandt:2018:MBO

[1141] Swetlana Herbrandt, Uwe Ligges, Manuel Pinho Ferreira, Michael Kansteiner, Dirk Biermann, Wolfgang Tillmann, and Claus Weihs. Model based optimization of a statistical simulation model for single diamond grinding. Computational Statistics, 33(3): 1127–1143, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0669-z.

Kuhn:2018:FSS

[1142] Matthias Kuhn, Thoralf Stange, Sylvia Herold, Christian Thiede, and Ingo Roeder. Finding small somatic structural variants in exome sequencing data: a machine learning approach. Computational Statistics, 33(3):1145– 1158, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0674-2.

Surmann:2018:PMU

[1143] Dirk Surmann, Uwe Ligges, and Claus Weihs. Predicting measurements at unobserved locations in an electrical transmission system. Computational Statistics, 33(3):1159–1172, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0734-2.

Villmann:2018:LVQ

[1144] T. Villmann, M. Kaden, W. Hermann, and M. Biehl. Learning vector quantization classifiers for ROC-optimization.

Computational Statistics, 33(3):1173– 1194, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-016-0678-y.

Seibold:2018:CIN

[1145] Heidi Seibold, Christoph Bernau, Anne-Laure Boulesteix, and Riccardo De Bin. On the choice and influence of the number of boosting steps for high-dimensional linear Cox-models. Computational Statistics, 33(3):1195– 1215, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0773-8.

Alizadeh:2018:OLL

[1146] Morad Alizadeh, Fazlollah Lak, Mahdi Rasekhi, Thiago G. Ramires, Haitham M. Yousof, and Emrah Altun. The odd log-logistic Topp–Leone G family of distributions: heteroscedastic regression models and applications. Computational Statistics, 33(3): 1217–1244, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0780-9.

Liu:2018:AMH

[1147] Wanrong Liu, Jianglin Fang, and Xuewen Lu. Additive-multiplicative hazards model with current status data. Computational Statistics, 33(3): 1245–1266, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0806-y.

Kumar:2018:DCR

[1148] M. Kumar and P. N. Bajeel. Design of component reliability test plan for a series system having time dependent testing cost with the presence of covariates. Computational Statistics, 33(3): 1267–1292, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0758-7.

Shih:2018:LBI

[1149] Jia-Han Shih and Takeshi Emura.

Likelihood-based inference for bivariate latent failure time models with competing risks under the generalized FGM copula. Computational Statistics, 33(3):1293–1323, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0804-0.

Mahdizadeh:2018:IER

[1150] M. Mahdizadeh and Ehsan Zamanzade.

Interval estimation of P (X&lt; Y) in ranked set sampling. Computational Statistics, 33(3):1325–1348, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0795-x.

Zamanzade:2018:PER

[1151] Ehsan Zamanzade and Xinlei Wang.

Proportion estimation in ranked set sampling in the presence of tie information. Computational Statistics, 33(3): 1349–1366, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0807-x.

Malekzadeh:2018:ICM

[1152] Ahad Malekzadeh and Mahmood Kharrati-Kopaei. Inferences on the common mean of several normal populations under heteroscedasticity. Computational Statistics, 33(3):1367– 1384, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0789-0.

Wyszynski:2018:SSM

[1153] Karol Wyszynski and Giampiero Marra. Sample selection models for count data in R. Computational Statistics, 33(3):1385–1412, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0762-y;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        017-0762-y.pdf.

Lee:2018:ECM

[1154] David Lee and Harry Joe. Efficient computation of multivariate empirical distribution functions at the observed values. Computational Statistics, 33(3): 1413–1428, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0771-x.

Ker:2018:SEL

[1155] Alan P. Ker and Abdoul G. Sam. Semiparametric estimation of the link function in binary-choice single-index models. Computational Statistics, 33(3): 1429–1455, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0779-2.

Okada:2018:PPM

[1156] Kensuke Okada and Shin ichi Mayekawa.

Post-processing of Markov chain Monte Carlo output in Bayesian latent variable models with application to multidimensional scaling. Computational Statistics, 33(3):1457–1473, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0759-6.

Chekouo:2018:HDV

[1157] Thierry Chekouo and Alejandro Murua. High-dimensional variable selection with the plaid mixture model for clustering. Computational Statistics, 33(3):1475–1496, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-018-0818-
7.

Salmeron:2018:TVC

[1158] Román Salmerón, José García, Catalina García, and María del Mar López. Transformation of variables and the condition number in ridge estimation. Computational Statistics, 33(3): 1497–1524, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0769-4.

Arashi:2018:RBL

[1159] Mohammad Arashi, Mina Norouzirad, S. Ejaz Ahmed, and Bahadır Yüzbaşı. Rank-based Liu regression. Computational Statistics, 33(3): 1525–1561, September 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0809-8.

Tian:2018:JMM

[1160] Yuzhu Tian, Manlai Tang, and Maozai Tian. Joint modeling for mixed-effects quantile regression of longitudinal data with detection limits and covariates measured with error, with application to AIDS studies. Computational Statistics, 33(4):1563–1587, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0812-0.

Cheng:2018:FIA

[1161] Hao Cheng and Ying Wei. A fast imputation algorithm in quantile regression. Computational Statistics, 33(4): 1589–1603, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0813-z.

Tian:2018:QRL

[1162] Yuzhu Tian, Manlai Tang, Yanchao Zang, and Maozai Tian. Quantile regression for linear models with autoregressive errors using EM algorithm. Computational Statistics, 33(4): 1605–1625, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0811-1.

Suesse:2018:ESA

[1163] Thomas Suesse. Estimation of spatial autoregressive models with measurement error for large data sets. Computational Statistics, 33(4):1627– 1648, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-017-0774-7.

Massing:2018:SSL

[1164] Till Massing. Simulation of Student-Lévy processes using series representations. Computational Statistics, 33(4): 1649–1685, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0814-y.

Simos:2018:BIF

[1165] Theodore Simos and Mike Tsionas.

Bayesian inference of the fractional Ornstein–Uhlenbeck process under a flow sampling scheme. Computational Statistics, 33(4):1687–1713, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0799-6.

Dey:2018:BTS

[1166] Tanujit Dey, Kun Ho Kim, and Chae Young Lim. Bayesian time series regression with nonparametric modeling of autocorrelation. Computational Statistics, 33(4):1715–1731, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0796-9.

Vosseler:2018:FST

[1167] Alexander Vosseler and Enzo Weber. Forecasting seasonal time series data: a Bayesian model averaging approach. Computational Statistics, 33(4):1733– 1765, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0801-3.

Micheas:2018:PSP

[1168] Athanasios C. Micheas and Jiaxun Chen. sppmix: Poisson point process modeling using normal mixture models. Computational Statistics, 33(4): 1767–1798, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0805-z.

Chavent:2018:CRP

[1169] Marie Chavent, Vanessa Kuentz-Simonet, Amaury Labenne, and Jérôme Saracco. ClustGeo: an R package for hierarchical clustering with spatial constraints. Computational Statistics, 33(4):1799–1822, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0791-1.

Lin:2018:FEA

[1170] Jie Lin, Donald A. Adjeroh, Bing-Hua Jiang, and Yue Jiang. fastWKendall: an efficient algorithm for weighted Kendall correlation. Computational Statistics, 33(4):1823–1845, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-017-
0775-6.

Samuh:2018:ACP

[1171] Monjed H. Samuh and Fortunato Pesarin. Applications of conditional power function of two-sample permutation test. Computational Statistics, 33(4):1847–1862, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-018-0803-
1.

Baringhaus:2018:SID

[1172] L. Baringhaus, D. Gaigall, and J. P.

Thiele. Statistical inference for L^{2}- distances to uniformity. Computational Statistics, 33(4):1863–1896, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0820-0.

Kleyn:2018:PTE

[1173] J. Kleyn, M. Arashi, and S. Millard.

Preliminary test estimation in system regression models in view of asymmetry. Computational Statistics, 33(4): 1897–1921, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0794-y.

Song:2018:SAS

[1174] Jiyeon Song and Seung Jun Shin. Stability approach to selecting the number of principal components. Computational Statistics, 33(4):1923–1938, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0826-7.

Benasseni:2018:CAU

[1175] Jacques Bénasséni. A correction of approximations used in sensitivity study of principal component analysis. Computational Statistics, 33(4): 1939–1955, December 2018. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-017-0790-7.

Usami:2019:PLG

[1176] Satoshi Usami, Ross Jacobucci, and Timothy Hayes. The performance of latent growth curve model-based structural equation model trees to uncover population heterogeneity in growth trajectories. Computational Statistics, 34(1):1–22, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0815-x;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-018-0815-x.pdf.

Garcia:2019:ERS

[1177] Oscar García. Estimating reducible stochastic differential equations by conversion to a least-squares problem. Computational Statistics, 34(1):23–46, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0837-4.

Manju:2019:SIC

[1178] Md Abu Manju, Math J. J. M. Candel, and Gerard J. P. van Breukelen. SamP2CeT: an interactive computer program for sample size and power calculation for two-level cost-effectiveness trials. Computational Statistics, 34 (1):47–70, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0829-4;   http:
//link.springer.com/content/pdf/
10.1007/s00180-018-0829-4.pdf.

Jiang:2019:PBO

[1179] Hong-Yan Jiang and Rong-Xian Yue.

Pseudo-Bayesian D-optimal designs for longitudinal Poisson mixed models with correlated errors. Computational Statistics, 34(1):71–87, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0834-7.

Muns:2019:IAB

[1180] Sander Muns. An iterative algorithm to bound partial moments. Computational Statistics, 34(1):89–122, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0825-8;   http://link.springer.
com/content/pdf/10.1007/s00180-
018-0825-8.pdf.

Lachos:2019:FRM

[1181] Víctor H. Lachos, Celso R. B. Cabral, Marcos O. Prates, and Dipak K. Dey. Flexible regression modeling for censored data based on mixtures of Student-t distributions. Computational Statistics, 34(1):123–152, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0856-1.

Osiewalski:2019:JMT

[1182] Jacek Osiewalski and Jerzy Marzec. Joint modelling of two count variables when one of them can be degenerate. Computational Statistics, 34 (1):153–171, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0828-5;   http:
                                        //link.springer.com/content/pdf/
                                        10.1007/s00180-018-0828-5.pdf.

Saboor:2019:MBM

[1183] Abdus Saboor, Muhammad Nauman Khan, Gauss M. Cordeiro, Marcelino A. R. Pascoa, Juliano Bortolini, and Shahid Mubeen. Modified beta modified-Weibull distribution. Computational Statistics, 34(1):173–199, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0822-y.

OHagan:2019:IMB

[1184] Adrian O’Hagan and Arthur White. Improved model-based clustering performance using Bayesian initialization averaging. Computational Statistics, 34(1):201–231, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0855-2.

Chacko:2019:BAW

[1185] Manoj Chacko and Rakhi Mohan.

Bayesian analysis of Weibull distribution based on progressive type-II censored competing risks data with binomial removals. Computational Statistics, 34(1):233–252, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0847-2.

Kobayashi:2019:ABC

[1186] Genya Kobayashi and Kazuhiko Kakamu. Approximate Bayesian computation for Lorenz curves from grouped data. Computational Statistics, 34(1):253–279, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-018-0831-
x.

Li:2019:NNG

[1187] Lingge Li, Andrew Holbrook, Babak Shahbaba, and Pierre Baldi. Neural network gradient Hamiltonian Monte Carlo. Computational Statistics, 34 (1):281–299, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-00861-z.

Badih:2019:AVI

[1188] Ghattas Badih, Michel Pierre, and Boyer Laurent. Assessing variable importance in clustering: a new method based on unsupervised binary decision trees. Computational Statistics, 34(1):301–321, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0857-0.

Tamandi:2019:SMS

[1189] Mostafa Tamandi, Ahad Jamalizadeh, and Tsung-I Lin. Shape mixtures of skew-t-normal distributions: characterizations and estimation. Computational Statistics, 34(1):323–347, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0835-6.

Yu:2019:BES

[1190] Han Yu, Brian Chapman, Arianna Di Florio, Ellen Eischen, David Gotz, Mathews Jacob, and Rachael Hageman Blair. Bootstrapping estimates of stability for clusters, observations and model selection. Computational Statistics, 34(1):349–372, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0830-y.

Melnykov:2019:EMA

[1191] Volodymyr Melnykov and Xuwen Zhu. An extension of the K-means algorithm to clustering skewed data. Computational Statistics, 34(1):373–394, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0821-z.

Tang:2019:FLA

[1192] Lu Tang, Ling Zhou, and Peter X. K.

Song. Fusion learning algorithm to combine partially heterogeneous Cox models. Computational Statistics, 34 (1):395–414, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0827-6.

Amini:2019:IPP

[1193] Morteza Amini and Mahdi Roozbeh.

Improving the prediction performance of the LASSO by subtracting the additive structural noises. Computational Statistics, 34(1):415–432, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0849-0.

Yu:2019:HTA

[1194] Lili Yu, Liang Liu, and Ding-Geng Chen. A homoscedasticity test for the accelerated failure time model. Computational Statistics, 34(1):433–446, March 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0840-9.

Aneiros:2019:ESI

[1195] Germán Aneiros, Ricardo Cao, and Philippe Vieu. Editorial on the special issue on functional data analysis and related topics. Computational Statistics, 34(2):447–450, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00892-0;   http://link.springer.
com/content/pdf/10.1007/s00180-
019-00892-0.pdf.

Bongiorno:2019:MFD

[1196] Enea G. Bongiorno, Aldo Goia, and Philippe Vieu. Modeling functional data: a test procedure. Computational Statistics, 34(2):451–468, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0816-9.

Febrero-Bande:2019:VSF

[1197] Manuel Febrero-Bande, Wenceslao González-Manteiga, and Manuel Oviedo de la Fuente. Variable selection in functional additive regression models. Computational Statistics, 34(2):469–487, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0844-5.

Burdejova:2019:DSP

[1198] Petra Burdejová and Wolfgang K. Härdle. Dynamic semi-parametric factor model for functional expectiles. Computational Statistics, 34 (2):489–502, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00883-1.

Yu:2019:RES

[1199] Ping Yu, Zhongyi Zhu, and Zhongzhan Zhang. Robust exponential squared loss-based estimation in semi-functional linear regression models. Computational Statistics, 34(2):503–525, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0810-2.

Zambom:2019:FDC

[1200] Adriano Zanin Zambom, Julian A. A.

Collazos, and Ronaldo Dias. Functional data clustering via hypothesis testing k-means. Computational Statistics, 34(2):527–549, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0808-9.

Sottile:2019:CEC

[1201] Gianluca Sottile and Giada Adelfio.

Clusters of effects curves in quantile regression models. Computational Statistics, 34(2):551–569, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0817-8.

Gorecki:2019:FRS

[1202] Tomasz Górecki and Lukasz Smaga.

fdANOVA: an R software package for analysis of variance for univariate and multivariate functional data. Computational Statistics, 34(2):571–597, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-018-
                                        0842-7;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        018-0842-7.pdf.

Tsukada:2019:HDT

[1203] Shin ichi Tsukada. High dimensional two-sample test based on the inter-point distance. Computational Statistics, 34(2):599–615, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-017-0777-
                                        4.

Kolacek:2019:BFP

[1204] Jan Kolácek, Ondrej Pokora, Daniela Kuruczová, and Tzai-Wen Chiu. Benefits of functional PCA in the analysis of single-trial auditory evoked potentials. Computational Statistics, 34(2):617–629, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0819-6.

Traore:2019:CAE

[1205] O. I. Traore, P. Cristini, N. Favretto-Cristini, L. Pantera, P. Vieu, and S. Viguier-Pla. Clustering acoustic emission signals by mixing two stages dimension reduction and nonparametric approaches. Computational Statistics, 34(2):631–652, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        00864-w.

Huang:2019:BSS

[1206] Yu-Min Huang. Binary surrogates with stratified samples when weights are unknown. Computational Statistics, 34(2):653–682, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0838-3.

Ozkale:2019:FOA

[1207] M. Revan Özkale and Engin Arican.

A first-order approximated jackknifed ridge estimator in binary logistic regression. Computational Statistics, 34(2):683–712, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0851-6.

Guo:2019:SLS

[1208] Xiao Guo, Hai Zhang, Yao Wang, and Yong Liang. Structure learning of sparse directed acyclic graphs incorporating the scale-free property. Computational Statistics, 34(2):713–742, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0841-8.

Choi:2019:PCP

[1209] Byeong Yeob Choi, Chen-Pin Wang, Joel Michalek, and Jonathan Gelfond. Power comparison for propensity score methods. Computational Statistics, 34(2):743–761, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0852-5.

Dvorak:2019:CTS

[1210] Jakub Dvorák. Classification trees with soft splits optimized for ranking. Computational Statistics, 34(2):763–786, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-019-
                                        00867-1.

DAmbrosio:2019:MCB

[1211] Antonio D’Ambrosio, Carmela Iorio, Michele Staiano, and Roberta Siciliano. Median constrained bucket order rank aggregation. Computational Statistics, 34(2):787–802, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0858-z.

Lin:2019:PMT

[1212] Chien-Tai Lin, Ying-Chen Lee, and Narayanaswamy Balakrishnan. Package mTEXO for testing the presence of outliers in exponential samples. Computational Statistics, 34(2):803–818, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0843-6.

Wang:2019:PML

[1213] Shaoxin Wang, Hu Yang, and Chaoli Yao. On the penalized maximum likelihood estimation of high-dimensional approximate factor model. Computational Statistics, 34(2):819–846, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-019-
00869-z.

Sun:2019:PSM

[1214] Libo Sun, Chihoon Lee, and Jennifer A.

Hoeting. A penalized simulated maximum likelihood method to estimate parameters for SDEs with measurement error. Computational Statistics, 34(2):847–863, June 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0846-3.

Schmid:2019:PR

[1215] Matthias Schmid, Bernd Bischl, and Hans A. Kestler. Proceedings of Reisensburg 2016–2017. Computational Statistics, 34(3):943–944, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00907-w;   http://link.springer.
com/content/pdf/10.1007/s00180-
019-00907-w.pdf.

Beggel:2019:TSA

[1216] Laura Beggel, Bernhard X. Kausler, Martin Schiegg, Michael Pfeiffer, and Bernd Bischl. Time series anomaly detection based on shapelet learning. Computational Statistics, 34(3): 945–976, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0824-9;   http:
//link.springer.com/content/pdf/
10.1007/s00180-018-0824-9.pdf.

Casalicchio:2019:ORP

[1217] Giuseppe Casalicchio, Jakob Bossek, Michel Lang, Dominik Kirchhoff, Pascal Kerschke, Benjamin Hofner, Heidi Seibold, Joaquin Vanschoren, and Bernd Bischl. OpenML: An R package to connect to the machine learning platform OpenML. Computational Statistics, 34(3):977–991, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-017-
                                        0742-2.

Welchowski:2019:SKD

[1218] Thomas Welchowski and Matthias Schmid. Sparse kernel deep stacking networks. Computational Statistics, 34(3):993–1014, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL http://link.springer.com/
                                        article/10.1007/s00180-018-0832-
                                        9.

Costilla:2019:BMB

[1219] Roy Costilla, Ivy Liu, Richard Arnold, and Daniel Fernández. Bayesian model-based clustering for longitudinal ordinal data. Computational Statistics, 34(3):1015–1038, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-019-
                                        00872-4.

Maleki:2019:MMR

[1220] Mohsen Maleki and Darren Wraith. Mixtures of multivariate restricted skew-normal factor analyzer models in a Bayesian framework. Computational Statistics, 34(3):1039–1053, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00870-6.

Arellano-Valle:2019:SND

[1221] Reinaldo B. Arellano-Valle, Javier E.

Contreras-Reyes, Freddy O. López Quintero, and Abel Valdebenito. A skew-normal dynamic linear model and Bayesian forecasting. Computational Statistics, 34(3):1055–1085, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0848-1.

Bauer:2019:NPS

[1222] Verena Bauer, Karl Fürlinger, and Göran Kauermann. A note on parallel sampling in Markov graphs. Computational Statistics, 34(3):1087– 1107, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00880-4.

Luo:2019:MTM

[1223] Xin Luo and Håkon Tjelmeland.

A multiple-try Metropolis–Hastings algorithm with tailored proposals. Computational Statistics, 34(3):1109– 1133, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00878-y.

Mbah:2019:RBS

[1224] Chamberlain Mbah, Kris Peremans, Stefan Van Aelst, and Dries F. Benoit. Robust Bayesian seemingly unrelated regression model. Computational Statistics, 34(3):1135–1157, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        0854-3.

Kim:2019:OBT

[1225] Dal Ho Kim, Woo Dong Lee, Sang Gil Kang, and Yongku Kim. Objective Bayesian tests for fieller–creasy problem. Computational Statistics, 34(3): 1159–1182, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-018-0853-4.

Manouchehri:2019:PAM

[1226] T. Manouchehri and A. R. Nematollahi. Periodic autoregressive models with closed skew-normal innovations. Computational Statistics, 34(3): 1183–1213, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-019-00893-z.

Altun:2019:TSE

[1227] Emrah Altun. Two-sided exponentialgeometric distribution: inference and volatility modeling. Computational Statistics, 34(3):1215–1245, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.

com/article/10.1007/s00180-019-
00873-3.

Cankaya:2019:RES

[1228] Mehmet Niyazi Çankaya, Abdullah Yalçınkaya, Ömer AltındaÇ§, and Olcay Arslan. On the robustness of an epsilon skew extension for Burr III distribution on the real line. Computational Statistics, 34(3):1247–1273, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
0859-y.

Khan:2019:IGB

[1229] N. Mamode Khan, Y. Sunecher, V. Jowaheer, M. M. Ristic, and M. Heenaye-Mamode Khan. Investigating GQL-based inferential approaches for non-stationary BI-NAR(1) model under different quantum of over-dispersion with application. Computational Statistics, 34(3): 1275–1313, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0836-5.

Britos:2019:RES

[1230] Grisel Maribel Britos and Silvia María Ojeda. Robust estimation for spatial autoregressive processes based on bounded innovation propagation representations. Computational Statistics, 34(3):1315–1335, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-018-0845-

Vaiciulyte:2019:REM

[1231] Jurate Vaiciulyte and Leonidas Sakalauskas. Recursive estimation of multivariate hidden Markov model parameters. Computational Statistics, 34(3): 1337–1353, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-019-00877-z.

El-Shorbagy:2019:EGA

[1232] M. A. El-Shorbagy, A. Y. Ayoub, A. A. Mousa, and I. M. El-Desoky. An enhanced genetic algorithm with new mutation for cluster analysis. Computational Statistics, 34(3):1355– 1392, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-019-00871-5.

Tian:2019:SNS

[1233] Guo-Liang Tian, Xiqian Ding, Yin Liu, and Man-Lai Tang. Some new statistical methods for a class of zero-truncated discrete distributions with applications. Computational Statistics, 34(3):1393–1426, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-018-
                                        00860-0.

Zhu:2019:PMM

[1234] Xuwen Zhu. Probability of misclassification in model-based clustering. Computational Statistics, 34(3): 1427–1442, September 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00868-0.

Hofmann:2019:DEA

[1235] Heike Hofmann, Hadley Wickham, and Dianne Cook. The 2013 Data Expo of the American Statistical Association. Computational Statistics, 34(4): 1443–1447, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00923-w;   http:
//link.springer.com/content/pdf/
10.1007/s00180-019-00923-w.pdf.

Kaplan:2019:PRG

[1236] Andee J. Kaplan and Eric R. Hare.

Putting down roots: a graphical exploration of community attachment. Computational Statistics, 34(4):1449– 1464, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-0850-7.

Maurer:2019:TFC

[1237] Karsten Maurer, Dave Osthus, and Adam Loy. A tale of four cities: exploring the soul of State College, Detroit, Milledgeville and Biloxi. Computational Statistics, 34(4):1465–1487, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-018-
00863-x.

Ackerman:2019:CSO

[1238] Samuel Ackerman. Consistency of survey opinions and external data.

Computational Statistics, 34(4):1489– 1509, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00882-2.

McNamara:2019:CES

[1239] Amelia A. McNamara. Community engagement and subgroup metaknowledge: some factors in the soul of a community. Computational Statistics, 34(4):1511–1535, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00879-x.

daSilva:2019:CCE

[1240] Natalia da Silva and Ignacio Alvarez-Castro. Clicks and cliques: exploring the soul of the community. Computational Statistics, 34(4):1537–1563, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00881-3.

Quach:2019:SCA

[1241] Anna Quach, Jürgen Symanzik, and Nicole Forsgren. Soul of the community: an attempt to assess attachment to a community. Computational Statistics, 34(4):1565–1589, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00866-2.

Orth:2019:DCA

[1242] Jessica M. Orth. Drivers of community attachment: an interactive analysis. Computational Statistics, 34(4): 1591–1611, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-018-00862-y;   http:
//link.springer.com/content/pdf/
10.1007/s00180-018-00862-y.pdf.

Poterie:2019:CTA

[1243] A. Poterie, J.-F. Dupuy, V. Monbet, and L. Rouvière. Classification tree algorithm for grouped variables. Computational Statistics, 34(4): 1613–1648, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00894-y.

Chang:2019:MEO

[1244] Chung Chang, Meng-Ke Hsieh, An Jen Chiang, Yi-Hsuan Tsai, Chia-Chiung Liu, and Jiabin Chen. Methods for estimating the optimal number and location of cut points in multivariate survival analysis: a statistical solution to the controversial effect of BMI. Computational Statistics, 34(4): 1649–1674, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00908-9.

Johnson:2019:SAM

[1245] Nels G. Johnson and Inyoung Kim.

Semiparametric approaches for matched case-control studies with error-in-covariates. Computational Statistics, 34(4):1675–1692, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL http://link.springer.com/
article/10.1007/s00180-019-00888-
w.

Bhuyan:2019:ERE

[1246] Prajamitra Bhuyan. Estimation of random-effects model for longitudinal data with nonignorable missingness using Gibbs sampling. Computational Statistics, 34(4):1693–1710, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00887-x;   http://link.springer.
com/content/pdf/10.1007/s00180-
019-00887-x.pdf.

Liu:2019:WCQ

[1247] Huilan Liu, Hu Yang, and Changgen Peng. Weighted composite quantile regression for single index model with missing covariates at random. Computational Statistics, 34(4):1711– 1740, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00886-y.

Ramosaj:2019:PMV

[1248] Burim Ramosaj and Markus Pauly. Predicting missing values: a comparative study on non-parametric approaches for imputation. Computational Statistics, 34(4):1741–1764, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00900-3.

Wang:2019:BIE

[1249] Bei Wang, Yi Zheng, Kyle M. Irimata, and Jeffrey R. Wilson. Bootstrap ICC estimators in analysis of small clustered binary data. Computational Statistics, 34(4):1765–1778, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00885-z.

OHagan:2019:IPU

[1250] Adrian O’Hagan, Thomas Brendan Murphy, Luca Scrucca, and Isobel Claire Gormley. Investigation of parameter uncertainty in clustering using a Gaussian mixture model via jackknife, bootstrap and weighted likelihood bootstrap. Computational Statistics, 34(4):1779–1813, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00897-9.

Bisaglia:2019:MBI

[1251] Luisa Bisaglia and Margherita Gerolimetto.

Model-based INAR bootstrap for forecasting INAR(p) models. Computational Statistics, 34(4):1815–1848, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00902-1.

Bourel:2019:BDE

[1252] Mathias Bourel and Jairo Cugliari.

Bagging of density estimators. Computational Statistics, 34(4):1849– 1869, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00889-9.

Konecna:2019:MLM

[1253] Katerina Konecná and Ivanka Horová. Maximum likelihood method for bandwidth selection in kernel conditional density estimate. Computational Statistics, 34(4):1871–1887, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00884-0.

Blanchet-Scalliet:2019:FAC

[1254] Christophette Blanchet-Scalliet, Céline Helbert, Mélina Ribaud, and Céline Vial. Four algorithms to construct a sparse kriging kernel for dimensionality reduction. Computational Statistics, 34(4):1889–1909, December 2019. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00874-2.

Renaux:2020:HIG

[1255] Claude Renaux, Laura Buzdugan, Markus Kalisch, and Peter Bühlmann. Hierarchical inference for genomewide association studies: a view on methodology with software. Computational Statistics, 35(1):1–40, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-

00939-2. See comments [1256, 1257, 1258, 1259] and rejoinder [1260].

Goeman:2020:CHI

[1256] Jelle J. Goeman and Stefan Böhringer.

Comments on: Hierarchical inference for genome-wide association studies. Computational Statistics, 35(1):41–45, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-

00943-6. See [1255].

Frommlet:2020:CHI

[1257] Florian Frommlet and Michael G.

Schimek. Comments on: Hierarchical inference for genome-wide association studies: a view on methodology with software by F. Frommlet and M. G. Schimek. Computational Statistics, 35(1):47–49, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-

00945-4. See [1255].

Heller:2020:CHI

[1258] Ruth Heller. Comments on: Hierarchical inference for genome-wide association studies: a view on methodology with software by F. Frommlet and M. G. Schimek. Computational Statistics, 35(1):51–55, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-

00942-7. See [1255].

Rodrigues:2020:CHI

[1259] Paulo C. Rodrigues and Vanda M.

Lourenço. Comments on: Hierarchical Inference for genome-wide association studies: a view on methodology with software. Computational Statistics, 35(1):57–58, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00941-8.  See

[1255].

Renaux:2020:RHI

[1260] Claude Renaux, Laura Buzdugan, Markus Kalisch, and Peter Bühlmann. Rejoinder on: Hierarchical inference for genome-wide association studies: a view on methodology with software. Computational Statistics, 35(1):59–67, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-

00948-1. See [1255].

Tang:2020:SCB

[1261] Fengqin Tang, Chunning Wang, Jinxia Su, and Yuanyuan Wang. Spectral clustering-based community detection using graph distance and node attributes. Computational Statistics, 35(1):69–94, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00909-8.

Tang:2020:FIA

[1262] Zhou Tang, Zhangsheng Yu, and Cheng Wang. A fast iterative algorithm for high-dimensional differential network. Computational Statistics, 35 (1):95–109, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00915-w.

Perrin:2020:AMI

[1263] Guillaume Perrin and Christian Soize.

Adaptive method for indirect identification of the statistical properties of random fields in a Bayesian framework. Computational Statistics, 35 (1):111–133, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00936-5.

Bakar:2020:IDR

[1264] K. Shuvo Bakar. Interpolation of daily rainfall data using censored Bayesian spatially varying model. Computational Statistics, 35(1):135–152, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00911-0.

Barzegar:2020:SBN

[1265] Zahra Barzegar and Firoozeh Rivaz. A scalable Bayesian nonparametric model for large spatio-temporal data. Computational Statistics, 35(1):153–173, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00905-y.

Magnusson:2020:DRS

[1266] Måns Magnusson, Leif Jonsson, and Mattias Villani. DOLDA: a regularized supervised topic model for high-dimensional multi-class regression. Computational Statistics, 35 (1):175–201, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00891-1;   http:
//link.springer.com/content/pdf/
10.1007/s00180-019-00891-1.pdf.

Shao:2020:CHD

[1267] Wei Shao and Yijun Zuo. Computing the halfspace depth with multiple try algorithm and simulated annealing algorithm. Computational Statistics, 35(1):203–226, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00906-x.

Yang:2020:BVS

[1268] Mingan Yang, Min Wang, and Guanghui Dong. Bayesian variable selection for mixed effects model with shrinkage prior. Computational Statistics, 35(1):227–243, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00895-x.

Yang:2020:SBV

[1269] Aijun Yang, Yuzhu Tian, Yunxian Li, and Jinguan Lin. Sparse Bayesian variable selection in kernel probit model for analyzing high-dimensional data. Computational Statistics, 35(1):245–258, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00917-8.

Altun:2020:UIS

[1270] Emrah Altun and Gauss M. Cordeiro. The unit-improved second-degree Lindley distribution: inference and regression modeling. Computational Statistics, 35(1):259–279, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00921-y.

Alizadeh:2020:OLL

[1271] Morad Alizadeh, Ahmed Z. Afify, M. S. Eliwa, and Sajid Ali. The odd log-logistic Lindley-g family of distributions: properties, Bayesian and non-Bayesian estimation with applications. Computational Statistics, 35 (1):281–308, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00932-9.

Asgharzadeh:2020:OCR

[1272] A. Asgharzadeh, S. F. Bagheri, N. A. Ibrahim, and M. R. Abubakar. Optimal confidence regions for the two-parameter exponential distribution based on records. Computational Statistics, 35(1):309–326, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00914-x.

Hu:2020:DDE

[1273] Chaoran Hu, Vladimir Pozdnyakov, and Jun Yan. Density and distribution evaluation for convolution of independent gamma variables. Computational Statistics, 35(1):327–342, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.

                                        com/article/10.1007/s00180-019-
                                        00924-9.

Saengkyongam:2020:ECS

[1274] Sorawit Saengkyongam, Anthony Hayter, Seksan Kiatsupaibul, and Wei Liu. Efficient computation of the stochastic behavior of partial sum processes. Computational Statistics, 35 (1):343–358, March 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

                                        //link.springer.com/article/10.
                                        1007/s00180-019-00920-z.

Wu:2020:SDA

[1275] Desheng Dash Wu and Wolfgang Karl Härdle. Service data analytics and business intelligence 2017. Computational Statistics, 35(2):423–426, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-020-
                                        00968-2;   http://link.springer.
                                        com/content/pdf/10.1007/s00180-
                                        020-00968-2.pdf.

Chen:2020:EDC

[1276] Shiyi Chen, Wolfgang K. Härdle, and Li Wang. Estimation and determinants of Chinese banks’ total factor efficiency: a new vision based on unbalanced development of Chinese banks and their overall risk. Computational Statistics, 35(2):427–468, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

                                        tronic). URL http://link.springer.
                                        com/article/10.1007/s00180-019-
                                        00951-6.

Fan:2020:ASO

[1277] Xuanzhu Fan, Jiafu Tang, and Chongjun Yan. Appointment scheduling optimization with two stages diagnosis for clinic outpatient. Computational Statistics, 35(2):469–490, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00876-0.

Guan:2020:IAF

[1278] Rong Guan, Huiwen Wang, and Haitao Zheng. Improving accuracy of financial distress prediction by considering volatility: an interval-data-based discriminant model. Computational Statistics, 35(2):491–514, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00916-9.

Kong:2020:RMO

[1279] Weichang Kong, Fei Qiao, and Qidi Wu. Real-manufacturing-oriented big data analysis and data value evaluation with domain knowledge. Computational Statistics, 35(2):515–538, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00919-6.

Yan:2020:LBE

[1280] Tianshun Yan, Yanyong Zhao, and Wentao Wang. Likelihood-based estimation of a semiparametric timedependent jump diffusion model of the short-term interest rate. Computational Statistics, 35(2):539–557, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00875-1.

Smallman:2020:SPP

[1281] Luke Smallman, William Underwood, and Andreas Artemiou. Simple Poisson PCA: an algorithm for (sparse) feature extraction with simultaneous dimension determination. Computational Statistics, 35(2):559–577, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00903-0;   http://link.springer.
com/content/pdf/10.1007/s00180-
019-00903-0.pdf.

Carcenac:2020:ASM

[1282] Manuel Carcenac and Soydan Redif. Application of the sequential matrix diagonalization algorithm to highdimensional functional MRI data. Computational Statistics, 35(2):579– 605, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://link.

springer.com/article/10.1007/s00180-
019-00925-8.

Chiapino:2020:MEV

[1283] Maël Chiapino, Stephan Clémençon, Vincent Feuillard, and Anne Sabourin. A multivariate extreme value theory approach to anomaly clustering and visualization. Computational Statistics, 35(2):607–628, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00913-y.

Peeters:2020:SCN

[1284] Carel F. W. Peeters, Mark A. van de Wiel, and Wessel N. van Wieringen. The spectral condition number plot for regularization parameter evaluation. Computational Statistics, 35 (2):629–646, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00912-z;   http:
//link.springer.com/content/pdf/
10.1007/s00180-019-00912-z.pdf.

Salmeron-Gomez:2020:DQN

[1285] Román Salmerón-Gómez, Ainara Rodríguez-Sánchez, and Catalina García-García. Diagnosis and quantification of the non-essential collinearity. Computational Statistics, 35 (2):647–666, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00922-x.

Jadhav:2020:LRL

[1286] N. H. Jadhav. On linearized ridge logistic estimator in the presence of multicollinearity. Computational Statistics, 35(2):667–687, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00935-6.

Brouste:2020:CFM

[1287] Alexandre Brouste, Christophe Dutang, and Tom Rohmer. Closedform maximum likelihood estimator for generalized linear models in the case of categorical explanatory variables: application to insurance loss modeling. Computational Statistics, 35(2):689–724, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00918-7.

Lee:2020:EZI

[1288] Shen-Ming Lee, T. Martin Lukusa, and Chin-Shang Li. Estimation of a zeroinflated Poisson regression model with missing covariates via nonparametric multiple imputation methods. Computational Statistics, 35(2):725–754, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00930-x.

Garcia-Barzana:2020:MLR

[1289] Marta García-Bárzana, Ana Belén Ramos-Guajardo, Ana Colubi, and Erricos J. Kontoghiorghes. Multiple linear regression models for random intervals: a set arithmetic approach. Computational Statistics, 35 (2):755–773, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00910-1.

Filova:2020:AQA

[1290] Lenka Filová and Radoslav Harman. Ascent with quadratic assistance for the construction of exact experimental designs. Computational Statistics, 35(2):775–801, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-020-00961-9.

Bos:2020:NOT

[1291] Len Bos, Federico Piazzon, and Marco Vianello. Near G-optimal Tchakaloff designs. Computational Statistics, 35 (2):803–819, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00933-8.

Pooladsaz:2020:AFE

[1292] Saeid Pooladsaz and Mahboobeh Doosti-Irani. An algorithm for finding efficient test-control block designs with correlated observations. Computational Statistics, 35(2):821–836, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00904-z.

DeSalvo:2020:RSC

[1293] Stephen DeSalvo and James Zhao.

Random sampling of contingency tables via probabilistic divide-and-conquer. Computational Statistics, 35(2):837–869, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http:

//link.springer.com/article/10.
1007/s00180-019-00899-7.

Powell:2020:CEM

[1294] Ben Powell and Paul A. Smith. Computing expectations and marginal likelihoods for permutations. Computational Statistics, 35(2):871–891, June 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s00180-019-
00901-2;   http://link.springer.
com/content/pdf/10.1007/s00180-
019-00901-2.pdf.

Sambasivan:2020:BPS

[1295] Rajiv Sambasivan, Sourish Das, and Sujit K. Sahu. A Bayesian perspective of statistical machine learning for big data. Computational Statistics, 35 (3):893–930, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00970-8.

Li:2020:SVM

[1296] Peizhi Li, Yingwei Peng, and Qingli Dong. A support vector machine based semiparametric mixture cure model. Computational Statistics, 35 (3):931–945, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00931-w.

Lux:2020:DDV

[1297] Marius Lux, Wolfgang Karl Härdle, and Stefan Lessmann. Data driven value-at-risk forecasting using a SVR-GARCH-KDE hybrid. Computational Statistics, 35(3):947–981, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-019-00934-
7.

Calhoun:2020:RFA

[1298] Peter Calhoun, Melodie J. Hallett, and Juanjuan Fan. Random forest with acceptance-rejection trees. Computational Statistics, 35(3):983– 999, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00929-4.

Beaulac:2020:BDT

[1299] Cédric Beaulac and Jeffrey S. Rosenthal. BEST: a decision tree algorithm that handles missing values. Computational Statistics, 35(3): 1001–1026, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00987-z.

Turner:2020:MRR

[1300] Heather L. Turner, Jacob van Etten, and Ioannis Kosmidis. Modelling rankings in R: the PlackettLuce package. Computational Statistics, 35(3): 1027–1057, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00959-3.

Marque-Pucheu:2020:EDR

[1301] Sophie Marque-Pucheu, Guillaume Perrin, and Josselin Garnier. An efficient dimension reduction for the Gaussian process emulation of two nested codes with functional outputs. Computational Statistics, 35(3): 1059–1099, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00926-7.

Schmutz:2020:CMF

[1302] Amandine Schmutz, Julien Jacques, and Pauline Martin. Clustering multivariate functional data in group-specific functional subspaces. Computational Statistics, 35(3):1101– 1131, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00958-4.

Larsson:2020:DFA

[1303] Rolf Larsson. Discrete factor analysis using a dependent Poisson model. Computational Statistics, 35(3):1133– 1152, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00960-w.

Wang:2020:UHD

[1304] Huiwen Wang, Ruiping Liu, and Gilbert Saporta. Ultra-high dimensional variable screening via Gram– Schmidt orthogonalization. Computational Statistics, 35(3):1153– 1170, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00963-7.

Laa:2020:UTV

[1305] Ursula Laa and Dianne Cook. Using tours to visually investigate properties of new projection pursuit indexes with application to problems in physics. Computational Statistics, 35(3):1171– 1205, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00954-8.

Palacios-Gonzalez:2020:FAE

[1306] Federico Palacios-González and Rosa M.

García-Fernández. A faster algorithm to estimate multiresolution densities. Computational Statistics, 35(3): 1207–1230, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00952-w.

Diaz-Coto:2020:SRP

[1307] Susana Díaz-Coto, Pablo Martínez-Camblor, and Sonia Pérez-Fernández. smoothROCtime: an R package for time-dependent ROC curve estimation. Computational Statistics, 35(3): 1231–1251, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00955-7.

Savchuk:2020:OSC

[1308] Olga Y. Savchuk. One-sided crossvalidation for nonsmooth density functions. Computational Statistics, 35(3): 1253–1272, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00938-3.

Yavuz:2020:PCL

[1309] Fulya Gokalp Yavuz and Barret Schloerke. Parallel computing in linear mixed models. Computational Statistics, 35(3):1273–1289, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-019-00950-
                                        7.

Noghrehchi:2020:MIF

[1310] Firouzeh Noghrehchi, Jakub Stoklosa, and Spiridon Penev. Multiple imputation and functional methods in the presence of measurement error and missingness in explanatory variables. Computational Statistics, 35(3): 1291–1317, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00976-2.

Henderson:2020:EIS

[1311] Donna Henderson and Gerton Lunter. Efficient inference in state-space models through adaptive learning in online Monte Carlo expectation maximization. Computational Statistics, 35(3): 1319–1344, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-019-00937-4.

Jose:2020:TDS

[1312] Joby K. Jose and M. Drisya. Timedependent stress-strength reliability models based on phase type distribution. Computational Statistics, 35(3): 1345–1371, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00991-3.

Ding:2020:SMI

[1313] Dong Ding, Axel Gandy, and Georg Hahn. A simple method for implementing Monte Carlo tests. Computational Statistics, 35(3):1373– 1392, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00927-6.

Battaglia:2020:DEA

[1314] Francesco Battaglia, Domenico Cucina, and Manuel Rizzo. Detection and estimation of additive outliers in seasonal time series. Computational Statistics, 35(3):1393– 1409, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00928-5.

Garcia-Jorcano:2020:DAC

[1315] Laura Garcia-Jorcano and Alfonso Novales. A dominance approach for comparing the performance of VaR forecasting models. Computational Statistics, 35(3):1411– 1448, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00990-4.

Diriba:2020:MDE

[1316] Tadele Akeba Diriba and Legesse Kassa Debusho. Modelling dependency effect to extreme value distributions with application to extreme wind speed at Port Elizabeth, South Africa: a frequentist and Bayesian approaches. Computational Statistics, 35(3):1449– 1479, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00947-2.

Huang:2020:BAE

[1317] Hanwen Huang, Andreas Handel, and Xiao Song. A Bayesian approach to estimate parameters of ordinary differential equation. Computational Statistics, 35(3):1481– 1499, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00962-8.

Shaochuan:2020:BMC

[1318] Lu Shaochuan. Bayesian multiple changepoints detection for Markov jump processes. Computational Statistics, 35(3):1501–1523, September 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-020-00956-
                                        6.

Wu:2020:MAI

[1319] Bo-Hong Wu, Hirofumi Michimae, and Takeshi Emura. Meta-analysis of individual patient data with semicompeting risks under the Weibull joint frailty-copula model. Computational Statistics, 35(4):1525– 1552, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00977-1.

Khalifeh:2020:SFA

[1320] Ashkan Khalifeh, Eisa Mahmoudi, and Ajit Chaturvedi. Sequential fixed-accuracy confidence intervals for the stress-strength reliability parameter for the exponential distribution: two-stage sampling procedure. Computational Statistics, 35(4): 1553–1575, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00957-5.

Guo:2020:PQL

[1321] Guangbao Guo, Yue Sun, and Xuejun Jiang. A partitioned quasi-likelihood for distributed statistical inference. Computational Statistics, 35(4):1577– 1596, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00974-4.

Daneshgar:2020:WAR

[1322] Neda Daneshgar and Majid Sar-

mad.  word.alignment: an R

package for computing statistical word alignment and its evaluation. Computational Statistics, 35(4):1597– 1619, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00979-z.

Sun:2020:IEC

[1323] Jing Sun. An improvement on the efficiency of complete-case-analysis with nonignorable missing covariate data. Computational Statistics, 35(4): 1621–1636, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00964-6.

Xiao:2020:EPL

[1324] Yan-Ting Xiao and Fu-Xiao Li. Estimation in partially linear varyingcoefficient errors-in-variables models with missing response variables. Computational Statistics, 35(4):1637– 1658, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00967-3.

Mostel:2020:SIM

[1325] Linda Möstel, Marius Pfeuffer, and Matthias Fischer. Statistical inference for Markov chains with applications to credit risk. Computational Statistics, 35(4):1659– 1684, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00978-0.

Zhang:2020:MTI

[1326] Lingyue Zhang, Dawei Lu, and Xiaoguang Wang. Measuring and testing interdependence among random vectors based on Spearman’s ρ and Kendall’s τ. Computational Statistics, 35(4):1685–1713, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-00973-
5.

Kim:2020:MAP

[1327] Hee-Young Kim, Christian H. Weiß, and Tobias A. Möller. Models for autoregressive processes of bounded counts: How different are they? Computational Statistics, 35(4):1715– 1736, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00980-6.

Shamma:2020:TSM

[1328] Nisreen Shamma, Mehrnaz Mohammadpour, and Masoumeh Shirozhan. A time series model based on dependent zero inflated counting series. Computational Statistics, 35(4): 1737–1757, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00982-4.

Zhang:2020:RJM

[1329] Weiping Zhang, Feiyue Xie, and Jiaxin Tan. A robust joint modeling approach for longitudinal data with informative dropouts. Computational Statistics, 35(4):1759– 1783, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00972-6.

Mahdizadeh:2020:EAV

[1330] M. Mahdizadeh and Ehsan Zamanzade. Estimating asymptotic variance of M-estimators in ranked set sampling. Computational Statistics, 35(4): 1785–1803, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00946-3.

Taconeli:2020:PEM

[1331] Cesar Augusto Taconeli and Wagner Hugo Bonat. On the performance of estimation methods under ranked set sampling. Computational Statistics, 35(4):1805– 1826, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00953-9.

Taconeli:2020:MLE

[1332] Cesar Augusto Taconeli and Suely Ruiz Giolo. Maximum likelihood estimation based on ranked set sampling designs for two extensions of the Lindley distribution with uncensored and right-censored data. Computational Statistics, 35(4):1827– 1851, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00984-2.

Liang:2020:RSD

[1333] Baosheng Liang, Peng Wu, and Yanping Qiu. Regression and subgroup detection for heterogeneous samples. Computational Statistics, 35(4):1853– 1878, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00965-5.

Haslbeck:2020:ENC

[1334] Jonas M. B. Haslbeck and Dirk U. Wulff. Estimating the number of clusters via a corrected clustering instability. Computational Statistics, 35(4): 1879–1894, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00981-5.

Ozsoy:2020:UHO

[1335] Volkan Soner Özsoy, Mehmet Güray Ünsal, and H. Hasan Örkcü. Use of the heuristic optimization in the parameter estimation of generalized gamma distribution: comparison of GA, DE, PSO and SA methods. Computational Statistics, 35(4):1895– 1925, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00966-4.

Jafari:2020:CSP

[1336] Ali Akbar Jafari and Javad Shaabani. Comparing scale parameters in several gamma distributions with known shapes. Computational Statistics, 35(4):1927– 1950, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00983-3.

Furlan:2020:CAS

[1337] Claudia Furlan and Cinzia Mortarino. Comparison among simultaneous confidence regions for nonlinear diffusion models. Computational Statistics, 35(4):1951– 1991, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-019-00949-0.

Kim:2020:IGF

[1338] Jiwoong Kim. Implementation of a goodness-of-fit test through Khmaladze martingale transformation. Computational Statistics, 35(4):

1993–2017, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00971-7.

Stewart:2020:MEL

[1339] Patrick Stewart and Wei Ning. Modified empirical likelihood-based confidence intervals for data containing many zero observations. Computational Statistics, 35(4):2019– 2042, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00993-1.

Spade:2020:GEM

[1340] David A. Spade. Geometric ergodicity of a Metropolis–Hastings algorithm for Bayesian inference of phylogenetic branch lengths. Computational Statistics, 35(4):2043– 2076, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00969-1.

Shan:2020:BRE

[1341] Guodong Shan, Yiheng Hou, and Baisen Liu. Bayesian robust estimation of partially functional linear regression models using heavy-tailed distributions. Computational Statistics, 35(4): 2077–2092, December 2020. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00975-3.

Jouvin:2021:GCC

[1342] Nicolas Jouvin, Pierre Latouche, and Alain Livartowski. Greedy clustering of count data through a mixture of multinomial PCA. Computational Statistics, 36(1):1–33, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01008-9.

Liverani:2021:CMC

[1343] Silvia Liverani, Lucy Leigh, and Julie E. Byles. Clustering method for censored and collinear survival data. Computational Statistics, 36 (1):35–60, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01000-3.

Melnykov:2021:TMM

[1344] Yana Melnykov, Xuwen Zhu, and Volodymyr Melnykov. Transformation mixture modeling for skewed data groups with heavy tails and scatter. Computational Statistics, 36 (1):61–78, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01009-8.

Slaoui:2021:RNP

[1345] Yousri Slaoui. Recursive nonparametric kernel classification rule estimation for independent functional data. Computational Statistics, 36 (1):79–112, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01024-9.

Shang:2021:TGN

[1346] Zongyuan Shang and Alan Ker. Two generalized nonparametric methods for estimating like densities. Computational Statistics, 36(1):113–126, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01007-
w.

Estevez-Perez:2021:NWR

[1347] Graciela Estévez-Pérez and Philippe Vieu. A new way for ranking functional data with applications in diagnostic test. Computational Statistics, 36(1):127–154, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01020-z.

Mohanty:2021:ASF

[1348] Soumya D. Mohanty and Ethan Fahnestock. Adaptive spline fitting with particle swarm optimization. Computational Statistics, 36 (1):155–191, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01022-x.

Nodehi:2021:EPM

[1349] Anahita Nodehi, Mousa Golalizadeh, and Claudio Agostinelli. Estimation of parameters in multivariate wrapped models for data on a ptorus. Computational Statistics, 36 (1):193–215, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01006-x.

Genc:2021:UGE

[1350] Murat Genç and M. Revan Özkale. Usage of the GO estimator in high dimensional linear models. Computational Statistics, 36(1):217–239, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01001-
2.

Biswas:2021:BQR

[1351] Jayabrata Biswas and Kiranmoy Das.

A Bayesian quantile regression approach to multivariate semi-continuous longitudinal data. Computational Statistics, 36(1):241–260, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01002-
1.

Chen:2021:BIN

[1352] Cathy W. S. Chen, Sangyeol Lee, and K. Khamthong. Bayesian inference of nonlinear hysteretic integervalued GARCH models for disease counts. Computational Statistics, 36 (1):261–281, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01018-7.

Zheng:2021:KKL

[1353] Songfeng Zheng. KLERC: kernel Lagrangian expectile regression calculator. Computational Statistics, 36 (1):283–311, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01003-0.

Kabaila:2021:CEV

[1354] Paul Kabaila and Nishika Ranathunga. Computation of the expected value of a function of a chi-distributed random variable. Computational Statistics, 36 (1):313–332, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01005-y.

Pietrosanu:2021:AAP

[1355] Matthew Pietrosanu, Jueyu Gao, and Di Niu. Advanced algorithms for penalized quantile and composite quantile regression. Computational Statistics, 36(1):333–346, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01010-1.

Ramirez-Padron:2021:RWG

[1356] Ruben Ramirez-Padron, Boris Mederos, and Avelino J. Gonzalez. Robust weighted Gaussian processes. Computational Statistics, 36(1):347–373, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https://link.

                                        springer.com/article/10.1007/s00180-
                                        020-01011-0.

Dorre:2021:LBA

[1357] Achim Dörre, Chung-Yan Huang, and Takeshi Emura. Likelihoodbased analysis of doubly-truncated data under the location-scale and AFT model. Computational Statistics, 36 (1):375–408, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01027-6.

Yuan:2021:PFB

[1358] Tianlu Yuan. The 8-parameter Fisher–Bingham distribution on the sphere. Computational Statistics, 36 (1):409–420, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01023-w.

Yuan:2021:CPF

[1359] Tianlu Yuan. Correction to: The 8-parameter Fisher-Bingham distribution on the sphere. Computational Statistics, 36(1):421, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01035-
6.

Dogru:2021:FMS

[1360] Fatma Zehra Dogru and Olcay Arslan.

Finite mixtures of skew Laplace normal distributions with random skewness. Computational Statistics, 36 (1):423–447, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01025-8.

Bhushan:2021:OIM

[1361] Shashi Bhushan and Abhay Pratap Pandey. Optimal imputation of the missing data using multi auxiliary information. Computational Statistics, 36(1):449–477, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01016-9.

Ren:2021:IOC

[1362] Junru Ren and Wenhao Gui. Inference and optimal censoring scheme for progressively Type-II censored competing risks model for generalized Rayleigh distribution. Computational Statistics, 36(1):479–513, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01021-y.

Siddiqa:2021:MRC

[1363] Hajra Siddiqa, Sajid Ali, and Ismail Shah. Most recent changepoint detection in censored panel data. Computational Statistics, 36 (1):515–540, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01028-5.

Jin:2021:PWC

[1364] Jun Jin, Tiefeng Ma, and Shuangzhe Liu. Penalized weighted composite quantile regression for partially linear varying coefficient models with missing covariates. Computational Statistics, 36(1):541–575, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01012-z.

Arbel:2021:DPM

[1365] Julyan Arbel, Riccardo Corradin, and Bernardo Nipoti. Dirichlet process mixtures under affine transformations of the data. Computational Statistics, 36(1):577–601, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01013-y.

Hanafi:2021:GPH

[1366] Mohamed Hanafi, Pasquale Dolce, and Zouhair El Hadri. Generalized properties for hanafi-wold’s procedure in partial least squares path modeling. Computational Statistics, 36 (1):603–614, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01015-w.

Romano:2021:PPR

[1367] Rosaria Romano and Francesco Palumbo.

Partial possibilistic regression path modeling: handling uncertainty in path modeling. Computational Statistics, 36(1):615–639, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01026-7.

Qin:2021:MCE

[1368] Xu Qin. A modified canny edge detector based on weighted least squares. Computational Statistics, 36 (1):641–659, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01017-8.

Ahmadi-Javid:2021:EDM

[1369] Amir Ahmadi-Javid and Mohsen Ebadi. Economic design of memorytype control charts: The fallacy of the formula proposed by Lorenzen and Vance (1986). Computational Statistics, 36(1):661–690, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01019-
6.

Simone:2021:AEA

[1370] Rosaria Simone. An accelerated EM algorithm for mixture models with uncertainty for rating data. Computational Statistics, 36 (1):691–714, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01004-z.

Niu:2021:RPS

[1371] Mu Niu, Joe Wandy, and Dirk Husmeier. R package for statistical inference in dynamical systems using kernel based gradient matching: KGode. Computational Statistics, 36 (1):715–747, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01014-x.

Raim:2021:STC

[1372] Andrew M. Raim, Scott H. Holan, and Christopher K. Wikle. Spatiotemporal change of support modeling with R. Computational Statistics, 36 (1):749–780, March 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01029-4.

Fu:2021:EDR

[1373] Liya Fu, Zhuoran Yang, and You-Gan Wang. Efficient and doubly-robust methods for variable selection and parameter estimation in longitudinal data analysis. Computational Statistics, 36(2):781–804, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01038-3.

Guney:2021:REV

[1374] Yesim Güney, Yetkin Tuaç, and Olcay Arslan. Robust estimation and variable selection in heteroscedastic regression model using least favorable distribution. Computational Statistics, 36(2):805–827, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01036-5.

Afzal:2021:VSP

[1375] Arfan Raheen Afzal, Jing Yang, and Xuewen Lu. Variable selection in partially linear additive hazards model with grouped covariates and a diverging number of parameters. Computational Statistics, 36 (2):829–855, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01062-3.

Chen:2021:FSB

[1376] Li-Pang Chen. Feature screening based on distance correlation for ultrahigh-dimensional censored data with covariate measurement error. Computational Statistics, 36 (2):857–884, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01039-2.

Chen:2021:EAJ

[1377] Xiaolin Chen, Catherine Chunling Liu, and Sheng Xu. An efficient algorithm for joint feature screening in ultrahigh-dimensional Cox’s model. Computational Statistics, 36 (2):885–910, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01032-9.

Beyaztas:2021:PLS

[1378] Ufuk Beyaztas and Han Lin Shang. A partial least squares approach for function-on-function interaction regression. Computational Statistics, 36 (2):911–939, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01058-z.

Li:2021:SBT

[1379] Zhengbang Li, Fuxiang Liu, and Guoxin Zuo. A stationary bootstrap test about two mean vectors comparison with somewhat dense differences and fewer sample size than dimension. Computational Statistics, 36(2):941–960, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01030-x.

Pan:2021:ODU

[1380] Yingli Pan, Zhan Liu, and Guangyu Song. Outlier detection under a covariate-adjusted exponential regression model with censored data. Computational Statistics, 36(2):961–976, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-01052-
5.

Cockeran:2021:GFT

[1381] M. Cockeran, S. G. Meintanis, and J. S. Allison. Goodness-of-fit testing of survival models in the presence of Type-II right censoring. Computational Statistics, 36 (2):977–1010, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01050-7.

Sanchez:2021:OTS

[1382] Ainara Rodríguez Sánchez, Román Salmerón Gómez, and Catalina García García. Obtaining a threshold for the Stewart index and its extension to ridge regression. Computational Statistics, 36(2):1011–1029, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01047-2.

Machalova:2021:CSR

[1383] Jitka Machalová, Renáta Talská, and Ales Gába. Compositional splines for representation of density functions. Computational Statistics, 36 (2):1031–1064, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01042-7.

Hoang-Nguyen-Thuy:2021:MCT

[1384] Ngan Hoang-Nguyen-Thuy and K. Krishnamoorthy. A method for computing tolerance intervals for a location-scale family of distributions. Computational Statistics, 36 (2):1065–1092, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01031-w.

Koyuncu:2021:DRM

[1385] Nursel Koyuncu and Derya Karagöz. Designing robust modified R control charts for asymmetric distributions under ranked set and median ranked set sampling. Computational Statistics, 36 (2):1093–1121, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01051-6.

Kharrati-Kopaei:2021:EDL

[1386] Mahmood Kharrati-Kopaei. On the exact distribution of the likelihood ratio test statistic for testing the homogeneity of the scale parameters of several inverse Gaussian distributions. Computational Statistics, 36 (2):1123–1138, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01053-4.

Zhang:2021:CWU

[1387] Yang Zhang, Yu Xiao, and Xin Lu. Comprehensive world university ranking based on ranking aggregation. Computational Statistics, 36 (2):1139–1152, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01033-8.

Soh:2021:ECP

[1388] Patrice Takam Soh, Eugene Kouassi, and Martin Kegnenlezom. Estimation of a CIR process with jumps using a closed form approximation likelihood under a strong approximation of order 1. Computational Statistics, 36 (2):1153–1176, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01040-9.

Bagnato:2021:URO

[1389] Luca Bagnato and Antonio Punzo.

Unconstrained representation of orthogonal matrices with application to common principal components. Computational Statistics, 36 (2):1177–1195, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01041-8.

Zhong:2021:SCB

[1390] Chen Zhong and Lijian Yang. Simultaneous confidence bands for comparing variance functions of two samples based on deterministic designs. Computational Statistics, 36 (2):1197–1218, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01043-6.

Kilic:2021:GAA

[1391] Muhammet Burak Kiliç, Yusuf Sahin, and Melih Burak Koca. Genetic algorithm approach with an adaptive search space based on EM algorithm in two-component mixture Weibull parameter estimation. Computational Statistics, 36 (2):1219–1242, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01044-5.

Burkner:2021:ELO

[1392] Paul-Christian Bürkner, Jonah Gabry, and Aki Vehtari. Efficient leaveone-out cross-validation for Bayesian non-factorized normal and Student-t models. Computational Statistics, 36 (2):1243–1261, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01045-4.

Kelter:2021:ATI

[1393] Riko Kelter. Analysis of type I and II error rates of Bayesian and frequentist parametric and nonparametric two-sample hypothesis tests under preliminary assessment of normality. Computational Statistics, 36 (2):1263–1288, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01034-7.

Tian:2021:BBR

[1394] Yu-Zhu Tian, Man-Lai Tang, and Mao-Zai Tian. Bayesian bridgerandomized penalized quantile regression for ordinal longitudinal data, with application to firm’s bond ratings. Computational Statistics, 36 (2):1289–1319, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01037-4.

Bayati:2021:BAR

[1395] Mahdieh Bayati, Seyed Kamran Ghoreishi, and Jingjing Wu. Bayesian analysis of restricted penalized empirical likelihood. Computational Statistics, 36(2):1321–1339, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01046-3.

Sousa:2021:BWS

[1396] Alex Rodrigo dos S. Sousa, Nancy L.

Garcia, and Brani Vidakovic. Bayesian wavelet shrinkage with beta priors. Computational Statistics, 36 (2):1341–1363, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01048-1.

Kang:2021:BMC

[1397] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Bayesian multiple change-points detection in a normal model with heterogeneous variances. Computational Statistics, 36 (2):1365–1390, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01054-3.

Wu:2021:VSL

[1398] Chien-Wei Wu, Ming-Hung Shu, and Bi-Min Hsu. Variables skiplot sampling plans on the basis of process capability index for products with a low fraction of defectives. Computational Statistics, 36 (2):1391–1413, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01049-0.

Grabchak:2021:TLE

[1399] Michael Grabchak. On the transition laws of p-tempered α-stable OUprocesses. Computational Statistics, 36 (2):1415–1436, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01055-2.

Gooijer:2021:AVM

[1400] Jan G. De Gooijer. Asymmetric vector moving average models: estimation and testing. Computational Statistics, 36(2):1437–1460, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-01056-1.

Cao:2021:EPG

[1401] Chunzheng Cao, Ming He, and Xin Liu. Estimation and prediction of a generalized mixed-effects model with tprocess for longitudinal correlated binary data. Computational Statistics, 36(2):1461–1479, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01057-0.

Hsiao:2021:FDD

[1402] Cheng Hsiao, Yimeng Xie, and Qiankun Zhou. Factor dimension determination for panel interactive effects models: an orthogonal projection approach. Computational Statistics, 36 (2):1481–1497, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01059-y.

Deng:2021:IAM

[1403] Shirong Deng, Jie Chen, and Huidong Shi. Integrative analysis of multiple types of genomic data using an accelerated failure time frailty model. Computational Statistics, 36 (2):1499–1532, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-01060-5.

Dmitrieva:2021:IAB

[1404] Tatiana Dmitrieva, Kristin McCullough, and Nader Ebrahimi. Improved approximate Bayesian computation methods via empirical likelihood. Computational Statistics, 36 (2):1533–1552, June 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00985-1.

Amjadi:2021:EDC

[1405] Roya Amjadi and Wendy Martinez. The 2016 Data Challenge of the American Statistical Association. Computational Statistics, 36(3):

1553–1560, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01076-5.

Alkan:2021:PMM

[1406] Gunes Alkan, Robert Farrow, and Yifan Zhong. Predictive modeling of maximum injury severity and potential economic cost in a car accident based on the General Estimates System data. Computational Statistics, 36(3): 1561–1575, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01074-7.

Auerbach:2021:HBA

[1407] Jonathan Auerbach, Christopher Eshleman, and Rob Trangucci. A hierarchical Bayes approach to adjust for selection bias in beforeafter analyses of vision zero policies. Computational Statistics, 36(3): 1577–1604, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01070-x.

Coyle:2021:ADD

[1408] Patrick Coyle, Chen Chen, and Nooreen Dabbish. Analysis of drowsy driving: exploring subpopulation risk with weighted contingency table tools. Computational Statistics, 36(3):1605– 1620, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01071-w.

Roy:2021:CBT

[1409] Dooti Roy, Ved Deshpande, and M. Henry Linder. A cluster-based taxonomy of bus crashes in the United States. Computational Statistics, 36(3): 1621–1638, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01073-8.

Philips:2021:ACS

[1410] Cody R. Philips, Robert C. Garrett, and Thomas J. Fisher. An analysis of crash-safety ratings and the true assessment of injuries by vehicle. Computational Statistics, 36(3): 1639–1660, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01072-9.

Zhong:2021:BLV

[1411] Wenyan Zhong, Xuewen Lu, and Jingjing Wu. Bi-level variable selection in semiparametric transformation models with right-censored data. Computational Statistics, 36(3): 1661–1692, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01075-6.

Moreira:2021:NED

[1412] Carla Moreira, Jacobo de Uña-Álvarez, and Roel Braekers. Nonparametric estimation of a distribution function from doubly truncated data under dependence. Computational Statistics, 36(3): 1693–1720, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01085-4.

Schwendinger:2021:COS

[1413] Florian Schwendinger, Bettina Grün, and Kurt Hornik. A comparison of optimization solvers for log binomial regression including conic programming. Computational Statistics, 36(3): 1721–1754, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01084-5.

Reschenhofer:2021:FCP

[1414] Erhard Reschenhofer and Manveer K. Mangat. Fast computation and practical use of amplitudes at non-Fourier frequencies. Computational Statistics, 36(3):1755–1773, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-020-01061-
                                        4.

Klaschka:2021:MCI

[1415] Jan Klaschka and Jenő Reiczigel. On matching confidence intervals and tests for some discrete distributions: methodological and computational aspects. Computational Statistics, 36(3): 1775–1790, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00986-0.

Al-Labadi:2021:TSP

[1416] Luai Al-Labadi. The two-sample problem via relative belief ratio.

Computational Statistics, 36(3):1791– 1808, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00988-y.

Gur:2021:EEB

[1417] Sercan Gür and Klaus Pötzelberger.

On the empirical estimator of the boundary in inverse first-exit problems. Computational Statistics, 36(3): 1809–1820, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00989-x.

Kim:2021:BSP

[1418] Seong W. Kim, Sabina Shahin, and Jinheum Kim. Binary segmentation procedures using the bivariate binomial distribution for detecting streakiness in sports data. Computational Statistics, 36(3):1821– 1843, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00992-2.

Jaser:2021:TSR

[1419] Miriam Jaser and Aleksey Min. On tests for symmetry and radial symmetry of bivariate copulas towards testing for ellipticity. Computational Statistics, 36(3):1–26, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00994-0.

Su:2021:CMN

[1420] Chun-Lung Su and Pao sheng Shen. On consistency of the monotone NPMLE of survival function under the mixed case interval-censored model with left truncation. Computational Statistics, 36(3):1871– 1883, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00995-z.

Rolke:2021:CSG

[1421] Wolfgang Rolke and Cristian Gutierrez Gongora. A chi-square goodnessof-fit test for continuous distributions against a known alternative. Computational Statistics, 36(3):1885– 1900, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-020-00997-x.

Su:2021:VAI

[1422] Xiao Su and Yuguo Chen. Variational approximation for importance sampling. Computational Statistics, 36(3): 1901–1930, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01063-w.

Opheim:2021:LMM

[1423] Timothy Opheim and Anuradha Roy. Linear models for multivariate repeated measures data with block exchangeable covariance structure. Computational Statistics, 36(3):1931– 1963, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01064-9.

Elshahhat:2021:BSA

[1424] Ahmed Elshahhat and Mazen Nassar. Bayesian survival analysis for adaptive Type-II progressive hybrid censored Hjorth data. Computational Statistics, 36(3):1965– 1990, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01065-8.

Jin:2021:BSB

[1425] Shiqiang Jin and Gyuhyeong Goh.

Bayesian selection of best subsets via hybrid search. Computational Statistics, 36(3):1991– 2007, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00996-y.

Marcot:2021:WOV

[1426] Bruce G. Marcot and Anca M.

Hanea. What is an optimal value of k in k-fold cross-validation in discrete Bayesian network analysis? Computational Statistics, 36(3):2009– 2031, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-020-00999-9.

Hu:2021:BJQ

[1427] Yingying Hu, Huixia Judy Wang, and Jianhua Guo. Bayesian jointquantile regression. Computational Statistics, 36(3):2033–2053, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-020-00998-
w.

Maturana-Russel:2021:BSD

[1428] Patricio Maturana-Russel and Renate Meyer. Bayesian spectral density estimation using p-splines with quantile-based knot placement. Computational Statistics, 36(3):2055– 2077, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01066-7.

Kang:2021:OBA

[1429] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective Bayesian analysis for generalized exponential stress-strength model. Computational Statistics, 36(3):2079– 2109, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01083-6.

Chang:2021:EPL

[1430] Pei-Chieh Chang, Kim-Hung Pho, and Chin-Shang Li. Estimation of parameters of logistic regression for two-stage randomized response technique. Computational Statistics, 36(3): 2111–2133, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01068-5.

Shao:2021:DGI

[1431] Bin Shao. Decomposition of the Gini index by income source for aggregated data and its applications. Computational Statistics, 36(3):2135– 2159, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01069-4.

Chen:2021:UQE

[1432] Luming Chen and Sujit K. Ghosh. Uncertainty quantification and estimation of closed curves based on noisy data. Computational Statistics, 36(3): 2161–2176, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01077-4.

Bee:2021:TPR

[1433] Marco Bee, Julien Hambuckers, and Luca Trapin. Testing a parameter restriction on the boundary for the g-and-h distribution: a simulated approach. Computational Statistics, 36(3):2177– 2200, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01078-3.

Mahdavi:2021:MLE

[1434] Abbas Mahdavi, Vahid Amirzadeh, and Tsung-I Lin. Maximum likelihood estimation for scale-shape mixtures of flexible generalized skew normal distributions via selection representation. Computational Statistics, 36(3): 2201–2230, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01079-2.

Ferreira:2021:DTR

[1435] Arthur De Sá Ferreira, Ney Meziat-Filho, and Ana Paula Antunes Ferreira. Double threshold receiver operating characteristic plot for three-modal continuous predictors. Computational Statistics, 36(3): 2231–2245, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01080-9.

Alshammri:2021:MDP

[1436] Fayed Alshammri and Jiazhu Pan. Moving dynamic principal component analysis for non-stationary multivariate time series. Computational Statistics, 36(3):2247– 2287, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01081-8.

Miljkovic:2021:NCA

[1437] Tatjana Miljkovic and Ying-Ju Chen. A new computational approach for estimation of the Gini index based on grouped data. Computational Statistics, 36(3):2289–2311, September 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-021-01082-
                                        7.

Robledo:2021:NAF

[1438] Kristy P. Robledo and Ian C. Marschner. A new algorithm for fitting semi-parametric variance regression models. Computational Statistics, 36(4):2313– 2335, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01067-6.

Masci:2021:ECS

[1439] Chiara Masci, Francesca Ieva, and Anna Maria Paganoni. Evaluating class and school effects on the joint student achievements in different subjects: a bivariate semiparametric model with random coefficients. Computational Statistics, 36(4): 2337–2377, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01107-1.

Siebenborn:2021:MPT

[1440] Martin Siebenborn and Julian Wagner. A multigrid preconditioner for tensor product spline smoothing. Computational Statistics, 36(4): 2379–2411, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01104-4.

Qin:2021:PSE

[1441] Fei Qin and Zhangsheng Yu. Penalized spline estimation for panel count data model with time-varying coefficients. Computational Statistics, 36(4): 2413–2434, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01109-z.

Oliveira:2021:AMA

[1442] Rodrigo A. Oliveira and Gilberto A. Paula. Additive models with autoregressive symmetric errors based on penalized regression splines. Computational Statistics, 36(4):2435– 2466, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01106-2.

Xie:2021:MCR

[1443] Yujing Xie and Zhangsheng Yu. Mixture cure rate models with neural network estimated nonparametric components. Computational Statistics, 36(4): 2467–2489, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01086-3.

Arenas:2021:CME

[1444] G. Y. Arenas, J. A. Villaseñor, and F. Tajonar. A computational method for estimating a change point in the Cox hazard model. Computational Statistics, 36(4):2491– 2506, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01087-2.

Novais:2021:CEC

[1445] Luísa Novais and Susana Faria. Comparison of the EM, CEM and SEM algorithms in the estimation of finite mixtures of linear mixed models: a simulation study. Computational Statistics, 36(4):2507– 2533, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01088-1.

Zuo:2021:OSS

[1446] Lulu Zuo, Haixiang Zhang, and Liuquan Sun. Optimal subsample selection for massive logistic regression with distributed data. Computational Statistics, 36(4):2535– 2562, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01089-0.

Santi:2021:RBE

[1447] Flavio Santi, Maria Michela Dickson, and Giuseppe Espa. Reduced-bias estimation of spatial autoregressive models with incompletely geocoded data. Computational Statistics, 36(4): 2563–2590, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01090-7.

Bar-Hen:2021:SCC

[1448] Avner Bar-Hen, Servane Gey, and Jean-Michel Poggi. Spatial CART classification trees. Computational Statistics, 36(4):2591–2613, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01091-
6.

Cavicchioli:2021:SIM

[1449] Maddalena Cavicchioli. Statistical inference for mixture GARCH models with financial application.

Computational Statistics, 36(4):2615– 2642, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01092-5.

Zhang:2021:BAB

[1450] Jiwei Zhang, Zhaoyuan Zhang, and Jian Tao. A Bayesian algorithm based on auxiliary variables for estimating GRM with non-ignorable missing data. Computational Statistics, 36(4): 2643–2669, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01100-8.

Shimamura:2021:BSC

[1451] Kaito Shimamura and Shuichi Kawano. Bayesian sparse convex clustering via global-local shrinkage priors. Computational Statistics, 36(4):2671– 2699, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01101-7.

Vazquez:2021:PMV

[1452] Elsa Vazquez and Jeffrey R. Wilson. Partitioned method of valid moment marginal model with Bayes interval estimates for correlated binary data with time-dependent covariates. Computational Statistics, 36(4): 2701–2718, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01105-3.

Viscardi:2021:WAB

[1453] Cecilia Viscardi, Michele Boreale, and Fabio Corradi. Weighted approximate Bayesian computation via Sanov’s theorem. Computational Statistics, 36(4): 2719–2753, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01093-4.

Hasan:2021:EAP

[1454] Mirza Nazmul Hasan and Roel Braekers. Estimation of the association parameters in hierarchically clustered survival data by nested Archimedean copula functions. Computational Statistics, 36(4):2755– 2787, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01094-3.

Rosenthal:2021:JMC

[1455] Jeffrey S. Rosenthal, Aki Dote, and Ali Sheikholeslami. Jump Markov chains and rejection-free Metropolis algorithms. Computational Statistics, 36(4):2789–2811, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01095-
2.

Chmielowiec:2021:AEF

[1456] Andrzej Chmielowiec. Algorithm for error-free determination of the variance of all contiguous subsequences and fixed-length contiguous subsequences for a sequence of industrial measurement data. Computational Statistics, 36(4):2813– 2840, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01096-1.

Altun:2021:NTP

[1457] Emrah Altun. A new two-parameter discrete Poisson-generalized Lindley distribution with properties and applications to healthcare data sets. Computational Statistics, 36(4):2841– 2861, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01097-0.

Sherlock:2021:DSI

[1458] Chris Sherlock. Direct statistical inference for finite Markov jump processes via the matrix exponential. Computational Statistics, 36(4):2863– 2887, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01102-6.

Lopez-Blazquez:2021:ADM

[1459] Fernando López-Blázquez and Begoña Salamanca-Miño. Automatic differentiation and maximal correlation of order statistics from discrete parents. Computational Statistics, 36(4): 2889–2915, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01103-5.

Parrella:2021:BJP

[1460] Maria Lucia Parrella, Giuseppina Albano, and Michele La Rocca. Bootstrap joint prediction regions for sequences of missing values in spatio-temporal datasets. Computational Statistics, 36(4):2917– 2938, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01099-y.

Meilan-Vila:2021:CVN

[1461] A. Meilán-Vila, R. Fernández-Casal, and M. Francisco-Fernández. A computational validation for nonparametric assessment of spatial trends. Computational Statistics, 36(4):2939– 2965, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01108-0.

Tian:2021:BJI

[1462] Yu-Zhu Tian, Man-Lai Tang, and Mao-Zai Tian. Bayesian joint inference for multivariate quantile regression model with L_{1/2} penalty. Computational Statistics, 36(4):2967– 2994, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01158-4. See cor-

rection [1463].

Tian:2021:CBJ

[1463] Yu-Zhu Tian, Man-Lai Tang, and Mao-Zai Tian. Correction to: Bayesian joint inference for multivariate quantile regression model with L_{1/2} penalty. Computational Statistics, 36 (4):2995, December 2021. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01168-2.  See

[1462].

Zhu:2022:LHT

[1464] Tianming Zhu and Jin-Ting Zhang. Linear hypothesis testing in highdimensional one-way MANOVA: a new normal reference approach. Computational Statistics, 37(1):1–27, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-021-01110-
                                        6.

Shen:2022:NEC

[1465] Pao sheng Shen. Nonparametric estimation for competing risks survival data subject to left truncation and interval censoring. Computational Statistics, 37(1):29–42, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01111-5.

Ballante:2022:NAM

[1466] Elena Ballante, Silvia Figini, and Pierpaolo Uberti. A new approach in model selection for ordinal target variables. Computational Statistics, 37(1):43–56, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01112-4.

Zhang:2022:DIH

[1467] Shuang Zhang and Xingdong Feng.

Distributed identification of heterogeneous treatment effects. Computational Statistics, 37(1):57–89, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01114-
2.

Paradis:2022:RMS

[1468] Emmanuel Paradis. Reduced multidimensional scaling. Computational Statistics, 37(1):91–105, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01116-
0.

Xiaohui:2022:AGR

[1469] Zhou Xiaohui and Gu Guiding. An algorithm of generating random number by wavelet denoising method and its application. Computational Statistics, 37(1):107–124, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01117-z.

Amiri:2022:ERC

[1470] Mehdi Amiri, Yaser Mehrali, and Ahad Jamalizadeh. Efficient recursive computational algorithms for multivariate t and multivariate unified skew-t distributions with applications to inference. Computational Statistics, 37 (1):125–158, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01119-x.

Genc:2022:NDR

[1471] Murat Genç. A new double-regularized regression using Liu and lasso regularization. Computational Statistics, 37(1):159–227, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01120-4.

Ramos-Guajardo:2022:HCM

[1472] Ana Belén Ramos-Guajardo. A hierarchical clustering method for random intervals based on a similarity measure. Computational Statistics, 37 (1):229–261, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01121-3.

Goel:2022:SIT

[1473] Rajni Goel and Hare Krishna. Statistical inference for two Lindley populations under balanced joint progressive type-II censoring scheme. Computational Statistics, 37(1):263–286, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-021-01122-
                                        2.

Li:2022:BGS

[1474] Weibing Li and Thierry Chekouo. Bayesian group selection with non-local priors. Computational Statistics, 37 (1):287–302, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01115-1.

Ariyo:2022:SBM

[1475] Oludare Samuel Ariyo and Matthew Adekunle Adeleke. Simultaneous Bayesian modelling of skew-normal longitudinal measurements with non-ignorable dropout. Computational Statistics, 37 (1):303–325, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01118-y.

Chen:2022:BAP

[1476] Zhiyong Chen and Jianbao Chen.

Bayesian analysis of partially linear, single-index, spatial autoregressive models. Computational Statistics, 37(1):327–353, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01123-1.

Rodrigues:2022:ECR

[1477] Agatha Rodrigues, Pascal Kerschke, and Adriano Polpo. Estimation of component reliability from superposed renewal processes by means of latent variables. Computational Statistics, 37(1):355–379, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01124-0. See cor-

rection [1478].

Rodrigues:2022:CEC

[1478] Agatha Rodrigues, Pascal Kerschke, and Adriano Polpo. Correction to: Estimation of component reliability from superposed renewal processes by means of latent variables. Computational Statistics, 37 (1):381, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01138-8.  See

[1477].

Wu:2022:SRK

[1479] Yi Wu, Wei Yu, and Xuejun Wang. Strong representations of the Kaplan– Meier estimator and hazard estimator with censored widely orthant dependent data. Computational Statistics, 37(1):383–402, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01125-z.

Allison:2022:DFG

[1480] James Allison, Bojana Milosević, and Marius Smuts. Distribution-free goodness-of-fit tests for the Pareto distribution based on a characterization. Computational Statistics, 37 (1):403–418, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01126-y.

Choi:2022:EPB

[1481] Young-Geun Choi, Seunghwan Lee, and Donghyeon Yu. An efficient parallel block coordinate descent algorithm for large-scale precision matrix estimation using graphics processing units. Computational Statistics, 37 (1):419–443, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01127-x.

Ferreira:2022:EDP

[1482] Clécio da Silva Ferreira, Gilberto A.

Paula, and Gustavo C. Lana. Estimation and diagnostic for partially linear models with first-order autoregressive skew-normal errors. Computational Statistics, 37(1):445–468, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01130-
2.

Teimouri:2022:BRP

[1483] Mahdi Teimouri. bccp: an R package for life-testing and survival analysis. Computational Statistics, 37 (1):469–489, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01129-9.

Benedetti:2022:SAB

[1484] Roberto Benedetti, Maria Michela Dickson, and Federica Piersimoni. A simulated annealing-based algorithm for selecting balanced samples. Computational Statistics, 37 (1):491–505, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01113-3.

Huang:2022:GFN

[1485] Jian Huang, Yuling Jiao, and Xiliang Lu. GSDAR: a fast Newton algorithm for ℓ_{0} regularized generalized linear models with statistical guarantee. Computational Statistics, 37 (1):507–533, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01098-z.

AguiareOliveira:2022:DSU

[1486] Hime Aguiar e Oliveira Jr. Deterministic sampling from uniform distributions with Sierpiński space-filling curves. Computational Statistics, 37 (1):535–549, March 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01128-w.

Huang:2022:CSO

[1487] Alan Huang, Lucas Sippel, and Thomas Fung. Consistent second-order discrete kernel smoothing using dispersed Conway–Maxwell–Poisson kernels. Computational Statistics, 37 (2):551–563, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01144-w.

Zhang:2022:AES

[1488] Shibin Zhang. Automatic estimation of spatial spectra via smoothing splines. Computational Statistics, 37(2):565–590, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01141-z.

Fan:2022:SFL

[1489] Ruzong Fan and Hong-Bin Fang. Stochastic functional linear models and Malliavin calculus. Computational Statistics, 37(2):591–611, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01142-
y.

ONeill:2022:CHD

[1490] Ben O’Neill. Computing highest density regions for continuous univariate distributions with known probability functions. Computational Statistics, 37(2):613–649, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01133-z.

Cappello:2022:CAS

[1491] Claudia Cappello, Sandra De Iaco, and Monica Palma. Computational advances for spatio-temporal multivariate environmental models. Computational Statistics, 37(2):651–670, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01132-
0.

Dvorak:2022:GTI

[1492] Jirí Dvorák and Tomás Mrkvicka.

Graphical tests of independence for general distributions. Computational Statistics, 37(2):671–699, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01134-
y.

Greengard:2022:FRS

[1493] Philip Greengard, Andrew Gelman, and Aki Vehtari. A fast regression via SVD and marginalization. Computational Statistics, 37 (2):701–720, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01135-x.

Du:2022:IES

[1494] Yusong Du, Baoying Fan, and Baodian Wei. An improved exact sampling algorithm for the standard normal distribution. Computational Statistics, 37(2):721–737, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01136-w.  See

[1636].

Johnson:2022:PTM

[1495] E. N. Johnson and S. J. Richter. Permutation tests for mixed paired and two-sample designs. Computational Statistics, 37(2):739–750, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01137-
9.

Soutinho:2022:MCM

[1496] Gustavo Soutinho and Luís Meira-Machado. Methods for checking the Markov condition in multi-state survival data. Computational Statistics, 37(2):751–780, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01139-7.

Hasan:2022:MAB

[1497] Mirza Nazmul Hasan and Roel Braekers. Modelling the association in bivariate survival data by using a Bernstein copula. Computational Statistics, 37(2):781–815, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01154-8.

Kosovalic:2022:HEA

[1498] Nemanja Kosovalić and Sandip Barui.

A hard EM algorithm for prediction of the cured fraction in survival data. Computational Statistics, 37 (2):817–835, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01140-0.

Das:2022:SPB

[1499] Kiranmoy Das, Bhuvanesh Pareek, and Pulak Ghosh. A semi-parametric Bayesian dynamic hurdle model with an application to the health and retirement study. Computational Statistics, 37(2):837–863, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01143-x.

Long:2022:MIG

[1500] Quan Long. Multimodal information gain in Bayesian design of experiments. Computational Statistics, 37(2):865–885, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01145-9.

Araya:2022:BIF

[1501] Héctor Araya, Meryem Slaoui, and Soledad Torres. Bayesian inference for fractional oscillating Brownian motion. Computational Statistics, 37 (2):887–907, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01146-8.

Henn:2022:LPT

[1502] L. L. Henn. Limitations and performance of three approaches to Bayesian inference for Gaussian copula regression models of discrete data. Computational Statistics, 37 (2):909–946, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01131-1.

Salvador:2022:BTS

[1503] Sara Salvador and Riccardo Gatto. Bayesian tests of symmetry for the generalized von Mises distribution. Computational Statistics, 37 (2):947–974, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01147-7.

Munezero:2022:EPS

[1504] Parfait Munezero. Efficient particle smoothing for Bayesian inference in dynamic survival models. Computational Statistics, 37 (2):975–994, April 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01155-7.

Bichat:2022:HCP

[1505] Antoine Bichat, Christophe Ambroise, and Mahendra Mariadassou. Hierarchical correction of p-values via an ultrametric tree running Ornstein–Uhlenbeck process. Computational Statistics, 37 (3):995–1013, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01148-6.

Li:2022:KSS

[1506] Jie Li, Jiangyan Wang, and Lijian Yang. Kolmogorov–Smirnov simultaneous confidence bands for time series distribution function. Computational Statistics, 37(3):1015–1039, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01149-
5.

Yazdani:2022:LRC

[1507] Akram Yazdani, Hojjat Zeraati, Mehdi Yaseri, Shahpar Haghighat, and Ahmad Kaviani. Laplace regression with clustered censored data. Computational Statistics, 37 (3):1041–1068, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01151-x.

Sazak:2022:MML

[1508] Hakan Savas Sazak and Melis Zeybek. The modified maximum likelihood estimators for the parameters of the regression model under bivariate median ranked set sampling. Computational Statistics, 37 (3):1069–1109, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01152-w.

Ayyala:2022:CMT

[1509] Deepak Nag Ayyala, Santu Ghosh, and Daniel F. Linder. Covariance matrix testing in high dimension using random projections. Computational Statistics, 37(3):1111–1141, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01166-4.

Ma:2022:SIM

[1510] Xuejun Ma, Shaochen Wang, and Wang Zhou. Statistical inference in massive datasets by empirical likelihood. Computational Statistics, 37 (3):1143–1164, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01153-9.

Signorelli:2022:CSV

[1511] Mirko Signorelli and Luisa Cutillo. On community structure validation in real networks. Computational Statistics, 37 (3):1165–1183, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01156-6.

Lu:2022:NEI

[1512] Feilong Lu and Dehui Wang. A new estimation for INAR(1) process with Poisson distribution. Computational Statistics, 37(3):1185–1201, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01157-
5.

DAmbrosio:2022:DSD

[1513] Federico D’Ambrosio, Hans L. Bodlaender, and Gerard T. Barkema. Dynamic sampling from a discrete probability distribution with a known distribution of rates. Computational Statistics, 37(3):1203–1228, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01159-
3.

ONeill:2022:SCR

[1514] Ben O’Neill. Smallest covering regions and highest density regions for discrete distributions. Computational Statistics, 37(3):1229–1254, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01172-
6.

Conceicao:2022:CGN

[1515] Katiane S. Conceição, Marinho G. Andrade, Francisco Louzada, and Nalini Ravishanker. Characterizations and generalizations of the negative binomial distribution. Computational Statistics, 37(3):1255–1286, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01150-y.

Kang:2022:OBG

[1516] Sang Gil Kang, Woo Dong Lee, and Yongku Kim. Objective Bayesian group variable selection for linear model. Computational Statistics, 37 (3):1287–1310, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01160-w.

Wei:2022:BAA

[1517] Zheng Wei, Daeyoung Kim, and Erin M. Conlon. A Bayesian approach to the analysis of asymmetric association for two-way contingency tables. Computational Statistics, 37 (3):1311–1338, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01161-9.

Dao:2022:BVS

[1518] Mai Dao, Min Wang, Souparno Ghosh, and Keying Ye. Bayesian variable selection and estimation in quantile regression using a quantile-specific prior. Computational Statistics, 37 (3):1339–1368, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-021-01181-5.

McCullough:2022:NAB

[1519] Kristin McCullough, Tatiana Dmitrieva, and Nader Ebrahimi. New approximate Bayesian computation algorithm for censored data. Computational Statistics, 37(3):1369–1397, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01167-
3.

Ravagli:2022:BAM

[1520] Davide Ravagli and Georgi N. Boshnakov. Bayesian analysis of mixture autoregressive models covering the complete parameter space. Computational Statistics, 37(3):1399–1433, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01162-
8.

Fabio:2022:HMR

[1521] Lizandra C. Fabio, Francisco J. A. Cysneiros, Gilberto A. Paula, and Jalmar M. F. Carrasco. Hierarchical and multivariate regression models to fit correlated asymmetric positive continuous outcomes. Computational Statistics, 37(3):1435–1459, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01163-7.

Bruch:2022:ARB

[1522] Christian Bruch. Applying the rescaling bootstrap under imputation for a multistage sampling design. Computational Statistics, 37 (3):1461–1494, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01164-6.

Yang:2022:REN

[1523] Shuquan Yang, Nengxiang Ling, and Yulin Gong. Robust estimation of the number of factors for the pair-elliptical factor models. Computational Statistics, 37 (3):1495–1522, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01165-5.

Scagliarini:2022:STS

[1524] Michele Scagliarini. A sequential test and a sequential sampling plan based on the process capability index C_{pmk}. Computational Statistics, 37 (3):1523–1550, July 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01169-1.

Arslan:2022:MMD

[1525] Talha Arslan, Sukru Acitas, and Birdal Senoglu. Modified minimum distance estimators: definition, properties and applications. Computational Statistics, 37(4):1551– 1568, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01170-8.

Khan:2022:AFT

[1526] Shahedul A. Khan and Nyla Basharat. Accelerated failure time models for recurrent event data analysis and joint modeling. Computational Statistics, 37(4):1569– 1597, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01171-7.

Benlakhdar:2022:SMD

[1527] Said Benlakhdar, Mohammed Rziza, and Rachid Oulad Haj Thami. Statistical modeling of directional data using a robust hierarchical von Mises distribution model: perspectives for wind energy. Computational Statistics, 37(4): 1599–1619, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01173-5.

Amato:2022:PWE

[1528] Umberto Amato, Anestis Antoniadis, Italia De Feis, and Irène Gijbels. Penalized wavelet estimation and robust denoising for irregular spaced data. Computational Statistics, 37(4): 1621–1651, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01174-4.

Dobi:2022:MRP

[1529] Balázs Dobi and András Zempléni.

Markovchart: an R package for costoptimal patient monitoring and treatment using control charts. Computational Statistics, 37(4):1653– 1693, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01175-3.

Sevil:2022:GBE

[1530] Yusuf Can Sevil and Tugba Ozkal Yildiz. Gumbel’s bivariate exponential distribution: estimation of the association parameter using ranked set sampling. Computational Statistics, 37(4):

1695–1726, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01176-2.

Zhou:2022:EIV

[1531] Xiaoshuang Zhou and Peixin Zhao. Estimation and inferences for varying coefficient partially nonlinear quantile models with censoring indicators missing at random. Computational Statistics, 37(4):1727– 1750, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01192-2.

Bothma:2022:NCT

[1532] E. Bothma, J. S. Allison, and I. J. H. Visagie. New classes of tests for the Weibull distribution using Stein’s method in the presence of random right censoring. Computational Statistics, 37(4):1751– 1770, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01178-0.

Bee:2022:TGH

[1533] Marco Bee. The truncated g-and-h distribution: estimation and application to loss modeling. Computational Statistics, 37(4):1771– 1794, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01179-z.

Saini:2022:RIM

[1534] Shubham Saini and Renu Garg.

Reliability inference for multicomponent stress-strength model from Kumaraswamy-G family of distributions based on progressively first failure censored samples. Computational Statistics, 37(4):1795– 1837, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01180-6.

Ferreira:2022:MUW

[1535] José A. Ferreira. Models under which random forests perform badly; consequences for applications. Computational Statistics, 37(4):1839– 1854, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01182-4.

Feng:2022:RMC

[1536] Kexin Feng, Lawrence M. Leemis, and Heather Sasinowska. RMSEminimizing confidence intervals for the binomial parameter. Computational Statistics, 37(4):1855– 1885, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01183-3.

Wang:2022:UMV

[1537] Liming Wang, Xingxiang Li, Xiaoqing Wang, and Peng Lai. Unified mean-variance feature screening for ultrahigh-dimensional regression.

Computational Statistics, 37(4):1887– 1918, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01184-2.

Baek:2022:CEA

[1538] Seungchul Baek and Junyong Park. A computationally efficient approach to estimating species richness and rarefaction curve. Computational Statistics, 37(4):1919–1941, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01185-
1.

Tavassoli:2022:HMI

[1539] A. Tavassoli, Y. Waghei, and A. Nazemi. Hybrid MLP–IDW approach based on nearest neighbor for spatial prediction. Computational Statistics, 37(4): 1943–1962, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01186-0.

Guzel:2022:NNA

[1540] Ismail Güzel and Atabey Kaygun. A new non-Archimedean metric on persistent homology. Computational Statistics, 37(4):1963– 1983, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01187-z.

Gonzalez-Estrada:2022:SWT

[1541] Elizabeth González-Estrada, José A. Villaseñor, and Rocío Acosta-Pech.

Shapiro–Wilk test for multivariate skew-normality. Computational Statistics, 37(4):1985–2001, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-021-01188-
y.

Silva:2022:BER

[1542] Wagner J. F. Silva, Renata M.

C. R. Souza, and F. J. A. Cysneiros. Bivariate elliptical regression for modeling interval-valued data. Computational Statistics, 37(4):2003– 2028, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01189-x.

Allerbo:2022:FNP

[1543] Oskar Allerbo and Rebecka Jörnsten.

Flexible, non-parametric modeling using regularized neural networks. Computational Statistics, 37(4):2029– 2047, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01190-4.

Ahmadi:2022:PTF

[1544] J. Ahmadi, B. Khatib Astaneh, M. Rezaie, and S. Ameli. Prediction of times to failure of censored units under generalized progressive hybrid censoring scheme. Computational Statistics, 37(4):2049– 2086, September 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-021-01191-3.

Garner:2022:EDC

[1545] Thesia I. Garner and Wendy Martinez. The 2017 Data Challenge of the American Statistical Association. Computational Statistics, 37(5): 2087–2094, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01257-w.

Meechai:2022:HEE

[1546] Joyance Meechai and Manel Wijesinha. Household energy expenditure and consumption patterns in the United States. Computational Statistics, 37(5): 2095–2127, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01255-y.

Hu:2022:MUI

[1547] Mingzhao Hu. Multivariate understanding of income and expenditure in United States households with statistical learning. Computational Statistics, 37(5):2129– 2160, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01251-2.

Sulc:2022:NRP

[1548] Zdenek Sulc, Jana Cibulkova, and Hana Rezankova. Nomclust 2.0: an R package for hierarchical clustering of objects characterized by nominal variables. Computational Statistics, 37(5): 2161–2184, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01209-4.

Cai:2022:NIS

[1549] Mingyang Cai and Gerko Vink.

A note on imputing squares via polynomial combination approach. Computational Statistics, 37(5):2185– 2201, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01194-8.

Bollwein:2022:ODT

[1550] Ferdinand Bollwein and Stephan Westphal. Oblique decision tree induction by cross-entropy optimization based on the von Mises–Fisher distribution. Computational Statistics, 37(5): 2203–2229, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01195-7.

Yu:2022:TST

[1551] Xinyang Yu, Cheng Wang, Zhongqing Yang, and Binyan Jiang. Tuning selection for two-scale kernel density estimators. Computational Statistics, 37(5): 2231–2247, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01196-6.

Pakyari:2022:GFT

[1552] Reza Pakyari and Ayman Baklizi. On goodness-of-fit testing for Burr type X distribution under progressively type-II censoring.

Computational Statistics, 37(5):2249– 2265, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01197-5.

Taconeli:2022:ICI

[1553] Cesar Augusto Taconeli and Idemauro Antonio Rodrigues de Lara. Improved confidence intervals based on ranked set sampling designs within a parametric bootstrap approach. Computational Statistics, 37(5):2267– 2293, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01198-4.

Zhang:2022:ASL

[1554] Boyao Zhang, Tobias Hepp, Sonja Greven, and Elisabeth Bergherr. Adaptive step-length selection in gradient boosting for Gaussian location and scale models. Computational Statistics, 37(5):2295– 2332, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01199-3.

Abdallah:2022:EDF

[1555] Mohamed S. Abdallah, Amer I. Al-Omari, Naif Alotaibi, Ghadah A. Alomani, and A. S. Al-Moisheer. Estimation of distribution function using L ranked set sampling and robust extreme ranked set sampling with application to reliability. Computational Statistics, 37(5): 2333–2362, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01201-y.

Loe:2022:EUC

[1556] Margrethe Kvale Loe and Håkon Tjelmeland. Ensemble updating of categorical state vectors. Computational Statistics, 37(5):2363– 2397, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01202-x.

Hua:2022:ICB

[1557] Rui Hua and Wenhao Gui. Inference for copula-based dependent competing risks model with stepstress accelerated life test under generalized progressive hybrid censoring. Computational Statistics, 37(5): 2399–2436, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01203-w.

Hanafi:2022:OCP

[1558] Mohamed Hanafi, Zouhair El Hadri, Abderrahim Sahli, and Pasquale Dolce. Overcoming convergence problems in PLS path modelling. Computational Statistics, 37(5):2437– 2470, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01204-9.

Clairon:2022:OCP

[1559] Quentin Clairon and Adeline Samson. Optimal control for parameter estimation in partially observed hypoelliptic stochastic differential equations. Computational Statistics, 37(5): 2471–2491, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01212-9.

Lee:2022:VIL

[1560] Youngseon Lee, Seongil Jo, and Jaeyong Lee. A variational inference for the Lévy adaptive regression with multiple kernels. Computational Statistics, 37(5):2493– 2515, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01200-z.

Sabino:2022:FST

[1561] Piergiacomo Sabino and Nicola Cufaro Petroni. Fast simulation of tempered stable Ornstein–Uhlenbeck processes. Computational Statistics, 37(5): 2517–2551, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01205-8.

Frascolla:2022:TST

[1562] Cindy Frascolla, Guillaume Lecuelle, Pascal Schlich, and Hervé Cardot. Two sample tests for semi-Markov processes with parametric sojourn time distributions: an application in sensory analysis. Computational Statistics, 37(5): 2553–2580, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01210-x.

Gogebakan:2022:NPS

[1563] Kemal Çaglar Gögebakan and Burak Alparslan Eroglu. Non-parametric seasonal unit root tests under periodic non-stationary volatility. Computational Statistics, 37(5):2581– 2636, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01211-w.

Lu:2022:ISC

[1564] Jun Lu, Dan Wang, and Qinqin Hu. Interaction screening via canonical correlation. Computational Statistics, 37(5): 2637–2670, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01206-7.

Pargent:2022:RTE

[1565] Florian Pargent, Florian Pfisterer, Janek Thomas, and Bernd Bischl. Regularized target encoding outperforms traditional methods in supervised machine learning with high cardinality features. Computational Statistics, 37(5):2671– 2692, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01207-6.

Lee:2022:RSB

[1566] Jaejoon Lee, Seongil Jo, and Jaeyong Lee. Robust sparse Bayesian infinite factor models. Computational Statistics, 37(5):2693–2715, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01208-
                                        5.

Luo:2022:UDP

[1567] Biao Luo, Hongyi Li, Yingying Wei, and Zujun Ou. Uniform design with prior information of factors under weighted wrap-around (L_{2})- discrepancy. Computational Statistics, 37(5):2717–2739, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01193-
                                        9.

Michimae:2022:BRE

[1568] Hirofumi Michimae and Takeshi Emura. Bayesian ridge estimators based on copula-based joint prior distributions for regression coefficients. Computational Statistics, 37(5):2741– 2769, November 2022. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01213-8.

Ozenne:2023:CFE

[1569] Brice Ozenne, Esben Budtz-Jørgensen, and Sebastian Elgaard Ebert. Controlling the familywise error rate when performing multiple comparisons in a linear latent variable model. Computational Statistics, 38(1):1–23, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01214-7.

Phamtoan:2023:FCA

[1570] Dinh Phamtoan and Tai Vovan.

The fuzzy cluster analysis for interval value using genetic algorithm and its application in image recognition. Computational Statistics, 38 (1):25–51, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01215-6.

Yoshikawa:2023:SRR

[1571] Kohei Yoshikawa and Shuichi Kawano.

Sparse reduced-rank regression for simultaneous rank and variable selection via manifold optimization. Computational Statistics, 38(1):53–75, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01216-

5. See correction [1572].

Yoshikawa:2023:CSR

[1572] Kohei Yoshikawa and Shuichi Kawano.

Correction: Sparse reduced-rank regression for simultaneous rank and variable selection via manifold optimization. Computational Statistics, 38(1):77–78, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01274-9.  See

[1571].

Xu:2023:ESF

[1573] Tengteng Xu, Riquan Zhang, and Xiuzhen Zhang. Estimation of spatial-functional based-line logit model for multivariate longitudinal data. Computational Statistics, 38 (1):79–99, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01217-4.

Langthaler:2023:PTT

[1574] Patrick B. Langthaler, Riccardo Ceccato, Luigi Salmaso, Rosa Arboretti, and Arne C. Bathke. Permutation testing for thick data when the number of variables is much greater than the sample size: recent developments and some recommendations. Computational Statistics, 38 (1):101–132, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01218-3.

Shiu:2023:ECT

[1575] Daniel Shiu. Efficient computation of tight approximations to Chernoff bounds. Computational Statistics, 38 (1):133–147, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01219-2.

Annas:2023:SSA

[1576] Monir El Annas, Badreddine Benyacoub, and Mohamed Ouzineb. Semisupervised adapted HMMs for P2P credit scoring systems with reject inference. Computational Statistics, 38 (1):149–169, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01220-9.

Nguyen:2023:SPO

[1577] Thuan Nguyen and Jiming Jiang.

Simplified partially observed quasiinformation matrix. Computational Statistics, 38(1):171–189, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01221-
8.

Centofanti:2023:ASS

[1578] Fabio Centofanti, Antonio Lepore, Alessandra Menafoglio, Biagio Palumbo, and Simone Vantini. Adaptive smoothing spline estimator for the function-on-function linear regression model. Computational Statistics, 38 (1):191–216, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01223-6.

Lee:2023:EGP

[1579] Seunghwan Lee, Sang Cheol Kim, and Donghyeon Yu. An efficient GPU-parallel coordinate descent algorithm for sparse precision matrix estimation via scaled lasso. Computational Statistics, 38 (1):217–242, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01224-5.

Kuroda:2023:SCA

[1580] Masahiro Kuroda, Yuichi Mori, and Masaya Iizuka. Speeding up the convergence of the alternating least squares algorithm using vector ε acceleration and restarting for nonlinear principal component analysis. Computational Statistics, 38 (1):243–262, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01225-4.

Qi:2023:KRC

[1581] Xiaomeng Qi and Zhangsheng Yu. Kernel regression for cause-specific hazard models with time-dependent coefficients. Computational Statistics, 38 (1):263–283, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01227-2.

Alcantara:2023:MSU

[1582] Ida Marie Alcantara, Joshua Naranjo, and Yanda Lang. Model selection using PRESS statistic. Computational Statistics, 38(1):285–298, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01228-
                                        1.

Zhang:2023:RTE

[1583] Rui Zhang, Guoyou Qin, and Dongsheng Tu. A robust threshold t linear mixed model for subgroup identification using multivariate T distributions. Computational Statistics, 38 (1):299–326, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01229-0.

Sundqvist:2023:AAR

[1584] Martina Sundqvist, Julien Chiquet, and Guillem Rigaill. Adjusting the adjusted Rand Index. Computational Statistics, 38(1):327–347, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01230-
7.

Pavone:2023:URM

[1585] Federico Pavone, Juho Piironen, Paul-Christian Bürkner, and Aki Vehtari. Using reference models in variable selection. Computational Statistics, 38 (1):349–371, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01231-6.

Fang:2023:PME

[1586] Xiaoqiong Fang, Andy W. Chen, and Derek S. Young. Predictors with measurement error in mixtures of polynomial regressions. Computational Statistics, 38(1):373–401, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01232-
5.

Cancho:2023:SRC

[1587] Vicente G. Cancho, Elizbeth C. Bedia, Gauss M. Cordeiro, Fábio Prataviera, Edwin M. M. Ortega, and Ana P. J. E. Santo. A survival regression with cure fraction applied to cervical cancer. Computational Statistics, 38(1):403–418, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01233-4.

Maturo:2023:SCC

[1588] Fabrizio Maturo and Rosanna Verde. Supervised classification of curves via a combined use of functional data analysis and tree-based methods. Computational Statistics, 38 (1):419–459, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01236-1.

Hocking:2023:LOP

[1589] Toby Dylan Hocking and Anuraag Srivastava. Labeled optimal partitioning. Computational Statistics, 38 (1):461–480, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01238-z.

Dey:2023:APT

[1590] Sanku Dey, Ahmed Elshahhat, and Mazen Nassar. Analysis of progressive type-II censored gamma distribution. Computational Statistics, 38 (1):481–508, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01239-y.

He:2023:MAD

[1591] Baihua He, Yanyan Liu, Guosheng Yin, and Yuanshan Wu. Model aggregation for doubly divided data with large size and large dimension. Computational Statistics, 38 (1):509–529, March 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01242-3.

Karabatsos:2023:ABC

[1592] George Karabatsos. Approximate Bayesian computation using asymptotically normal point estimates. Computational Statistics, 38(2):531–568, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01226-
3.

Teimouri:2023:FBI

[1593] Mahdi Teimouri. Fast Bayesian inference for Birnbaum–Saunders distribution. Computational Statistics, 38(2):569–601, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01234-3.

Garcia:2023:BSQ

[1594] Carlos García, Zaida Quiroz, and Marcos Prates. Bayesian spatial quantile modeling applied to the incidence of extreme poverty in Lima, Peru. Computational Statistics, 38 (2):603–621, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01235-2.

Brisco:2023:BFB

[1595] Agnese Maria Di Brisco, Enea Giuseppe Bongiorno, Aldo Goia, and Sonia Migliorati. Bayesian flexible beta regression model with functional covariate. Computational Statistics, 38 (2):623–645, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01240-5.

Weisser:2023:PDS

[1596] Christoph Weisser, Christoph Gerloff, Anton Thielmann, Andre Python, Arik Reuter, Thomas Kneib, and Benjamin Säfken. Pseudo-document simulation for comparing LDA, GS-DMM and GPM topic models on short and sparse text using Twitter data. Computational Statistics, 38(2):647–674, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01246-z.

Jokiel-Rokita:2023:BEV

[1597] Alicja Jokiel-Rokita and Ryszard Magiera. Bayesian estimation versus maximum likelihood estimation in the Weibull-power law process. Computational Statistics, 38 (2):675–710, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01241-4.

Kumar:2023:BEC

[1598] Pushkal Kumar, Manas Ranjan Tripathy, and Somesh Kumar. Bayesian estimation and classification for two logistic populations with a common location. Computational Statistics, 38(2):711–748, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01247-y.

Zhu:2023:SLA

[1599] Wanchuang Zhu and Yanan Fan.

A synthetic likelihood approach for intractable Markov random fields. Computational Statistics, 38(2):749– 777, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https://link.

springer.com/article/10.1007/s00180-
022-01256-x.

Salam:2023:MAS

[1600] Abdul Salam and Marco Grzegorczyk. Model averaging for sparse seemingly unrelated regression using Bayesian networks among the errors. Computational Statistics, 38 (2):779–808, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01258-9.

Sen:2023:VSC

[1601] Sweata Sen, Damitri Kundu, and Kiranmoy Das. Variable selection for categorical response: a comparative study. Computational Statistics, 38 (2):809–826, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01260-1.

Mai:2023:EAM

[1602] The Tien Mai. An efficient adaptive MCMC algorithm for pseudo-Bayesian quantum tomography. Computational Statistics, 38(2):827–843, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01264-
                                        x.

Hu:2023:BAS

[1603] Guanyu Hu, Ming-Hui Chen, and Nalini Ravishanker. Bayesian analysis of spherically parameterized dynamic multivariate stochastic volatility models. Computational Statistics, 38(2):845–869, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01266-9.

Li:2023:GTS

[1604] Peili Li, Min Liu, and Zhou Yu. A global two-stage algorithm for nonconvex penalized high-dimensional linear regression problems. Computational Statistics, 38(2):871–898, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01249-
                                        w.

Lee:2023:ELR

[1605] Shen-Ming Lee, Truong-Nhat Le, Phuoc-Loc Tran, and Chin-Shang Li. Estimation of logistic regression with covariates missing separately or simultaneously via multiple imputation methods. Computational Statistics, 38(2):899–934, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01250-3.

Liu:2023:LSC

[1606] Yu Liu and Chin-Shang Li. A linear spline Cox cure model with its applications. Computational Statistics, 38(2):935–954, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01252-1.

Chen:2023:MMP

[1607] Zezhun Chen, Angelos Dassios, and George Tzougas. Multivariate mixed Poisson generalized inverse Gaussian INAR(1) regression. Computational Statistics, 38(2):955–977, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01253-
0.

Liu:2023:SPA

[1608] Yuyang Liu, Pengfei Pi, and Shan Luo. A semi-parametric approach to feature selection in high-dimensional linear regression models. Computational Statistics, 38(2):979–1000, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01254-
z.

Kim:2023:SQR

[1609] Kyu Hyun Kim, Daniel J. Caplan, and Sangwook Kang. Smoothed quantile regression for censored residual life. Computational Statistics, 38 (2):1001–1022, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01262-z.

Khodamoradi:2023:EMC

[1610] Tahereh Khodamoradi and Maziar Salahi. Extended mean-conditional value-at-risk portfolio optimization with PADM and conditional scenario reduction technique. Computational Statistics, 38(2):1023–1040, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-022-01263-
                                        y.

Huang:2023:IFB

[1611] Zhonglu Huang and Gengsheng Qin. Influence function-based confidence intervals for the Kendall rank correlation coefficient. Computational Statistics, 38(2):1041–1055, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01267-8.

Zhang:2023:DTD

[1612] Lu ning Zhang, Jian wei Liu, and Xin Zuo. Doubly timedependent Hawkes process and applications in failure sequence analysis. Computational Statistics, 38 (2):1057–1093, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-022-01269-6.

Guo:2023:SOP

[1613] Guangbao Guo, Chunjie Wei, and Guoqi Qian. Sparse online principal component analysis for parameter estimation in factor model. Computational Statistics, 38(2):1095–1116, June 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01270-
z.

Cetinkaya-Rundel:2023:EDC

[1614] Mine Cetinkaya-Rundel and Wendy Martinez. The 2018 Data Challenge Expo of the American Statistical Association. Computational Statistics, 38(3):1117–1122, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-023-01363-
3.

Keeley:2023:APF

[1615] Darren Keeley and Eric A. Suess.

Accuracy of precipitation forecasts: finding the right threshold for what is considered rain. Computational Statistics, 38(3):1123–1134, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-023-01337-
5.

Lundell:2023:LTA

[1616] Jill F. Lundell, Brennan Bean, and Jürgen Symanzik. Let’s talk about the weather: a cluster-based approach to weather forecast accuracy. Computational Statistics, 38(3):1135– 1155, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-023-01339-3.

Schweitzer:2023:SEW

[1617] Benjamin Schweitzer, Robert C. Garrett, Nichole Rook, and Thomas J. Fisher. A spatial extension of weather forecasts. Computational Statistics, 38(3):1157–1171, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

                                        URL https://link.springer.com/
                                        article/10.1007/s00180-023-01336-
                                        6.

Roy:2023:ENW

[1618] Dooti Roy, Gregory Vaughan, Jianan Hui, and Junxian Geng. An exploration of National Weather Service daily forecasts using R Shiny. Computational Statistics, 38(3):1173– 1191, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-023-01341-9.

Lin:2023:ULU

[1619] Chuyuan Lin, Ying Yu, Lucas Y. Wu, and Jiguo Cao. Unsupervised learning on U.S. weather forecast performance. Computational Statistics, 38(3):1193– 1213, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

                                        //link.springer.com/article/10.
                                        1007/s00180-023-01340-w.

Jang:2023:SCW

[1620] Phillip A. Jang and David S. Matteson. Spatial correlation in weather forecast accuracy: a functional time series approach. Computational Statistics, 38(3):1215– 1229, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-023-01338-4.

Hernandez:2023:GGF

[1621] Alberto J. Hernández and Maikol Solís.

Geometric goodness of fit measure to detect patterns in data point clouds. Computational Statistics, 38(3):1231– 1253, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01244-1.

Yan:2023:CDN

[1622] Yanan Yan and Yuehan Yang. Community detection for New York stock market by SCORE–CCD. Computational Statistics, 38(3):1255– 1282, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01245-0.

Amini:2023:HRP

[1623] Morteza Amini, Afarin Bayat, and Reza Salehian. hhsmm: an R package for hidden hybrid Markov/semi-Markov models. Computational Statistics, 38(3):1283–1335, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01248-
x.

Zhang:2023:SPM

[1624] Ning Zhang and Jin Yang. Sparse precision matrix estimation with missing observations. Computational Statistics, 38(3):1337–1355, September 2023.

CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01265-
w.

Riva-Palacio:2023:EPO

[1625] Alan Riva-Palacio, Ramsés H. Mena, and Stephen G. Walker. On the estimation of partially observed continuous-time Markov chains. Computational Statistics, 38(3):1357– 1389, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01273-w.

Brusa:2023:TEM

[1626] Luca Brusa, Francesco Bartolucci, and Fulvia Pennoni. Tempered expectation-maximization algorithm for the estimation of discrete latent variable models. Computational Statistics, 38(3):1391–1424, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01276-

7. See correction [1627].

Brusa:2023:CTE

[1627] Luca Brusa, Francesco Bartolucci, and Fulvia Pennoni. Correction to: Tempered expectation-maximization algorithm for the estimation of discrete latent variable models. Computational Statistics, 38(3):1425, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01321-

5. See [1626].

Gillard:2023:PWH

[1628] Jonathan Gillard, Emily O’Riordan, and Anatoly Zhigljavsky. Polynomial whitening for high-dimensional data. Computational Statistics, 38(3): 1427–1461, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01277-6.

Miranda-Soberanis:2023:TPL

[1629] V. F. Miranda-Soberanis and Thomas W.

Yee. Two-parameter link functions, with applications to negative binomial, Weibull and quantile regression. Computational Statistics, 38(3): 1463–1485, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01279-4.

Pandolfo:2023:CDD

[1630] Giuseppe Pandolfo and Antonio D’Ambrosio. Clustering directional data through depth functions. Computational Statistics, 38(3):1487– 1506, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01281-w.

Elias:2023:DBR

[1631] Antonio Elías, Raúl Jiménez, and Han Lin Shang. Depth-based reconstruction method for incomplete functional data. Computational Statistics, 38(3):1507–1535, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic).

URL https://link.springer.com/
article/10.1007/s00180-022-01282-
9.

Basol:2023:CDP

[1632] Merve Basol, Dincer Goksuluk, and Ergun Karaagaoglu. Comparing the diagnostic performance of methods used in a full-factorial design multi-reader multi-case studies. Computational Statistics, 38(3):1537– 1553, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01309-1.

Aldeni:2023:PSE

[1633] Mahmoud Aldeni, John Wagaman, Mohamed Amezziane, and S. Ejaz Ahmed. Pretest and shrinkage estimators for log-normal means. Computational Statistics, 38(3):1555– 1578, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01286-5.

Numminen:2023:QLP

[1634] Riikka Numminen, Ileana Montoya Perez, Ivan Jambor, Tapio Pahikkala, and Antti Airola. Quicksort leave-pair-out cross-validation for ROC curve analysis. Computational Statistics, 38(3):1579– 1595, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01288-3.

Gallaugher:2023:PWC

[1635] M. P. B. Gallaugher, C. Biernacki, and P. D. McNicholas. Parameterwise co-clustering for high-dimensional data. Computational Statistics, 38(3): 1597–1619, September 2023. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL https:

//link.springer.com/article/10.
1007/s00180-022-01289-2.

Karney:2016:SEN

[1636] Charles F. F. Karney. Sampling exactly from the normal distribution. ACM Transactions on Mathematical Software, 42(1):3:1–3:14, February 2016. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See improvement in [?].